A Complete Bibliography of Publications in the Journal of Mathematical Physics: 2000–2004

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.12

Title word cross-reference

(1 + 1) [278, 1699, 1816, 365]. (2 + 1) [271, 507, 33, 487, 1110, 1492]. (2m, 2n) [1307]. (2N) [1932]. (3/2) [591]. (5) [777]. (: ϕ^{4}: )_{2} [1788]. (d^{2}X/dx^{2}) + λ^{2}X = 0 [1031]. (λ/4!)(ϕ_{1}^{4} + ϕ_{2}^{4})_{d} [813]. (M + N) [1110]. (n) [518]. (N + 1) [1932]. (n ≥ 2) [413]. (p, q) [1744]. (q, h) [573]. 1 + 1 [1100, 759]. 1/2 [699, 642]. 1/N [152]. 1/r [1160]. 1: 1: 2 [1943]. 2 [936, 1777, 1669, 1733, 1636]. 2 + 1 [714, 902, 1290, 906, 381, 759, 926, 795, 124, 1269, 564]. 27 ⊗ \overline{27} [277]. 2N [1317]. 3 [1786]. 3 + 1 [297]. 3nj [1094]. 5 [1091, 1278, 1717]. 78 ⊗ 78 [479]. _{−−} [1169]. _{2} [255]. _{2}F_{1}(a, b; c; x) [827, 611]. _{N} [889]. _{q} [97, 1300, 388]. A [676, 985]. A_{M}^{(1)} [509]. A_{n}^{q} [1118]. A_{4} [1020]. A_{∞} [1858]. A_{n−1} [1753, 1657]. A_{n−1}^{(1)} [1898]. AB [83]. ADE [1482]. AdS_{3} [677]. AdS_{3} × S^{3} [679]. AdS_{7/4}/CFT_{6/3} [680]. B_{q,λ}(A_{2}^{(2)}) [1821]. ¯∂ [37, 1093, 365]. bc [148, 770]. BC_{n} [476, 792]. β [1537]. BF [240]. c = 1 [1637]. C^{∗} [1387, 143, 1806]. C^{n} [189, 405]. C_{2} [1233]. C_{n} [476, 792]. C^{2} [455]. : ϕ_{4}^{4}: [58]. CP^{2} [1063]. CP^{d} + 1 [490]. CP^{N} [1275, 1301, 1420]. CP^{N−1} [1454]. D [894, 669, 1291, 1664, 1035, 1479, 589, 272, 545, 469, 386, 1390, 1180, 676, 1829, 490]. D + 1 [1516, 425].

D < 2 [613]. d = 2 + 1 [1378]. D = 4 [685]. D = 5 [389]. d^{2} = 0 [1771]. d^{3} = 0 [1771]. D_{4} [1547]. δ [1844, 16, 104, 496, 791, 1235, 1255, 1231]. Δ = −1/2 [1113]. δ′ [103, 496]. d ≥ 4 [316]. E(1, 1) [1813]. E^{3} [1901]. E_{6} [277, 479]. E_{7} [1020]. E_{8}^{(1)} [1336]. E_{κ}(2) [195]. E_{q}(2) [98]. ϵ [82]. η [3, 637]. F [258, 1183]. F_{1} [1339]. F_{2} [1339]. F_{3} [1339]. F_{4} [1443]. G [1752, 1406, 452, 1529]. G_{2} [136, 1758]. G_{r,s} [145]. gc_{N} [1654]. gl [383]. GL(2) ⊗ GL(1) [831]. GL_{h}(1|1) [575]. h(1) ⊕ su(2) [989]. H^{2} [891, 1384]. J [937, 338, 444, 962, 830]. K [1565, 982, 621]. k = −1 [514]. K′(4) [525]. κ [1404]. l [1026]. L^{2} [341, 34]. L^{p} [935]. l_{p} [943]. λ [1874]. λ_{l} [1645]. ≤ 4 [1918]. LS [844]. M [1613, 691, 1893, 317, 1428]. M + 1 [1893]. M_{4} × X_{2} [353]. CP^{n} [1049]. R [1785]. R^{2d} [1361]. R^{3} [1594, 1785]. R^{4} [1519]. R^{n} [34]. T^{4}/Z_{2} [659]. Z^{d} [1696, 82]. C^{∈} [554]. F_{√}(η) = ∞/−(√ + ∞) ∫_{′}^{∞}[ϵ^{√}/(∞ + ⌉^{ϵ−η})] ⌈ϵ [605]. N

[759, 1072, 1367]. N = △ [1156]. P(ϕ)_{∈} [59]. U_{⨿}(A_{∈\−∞}) [138]. W_{∞} [1217]. ⌊(∞) [1619]. SU(∞, ∞) [999]. SU(N) [988]. osp(2 /2 ) [140]. SL(2 , R) [22]. sl(n, C) [208]. SO(2n) [286, 327]. SO(3 ) [384]. Sp(2n, R) [286]. Sp(2n, R) [327]. SU (1 , 1 ) [41]. SU (2 ) [256, 24, 384, 404]. SU (3 ) [8, 384]. SU (N ) [19]. U_{q}[sl(ˆ2|)] [304]. GL_{p,q}(1|1) [408]. SL(2, c) [44]. SL(2, c)/GL(1, c) [44]. SL(3, C) [280]. SU_{μ}(2) [195]. U(1) [262]. U(∞) [71]. U_{pq}[gl(2/2)] [380]. N = 1 [669]. N = 4 [674]. PT [627]. PT [648]. GL_{p,q}(2) [1360]. o(6, C) [1253]. osp(1|2) [782, 1367]. osp(N, M) [769]. sl(2, R) [984]. sl(4, C) [1253]. so(M) [802]. SO(N) [1046]. SO(N, 1) [1256]. SO_{q}(3) [963]. Sp(6, R) [1413]. su(1, 1) [1094]. SU(2) [997, 1491]. SU(N) [977]. su(n + 1) [1069]. UOSp(k_{1}/2k_{2}) [1506]. U′(so_{4}) [829]. N [1335, 970, 888, 31, 1343, 691, 1645, 1300, 1606, 450, 675, 942, 1421, 1818, 1, 1256, 1250, 1369, 214, 1042, 701, 1193, 571, 820, 1234, 1425, 751, 1046, 949, 668, 511, 1217, 1781, 601, 1535, 1450, 658, 1094, 786, 1065, 947, 1553, 713, 1229, 789, 388, 1069]. N = 2 [42, 710, 255, 352, 256, 127]. N = 2, 3, 4 [547]. N = 3 [1250, 1369]. nj [190]. O(4) [954]. o(4, C) [1764]. o(5, C) [729]. O^{∗} [1059]. ω [1864]. osp(1|2) [1295]. OSp(3, 1|2) [250]. OSp_{h}(2/1) [1724]. P [1874, 10, 371, 40, 835]. p_{1} [316]. p_{2} [316]. P_{N} [64]. ⊥ [609]. ϕ [243]. ϕ^{4} [497]. ϕ_{2}^{4} [1632]. P T [877, 1029, 1102, 1314, 1906]. q [640, 626, 449, 1861, 829, 578, 1902, 46, 801, 431, 333]. QED_{4} [693]. R [1272, 634, 663, 1579, 35, 140, 1296, 129]. R^{8} [1063]. S [1026]. S(g) [1788]. S^{1} × R^{+} [153]. S^{2} [1460, 891, 1384]. S^{2} ⟶ Gr(2, N) [977]. S_{3} [676]. S_{f} ⊃ S_{f−1} [1248]. sech^{2} [1779]. sl [449]. sl(2) [923]. SL(2, R) [1744, 1636]. SL(2, R) [677, 676]. sl(m/n) [48, 289, 832]. sl(n + 1) [1095]. sl(n, C) [932, 13]. sl_{2} [626]. sl_{n} [1568]. so(2, 1) [1795, 1947]. SO(2, 2) [526, 1704]. SO(3) [694, 1223, 1799]. so(4) [661]. SO(5) [1799]. SO(8) [1443]. SO(9) [1443]. SO(N) [949]. SO(N, 1) [1]. SO_{o}(2p, 2) [47]. sp(4, C) [729]. ⋆ [972]. SU(1, 1) [1701]. SU(1|3) [1610]. SU(2) [556, 1811, 430, 570, 1639, 636, 444, 962]. SU(2n) [1708]. SU(3) [73, 1189, 1190, 556, 1811, 750, 636, 1410, 1837]. SU(3|1) [1610]. SU(4) [1231]. SU(N) [9, 49, 841, 563, 644, 1759]. SU_{q}(2) [1698]. T [1482, 434, 338]. T^{2}/Z_{4} [1688]. T^{4} [1650]. p̃ [243]. U(SU(∈)) [1262]. U(n) [1814]. Ures(sl_{2}) [801]. U_{q} [1568, 449, 1026, 383]. u_{q}(2) [444, 962].

U_{q}(C_{n}(1)) [1935]. U_{q}(ŝl_{n}) [247]. U_{q}(sl_{n+1 ,tor}) [413]. U_{q}(sl_{2,tor}) [634]. U′_{q}(so_{3}) [522]. U_{q}[gl(m|n)] [1274]. U_{q}[sl(N̂ |1 )] [338]. U_{q}[gl(ˆ2|2)] [149]. U_{q}[gl(N̂|N)] [149]. U_{q}[sl(n + 1|m)] [967]. U_{q}sl(2|2)^{(1)} [1579]. U_{q→0}(sl(2)) [331]. U_{q}so(5) [1305]. UOSp(k_{1}/2k_{2}) [1852]. V [514]. V (r_{ij}) = g(r_{ij}^{2}) [1818]. W [133]. W^{2} [1912]. W_{∞} [1388]. XXZ [1475, 247]. Y (osp(m|2n)) [1272]. Y (sl_{n}) [1648]. ζ [693]. z → (αz + β)/(γz + δ) [1902].

-adic [1874]. -algebra [1864]. -algebras [133, 1059, 1217]. -Bernoulli [82]. -body [970, 888, 1606, 214]. -boson [1421, 942, 1818]. -branes [243, 10, 371, 40, 835, 317, 669]. -bundles [1455]. -categories [143, 1858]. -channel [1893]. -coefficients [1094]. -component [1450, 31, 1110]. -conformal [1861]. -constraint [982]. -cosymplectic [621]. -deformation [1902]. -deformations [573]. -deformed [626, 829, 333]. -dimensional [507, 425, 936, 1180, 33, 658, 124, 487, 1110, 297, 1065, 1816, 1829, 1492, 1229, 365, 564, 1932, 1291, 571, 469, 386]. -dimensions [1269]. -distributed [1565]. -dressing [1093, 365]. -extended [820]. -Fock [801, 333]. -fold [490, 789, 450, 1072]. -function [791, 1231, 1613]. -generalized [452]. -geometries [1669, 1786]. -Hahn [46]. -Hamiltonian [511]. -harmonic [578]. -hypergeometric [449]. -invariant [1529, 262]. -invariants [3, 637]. -Jacobi [1537]. -Laplace [578]. -manifolds [1091, 1278]. -mapping [243]. -matrices [140, 1579]. -matrix [1272, 35, 1296, 129, 937, 434]. -model [1874]. -modules [1387]. -operators [663]. -order [609]. -oscillators [640]. -particle [1234]. -percolation [83]. -piece [1317]. -plane [1861]. -point [1335, 1367]. -potential [1160]. -product [972]. -quantum [1632]. -qubit [1781]. -sphere [1256, 103, 104, 1235, 1255, 1]. -symmetric [627, 954, 648]. -symmetry [1029, 1102, 1906]. -symplectic [786, 1752]. -system [770]. -systems [148]. -theory [497, 691, 240, 1428]. -torus [1553, 1069]. -type [496]. -wave [713, 589, 759].

1/2 [1294]. 12- [962]. 1D [1352].

2-dimensional [567]. 28 [1256]. 2D [1271, 57].

33 [492]. 35 [637]. 36 [1580]. 38 [1810]. 39 [207, 73, 293]. 3d [1207].

40 [49, 737, 835, 738, 208, 209]. 41 [552, 491, 459, 339, 340, 1254, 774, 553, 638, 554, 664, 1729, 493, 206]. 42 [1811, 936, 739, 1400]. 43 [1757, 1756, 1096, 1676, 1542, 1421, 1948, 1369, 1278, 1066, 1310, 1340, 1954, 1945, 1067]. 44 [1951, 1420, 1399, 1944, 1950, 1444]. 45 [1755, 1952, 1955, 1953, 1949, 1946, 1947].

5D [814, 1010].

9- [830].

= [1390, 1535, 1597, 1425].

Abelian [420, 672, 240, 1033, 1647, 675, 667, 1287]. Ablowitz [1521, 1926]. Abrikosov [793]. Absence [79, 164, 983]. Absolute [348, 742, 414]. absolutely [1712, 468]. absorption [482]. according [1718]. account [1415]. accurate [1426]. acoustic [1366, 1212]. Acoustics [544]. across [1281, 1195]. acting [1154]. action [693, 316, 1411, 696, 344, 1825]. actions [625, 819, 671]. Addendum [1340, 1887, 1444]. Adding [959, 82]. Additional [1451]. additive [1186, 1187, 1677]. Additivity [1147, 1127, 1129, 1128]. Adiabatic [1751, 1068, 866, 1208, 1474, 246]. adic [1874]. Adjoint [438, 1318, 174, 1738, 1511, 485, 515, 335]. adjointness [658, 768]. admit [1079]. AdS [1634, 675]. AdS/CFT [675]. AdS3 [678]. advection [1327]. affine [626, 1471, 1276, 1935, 373, 1210, 472, 1346]. Aharonov [612, 1511, 993, 557, 1626, 298, 562, 1433]. AKNS [1048, 1489, 1110, 1524, 511]. AKS [1805]. algebra [989, 587, 42, 1806, 723, 1390, 824, 1544, 1205, 1233, 1598, 522, 829, 10, 835, 138, 728, 1864, 1935, 1619, 633, 661, 1673, 413, 634, 1637, 786, 1547, 525, 1829, 662, 1067, 880, 1236, 290, 950, 1795, 1947, 769, 1115]. Algebraic [1059, 266, 481, 778, 34, 284, 1807, 1510, 507, 895, 1106, 52, 1412, 426, 343, 377, 1922, 1335, 1702, 966, 22, 1674, 1439, 1695, 1808, 803, 1296, 1410, 586, 1528, 1703]. algebraisations [1348]. algebras [1568, 625, 1889, 278, 1059, 1297, 133, 1217, 1232, 407, 1298, 1389, 852, 1299, 475, 547, 52, 565, 1471, 1865, 680, 794, 632, 45, 102, 138, 581, 1118, 1826, 287, 198, 553, 1027, 609, 500, 985, 165, 166, 1867, 454, 1058, 1183, 456, 1804, 771, 1917, 1655, 1654, 772, 1610, 458, 1909]. Algebro [902, 33, 1521]. Algebro-geometric [902, 33]. algebroid [1210]. algorithm [497, 1830, 1890, 1935]. Algorithms [748, 93]. alignment [1077]. Almost [1888, 1941, 711, 1855, 206, 28, 945]. almost-intertwining [711]. Aloff [1455]. Alternative [563]. Amalgamations [1618]. ambiguity [1040]. among [461, 1286]. amplitude [866, 459, 241, 1565, 994]. amplitudes [1570]. Analog [1218, 145]. Analogies [286, 327, 1670]. analogs [558, 934, 830, 437]. analogue [247]. analyses [998, 1072]. Analysis [323, 807, 1345, 1371, 74, 710, 1679, 716, 1455, 1843, 1649, 1357, 381, 329, 412, 921, 1071, 616, 1553, 1055, 1324, 1203, 1956, 1936, 1946, 357, 276, 1641]. Analytic [1895, 1089, 764, 1837, 968, 937, 1088, 934, 1787, 1675, 500, 1203]. Analytical [1248, 557, 1358, 1444]. Analyticity [1827]. Analyzing [95]. Anderson [1548]. Angle [930]. angles [1458]. Angular [1927, 1506, 1262, 765, 583, 662]. anharmonic [780, 7, 293]. Anisotropic [84, 1737, 272, 1164, 1925, 1083, 813, 592, 593, 624, 959, 164, 351, 1054]. Annihilation [1019, 1621, 999]. Annihilation-diffusion [1019]. annular [557]. annulus [851]. Anomalies [783, 1470]. Anomalous [1386, 886]. anomaly [136, 426, 1700, 1165, 168]. Anosov [1827]. ansatz [867, 1216, 1657, 1502, 1057, 644, 1231, 1528]. anti [1931, 1539, 795, 564]. anti-de [1931, 795, 564]. anti-self-dual [1539]. anticommutator [1368]. antifermion [1715]. antilinear [1102, 1197]. antimonopole [1639]. antisymmetric [1316]. antisymmetrization [905]. antithesis [1822, 1952]. any [1219, 266, 1201, 1673, 1378, 457]. anyon [1860, 583]. Aperiodic [290]. appear [1395]. appearance [1759]. Appell [1339]. Application [1413, 1089, 722, 1841, 1452, 546, 695, 1, 1256, 606, 370, 161, 1609, 550, 978, 1553, 111, 533, 1230, 1304]. Applications [1750, 1107, 403, 8, 73, 1572, 211, 1521, 899, 1250, 1369, 549, 146, 221, 1425, 38, 1823, 458]. applied [126]. approach [1623, 1354, 1059, 1217, 971, 507, 1389, 867, 1663, 60, 1680, 1516, 1606, 44, 343, 1355, 1863, 1478, 589, 451, 1905, 1682, 162, 661, 1358, 1487, 561, 803, 1775, 1195, 703, 1930, 307, 755, 7, 308, 293, 664, 1492, 671, 1444, 1670, 1177, 1017]. approaches [1857]. approximate [1675, 334]. approximately [1270]. Approximation [1899, 1474, 1012, 1427, 1459, 350, 702, 1748, 978, 1751, 703, 704, 705, 434, 64]. approximations [1632, 1511, 725, 1135, 571]. arbitrary [1232, 420, 520, 102, 1937, 378, 36, 1108, 700, 1691, 1158, 1566, 1172, 304]. area [199]. area-preserving [199]. areas [913]. argument [596]. arguments [706]. arising [138]. Arnold [118, 826]. array [1028]. arrays [453]. Askey [448]. aspect [378, 1128]. aspects [1857, 601, 1275, 1301, 1420, 500, 213]. assignments [1141]. assisted [1126]. Associated [984, 1748, 31, 1772, 1238, 824, 134, 475, 1573, 1455, 390, 932, 1657, 1399, 1303, 845, 37, 1169, 1619, 256, 1170, 1692, 999, 772]. associative [565]. asymmetric [270]. asymmetry [910]. Asymptotic [1359, 1892, 1923, 827, 1498, 1843, 248, 197, 125, 636, 1447, 1878, 605, 1180, 250, 1396, 945, 611, 1368, 708]. asymptotically [1261, 187]. Asymptotics [1258, 270, 849, 288, 878, 566, 236, 775, 720, 850, 6, 207, 1839, 1750, 592, 593, 1022, 1056, 922, 1520, 1944]. atom [1312, 1150, 1340, 215]. atomic [461, 1795, 1947]. atoms [419, 842, 1795, 1947]. attractive [1347]. attractor [283, 1599, 1840]. Attractors [892]. atypical [48]. automata [509]. Automatic [695]. automorphism [715]. Automorphisms [932, 609, 862]. average [1066, 875]. averaged [26]. averages [1500]. averaging [1084, 1744, 1328, 308, 664]. averaging-theory [308, 664]. Avez [1881]. Avoided [779, 27]. Avoiding [1912, 84, 93]. Avossevou [1255]. axial [106, 308, 664]. axiomatic [751]. axis [1156]. axisymmetric [1355, 187].

B [1845, 29]. background [836, 50, 527, 762, 810, 1041, 678, 1100, 131, 1542, 1166, 679]. backgrounds [316, 1294, 515]. Bäcklund [1271, 995, 1050]. bag [1100]. Baker [912, 147]. Balian [1361, 1833]. ball [519, 1509, 1466]. Bañados [1749]. band [940, 845, 646, 307]. band-gap [845]. bands [878]. bandwidths [141]. Bardeen [350, 1376]. Bargmann [189, 724, 747, 1000]. barrier [1678, 1373]. based [989, 191, 1837, 450, 307, 755, 148, 650, 1610]. bases [888, 1116, 1667, 1935, 1065, 13, 208, 384]. basic [830, 1675, 1799]. basin [1078]. basis [449, 1799, 280]. Batalin [815]. Bateman [520]. Baxter [1331, 1863, 917, 383]. Baxterizations [1300]. Baxterized [1572]. Baym [434]. BC [889]. be [1822, 1952, 199, 1131]. beams [1266]. Becchi [427, 698, 1157, 109]. Beethoven [141]. Behavior [1142, 1117, 366, 232, 1878, 1498, 900, 570, 1076, 1351]. Belinskii [274]. Bell [1422, 102, 107]. Belov [730, 1306]. Beltrami [1385]. Benjamin [1674]. Berezin [1509]. Bergman [930]. Bering [309]. Bernoulli [82, 1496]. Berry [808, 1208, 617]. Bessel [1273, 1393, 863, 1748]. beta [1222]. Bethe [1243, 1645, 867, 1216, 1657, 1502, 1057, 782, 644, 1231, 1528, 1113]. between [1583, 1186, 1187, 1215, 311, 1190, 136, 886, 102, 1353, 286, 327, 1094, 125, 622, 1313, 1638, 1670, 1288, 1325, 795]. beyond [1022, 733]. Bi [1011, 1374, 450, 127, 123, 1805, 953]. Bi-Hamiltonian [1374, 450, 127, 123, 1805, 953]. Bi-quasi-Hamiltonian [1011]. bialgebras [1245, 1755, 282, 1754, 861, 524]. Bianchi [401, 715, 514, 599, 274, 443]. bicategorical [1389]. bicovariant [1504]. bidirectional [652, 1168]. bifurcation [1733, 1021]. bilinear [965, 58]. Billiard [371, 557]. billiards [1909]. Binary [759, 1343, 857, 367, 623]. binding [1884, 1372]. Binegar [47]. binomial [142, 1794]. Biot [325, 546]. biparametric [831]. bipolar [1010]. Birkhoff [1274, 1922]. Bisognano [1154]. Bispectrality [294]. bistatic [119]. bit [1937]. Black [1233, 141, 1321, 1923, 438, 1318, 1938, 131, 684, 677, 1749, 275, 1592, 442]. Black-brane [1233]. black-holes [442]. Blaszak [730, 1306]. blistering [504]. Bloch [643, 1148, 1442, 1397, 1651, 1652, 1028, 847]. blow [179]. blow-up [179]. BMS [1689]. bodies [1620]. body [970, 888, 265, 1560, 1769, 1921, 1413, 339, 113, 1606, 1035, 1479, 1084, 214, 1512, 1495, 1722, 594, 1304, 213, 170]. Bogoliubov [1099, 1788, 1732]. Bohm [1511, 562, 612, 993, 557, 1626, 298, 1433]. Boltzmann [1876, 1564, 434]. Bondi [132]. Bondi-type [132]. Boost [961]. Borcea [188]. Borel [1377, 530, 1609]. Borg [1613]. Born [672, 1841, 1616, 1213]. Borromean [197]. Bose [1500, 791, 1646, 1847, 1897, 1723, 1528]. Boson [662, 1571, 1343, 942, 1421, 1818]. Bosonic [377, 687, 1773]. Bosonization [338]. bosons [538, 1042, 1103, 1159, 1267, 1046]. Bott [1720]. Bound [1280, 1098, 1709, 1710, 1683, 1347, 5, 1810, 1404, 177, 942, 1421, 1818, 1132, 1133, 1194, 878, 259, 1112, 1823]. boundaries [836, 971, 1078, 1145, 329, 1713, 1949]. Boundary [1117, 1585, 1351, 1236, 1081, 1082, 1204, 1757, 16, 1447, 6, 207, 1486, 1737, 1098, 993, 1449, 1500, 247, 1057, 486, 644, 1842, 536, 1319, 1005, 309, 25, 64, 204]. Boundary-layer [1117]. Bounded [549, 1204, 1105, 1757, 1756, 424, 862]. Bounds [1809, 1834, 1434, 1892, 539, 540, 818, 843, 1666, 92]. box [20]. boxes [1847]. bracket [435, 1778, 126, 309]. brackets [115]. Bradlow [1460]. braid [657, 963, 1572, 1300, 1287]. Braided [1866, 548, 969, 1332, 521, 1860, 143, 1199]. braiding [626]. Branch [27, 983, 158]. Branch-point [27]. Branching [313]. Brane [682, 1163, 1233, 838, 690]. Brane-world [682]. branes [685, 238, 317, 682, 669, 243, 10, 371, 40, 835, 667]. Brans [1096, 957]. breaking [1097, 695, 1427, 124, 705, 887, 1127]. Breit [1373]. broadened [843]. Broken [1403, 613, 220, 109]. Brown [569]. Brownian [81, 1198]. BRS [106]. Bruhat [1049]. building [277, 479]. bundle [770]. bundles [1872, 764, 1218, 396, 1808, 72, 1210, 148, 948, 458, 1455]. Burgers [902, 1060]. Buslaev [1423]. butterfly [846].

Calabi [188, 490]. Calapso [181]. calculable [536]. Calculation [1780, 55, 1775, 132, 1353, 1178, 1502, 1900]. calculi [1771]. calculus [1545, 1535, 1259, 870, 284, 1879, 1711, 23, 1745]. calculuses [521]. Calogero [265, 422, 1348, 266, 824, 794, 431, 1547, 585, 1491, 1211]. Camassa [508, 196, 486]. Campbell [147, 1220]. cannot [1131]. Canonical [4, 1580, 1781, 1572, 428, 888, 1920, 133, 1346, 1836, 429, 622, 1847, 1000]. Capacity [499, 1138, 1132, 1126, 943, 1133, 1127, 1128]. CAR [1115]. CAR-algebra [1115]. Cartan [964, 1438]. Cartesian [1690]. cascades [362]. case [805, 806, 424, 242, 1938, 1404, 1859, 1635, 917, 1683, 609, 1631, 1751, 879]. cases [928]. Casimir [553, 1583, 1687, 1291, 137, 45, 57, 198, 614, 1847, 518]. cat [118]. categories [669, 1061, 143, 821, 1858]. Category [1296, 579]. Cauchy [1200, 398]. causal [1354, 382, 14, 209]. Causality [1931]. caustic [1206]. caustics [211]. cavity [1895, 1687, 1073]. Cayley [1853, 1874]. CCR [1297]. celebrated [758]. celestial [60]. center [804, 805, 806, 994]. Centers [970]. Central [1244, 1260, 489, 1637, 1579]. centrally [633]. centroids [1498]. century [222, 223, 216]. certain [1771, 360, 1395, 1275, 1301, 1420, 1899, 1490, 1431, 1236, 205]. CFT [675]. chain [1475, 1645, 1608, 340, 156, 247, 728, 952, 1113]. Chains [802, 40, 1501, 825, 1499, 95, 1325]. Chaltikian [730, 1306]. Chandra [1862, 1852]. change [689]. channel [1893, 1133, 1670]. channeling [308, 664]. channels [1132, 499, 1147, 943, 1127, 1129, 1128]. Chaos [1938, 714, 912]. Chaotic [232, 618, 1012, 1575]. Character [901, 41]. characteristic [1472, 1841, 1660, 198, 553]. characteristics [1596]. characterization [1029, 27]. Characterizations [429, 1337]. Characterizing [1119, 1324]. Characters [286, 1438, 1597]. Charge [1587, 837, 1160, 1469, 974, 1637, 1515]. charged [534, 1494, 642, 272, 699]. charges [1920, 1238, 569, 1176]. Checkerboard [828]. checkerboards [1016, 790]. Cherenkov [1493]. Chern [1200, 1660, 1589, 810, 1263, 1004, 160, 582, 849, 467, 1518, 299, 707]. Chernoff [1184]. Cherry [1819]. Chetaev [1800]. Chiral [101, 1114, 1158, 685, 358, 391, 267, 168, 1112, 1174]. Chirality [591, 348, 349]. Choi [1285]. choices [1489]. cholesteric [1780]. chord [1620]. Christoffel [1149]. chromodynamics [974, 1662]. Chronon [1153]. CHSH [346]. circle [1458, 723, 254, 1555]. circular [359, 614, 1729]. Cízek [1703]. class [1370, 1387, 1892, 1800, 757, 541, 1024, 1025, 34, 545, 370, 138, 251, 1080, 1826, 719, 660, 373, 400, 1252, 1310, 610, 822, 1552, 1210, 1112, 1554, 1668]. Classes [165, 401, 984, 1191, 518]. Classical [263, 923, 741, 723, 1494, 558, 140, 1044, 300, 217, 264, 865, 692, 120, 722, 990, 740, 1836, 809, 811, 565, 1604, 1605, 18, 227, 1627, 1227, 1790, 1134, 1126, 1326, 474, 1849, 581, 1439, 1695, 1731, 1759, 213, 1176, 397, 731, 532, 1127, 771, 950, 1128, 1903, 1111]. Classicality [1226]. Classification [1872, 130, 360, 282, 1774, 1718, 663, 48, 1655, 602, 43, 796, 1664, 1440, 1770, 522, 1941, 825, 954, 820, 1816, 1829, 1510]. classify [1752]. Clebsch [277, 479, 1701, 384]. Clifford [587, 950]. Cliffordons [697]. close [1583, 503, 213]. Closed [625, 715, 686, 1031]. closeness [406]. Closure [920]. clover [734, 735, 733]. Cluster [1234, 59, 81]. coaddition [548]. coadjoint [1413]. coalgebra [923]. coarse [719]. cocycles [1866]. Codazzi [134]. code [312, 313]. codes [1900]. codimension [1284, 1329]. codimension-one [1329]. coding [788]. Codon [312]. coefficient [654, 1738, 1838]. coefficients [1543, 444, 962, 277, 479, 182, 136, 69, 519, 1080, 830, 1094, 384, 1799]. coercivity [1459]. coframe [370]. Coherent [941, 1948, 1911, 750, 1954, 1196, 999, 847, 1069, 4, 1580, 989, 639, 1068, 1190, 1837, 911, 246, 1259, 1628, 1193, 1408, 166, 1741, 470, 1554, 1796, 472, 1582]. coherent-state [1259, 1741, 1582]. cohomological [919]. Cohomology [754, 1529, 1752, 1654, 908]. cokernel [1872]. Coleman [733]. Colliding [492, 493, 373]. collineations [130, 1845, 1886, 1718]. collision [608]. collisions [1707]. color [1655]. Colored [383, 145, 1670]. colorings [92]. Combinatorial [857, 160, 1863]. combinatorics [86]. combined [573, 1098, 993]. combs [1513]. Commensurate [1111]. Comment [491, 1755, 1254, 1542, 492, 737, 774, 738, 739, 1954, 1729, 1956, 493, 206]. Comments [1886]. commitment [1937]. commutant [1197]. commutative [1941]. commutativity [428]. commutator [1274, 1368]. commutators [1843]. Commuting [485, 767]. comodules [524]. compact [1279, 765, 1005, 1804, 617]. compactification [389, 659]. compactifications [685, 1428, 684, 688]. Companion [238]. comparison [95]. Compatibility [1141, 1030, 1313, 1308]. competing [1853]. Complete [605, 1079, 387, 805, 806, 715, 90, 1440, 1496, 1029, 1762, 1031]. completely [758, 1175, 1627]. completeness [1186, 895, 1106, 724, 586]. Complex [1154, 1873, 1217, 374, 1570, 495, 904, 520, 1302, 349, 1562, 1924, 1953, 1320]. complexity [1747]. component [31, 1271, 1440, 402, 1450, 1110]. components [536]. Composite [422, 1924, 1953, 1875]. composites [828]. compound [1288, 1736]. Comprehensive [616]. compressible [65]. compression [1134]. Computation [535]. computations [1781]. computer [1900]. Computing [594]. concentration [608]. concentric [104]. concept [51]. concepts [1544, 223]. concurrence [1781]. condensate [1471, 849]. condensates [589, 1528]. condensation [1897]. condition [1232, 88, 5, 1810, 661, 199, 877, 260, 26, 64]. Conditional [1124, 516, 487]. conditionally [616]. Conditions [1077, 1204, 1757, 805, 806, 16, 1186, 1486, 656, 1737, 1590, 1098, 993, 1449, 1500, 1469, 1057, 644, 785, 1130, 1945, 1823, 25]. conductance [18]. conducting [172]. conductivity [790, 706]. cone [519, 391, 679]. Cones [14, 209, 1195]. confined [1291]. confinement [1702, 1350, 1751, 291]. Confluence [1653]. Confluent [466, 827]. Conformal [1928, 1298, 1237, 135, 547, 695, 1861, 184, 39, 1777, 1395, 1585, 273, 1654, 170]. Conformally [624, 897]. Conjecture [374, 67, 790, 1914, 1129]. conjectures [870]. Conjugation [143]. connected [1674]. connectedness [1162]. connection [1186, 1187, 1094, 561, 1638, 1288]. connections [1612, 246]. Connes [416]. consequence [910, 1165]. Consequences [719]. Conservation [1080, 598, 181, 1208, 1239, 976]. conservative [89]. conserved [1041]. conserving [1043]. consistency [1232]. consistent [315, 623]. constant [1458, 602, 512, 604, 1261, 159, 474, 1294, 1883, 1882]. constants [132, 68, 1496]. constrained [475, 1489, 1169, 1485]. constraint [625, 401, 982]. Constraints [173, 971, 1800, 1574, 306, 759, 126, 1665, 1453, 885, 763]. Constructing [1271, 1417]. Construction [1806, 285, 1661, 56, 1058, 457, 507, 1803, 1189, 1190, 1024, 1025, 1507, 633, 58, 1685]. constructions [1572, 281, 1521, 774]. Constructive [221]. Contact [1091, 1278, 525]. content [155]. context [591]. Contextual [1407, 1682]. continuation [1089]. continuations [1237]. continued [1891]. Continuity [1645, 414]. Continuous [692, 1247, 1642, 1505, 382, 1782, 267, 1922, 1037, 1739, 1722, 1712, 468, 1045]. Continuous-time [692, 1045]. Continuum [776, 1346, 1707, 1207, 837, 1632, 1453, 885]. Contraction [32, 145, 195]. contribution [309]. control [991, 302]. Controllability [988, 1401, 765]. convection [539, 540, 1809]. convective [1797]. Convergence [523, 64, 1414, 192, 1315, 1433]. convergent [497]. conversion [593]. converting [1286]. convex [907, 1304]. Convexity [1228, 151]. convolutions [861]. Cooper [350, 1376]. Coordinate [644, 1156]. coordinates [1144, 1676, 855, 1690, 1506, 1437, 124, 1748]. copolymer [1880]. corner [745]. corotating [477]. correction [1790, 256, 1741, 1122]. corrections [543, 1153, 842]. correlated [1143, 1132]. correlation [1780, 1040, 102, 921, 1071, 72, 770, 304]. correlations [998, 922]. correspondence [680, 675, 585]. correspondences [980]. corresponding [1568, 1598, 802, 145]. Corrigendum [1223]. coset [1530, 531]. Cosgrove [597]. cosine [1782]. cosmic [158, 372]. Cosmological [814, 602, 512, 1008, 1009, 1740, 1936, 1946, 323]. cosmologies [401, 366, 715, 1161, 1758, 856]. cosmology [1096, 957, 371]. cosymplectic [621]. cotangent [826]. Coulomb [159, 804, 805, 806, 16, 1352, 867, 1516, 994, 108, 562, 708, 968, 920]. Coulomb-oscillator [159]. countable [1696]. Counterexample [1129]. Counterexamples [122]. counterspace [409]. Coupled [343, 836, 1316, 15, 1827, 896, 360, 245, 33, 1250, 1369, 1326, 1357, 1709, 1710, 600, 1194, 572, 878, 1728, 1050, 1159, 1267, 1668, 1908]. Coupling [865, 1117, 1246, 675, 690, 1086, 1598, 844, 488, 1662, 1799, 303, 1701]. couplings [356, 1172, 893]. covariance [344, 1198, 1199]. Covariant [1204, 1757, 1531, 909, 869, 1282, 396, 1804, 263, 324, 1724, 640, 1318, 1660, 1533, 1647, 1907, 1363, 1851, 426, 305, 606, 786, 946, 397, 756]. Covariants [1882]. Coxeter [266]. CP [849]. CPT} [1314]. CPT}-symmetries [1314]. CR [1278, 1091]. Crane [821]. Creation [782, 696, 1294, 1621, 999]. criteria [1226, 1124]. criterion [361]. Critical [1311, 328, 1040, 283, 497, 944, 270, 1416]. cross [1687, 419]. crossed [407, 1101]. crossing [556, 1811, 1293]. crossings [779, 1284, 27]. crossover [1680]. Crystal [1935, 74, 79, 495, 1503]. crystals [1780, 1414, 868, 509, 379]. Cubic [1412, 630, 36]. cubics [1338]. Cummings [1341, 1151]. curl [325, 288]. current [243, 1192, 1214, 63]. currents [850, 197, 359, 1729, 304]. Curvature [1749, 506, 1473, 1434, 130, 604, 1008, 1009, 159, 924, 474, 1882, 1510]. curvatures [246]. curve [764, 1491]. curved [1224, 1342, 1798, 529, 1517, 1203]. curves [721, 792, 377, 1276, 771]. cusp [1088]. cusps [367]. Cyclic [1092, 1365, 1784, 1866]. Cyclical [366]. cylinder [779, 254, 614, 1656]. cylindrical [1748, 1010]. cylindrically [797, 1886, 1719].

D [1636, 1717, 838, 1777, 1733, 667, 515]. D-brane [838]. D-branes [667]. damped [1476, 1678, 283, 1840]. damping [128, 892, 1678]. Darboux [96, 1362, 1363, 450, 267, 1600, 933, 1149, 600, 551, 385, 623]. Darboux-type [385]. Darbouxian [1173]. Dark [925]. data [1447, 1616, 568, 898, 1139, 1134, 1789, 1665]. Davey [1449, 179, 660]. decay [1243, 696, 34, 55]. Decoherence [1706, 1546, 347]. Decomposition [1698, 988, 405, 463, 436, 1922, 1260, 1251, 454]. decompositions [1781, 71, 1030]. deconfined [781]. decoupling [343]. defects [1290, 378, 1329]. Deficiency [1621]. defined [485, 58, 1882, 1554, 404, 518]. definition [742, 91]. definitions [1469]. Deformation [1871, 582, 154, 1857, 1870, 294, 1681, 54, 1335, 1902, 380, 431, 831, 1656]. Deformations [134, 607, 1917, 1270, 573, 923, 1217, 194, 181, 1524]. Deformed [1793, 626, 829, 333, 1199]. degeneracy [789]. Degenerate [1862, 760, 897, 1323, 179, 335]. degeneration [482]. degree [597, 912]. degrees [400]. delay [307]. delta [169]. Demazure [379]. Demidowitsch [361]. Democratic [858]. Dense [1513, 1343]. densities [461]. Density [1192, 1473, 461, 154, 1415, 1721, 1891, 1779, 843, 1735, 1304, 1587, 307, 533]. departures [701]. dependence [87]. dependent [865, 715, 1681, 652, 1168, 905, 1617, 1175, 427, 869, 1034, 1542, 106, 1838, 1727, 1166, 1241, 173, 165, 166, 1480, 300, 502]. Derivation [1852, 1819, 376, 426, 1518, 255, 1655, 1715]. derivation-simple [1655]. derivations [1655]. derivative [894, 271]. derivatives [485, 1551]. derived [1792, 287, 550, 415, 1174]. descent [721, 299]. describing [645, 181]. Description [1940, 721, 225, 1444]. deSitter [316]. Detecting [1486]. determinant [3, 637, 101]. determinants [819, 1113]. Determination [1482, 589, 1727]. determine [811]. deviation [781]. deviations [80]. diagonal [139, 1408, 533]. diagonalization [1572]. Diagram [579, 494, 793, 846]. diagrams [839, 390, 579, 1043, 454]. diatomic [825]. diblock [1880]. Dicke [1096, 957]. Dicke-like [1096, 957]. dielectric [1493, 1052]. diffeomorphism [633]. diffeomorphisms [1827, 715, 1829]. difference [43, 267, 1282, 767, 965, 1879, 125, 608]. differentiable [14, 209]. Differential [8, 73, 408, 575, 1364, 521, 506, 516, 1738, 421, 375, 723, 6, 207, 629, 1455, 631, 1391, 1435, 1535, 1771, 1916, 234, 994, 267, 177, 1869, 269, 195, 109, 353, 1666, 662, 505, 1324, 772, 1686, 708, 322, 950, 365, 834]. differentially [1834]. differentials [324]. differentiation [606, 701, 203]. diffraction [1513]. diffractive [592]. diffusion [1798, 886, 1817, 1739, 1386, 1838, 595, 1058, 1019]. diffusive [534]. Digitalizing [340, 156]. dihedral [481]. dilute [1646, 1020]. dimension [1219, 1205, 390, 520, 632, 538, 1201, 872, 1507, 283, 700, 1194, 1230]. dimensional [865, 1401, 1844, 1142, 1422, 1932, 462, 927, 507, 850, 1246, 495, 777, 1474, 232, 902, 1596, 1289, 425, 630, 565, 1640, 1291, 123, 567, 936, 642, 450, 1016, 101, 1885, 1180, 33, 77, 1438, 530, 282, 185, 1264, 1070, 817, 81, 522, 1002, 800, 699, 1779, 55, 451, 746, 378, 498, 658, 924, 791, 286, 327, 559, 56, 348, 349, 1849, 1398, 124, 644, 712, 1658, 798, 381, 487, 1110, 108, 1697, 701, 359, 571, 572, 609, 1637, 297, 985, 1150, 1340, 821, 24, 1065, 1816, 891, 1384, 1829, 1183, 110, 1170, 469, 1592, 442, 129, 713, 1729]. dimensional [48, 1654, 926, 1492, 1229, 1510, 1159, 1267, 386, 1007, 511, 1231, 365, 795, 564, 1174]. Dimensionally [905, 1617, 1891]. dimensions [894, 1217, 613, 79, 316, 601, 1280, 1411, 1418, 1290, 1047, 1220, 992, 1435, 1436, 1516, 1100, 9, 49, 19, 318, 1163, 1035, 1479, 860, 1526, 272, 497, 83, 1165, 1709, 1710, 759, 1269, 1683, 417, 947, 1894, 1158, 1784, 271, 1426, 26, 1699]. dimer [94]. dipole [1052]. dipoles [536]. Dirac [1944, 1768, 1530, 1352, 867, 1041, 1516, 1281, 642, 1100, 391, 131, 1153, 605, 1742, 1104, 658, 819, 1835, 1808, 297, 1513, 1883, 1807, 1520, 773, 1433, 1167, 533]. Direct [1178, 255, 178, 1502, 487, 1518, 893]. direction [1883]. directionally [652, 1168]. directions [1077]. directive [359, 1729]. director [1780]. Dirichlet [1297, 1449, 1500]. Discontinuities [773]. Discontinuous [510, 364]. Discrete [725, 1056, 1398, 876, 228, 1325, 1458, 418, 1782, 776, 1346, 1738, 902, 1642, 1521, 1790, 1607, 1135, 1037, 1080, 1522, 1879, 1133, 1336, 1748, 353, 1588, 1712, 1015, 337]. discrete-velocity [418]. discretization [1392]. discretized [1370]. disk [851, 968, 753]. disk-annulus [851]. disorder [781]. disordered [1670]. Dispersionless [1093, 1451, 311, 475, 1776, 600]. dispersive [358, 232]. Dissipative [1402, 1624, 1599, 1850, 1192, 1327, 584]. distance [503, 1076, 922, 146]. distances [1735]. distributed [1565]. distribution [1341, 1614, 1138, 1726, 1304, 608, 146, 1368]. Distributional [1377, 958, 1749]. distributions [1932, 346, 142, 63, 1620, 1631]. Distributive [1787]. distributivity [1288]. divergence [347, 1242]. divergences [11, 737, 918]. divergent [1312, 336]. Division [1865]. do [811]. does [568]. domain [358, 1290, 696, 745, 683, 1057, 864]. domains [1713, 1949, 1554]. dominated [192]. dominated-type [192]. Doppler [307]. Doppler-scale [307]. dot [1705]. Double [613, 1755, 1591, 1754, 1540, 1819, 630, 905, 484, 830]. Double-scaling [613]. doubles [1826]. doubling [270]. doubly [794]. down [1641]. dragging [1834, 1666]. Dressing [157, 1093, 365]. Drinfel’d [1648, 1531, 1821]. Driven [1241, 230, 1809, 505, 1015]. driving [1840]. Dual [114, 352, 1445, 1344, 810, 66, 853, 1004, 46, 849, 1539, 1898, 998]. Duality [1245, 1344, 691, 1115, 344, 159, 416, 1430, 1016]. Duffin [769]. dust [653, 718, 514, 187]. dyadic [1023]. Dym [1238, 1532, 1793, 1720, 1773]. dynamic [62]. Dynamical [888, 852, 1427, 689, 162, 1487, 1485, 117, 127, 276, 1371, 1769, 1921, 576, 1331, 1286, 35, 1146, 1599, 1079, 125, 760, 892, 920]. Dynamics [308, 664, 120, 1059, 1558, 1460, 990, 1018, 459, 241, 1792, 88, 546, 601, 266, 1413, 1634, 1565, 1084, 1459, 295, 248, 1850, 913, 1739, 126, 1684, 95, 1073, 1007, 89]. dynamo [1441]. dynamos [1053]. dyonic [1156]. dyons [694, 1223, 900]. Dyson [1541, 870].

each [1583]. early [366, 544]. Eckart [331]. edge [435, 1477]. Effect [356, 807, 82, 1144, 1676, 993, 1626, 609, 298, 1376]. Effective [693, 1633, 684, 894, 1574, 1658, 671, 619, 696]. effects [1888, 1633, 813, 797, 973, 1030, 1913]. Efficient [1188, 1135]. Eguchi [1508]. eigenelements [335]. eigenfunction [773]. eigenfunctions [1547, 387]. eigenproblems [648]. eigenstates [422, 191]. eigensubspaces [335]. Eigenvalue [1726, 1658, 1221, 1508, 337, 204]. Eigenvalues [462, 1657, 557, 873, 1830, 1426, 518]. eigenvectors [709, 1830, 200]. eight [1253]. eikonal [1873]. Einstein [51, 1590, 897, 319, 814, 1843, 513, 1635, 545, 900, 570, 1665, 1005, 321, 1381, 322, 404, 1528, 276, 1909]. Elbow [38]. electric [612, 1377, 1101, 536, 1052, 1294, 1587]. electrodynamics [237, 1075, 53, 973, 1556, 572, 1587, 1157, 1176]. Electromagnetic [856, 263, 1212, 419, 868, 558, 913, 117, 259, 1424, 566, 119, 1017]. electromagnetism [240, 1761]. electron [867, 1697, 1266, 1795, 1947, 779]. electrons [1474]. electrostatic [1213]. electrostatics [550]. element [1767]. Elementary [176, 343, 661, 355]. elements [465, 1673, 1659, 968, 280]. ellipsoid [146]. Ellipsoidal [1010, 1268]. ellipsoids [601]. Elliptic [1667, 1209, 805, 806, 764, 1348, 792, 1331, 984, 1597, 35, 1146, 1092, 1365, 1821, 768, 194]. elliptical [477]. embedded [852, 1716, 1955, 490]. Embedding [1163, 1047, 1220, 545, 1685]. Embeddings [1557]. Emden [410]. encircling [1592]. energies [1311, 650]. Energy [1841, 244, 818, 504, 1103, 1687, 236, 775, 15, 491, 1068, 628, 527, 850, 316, 30, 1839, 1372, 1356, 1680, 1291, 569, 1517, 1779, 1512, 55, 36, 1108, 1502, 252, 57, 701, 399, 614, 1327, 1666, 257, 320, 615, 1055, 1112, 1559, 1887, 1719, 26, 27, 502]. energy-level [1680]. energy-momentum [1887, 1719]. Engel [199].

Enhanced [1884, 1372]. Enriquez [1754]. Ensemble [914, 346, 1537, 393]. ensembles [1222, 1937, 873, 200, 230]. Enstrophy [1018]. entangled [1895, 1125]. Entanglement [1122, 1341, 1403, 1119, 1677, 1120, 1126, 1121, 822, 647, 1127, 1123, 1124]. entanglement-assisted [1126]. entanglement-breaking [1127]. Entropic [1431, 635, 1400]. entropies [1258, 921, 650, 1124]. Entropy [1608, 1501, 1892, 850, 1012, 1934, 1599, 91, 1439, 921, 918, 1130, 1945, 1823, 1747, 230]. envelope [890]. enveloping [138, 987]. environment [1131]. Epstein [1484]. Equal [1685]. equality [1130, 1945]. equation [1873, 150, 552, 96, 236, 775, 720, 418, 1097, 1738, 421, 1186, 1187, 929, 1447, 1352, 1876, 902, 1881, 1595, 657, 651, 981, 1564, 1573, 1545, 436, 1679, 1534, 1878, 508, 1436, 1640, 1516, 904, 175, 1168, 567, 936, 520, 642, 131, 1449, 1153, 267, 1669, 1786, 1863, 1742, 530, 410, 597, 1725, 818, 283, 178, 1104, 1326, 1649, 1502, 116, 1148, 1442, 1397, 1651, 1652, 196, 486, 933, 965, 549, 1386, 1838, 1879, 124, 712, 934, 1169, 917, 1109, 487, 383, 1842, 1508, 274, 412, 600, 925, 1674, 350, 488, 1942, 399, 1926, 702, 1337, 1831, 297, 299]. equation [1646, 1880, 1804, 1578, 505, 1763, 434, 1021, 334, 271, 833, 1622, 1950, 756, 1198, 301, 1177, 1320, 1898, 64, 1929, 1641, 502]. Equations [1691, 1445, 341, 506, 1458, 31, 1401, 1613, 655, 1594, 612, 1001, 238, 401, 516, 375, 628, 507, 1841, 424, 1271, 1603, 710, 316, 1238, 1793, 1033, 376, 656, 890, 134, 1616, 963, 1300, 1596, 1363, 811, 1418, 1455, 1590, 1060, 1391, 1412, 1861, 1436, 43, 1642, 652, 1574, 483, 123, 450, 758, 541, 780, 234, 1213, 360, 1440, 39, 319, 994, 814, 1216, 592, 593, 177, 33, 995, 1415, 1721, 269, 597, 1056, 1770, 817, 577, 800, 1250, 1369, 1334, 1890, 1357, 853, 1080, 997, 1209, 559, 598, 1618, 1879, 179, 57, 1840, 570, 1562, 1241, 759]. equations [1269, 198, 553, 661, 62, 212, 600, 1674, 1194, 1743, 1336, 466, 165, 166, 595, 362, 352, 1480, 1539, 1519, 1816, 1296, 181, 1381, 1448, 730, 1050, 1306, 1490, 826, 505, 1324, 434, 128, 1523, 713, 650, 515, 1917, 772, 457, 1715, 322, 854, 1240, 1211, 1167, 404, 323, 511, 387, 1325, 365, 834, 951, 564, 276]. Equilibrium [1723, 229, 74, 1414, 1415, 608, 1198, 1179]. equipartition [1644]. Equivalence [401, 320, 807, 1208, 1389, 541, 230, 719, 1882, 1102]. equivalent [344, 844, 437]. equivariant [528]. Ermanno [1496]. Erratum [552, 1580, 1757, 1756, 459, 207, 73, 1951, 1810, 1096, 637, 1676, 1811, 339, 936, 340, 49, 1952, 1955, 1421, 1948, 1256, 1369, 1278, 835, 1399, 1066, 553, 1310, 1953, 638, 554, 208, 209, 1400, 1945, 293, 664, 1949, 1067, 1944, 1950, 1947]. Error [1900, 1790, 1122, 1426]. error-correction [1790]. errors [1132]. Essential [658, 768, 1261]. estimates [504, 59, 1426]. Estimation [1611]. Euclidean [1096, 927, 957, 642, 1201, 677, 954, 1513, 1379]. Euler [1841, 1660, 57, 712, 661, 600, 415, 501]. Evaluation [1848, 1264, 292, 1795, 1947]. even [1526, 785]. every [206, 28]. everywhere [82]. evidence [627]. evolution [239, 981, 652, 123, 450, 1770, 997, 296, 549, 1748, 1296, 7, 293]. evolutionary [516]. Exact [1877, 567, 936, 642, 281, 410, 774, 997, 296, 431, 1430, 752, 1567, 1730, 721, 857, 1642, 674, 911, 1099, 994, 269, 817, 1334, 1649, 845, 381, 1110, 572, 1266, 610, 947, 1167, 1180, 1022, 21, 1320]. Exactly [103, 536, 1235, 1191, 1019, 616]. example [153]. examples [91]. Exceptional [112, 41]. exchange [556, 1811]. Excision [683]. excitation [708]. excitations [1602, 1492]. exciton [351]. excursion [81]. exhibiting [1021]. Existence [1803, 1589, 317, 810, 1725, 1722, 179, 1562, 1662, 1665, 1788, 5, 1810, 1200, 1616, 576, 131, 1850, 1005, 1675]. exotic [1245]. expanded [387]. expanding [117]. Expansion [503, 1794, 1359, 1819, 534, 627, 827, 1625, 391, 1396, 399, 1064, 863, 202, 59, 952, 611, 1185]. Expansions [1629, 901, 152, 605, 1653, 1552, 773, 708]. expectation [6, 207, 386]. Explicit [189, 574, 436, 59, 1857, 1572]. explicitly [758]. exploding [1015]. exponent [283, 1904]. Exponential [1173, 523, 772]. Exponentially [1426]. exponentials [724, 1673]. Exponentiation [766]. exponents [1012, 497]. expressed [805, 806]. expression [147]. expressions [1248]. Extended [1452, 1477, 688, 1492, 433, 649, 69, 1276, 633, 1775, 820, 1236, 170, 1699]. extended-object [1775]. Extension [1231, 150, 552, 655, 1782, 1921, 1220, 1695, 1553, 186, 734, 735, 1579, 1835]. Extensions [151, 135, 1872, 1511, 1532, 1635, 485, 1773, 733]. Extensive [1179]. extent [811]. Exterior [324, 1435, 1734, 1713, 1949]. external [78, 1377, 1838, 1432, 1566]. Extra [1418, 1741]. Extremal [1907, 1851, 1875].

facets [1116]. factor [1663, 1939, 749]. factored [588]. Factorization [236, 1468, 1572, 1700, 590, 252, 1306, 529]. factors [1248, 273, 1173]. Faddeev [1423]. Faddeev-type [1423]. False [696]. Families [572, 437, 834, 1538, 992, 455, 554, 29]. family [1498, 758, 1720, 1661, 1176, 854, 1172, 1865]. Fano [1455]. Far [570, 404]. Fast [497]. Fast-convergent [497]. faster [11, 737]. Fedosov [882]. feedback [1790]. Fejér [1066, 875]. Fermat [899]. Fermi [605, 1721, 615, 1230]. fermion [1100, 1629, 1715]. fermion-antifermion [1715]. fermionic [837, 1544, 101]. fermions [741, 317, 1290, 1042, 701, 1046, 1267, 1231]. ferroelectric [1828]. ferromagnet [1625]. ferromagnetic [356]. Feynman [1032, 1152, 596, 1603, 390, 100, 870, 192, 1043, 354, 1184, 29]. Fiber [948]. fibered [396]. fidelity [990]. Field [97, 263, 1895, 836, 237, 883, 938, 1558, 534, 239, 613, 1746, 527, 507, 424, 218, 596, 1765, 1633, 1377, 98, 1312, 1494, 1404, 1634, 17, 1061, 432, 1643, 993, 1051, 377, 1263, 114, 1714, 39, 1777, 101, 814, 1843, 1254, 1483, 1742, 345, 1395, 1707, 1927, 1828, 1429, 1086, 1607, 172, 1201, 102, 1002, 558, 699, 248, 1202, 105, 106, 498, 1042, 783, 570, 944, 749, 536, 1265, 396, 1294, 1076, 1731, 500, 786, 58, 1685, 1883, 221, 1767, 222, 1431, 1553, 919, 223, 1112, 975, 1515, 262, 1103, 1006, 1913, 1433, 1566, 26]. field [404, 1788, 1320, 1174, 112, 621]. field-dependent [106]. field-theoretic [1714]. fields [1632, 50, 1316, 1557, 612, 1001, 389, 78, 528, 1271, 1633, 325, 896, 867, 395, 588, 529, 1517, 1213, 344, 1860, 1154, 220, 1349, 842, 1585, 1635, 606, 1261, 913, 1101, 68, 349, 1911, 164, 63, 180, 23, 1808, 1783, 1550, 1044, 1005, 1807, 1324, 1432, 1379, 1745, 1203, 1936, 1946, 1820, 1908, 1017, 264, 795]. Fierz [1811, 556]. fifth [652]. fifth-order [652]. film [504]. Filtered [861]. finding [1334]. Fine [729, 1764, 932, 1253]. Finetti [1951, 1140]. Finite [1298, 316, 547, 813, 675, 1002, 1183, 713, 1913, 1854, 1401, 491, 628, 1633, 890, 123, 939, 450, 1477, 522, 1501, 106, 286, 327, 917, 985, 257, 1431, 129, 48, 1655, 457, 302, 511, 1174]. Finite-dimensional [1183, 713, 123, 450, 522, 286, 327, 129, 511]. finite-gap [917]. finitely [104]. First [1362, 1541, 1660, 432, 177, 1382, 1173, 1187, 1488]. first-order [432]. Fisher [1750]. five [1047, 1220, 185, 24, 1592]. five-dimensional [185, 24, 1592]. Fixed [1225, 1938, 36, 1108, 488, 703, 1875]. flat [1923, 716, 624, 187, 1455]. flatness [1344]. flavor [1877, 167, 638]. flights [1180, 1022]. flop [691]. flow [568, 508, 1720, 1470, 1170, 1692, 1215]. flows [1018, 1521, 1169, 1733, 1308, 1385]. Fluctuation [1905, 887]. fluctuations [1780, 1244, 1501, 346, 914, 876]. fluid [546, 1205, 1925, 1291, 1758, 230, 34, 272, 1164, 928, 1266, 1082, 756]. fluids [402, 983, 959, 584, 65]. Flux [1281, 1195]. Flux-across-surfaces [1281]. foam [1207]. Fock [724, 1860, 801, 1095, 967, 333]. focusing [1087, 281, 774]. Fokker [1386, 399, 505, 650, 1021]. fold [450, 1072, 490, 789]. foldability [755]. Folding [744, 340, 156]. Foldy [1432]. foliation [1202]. foliations [1529, 1008, 1009, 1171]. force [1583, 1881, 395, 1840, 1775, 1017]. force-free [395, 1017]. forced [1018, 1071]. forces [1838]. foreign [213]. form [970, 389, 1361, 1318, 715, 1939, 62, 199, 299, 58, 1179, 1242]. formalism [1356, 1643, 1323, 698, 1002, 1657, 1731]. formation [1403, 1503, 906]. forms [1297, 629, 729, 1381, 174]. formula [654, 1569, 557, 1373, 1527, 1149, 399, 803, 1184, 29]. formulas [444, 962, 189, 1423, 857, 449, 465, 269, 160, 830, 533]. formulation [120, 1660, 318, 964, 726, 427, 1869, 250, 953, 274, 109, 1490]. fouled [1074]. foundation [433]. Foundations [870]. Four [1243, 821, 1422, 1138, 777, 1411, 630, 390, 1435, 1436, 1640, 904, 632, 1479, 498, 986, 264]. four-body [1479]. Four-dimensional [821, 777, 630, 1640, 498]. four-manifolds [904]. four-metrics [986]. Four-particle [1243]. four-state [1138]. four-vector [264]. Fourier [1782, 1553, 1277]. fourth [1573, 597]. fractal [1078, 860]. Fractional [423, 629, 1393, 1394, 1469, 258, 505, 807, 1838, 1831, 128, 1386]. Fragile [1906]. frame [970, 645, 396, 307]. frames [888, 1833]. framework [1603, 606, 1380, 1484, 472]. Fredholm [494]. Free [30, 1634, 650, 1297, 395, 1248, 17, 780, 344, 1292, 983, 1502, 1629, 1748, 787, 1082, 1796, 1017]. free-fermion [1629]. free-particle [1748]. freedom [400]. freely [339, 113]. frequencies [1923]. frequency [358, 592, 593, 1641]. Friedman [1096, 957]. Friedmann [718, 514]. Friedrichs [21, 1006]. Frobenius [802, 1868]. Frölicher [115]. front [211, 1203, 1177]. FRW [1487]. Fuchs [826]. full [735]. function [1359, 1089, 1780, 1768, 1844, 1613, 1345, 445, 446, 693, 927, 1839, 1832, 1412, 44, 1625, 1466, 451, 791, 608, 399, 863, 1653, 1350, 24, 879, 1745, 611, 708, 1231]. Functional [120, 233, 701, 971, 154, 1427, 915, 1209, 288]. Functionals [386, 461]. Functions [203, 1224, 1342, 1040, 1088, 1370, 1671, 1273, 1570, 827, 1832, 641, 1145, 449, 1393, 1394, 1916, 799, 1742, 1024, 1025, 106, 1092, 1365, 1396, 1650, 997, 767, 379, 1576, 1697, 1766, 583, 1064, 863, 1748, 254, 72, 1940, 636, 1351, 770, 1103, 1339, 1277, 1172, 1813, 304, 1113]. Fundamental [1544, 1934, 670, 1463, 1765, 225, 1682]. further [1560]. Fusion [1116, 446, 1367]. future [231, 673]. fuzzy [972].

G [255]. G. [1255]. Gabor [1833]. Galerkin [1679]. Galilean [1322]. Galilei [278]. gap [88, 845, 917, 350]. gaps [462]. gas [1792, 1500, 791, 1722, 1697, 1646, 1847, 1897, 1723]. gases [1564, 1832, 469]. gate [206, 28]. gates [1135, 1039, 1128]. Gaudin [709, 1216, 1146, 782, 1295, 431, 1491]. Gauge [1561, 1316, 1606, 305, 321, 837, 389, 420, 1318, 1535, 1263, 675, 427, 1869, 544, 995, 428, 1927, 1489, 1002, 160, 1556, 679, 945, 1518, 109, 1005, 821, 1430, 1046, 949, 264]. Gauge-invariant [305, 544]. gauged [777, 694, 1223]. gauges [106]. Gauss [134, 974, 1410]. Gaussian [1427, 843, 329, 200, 608, 1431, 505]. Gegenbauer [386]. gelation [1085]. Gel’fand [449, 581, 1805, 1506, 1118]. genera [268, 1597]. General [1757, 1590, 1353, 1425, 1523, 335, 1105, 1756, 806, 534, 1806, 761, 1486, 242, 1404, 227, 1845, 886, 101, 1591, 427, 899, 1708, 624, 1164, 1652, 960, 222, 1654, 26, 1204, 804, 653, 301]. Generalization [1208, 551, 324, 1782, 1769, 463, 1800, 183, 52, 1149, 921, 455, 554, 770]. generalizations [1596, 1363, 1750, 361, 1211]. Generalized [980, 989, 695, 840, 911, 993, 1051, 1488, 451, 142, 819, 399, 1408, 299, 1381, 364, 148, 472, 1341, 189, 1370, 574, 1362, 1564, 390, 1634, 519, 1100, 344, 855, 137, 410, 860, 932, 1935, 452, 590, 1549, 1609, 1689, 1942, 561, 314, 1314, 1653, 109, 353, 730, 385, 1763, 434, 650, 1021, 1745, 1610, 979, 511, 1737]. generalizing [514]. generate [1219, 1538]. generated [1345, 1732]. Generating [445, 446, 641, 1740, 1565, 1522]. Generating-function [445, 446]. Generation [1447, 1402, 925]. generator [280]. generators [1640, 1095, 967]. Generic [102, 890]. genetic [312, 313]. genus [580, 1367]. Geodesic [1720, 1735, 732, 1308, 183, 508, 1162, 1459, 1164, 474, 124, 62, 1593]. geodesically [437]. Geometric [866, 1246, 1570, 425, 1392, 1175, 344, 1034, 589, 1337, 477, 86, 753, 1693, 1208, 902, 743, 964, 592, 33, 1521, 427, 869, 1478, 1275, 1301, 1420, 1358, 1588, 1453, 1444]. Geometrical [1406, 292, 903, 126, 1908, 1060]. geometries [1885, 1669, 1786, 1941, 1918]. Geometry [432, 1293, 1856, 23, 1724, 1866, 210, 1344, 1262, 406, 1603, 1870, 8, 73, 408, 575, 542, 224, 1268, 1364, 1504, 1805, 958, 996, 225, 717, 1536, 1109, 226, 876, 1695, 667, 109, 353, 1749, 416, 168, 986, 1858, 968, 950, 1320]. geophysical [1018]. Gepner [1597]. Gerdjikov [450]. Gevrey [251]. Gibbs [78, 1722]. Ginzburg [424, 1878, 915, 1562, 1329]. given [558]. Givental [1569, 1914]. Glaser [1484]. glass [1244, 1086, 755]. Glasser [879]. Glauber [95]. Global [910, 1616, 1008, 1009, 915, 1840, 1121, 1480, 1903, 527, 927, 890, 1200, 817, 1725, 283, 1599, 974, 978]. globally [1712]. gluing [960]. gluino [1471]. GMS [919]. go [1441]. Goldberg [368]. Golden [488, 1942]. goldfish [1769, 1921]. Goldstone [695, 1427]. Gordan [277, 479, 1701, 384]. Gordon [651, 1041, 1446, 1326, 381, 1855, 1339, 1167]. Gordon/massive [980]. Goto [1920, 1825, 1158]. Goursat [1614, 933]. Gowdy [1740]. GR [369]. graded [285, 1783, 1081]. Gradient [842]. grading [932]. gradings [729, 1253, 1764]. graining [719]. grammars [253]. graph [90, 1900]. Graphical [791, 120, 78]. graphs [1152, 100, 1260, 1734, 1726, 590, 75, 571, 1550]. Grassmann [1612, 1578]. Grassmannian [1406, 1042]. gratings [453]. gravitating [601, 567, 936, 928, 65]. gravitation [1891, 717, 1165]. Gravitational [2, 1558, 603, 1590, 1206, 1843, 1380, 1165, 1689, 1294, 1456]. graviting [1589]. gravitomagnetic [797]. gravity [894, 1354, 714, 1237, 1411, 1096, 957, 368, 1412, 184, 1801, 1591, 370, 763, 798, 373, 1487, 572, 403, 228, 186, 1010, 1908, 1207]. Gravoelectric [66]. Green [1768, 927, 1742, 1527, 106, 1766, 399, 24, 879, 1351]. grids [90]. Gröbner [1667]. Ground [527, 699, 53, 589, 1760]. Group [153, 464, 1516, 1429, 1752, 1744, 954, 1816, 625, 640, 1698, 1615, 1814, 448, 266, 1647, 1413, 517, 136, 630, 508, 1364, 1504, 43, 1939, 1331, 1263, 766, 1922, 1062, 392, 1720, 1, 1256, 1821, 296, 1496, 580, 633, 1691, 1181, 1689, 1743, 1336, 145, 1759, 1588, 1799, 1309, 1199, 1813, 205, 1699]. Group-cohomology [1752]. Group-invariant [1429]. groupoids [1872]. Groups [1856, 1889, 901, 1857, 1247, 1300, 481, 1417, 765, 1333, 521, 137, 411, 1443, 143, 226, 144, 1631, 1005, 145, 261, 617, 1229]. grow [568]. growth [1298, 1608, 1815]. Guichard [181]. guides [1709, 1710]. Günther [1731]. gyroscopes [441].

H [1542, 1255, 650]. H-theorems [650]. Hadamard [447, 260]. Hagedorn [665]. Hahn [46]. half [1423, 1261, 1405, 1601, 1794]. half-integer [1794]. half-line [1423]. Hall [1676, 494, 807, 1144, 18, 1477, 706]. Halphen [1362]. Hamilton [1186, 1187, 1271, 890, 1606, 1669, 1786]. Hamiltonian [1757, 340, 1204, 237, 1594, 923, 627, 507, 1048, 866, 1208, 1511, 32, 1356, 1011, 432, 1643, 123, 450, 758, 1323, 156, 1175, 177, 1374, 1805, 535, 1002, 844, 598, 953, 599, 173, 400, 954, 1695, 180, 314, 877, 363, 127, 1490, 1170, 826, 1324, 128, 975, 1184, 1081, 1082, 1659, 1491, 1171, 979, 511, 1174]. Hamiltonians [1310, 506, 620, 808, 1475, 1099, 305, 1690, 1228, 746, 162, 1252, 1029, 468, 1514, 976, 1074]. handling [11, 737]. Hanson [1508]. Harish [1862, 1852]. Harish-Chandra [1852]. Harmonic [583, 1936, 1946, 4, 1580, 1111, 1370, 79, 191, 1414, 343, 1275, 1301, 1454, 1420, 871, 1459, 872, 578, 1849, 1766, 403, 891, 1384, 1074, 977, 920]. harmonics [1, 1256, 1302, 1468, 1799]. harmony [537]. Harper [1534, 488, 1942]. Harry [1238, 1532, 1793, 1720, 1773]. Hartree [1097, 703, 704, 705]. Hartwig [1750]. Hasegawa [1725]. Hausdorff [1400, 147, 635]. having [165, 166]. Hawking [1320]. heat [1730, 929, 517, 1060, 519, 1797]. heat-kernel [519]. heating [1809]. Hegerfeldt [1956, 357]. Heiles [979]. Heisenberg [1067, 1190, 1625, 1178, 1502, 880, 1309]. helically [1622, 1950]. helically-reduced [1622, 1950]. helicity [325]. Helmholtz [774, 281]. Hénon [979]. hereditary [1525]. Herglotz [1488]. Hermite [1370]. Hermitian [1472, 627, 200, 877, 1029, 148, 1514, 1659]. Hermiticity [1310, 1441, 1252, 877, 1029, 1102, 1314, 1197]. Hessian [1459]. Hestenes [1808, 1807]. Hexagonal [1458]. hidden [1457]. Hiding [1139]. hierarchies [1525, 311, 475, 1093, 1014, 1917, 732, 365]. Hierarchy [1041, 1415, 1575, 1451, 1730, 1238, 1532, 1792, 123, 1721, 1805, 1598, 1489, 1450, 37, 1093, 1773, 1270, 511, 1172, 315, 623, 893]. Higgs [938, 1200, 695, 810, 559, 1639, 795, 564]. Higgs-Monopole [938]. High [1839, 543, 592, 593, 952, 1243, 1292, 1244, 81, 83, 335]. high-dimensional [81]. High-frequency [592, 593]. high-temperature [1243]. Higher [1423, 1869, 306, 324, 894, 516, 927, 1475, 1800, 1482, 995, 597, 56, 934, 798, 1487, 1912, 1367, 1240, 1007, 1732]. higher-derivative [894]. higher-dimensional [56, 798, 1007]. higher-genus [1367]. Higher-order [306, 516, 1487]. highest [338, 149]. Highly [116, 172, 359, 1729]. Hilbert [1387, 242, 1143, 1604, 1605, 1801, 763, 412, 609, 1523, 471, 1309]. Hilbertian [1507]. Historical [225]. Histories [1761, 50, 692, 1157, 1045]. history [1202, 1044]. Hitchin [268, 452]. HKT [1090]. HKT-structures [1090]. Hodge [332]. Hofstadter [846]. Hohenberg [125]. hole [131, 684, 677, 1749]. holes [1321, 1923, 438, 1318, 1938, 141, 275, 1592, 442]. Holm [508, 196, 486]. Holography [1636, 1777]. holomorphic [1218, 157, 1251, 986, 29]. holonomy [632, 1627, 144]. homogeneity [701]. homogeneous [1321, 808, 480, 993, 1505, 1449, 1440, 1148, 1651, 718, 65, 1103]. homogenization [868]. Homological [1870]. Homotopy [1717, 706]. Hooft [719]. Hopf [1567, 1873, 1568, 1544, 1648, 52, 45, 1826, 287, 1867]. hopping [1548]. horizon [1925, 1637]. horizons [1928]. horizontal [1052]. horns [1279]. hot [1927]. Hounkonnou [1255]. Hubbard [1070, 644, 789]. Hudson [1755]. Hughston [552, 150]. Hunter [1793]. Hurwitz [1089]. hybrid [573, 1522]. Hydrodynamic [1776, 1323, 577, 1499]. hydrodynamical [1107]. hydrogen [1150, 1340, 1312]. hydrogenic [386]. hydromagnetic [1053]. Hyper [904, 1457]. Hyper-complex [904]. hyper-Kähler [1457]. Hyperbolic [10, 835, 1909, 799, 584, 891, 1384, 856, 854]. hyperboloids [1224, 1065]. hypercube [1291]. Hyperelliptic [1692, 771]. hyperfinite [1623]. Hyperfunction [500]. Hypergeometric [72, 1359, 827, 449, 799, 830, 863, 1653, 466, 611]. Hyperspherical [351]. hypersurface [718, 1759, 510]. hypersurfaces [1891, 736].

Icosahedral [1464]. Ideal [469, 1832, 1214, 80, 885, 1082]. Ideally [1716, 1955]. ideas [740]. identical [1544, 1036, 1686]. identify [1769, 1921]. identities [857, 905, 1617, 1755, 1754, 1092, 1365, 198, 553]. identity [803]. II [1342, 479, 805, 1704, 1187, 656, 895, 1190, 715, 1678, 1605, 1394, 313, 593, 185, 1025, 81, 1009, 53, 327, 763, 1484, 349, 1710, 677, 1029, 166, 704, 1384, 1081, 1426, 734]. IIB [681]. III [806, 776, 1106, 274, 1102, 705, 1082, 735]. Ill [319, 1486]. Ill-posedness [319]. illustration [805, 806]. Image [1215]. Imaginary [1536]. immersed [1542, 1063, 1166]. impenetrable [1500, 1713, 1949]. implemented [1131]. implications [814]. implicit [1857]. Import [1651]. imposition [695]. imprimitivity [279]. Improved [1484, 92]. improvement [1040]. impurities [169]. impurity [1705, 1101]. incipient [81]. including [1522]. incomplete [222]. incompressible [1733]. increase [1372]. Indecomposable [456, 1417]. indefinite [1539, 1510]. independence [1931, 9, 49, 19, 935]. independent [1488, 702, 1055]. index [1671, 37]. indices [494, 1621]. indirect [21]. indistinguishability [1708]. Induced [278, 380, 1305, 991, 1189, 346, 197]. inducing [715]. Inelastic [339, 113]. Inequalities [1130, 1945, 461, 1422, 356]. inequality [1517, 107, 346]. inertia [594]. inertial [334]. Infeld [672, 1841, 1616, 1213]. Infinite [1239, 1067, 880, 460, 639, 1298, 539, 540, 754, 81, 286, 327, 614, 1797]. infinite-dimensional [286, 327]. infinite-order [754]. infinitely [1279, 1110]. infinity [132]. Information [155, 1344, 1258, 1478, 996]. informationally [1762]. infrared [347, 1067, 880, 1379, 1745]. inheriting [1205]. Inhomogeneous [1889, 1161, 1057, 171, 395, 915, 1442, 1397, 1652, 1522, 1673, 1766, 1611]. Initial [516, 1616, 898, 486, 1842, 1789, 1665]. Initial-value [516]. inner [803, 1196]. instant [970]. Instanton [256, 1819, 1259, 501]. Instantons [777, 1824, 55, 1689]. integer [1794]. Integrability [476, 1596, 1357, 656, 1776, 234, 1091, 1278, 1702, 1110, 661, 730, 1050, 1308, 1174]. Integrable [1445, 1746, 1859, 450, 529, 1753, 1070, 928, 559, 180, 1499, 1610, 506, 923, 507, 1048, 1792, 1475, 279, 134, 294, 778, 123, 758, 176, 1175, 1049, 360, 1440, 1374, 1627, 269, 597, 1598, 1467, 1283, 37, 1093, 196, 56, 1522, 660, 1269, 1013, 582, 1014, 1497, 955, 1453, 1712, 1490, 1692, 826, 956, 1523, 129, 713, 771, 1917, 1270, 1492, 854, 511, 1172, 387, 1325, 365, 893, 400]. Integral [1844, 804, 805, 806, 596, 465, 1869, 1852, 192, 1905, 163, 1181, 608, 562, 354, 1578, 847, 1339]. integral-difference [608]. Integrals [1062, 1032, 625, 805, 806, 1187, 971, 1362, 603, 725, 1755, 1259, 605, 177, 1395, 1227, 1690, 1264, 1628, 284, 484, 1754, 1353, 1382, 1064, 1184, 1795, 1947, 1173]. Integrated [81]. Integration [1182, 979, 315, 1615, 1671, 233, 701, 731]. integrations [1814]. integrators [1356, 1693]. intelligent [1837, 911]. Interacting [1290, 1566, 1571, 538, 1646, 1760]. interaction [16, 794, 1739, 1722, 759, 58, 1431, 713, 848, 2, 1231]. interactions [1561, 1844, 628, 1352, 1853, 103, 104, 496, 225, 1176, 1235, 1255, 1315, 25, 1015]. Interior [1733]. interlacing [374]. intermediate [925]. intermittency [230]. Internal [1038, 1561, 1809]. interpolation [161]. interpretation [427]. interrelated [165]. Intersecting [1891, 685, 10, 371, 40, 835]. intersection [1458]. intersite [1015]. Intertwined [700]. intertwining [711]. interval [424, 1031]. intricacies [1927]. introduced [1703]. Introduction [1856]. intuition [213]. invariance [1152, 642, 746, 250, 287, 1045]. Invariant [1615, 1814, 392, 966, 362, 731, 1529, 1920, 190, 1318, 1411, 897, 1861, 1608, 305, 544, 1429, 1419, 1825, 589, 1201, 370, 54, 1575, 767, 1484, 917, 1691, 954, 1912, 945, 1855, 109, 1182, 1882, 262, 881, 1277]. invariants [1218, 866, 1208, 1299, 895, 1106, 3, 637, 45, 71, 1079, 36, 1108, 160, 1727, 1695, 1518, 363, 501, 1882, 586, 987]. Inverse [461, 929, 171, 1791, 502, 265, 1452, 358, 175, 754, 1495, 1456, 298, 362, 38, 231, 1728, 1577, 119, 1794, 1177]. inverse-square [265]. inversion [1630]. Investigation [844, 1341, 1269, 271]. Involution [1324, 1436]. involutive [1169, 511]. involutivity [732]. involving [1092, 288]. ion [30]. Ionization [215]. Iper [885]. Iper-ideal [885]. irrationals [1534]. Irreducible [985, 1533, 522, 1941, 728, 286, 327, 581, 1743, 999, 1699]. irregular [1485]. irreversible [433, 649, 618]. isentropic [1792]. Ising [1243, 1853, 77, 922, 356, 1020]. isolated [1099]. isometries [796]. isomonodromic [194, 1491]. Isomonodromy [1524]. isomorphic [165, 166]. Isoperimetric [325]. isoscalar [1248]. Isospectral [54, 992, 700, 1656]. Isospectrality [1441]. isothermal [928]. isotope [1376]. isotropic [589, 1849, 920]. Isotropization [402]. isotypic [342]. IV [1346]. Ivanov [450]. Iwasawa [454].

J [552, 1580, 1757, 1756, 491, 459, 207, 73, 1951, 1810, 1096, 637, 1676, 1811, 339, 1755, 936, 340, 49, 1952, 1254, 1542, 1420, 1955, 492, 1421, 1948, 737, 1256, 1369, 1278, 835, 774, 1399, 738, 739, 1066, 553, 1310, 1953, 638, 1340, 554, 1954, 208, 209, 1400, 1945, 293, 664, 1949, 1067, 1944, 1729, 1956, 1946, 1950, 1444, 493, 206, 1947]. Jacobi [1186, 1187, 1271, 890, 1606, 1669, 1786, 800, 1276, 1092, 1365, 1537]. Jacobson [1095, 967]. Janet [1435]. Jaynes [1341, 1151]. jet [754]. joint [307, 1882]. Jordan [52]. Jordanian [1724]. Jost [1896]. Journal [1255]. jump [1073]. junction [1590].

K-property [118]. K-theory [666]. K3 [1409, 1335]. Kac [835, 10, 285, 454, 289, 832, 29, 1909]. Kac-modules [289, 832]. Kadanoff [434]. Kadomtsev [720, 475, 982, 436, 33, 37, 622, 1691, 1763]. Kähler [808, 1457, 490]. Kaluza [682, 855, 876]. kappa [672]. kappa-symmetry [672]. Kármán [504]. Kaup [1926, 1168]. Kawasaki [88]. KdV [360, 1770, 1357, 487, 38, 1448, 1668, 623]. KdV-type [1770, 360]. Keller [1083]. Kemmer [769]. Kepler [474, 1496, 1849, 430, 1549, 537, 24]. Keplerian [543]. kernel [1730, 1872, 405, 1243, 1737, 519]. kernels [189, 517]. Kerr [762, 66, 1162, 442, 1320]. Kersten [1357]. Ketoja [1942]. key [1138]. Killing [68, 1536, 1789, 1882, 1928]. kinematic [273]. kinematics [970]. Kinetic [433, 1564, 1779, 1512, 310, 701, 885]. kinetics [1670]. Kirkpatrick [1244, 1086]. Klein [1041, 682, 1326, 876, 1855, 1167]. Knizhnik [1216]. knot [546]. Kodaira [1217]. Koenigs [1392]. Kolakoski [1307]. Kolakoski- [1307]. Kolmogorov [1599, 107, 921]. Kolmogorov-like [921]. Koornwinder [803]. Korteweg [271, 655, 1730, 710, 1525, 1334, 600, 1050, 1240]. Kowalevski [1382]. KP [1451, 311, 1489, 1450, 1109, 1014]. KPI [1452]. Krasil’shchik [1357]. Kraus [1286]. Krein [1154]. Krichever [42]. Kubo [1527]. Kupershmidt [1168].

ladder [1859, 1610]. Ladik [1521]. Lagrange [726, 212]. Lagrangian [225, 757, 1643, 114, 326, 786, 510, 621]. Lagrangians [1525, 1633, 1642, 955, 881, 112]. Lamb [1337]. Lamé [1772, 984, 845]. Lampard [550]. Lanczos [182, 1435, 1436, 1640, 1574, 1438]. Landau [424, 1878, 817, 940, 843, 915, 1562, 1329, 833, 968, 1903]. Lane [410]. Langer [1819]. language [587, 740]. Laplace [578, 1302, 933, 385]. Laplacian [405, 1778, 1553]. Large [80, 751, 1046, 949, 691, 1117, 1018, 424, 1260, 1726, 1042, 359, 1052, 1729, 668]. Large- [751]. large-order [1117]. large-scale [1018]. larger [79]. Lascoux [1935]. laser [991]. laser-induced [991]. laterally [245, 878]. Lattice [1632, 453, 837, 1772, 507, 1827, 793, 1276, 1599, 622, 1787, 646, 1662, 821, 730, 1306, 1490, 1431, 879, 892, 1848, 833, 94]. lattice-field [507]. lattices [1474, 1364, 1504, 70, 859, 176, 1013, 926, 1919, 1288]. law [1541, 1495, 974, 110, 469, 1410, 1151]. laws [1080, 598, 181, 1239]. Lax [1525, 1595, 792, 652, 1357, 712, 1448, 471, 1006]. layer [1117, 1503, 1177]. layers [1672, 859, 939, 878]. leading [1563]. Leclerc [1935]. Lee [1584, 1815, 1006]. Legendre [1399, 1444, 1394, 1303, 1396, 1358]. Leibniz [1778]. Leibnizian [1322]. lemma [1274]. length [1765, 1645, 393, 1620]. lenses [367]. lensing [603, 1206, 1380]. Lenz [884]. Lerch [1832]. Level [338, 894, 419, 991, 459, 241, 1680, 55, 225, 1071, 968, 27, 304]. Level-one [338]. levels [491, 843, 257, 1103]. Levin [1703]. Levin-type [1703]. Lévy [1180, 505, 1670, 1022]. Lewis [440]. Lie [1755, 553, 293, 1872, 1232, 507, 723, 1298, 852, 1299, 601, 765, 43, 312, 313, 632, 282, 535, 1754, 606, 728, 1649, 1864, 140, 285, 633, 198, 661, 1673, 1382, 1027, 523, 1631, 1338, 1901, 1547, 769, 1759, 1829, 1183, 456, 7, 1050, 1804, 65, 1210, 731, 532, 771, 617, 48, 289, 832, 772, 337, 458]. Lie-algebra [1829]. Lie-series [293, 7]. Lieb [1541]. Liénard [1526]. Lifshits [1910]. Lifshitz [544, 817, 1038, 1366, 833]. liftings [560]. Light [391, 679]. Light-cone [391, 679]. lightlike [1820]. like [761, 1096, 957, 921, 525, 1240, 472, 736, 1258, 1077]. limit [837, 692, 1460, 613, 741, 1876, 1481, 1404, 101, 1191, 690, 1244, 81, 1260, 1042, 310, 47, 1855, 704, 1307, 1046, 949]. limit-periodic [1307]. Limitations [291]. limiting [482]. Limits [1329, 1632, 1040, 776, 1346, 1347, 69, 868, 267, 973, 718, 636, 1577, 1773, 1797]. Lindblad [1286]. line [775, 1087, 1423, 1601, 701, 1304, 1550, 1767, 1712]. Linear [412, 458, 1208, 808, 395, 234, 1890, 1849, 203, 772, 1936, 1946, 592]. linearity [1686]. linearization [266]. linearized [368, 57, 387]. linearly [1323]. lines [3, 637, 859]. link [886, 287, 987]. links [197, 1325]. Liouville [654, 1613, 374, 1048, 1893, 400, 826, 1031, 1174]. Liouville-integrable [400]. liquid [1780, 1503, 1721]. Liu [1542]. LKdV [38]. LKR [1602]. LLT [1935]. Local [39, 1005, 336, 102, 125, 1076, 500]. Locality [17, 1736, 1288]. localization [1371, 1474, 1548, 1378, 355, 1956, 357]. localizations [909]. localized [116, 1602, 1492]. locally [1712, 1655]. located [169]. loci [890]. locus [698]. Logarithmic [539]. logic [206, 28]. Long [85, 1423, 175, 1424]. Long-range [85, 1423, 1424]. long-wave [175]. longitudinal [1001, 264]. look [115, 1026, 1908]. loop [625, 696, 1801, 9, 49, 19, 794, 1598, 580, 1917]. loops [359, 1729]. Lorentz [263, 1434, 448, 1881, 1075, 1824, 766, 344, 106, 1484, 1181, 881, 1277]. Lorentz-violating [1075]. Lorentzian [761, 406, 1885, 1536, 1559, 1656]. Lorenz [966, 1497, 1173]. loss [1802]. Lotka [1276]. Low [1654, 268, 399, 1112, 1361, 1833]. low-energy [399]. Lower [1404, 1133, 1347, 1132, 944, 1823]. lowest [940, 968]. Lüders [1582]. lump [1459]. Lund [981]. Lyapunov [1012, 1904].

M [1161, 689, 687, 688]. M-theory [1161]. M. [1255]. M5 [690]. M5-brane [690]. Macdonald [803]. machines [714]. macromolecules [1609]. Magaard [1220]. Magnetic [588, 1229, 836, 405, 612, 1001, 1746, 1910, 596, 1839, 867, 1312, 1494, 779, 993, 529, 101, 1742, 1349, 172, 842, 699, 1261, 1101, 164, 1711, 180, 1550, 945, 169, 1883, 1553, 1103, 1796]. magnetic-solenoid [1742]. Magnetoencephalography [1268]. magnetohydrodynamics [1214, 362]. Magnification [603]. Magnus [202]. Mainardi [134]. Majorana [317, 193]. Mandula [733]. Manev [1054]. manifold [1406, 334, 291]. manifolds [188, 1612, 620, 406, 808, 6, 207, 1279, 757, 390, 904, 44, 727, 1752, 1091, 1278, 296, 1251, 396, 960, 14, 209, 1559, 111, 1510, 621]. many [1769, 1921, 1413, 104, 1512, 1722, 1110, 170]. many-body [1769, 1921, 1512, 1722, 170]. map [1088, 52, 840, 912, 622, 1731, 1170, 311]. mapping [243, 199]. maps [1873, 1413, 1012, 1063, 1275, 1301, 1454, 1420, 1459, 535, 906, 1575, 270, 118, 998, 332, 1692, 977]. Marchenko [1613]. Marciniak [730, 1306]. Marginal [1932, 1422]. marginals [1875]. Markov [931, 1501, 1527, 95]. Markovian [1401, 1297, 310, 434, 756]. Martin [120]. Maryland [1257, 249]. mas [739]. mass [970, 427, 1292, 658, 738, 12, 1658, 1662, 1020]. mass-dependent [427]. Massey [197]. Massive [1254, 1483, 345, 1320, 980, 527, 427, 1378]. Master [44, 1401, 451, 756, 1198]. Math [552, 1580, 1757, 1756, 491, 459, 207, 73, 1951, 1810, 1096, 637, 1676, 1811, 339, 1755, 936, 340, 49, 1952, 1254, 1542, 1420, 1955, 492, 1421, 1948, 737, 1256, 1369, 1278, 835, 774, 1399, 738, 739, 1066, 553, 1310, 1953, 638, 1340, 554, 1954, 208, 209, 1400, 1945, 293, 664, 1949, 1067, 1944, 1729, 1956, 1946, 1950, 1444, 493, 206, 1947]. Mathematical [1556, 1255, 1703, 1085, 740, 121, 550, 1675, 619]. Mathieu [1370]. matrices [1473, 1472, 1572, 556, 1811, 548, 1415, 1721, 932, 1735, 1830, 140, 711, 590, 634, 1523, 1368, 1579]. Matrix [1222, 659, 775, 937, 1698, 1272, 723, 1482, 1680, 390, 1500, 465, 1286, 35, 873, 138, 728, 1817, 1398, 1537, 200, 201, 524, 785, 1350, 1182, 1296, 1940, 1424, 434, 129, 1554, 968, 533, 280]. matrix-valued [1817]. Matroid [467, 1638]. matter [1343, 1877, 492, 167, 638, 1886, 1718, 2, 493]. Maximal [528]. Maximization [943]. Maximum [1012, 1205]. Maxwell [1944, 1903, 389, 1603, 1200, 1737, 592, 593, 513, 1635, 414, 1743, 1665, 515, 1520]. Maxwellians [1903]. mean [239, 1707, 1086, 1008, 1009, 1066, 875, 488, 1942]. mean-field [239, 1707]. means [804, 805, 806, 1271, 441, 1339]. measure [1827, 1677, 1349, 860, 1575, 1313, 297, 1747]. measurements [816, 499, 1762]. measures [1533, 1647, 1907, 1851, 1120, 392, 1284, 1182, 755, 822, 647]. Measuring [441]. mechanical [988, 1034, 306, 126, 820, 1713, 1949, 471]. Mechanics [409, 1563, 740, 152, 1836, 1476, 1678, 60, 473, 1604, 1605, 840, 939, 567, 936, 1051, 1100, 230, 869, 1227, 1107, 1607, 1282, 103, 496, 1830, 1682, 747, 173, 1609, 1358, 1439, 1695, 1631, 466, 945, 1924, 1953, 1675, 354, 787, 1234, 1195, 1930, 1453, 1425, 7, 293, 397, 300, 1072, 1235, 1255, 532, 86, 111, 885, 1444, 170, 1693]. mechanism [1563, 695]. media [171, 1083, 592, 593, 1304, 414, 1611]. medium [358]. Mel’nikov [60]. memory [1136, 1900]. memoryless [1133]. mesons [1662]. mesoscopic [779]. messages [1139]. metallic [1493]. metaplectic [607]. Metastability [89]. Method [11, 737, 1148, 188, 1780, 120, 721, 804, 805, 806, 445, 446, 1564, 529, 1162, 994, 273, 1334, 55, 1093, 487, 1468, 201, 702, 1456, 978, 1328, 1410, 129, 515, 1812, 893, 205, 1260]. Methodology [161]. methods [120, 517, 1438, 960, 1350, 320, 1555]. Metric [1055, 810, 440, 690, 1091, 1278, 853, 373, 1510, 472]. metric-affine [373]. Metric-independent [1055]. metric-like [472]. Metrics [1249, 1473, 1434, 478, 1457, 632, 441, 1539, 986, 437]. MHD [1644, 1441, 1242]. MIC [1549, 24]. MIC-Kepler [1549]. Microlocal [1203]. MICZ [1849]. mid [845]. mid-band [845]. Mills [49, 559, 1639, 795, 564, 1445, 1612, 1316, 389, 316, 1824, 674, 244, 9, 19, 900, 1003, 250, 853, 819, 570, 749, 954, 1156, 255, 352, 1005, 321, 1908, 404, 276]. Mima [1725]. minimal [1331, 285, 47, 303]. minimally [572]. Minimization [1823]. minimizer [915]. minimizers [288]. Minkowski [1507, 795]. Mirror [1858, 1870]. Miura [311, 622]. mixed [393, 1131]. Mixing [90, 88, 1753]. mixtures [1114]. MKdV [1169, 1357]. mobile [1541]. Möbius [848]. mode [743, 1728]. model [235, 1341, 277, 479, 1243, 1584, 1402, 1257, 78, 1348, 480, 294, 1925, 1706, 813, 1853, 1696, 709, 1216, 77, 815, 1828, 1070, 1244, 1086, 53, 973, 35, 744, 1791, 249, 504, 21, 1178, 1502, 1874, 1548, 1057, 514, 782, 1295, 748, 1586, 644, 1537, 676, 677, 572, 849, 1547, 256, 353, 1639, 1266, 952, 215, 1019, 1329, 1897, 1020, 1307, 1006, 1936, 1946, 671, 76, 338, 1898, 789, 1207, 1906, 1151, 94, 1835]. modeling [1766]. models [263, 865, 980, 1571, 889, 419, 151, 777, 694, 1223, 1311, 824, 476, 792, 425, 390, 1222, 44, 541, 841, 1714, 1591, 1753, 1597, 1429, 1825, 1146, 1657, 103, 451, 728, 107, 768, 718, 514, 56, 1522, 783, 328, 125, 274, 201, 1487, 1629, 1430, 1940, 755, 616, 1740, 1072, 1235, 1236, 1074, 1610, 563, 1528]. modes [784, 117]. modification [622]. Modified [963, 1300, 720, 657, 311, 475, 982, 1737, 436, 360, 33, 107, 1050, 1763, 457]. Modifying [1078]. Modular [498, 1378, 355, 218, 1482, 344, 500]. modulated [1848]. module [1236]. modules [1387, 1862, 135, 547, 985, 1829, 48, 289, 832]. Moduli [1918, 1933, 1462, 72, 501, 490, 1461]. modulus [1570]. molecular [825, 111]. Moment [1354, 363, 1564, 635, 1400]. momenta [1481]. moments [441]. momentum [1262, 1841, 1785, 765, 583, 1731, 297, 662, 1559, 1887, 1719, 386]. monads [1868]. Monge [1270]. Monge/Riemann [1270]. Monodromy [1943, 838]. monomer [94]. monomials [1062]. Monopole [61, 1639, 883, 527, 927, 1468, 1462, 1519, 938]. Monopole-antimonopole [1639]. Monopoles [694, 1223, 1471, 1463, 900, 1319]. Monotone [1473]. Moody [835, 10, 285, 454, 1909]. Morita [1389]. morphisms [403]. Morse [1768]. Moser [1211, 824, 794]. motion [981, 811, 543, 780, 1690, 172, 474, 124, 65, 1198]. movable [983]. move [11, 737]. moves [1497]. Multi [807, 1525, 955, 1819, 758]. multi-Hamiltonian [758]. multi-instanton [1819]. Multi-Lagrangians [1525, 955]. Multi-resolution [807]. multicomponent [1925, 956, 338]. Multidimensional [839, 1033, 365, 235, 640, 1782, 603, 1776, 371, 834]. multifractal [921]. multigroup [1056]. multikink [412]. multilinear [1492]. multiloop [1064]. multiparticle [1121, 1686]. Multipartite [463]. multiplane [1380]. Multiple [1348, 962, 1231]. multiplet [312, 313]. multiplets [261]. Multiplicative [1563, 1700, 1186, 1187, 1878]. multiplicativity [1747]. multiplicities [285]. Multiplicity [71, 1248, 1189]. multiplicity-free [1248]. Multiply [478]. Multipole [1185, 441, 864]. Multiresolution [1345]. multiscale [376, 656, 1064]. multisoliton [1674]. multistate [1855]. Multisymplectic [1109, 432, 457]. multivalued [1492]. Multivariable [1619]. multivortex [810, 849].

N [1255, 1597]. Nambu [1920, 1427, 1825, 1107, 1692, 1158]. narrow [90]. Natanzon [1704]. natural [620, 472]. nature [1208]. Navier [362]. navigation [838]. NCG [783]. near [1460, 990, 132, 371, 944, 329, 1416, 1903]. near-optimal [990]. nearby [646]. nearly [956]. necessary [877, 656]. necklaces [857]. negative [1525, 602, 1156, 87, 1712]. nematic [1503]. Nested [1064]. nets [1392, 1013]. networks [100, 972, 1333, 1905, 874]. Neumann [1389, 1363, 1500, 918]. Neutral [99, 632]. neutrino [1877, 167, 638]. neutron [1056, 64]. Neveu [135]. Newell [1926]. Newman [132, 68]. Newton [1241]. Newtonian [1322, 1876, 34]. Nijenhuis [115]. Nikodym [1551]. nilgeometric [1936, 1946]. nilpotent [1219, 47]. NLS [1596]. no [1441]. no-go [1441]. Noether [1488, 428]. Noether-type [1488]. noise [1563, 971, 1878, 1565, 1670]. noises [505]. Non [341, 672, 131, 107, 667, 1901, 1287, 1473, 1738, 627, 1263, 177, 34, 200, 1029, 186, 434, 420, 675, 877, 505, 1659]. Non- [341]. Non-Abelian [672, 667, 1287, 420, 675]. Non-existence [131]. non-Gaussian [505]. non-Hamiltonian [177]. non-Hermitian [627, 200, 1029, 877, 1659]. Non-Kolmogorov [107]. non-Markovian [434]. non-monotone [1473]. non-Newtonian [34]. Non-null [1901]. non-Riemannian [186]. non-self-adjoint [1738]. non-semisimple [1263]. Nonadiabatic [1627, 1099]. nonamenable [75]. nonarithmetic [1412]. nonassociative [1866]. nonasymptotically [1923]. nonautonomous [772, 1668]. nonbinary [1039]. Noncanonical [763, 1048]. noncentral [1683]. nonclassical [722]. noncocommutative [1826]. noncolliding [1817]. Noncommutative [1724, 1360, 1262, 1603, 643, 666, 225, 1856, 919, 1039, 1409, 1870, 98, 657, 963, 1300, 1411, 224, 1144, 1676, 1688, 1771, 665, 872, 295, 717, 659, 1650, 1080, 226, 876, 23, 1156, 667, 1519, 168, 1940, 1858, 1461, 1656]. Noncommuting [868]. noncompact [1333, 296, 570]. nonconfocal [1169]. nonconservative [757, 584]. nonconstant [924]. nonconvex [1620]. nondecaying [720]. nondegenerate [1047]. nondiagonalizable [1252, 1310, 1514]. nondynamical [35]. Nonequilibrium [230]. nonextensive [1563, 1179, 1747]. nonextensivity [886]. nonholonomic [1800, 1693]. nonhomogeneous [1919]. nonideal [1897]. Nonintegrability [599, 1694, 1919, 314]. Nonintegrable [853]. Nonlinear [1836, 762, 376, 656, 1481, 332, 128, 1021, 1097, 1447, 232, 1595, 1418, 652, 483, 123, 234, 995, 269, 1770, 1250, 1369, 1283, 178, 1601, 1080, 1838, 328, 310, 925, 1674, 1675, 1296, 363, 1453, 456, 505, 956, 1523, 650, 334, 854, 1015, 1919, 593]. nonlinearities [1816]. nonlinearized [337]. Nonlocal [420, 1238, 125, 1323]. nonmetricity [373]. nonminimally [896]. nonparabolic [1658]. nonperturbative [702]. nonradiating [63]. Nonrelativistic [686, 894, 888, 1584, 152, 1884, 1372, 869, 1860, 1429, 53, 973, 1353]. nonresonant [7, 293]. nonseparable [814]. Nonsingular [1758, 709]. nonsmooth [478, 1663]. Nonstandard [524, 573, 1701, 145, 1930]. nonstatic [50]. Nonsymmetric [717, 365]. nontrivial [1672, 817]. nonuniform [779]. nonvacuum [898]. nonzero [37]. Nordström [131]. norm [1375, 1433]. norm-1-property [1375]. normal [615, 262]. Normalized [449]. norms [1507, 943]. Note [246, 1420, 1946, 1040, 860, 1548, 1462, 455, 554, 1255]. Nöther [569]. notion [183]. Novel [1453]. Novikov [1586, 42]. nuclear [1791]. nucleation [89]. Null [318, 1338, 964, 369, 492, 513, 1155, 1293, 72, 1901, 1077, 1593, 26, 493, 187]. null-correlation [72]. nullvectors [1242]. number [1347, 1660, 539, 540, 939, 77, 501, 976, 1797]. numbers [1645, 1924, 1953]. Numerical [627, 805, 806]. numerically [994]. NUT [1041].

O [777]. object [1775]. objects [1766, 688]. Observable [973, 1375, 429, 500]. observables [1581, 1785, 342, 909, 242, 1282, 645, 862, 786, 946, 455, 554, 619, 692]. obstructed [1061]. obtain [273]. obtained [1825]. obtainment [389]. Obukhov [186]. occupancies [1231]. octonions [1867]. odd [201]. ODE [1885]. ODEs [1675]. Oehme [1006]. Off [139, 566, 533]. Off-diagonal [139, 533]. Old [1617, 223]. One [495, 865, 1844, 15, 1142, 850, 929, 1271, 232, 1280, 696, 1517, 530, 538, 1070, 872, 1779, 55, 451, 791, 644, 108, 701, 1194, 571, 610, 110, 1894, 1329, 1900, 1230, 338, 1231]. One-dimensional [495, 865, 1844, 1142, 850, 232, 530, 1070, 1779, 55, 451, 791, 644, 108, 701, 571, 110, 1231]. one-loop [696]. one-particle [15]. one-soliton [929]. one-way [1900]. Ono [1674]. onset [460, 745]. onto [1912]. open [665, 784, 644, 221, 670, 1167]. operations [1225, 1851, 1551, 1131]. operator [654, 655, 1343, 1738, 1423, 462, 154, 1803, 546, 1533, 1647, 1893, 1907, 1851, 1916, 1100, 281, 1626, 870, 137, 1334, 578, 1302, 1202, 774, 1192, 296, 819, 594, 1658, 608, 561, 414, 1748, 1621, 169, 1424, 515, 1433, 1796, 1788, 935]. operators [1569, 1530, 1402, 1187, 42, 1219, 723, 6, 207, 1525, 325, 438, 1318, 1417, 1041, 631, 556, 1811, 678, 930, 1393, 1394, 529, 1323, 1627, 1349, 1441, 577, 1261, 1405, 370, 1512, 249, 485, 139, 251, 658, 1038, 767, 1209, 37, 782, 198, 553, 164, 331, 571, 1684, 1408, 1366, 1621, 203, 887, 1883, 1712, 1306, 999, 663, 515, 216, 262, 1686, 335, 848, 149, 1812, 304, 518, 776]. Oppenheimer [323]. Optical [1791, 1198]. optics [592, 1283]. Optimal [991, 1214, 393, 990, 302]. Optimally [518]. optimization [336]. orbifold [1688]. orbifolds [1409, 675]. orbital [789]. orbits [938, 1413, 176, 1597, 47, 920]. orchid [1942]. Order [862, 324, 516, 1423, 1117, 6, 207, 1800, 43, 652, 432, 1916, 543, 841, 1885, 754, 995, 597, 1227, 1690, 306, 965, 549, 1691, 62, 1487, 609, 1183, 892, 457, 322, 1240]. Order-automorphisms [862]. ordered [202]. ordering [252, 1673]. orderings [382]. orders [1022, 335]. ordinary [1455, 43, 1024]. orientation [888]. orientation-shape [888]. origin [18, 914, 168]. Ornstein [398, 1661]. Orthogonal [201, 901, 1892, 191, 984, 1191, 1062, 143, 1094, 1619, 785]. orthomodular [1736]. orthonormal [1023]. oscillating [1899]. oscillation [1877]. oscillations [167, 638]. oscillator [4, 1580, 1370, 191, 343, 942, 1421, 872, 159, 1849, 1619, 22, 1744, 1074, 920]. oscillators [1111, 640, 425, 780, 871, 1904, 891, 1384, 7, 293, 648, 969, 1199]. oscillatory [1208]. OSP [388]. other [1583, 984, 1867]. Outer [1162]. output [1129].

P [1288]. P-lattices [1288]. packet [1066, 875, 110, 1894]. packets [910]. Padé [161]. Painlevé [1455, 1458, 710, 1573, 716, 1618, 934, 1110, 274, 1336, 1539, 585]. pair [30, 1595, 696, 942, 1421, 1818, 1357, 712, 998]. pairs [1525, 792, 652]. palindrome [1311]. papers [1255]. parabolic [1678, 978]. parabosonic [52]. paradox [1956, 357]. parafermionic [52]. parafermions [1137]. Parallel [144, 1125, 1215]. parameter [420, 1570, 503, 1579, 203]. parameters [534, 1034, 270, 300]. Parametric [1558, 482, 1904, 380]. parametrically [1015]. parametrization [1412, 1181, 1043]. parametrized [973]. paraparticle [52]. Parker [122]. part [785, 592, 593]. Partial [1829, 506, 516, 375, 234, 269, 1084, 485, 1751, 1328, 1324, 365, 834]. Partially [1825, 1374, 1419]. Particle [223, 980, 1844, 15, 237, 1343, 1889, 1243, 1558, 534, 1544, 1884, 1281, 1706, 642, 815, 699, 1326, 1817, 701, 1294, 1748, 787, 1646, 1234, 261, 847, 976, 756]. particle-field [237]. particles [422, 761, 1494, 99, 1415, 1036, 591, 1378, 1432, 1686, 1555]. particular [1072, 834]. Partition [850, 791]. parts [25]. Past [231]. patches [477]. Path [971, 1264, 163, 297, 562, 1032, 625, 596, 1259, 1628, 1905, 847, 1184]. Paths [379, 725]. pathwise [1195]. pattern [1503]. patterns [1458]. Pauli [462, 1511, 1041, 1626, 932, 1405, 951, 280]. PDEs [376, 656, 1776, 956]. peculiarity [1156]. Penning [823]. Penrose [132, 68]. peptide [340, 156]. percolating [82]. Percolation [77, 82, 81, 83, 1611]. perfect [1205, 1291, 1847, 1122, 1854]. period [270]. period-doubling [270]. periodic [1343, 1558, 495, 1078, 1428, 483, 18, 131, 794, 812, 1446, 1349, 162, 1840, 414, 362, 1713, 1054, 1307]. permutation [1287]. perpetuum [1541]. perspective [50]. Perspectives [221]. Perturbation [780, 646, 615, 1552, 956, 627, 438, 1312, 346, 163, 330, 945, 335]. Perturbations [1760, 401, 1318, 939, 439, 196, 953, 1883, 215, 204]. Perturbative [1371, 1059, 1518]. perturbed [929, 535, 178]. Petiau [769]. Petrov [545]. Petviashvili [720, 475, 982, 436, 33, 37, 622, 1691, 1763]. Phase [1032, 342, 793, 465, 994, 251, 75, 1766, 1076, 1555, 692, 1360, 804, 805, 806, 1422, 1932, 494, 839, 1208, 153, 741, 1570, 808, 1660, 909, 742, 851, 1282, 781, 429, 846, 1759, 946, 1741, 826, 617, 1848]. Phase-amplitude [994]. Phase-integral [465, 804, 805, 806]. phase-modulated [1848]. phase-shift [742]. Phase-space [1766]. Phases [1023, 828, 77, 1406, 477]. phenomena [1558, 1446, 328, 944, 1327, 1021]. phenomenology [1006]. phenomenon [1833]. Phillips [471, 1006]. photon [542]. Phys [552, 1580, 1757, 1756, 491, 459, 207, 73, 1951, 1810, 1096, 637, 1676, 1811, 339, 1755, 936, 340, 49, 1952, 1254, 1542, 1420, 1955, 492, 1421, 1948, 737, 1256, 1369, 1278, 835, 774, 1399, 738, 739, 1066, 553, 1310, 1953, 638, 1340, 554, 1954, 208, 209, 1400, 1945, 293, 664, 1949, 1067, 1944, 1729, 1956, 1946, 1950, 1444, 493, 206, 1947]. physical [1089, 740, 809, 211, 1899, 225, 763, 329, 213]. physics [219, 546, 3, 1791, 1675, 146, 223, 948, 1255]. Picard [826]. piece [1317]. Piezoelectricity [1016]. Piezomagnetism [1016]. pinned [79]. Pinning [944]. Planar [883, 867, 1215, 1503, 884, 536]. Planck [1481, 1386, 399, 505, 650, 1021, 1151]. plane [130, 1560, 98, 1861, 1535, 796, 1249, 1449, 1261, 1405, 1036, 853, 954, 1886, 891, 1384, 1718, 1185, 2, 1641, 864]. planes [657]. Plasma [1053, 546]. Plethysms [1571, 327]. plus [16, 1516]. Podolsky [51]. Poincaré [1045, 1247, 1274, 1261, 1405, 1743, 975, 1929, 1699]. Poinsot [1497]. point [1212, 534, 628, 761, 939, 1003, 1335, 1382, 488, 998, 703, 1367, 1513, 1315, 751, 1515, 25, 848, 27]. point-like [761]. points [1225, 983, 1562, 1918, 270]. Poisson [235, 1624, 435, 1413, 475, 565, 1680, 1049, 800, 329, 1751, 610, 333, 1804, 65, 309, 437, 337]. polar [1507]. polarization [1377]. polarized [392]. polaron [239]. poles [339, 113]. Polyakov [67]. Polyhedral [1463]. polymer [1625, 541]. polynomial [1727, 201, 1518, 662, 76, 1929]. polynomials [1370, 1472, 1892, 448, 1573, 191, 984, 1191, 578, 1399, 1303, 46, 1094, 1619, 287, 803, 1621, 386]. polytopes [1116]. population [991]. Pöschl [639]. posed [1486]. posedness [319, 1242]. Position [1785]. Positive [1533, 1647, 1349, 1525, 1907, 1851, 350, 320, 1622, 1950]. Potential [1515, 894, 1543, 1343, 491, 929, 495, 1347, 1160, 5, 742, 1810, 1678, 1516, 812, 360, 1228, 1779, 1495, 257, 215, 1424, 490, 1230, 920, 502]. potentials [96, 639, 526, 1704, 1423, 1772, 620, 1033, 778, 1258, 992, 18, 391, 794, 984, 1191, 1526, 942, 1421, 1818, 843, 55, 54, 845, 1038, 700, 917, 1241, 1697, 164, 466, 1816, 110, 1712, 1552, 1713, 1949, 469]. Potts [78, 1696, 76]. POVMs [1582]. power [1527, 1816, 110, 469, 336]. power-law [469]. powers [1273, 1732]. Prandtl [539, 540, 1797]. pre [1806]. pre-symplectic [1806]. Precession [339, 113]. prediction [674]. Preissman [1109]. preparation [1937]. prepotential [255, 256]. prescribed [1872, 1008, 1009]. presence [391, 1102]. present [225]. presentation [1272]. Preserving [1313, 199]. prespace [1682]. primary [1783]. Primitive [289, 832]. principal [1872, 1862, 448, 267, 1077, 1174]. principle [1561, 1800, 1488, 899, 653, 738, 12, 739]. principles [223]. probabilistic [74, 914, 1195]. probabilities [692]. Probability [459, 241, 1304, 146, 1142, 1570, 860, 1850, 107, 561, 217]. problem [970, 888, 804, 805, 806, 1085, 761, 1048, 1877, 358, 1560, 1769, 1921, 895, 1106, 1200, 1189, 1564, 1221, 1606, 482, 340, 156, 994, 1438, 1701, 754, 1454, 214, 1791, 1495, 1850, 474, 486, 1214, 1496, 1849, 430, 1398, 108, 1842, 1694, 1319, 298, 1005, 320, 24, 213, 635, 1400, 1728, 1054, 1523, 1882, 1031, 1854, 1211, 119, 1128, 337, 586]. problems [1354, 1142, 516, 374, 325, 1435, 1899, 46, 1658, 537, 288, 221, 231, 86, 1823, 204]. procedure [63, 1444]. procedures [1286]. process [90, 1798, 398, 158, 1285]. processes [558, 1527, 1817, 1815, 1019]. produce [63]. product [1698, 407, 448, 1663, 1417, 1755, 1332, 972, 71, 284, 484, 1754, 326, 499, 943, 590, 803]. production [684, 1501, 91]. Products [1472, 1671, 445, 1531, 478, 716, 1455, 1860, 286, 907, 197, 1619, 415, 202, 1196, 292]. profile [281, 1466, 774]. programmable [1128]. Progress [1456, 227]. projection [1219, 579]. projections [931]. Projective [513, 1287, 489]. projectors [907, 1912]. Prolongation [1438]. Proof [1292, 790, 1603, 361, 1285, 550, 1376]. proofs [1005]. propagation [1257, 779, 593, 1912, 978, 401]. Propagator [894, 1104, 1741, 470]. propagators [377, 856]. proper [177, 738, 12, 739]. Properties [1105, 1756, 1782, 1620, 150, 552, 1401, 1257, 574, 1317, 1705, 1931, 1544, 90, 1060, 1934, 428, 496, 1405, 912, 85, 934, 1562, 1787, 200, 1071, 595, 1234, 976, 1198, 1703, 1929, 618]. property [1292, 1375, 105, 118]. proposal [309]. protein [340, 156, 755]. protocol [1138]. Prym [452].

Pseudo [1441, 1252, 877, 1029, 1102, 1310, 1314, 1695, 562, 416, 181, 1514, 1882, 1197, 1230]. pseudo-Hermitian [1514]. Pseudo-Hermiticity [1310, 1441, 1252, 877, 1029, 1102, 1314, 1197]. pseudo-potential [1230]. pseudo-Riemannian [1695, 416, 1882]. pseudo-sphere [562]. pseudo-spherical [181]. pseudomonoids [1868]. pseudosymmetric [1774]. Pseudounitary [1684]. published [1255]. Publisher [1420, 1946]. Pulmannová [1755]. pure [1727, 1513, 1727]. purification [996, 1123]. purity [1129]. Putting [435].

Q [1542, 1466, 917, 1465]. Q-ball [1466]. Q-solitons [1465]. QCD [1292]. QED [1884, 1372, 1073, 1715]. Quadratic [824, 565, 1086, 1534, 281, 138, 774, 1659, 1524]. quadratics [1338, 1901]. quadrilateral [70]. quadruple [916]. Quantal [804, 805, 806, 263, 465]. quantities [1178]. quantity [1066, 875]. quantizability [1208]. Quantization [531, 1920, 866, 153, 1857, 1870, 811, 1681, 1755, 1175, 1822, 1952, 815, 869, 1034, 411, 392, 1754, 913, 763, 22, 1871, 749, 47, 1157, 25, 753, 882, 472, 805, 806]. quantizations [528]. quantized [18, 548, 521, 558, 138, 118, 706]. Quantizing [1003, 1767]. Quantum [50, 419, 421, 1857, 809, 98, 1476, 1678, 939, 1035, 1479, 1708, 996, 1826, 226, 1856, 253, 634, 1630, 58, 1646, 947, 707, 1713, 1949, 1073, 1956, 1074, 1198, 950, 357, 1199, 1131, 987, 865, 836, 1724, 1632, 988, 1401, 15, 237, 1225, 1888, 640, 1698, 1889, 278, 1354, 1932, 494, 1344, 807, 776, 1346, 990, 587, 613, 1581, 740, 1862, 1746, 152, 850, 1531, 218, 991, 1138, 1836, 155, 1188, 1765, 1705, 219, 1931, 448, 1781, 1672, 464, 242, 1140, 1951, 1300, 294, 1832, 1075, 1482, 1851, 565, 1136, 840, 1098, 1061, 1535, 911, 743, 576]. quantum [931, 1801, 44, 1331, 1608, 1625, 1100, 1517, 1249, 9, 49, 19, 245, 1139, 1627, 1943, 1478, 870, 1154, 1227, 1790, 411, 1828, 816, 102, 1937, 1132, 1135, 1260, 1286, 1134, 1282, 1375, 1002, 558, 1467, 53, 973, 248, 1283, 1146, 103, 496, 284, 484, 214, 912, 578, 1302, 1202, 1036, 1037, 1904, 1830, 1890, 1850, 498, 914, 1407, 1682, 974, 1935, 499, 1661, 1821, 1042, 747, 56, 560, 380, 143, 748, 80, 1285, 1118, 1556, 310, 1918, 1133, 645, 23, 413, 1076, 1313, 524, 1587, 1684, 1631, 431, 500, 985, 466, 945, 1662, 1924]. quantum [1953, 998, 1547, 846, 967, 145, 831, 820, 788, 1234, 1195, 1930, 1551, 1762, 1588, 222, 1425, 1875, 1130, 1945, 7, 293, 469, 1351, 300, 1072, 731, 1900, 223, 1122, 948, 1235, 1255, 532, 1509, 1127, 1379, 1745, 471, 1006, 217, 1203, 111, 1820, 756, 26, 1211, 1039, 206, 302, 28, 1129, 1128, 1788, 1732, 170, 1760, 388, 618]. Quantum-mechanical [1713, 1949]. quark [556, 1811]. quarkonium [161]. quarter [1449]. quartic [1108]. Quasi [1321, 1867, 1297, 1011, 1191, 45, 56, 287, 1747]. quasi-free [1297]. Quasi-homogeneous [1321]. Quasi-Hopf [1867, 45, 287]. quasi-multiplicativity [1747]. quasi-solvable [1191]. quasi-two- [56]. quasialgebras [1867]. quasiclassical [1093, 1066, 875, 1426]. quasiconfluent [1359]. quasicrystal [1027]. quasicrystals [308, 664]. quasidistribution [997]. Quasiexactly [46, 451]. quasifree [558].

Quasigraded [771]. quasilocal [569]. quasinormal [1923, 784]. Quasiperiodic [483, 1697]. quasiperiodically [1071]. quasiprobability [1341]. quasirevivals [110]. quasitriangular [1755, 1754]. quaternion [1830]. Quaternionic [631, 1221, 1443, 261, 1391, 1830, 1890, 489]. qubit [1895, 1781, 1147]. Qubits [1137, 394]. quest [219]. Quotient [580]. qutrit [1933].

Racah [69, 579, 518]. Radial [172, 236, 1024, 1025, 1104, 1353]. Radiating [798]. Radiation [1052, 263, 11, 1542, 737, 1166, 1176, 864, 566, 1320, 1151]. radiations [1493]. Radon [1551]. Ramond [678, 679]. Randall [681]. Random [1500, 1828, 1550, 94, 1472, 1910, 1311, 724, 1680, 843, 873, 1575, 1726, 1038, 1548, 861, 1304, 200, 1366, 333, 1368, 1670]. Random-field [1828]. random-matrix [873]. randomly [169]. range [1423, 1705, 85, 1424, 1713, 1949, 1431, 1015]. rank [1685]. rapidly [1312]. rate [55, 1834, 1666]. ratio [1543]. rational [824, 1236]. ratios [1472]. ray [857]. ray-splitting [857]. reaction [595, 1176]. real [656, 1249, 1446, 1029, 785, 729, 986]. reality [877]. realization [1531, 1390, 1648, 681, 1821, 581, 1425]. realizations [1181, 662, 1069]. recipe [1285]. Reciprocal [1323]. recombination [1402]. Reconstruction [307, 1581, 498]. Recoupling [1864, 579]. Recovering [1893]. Recovery [1543]. Recursion [577, 1306]. Recursive [1024, 1025]. Redefining [1075]. reduced [1535, 1415, 1721, 1622, 1950]. Reduction [1143, 727, 1090, 1037, 1849, 933, 212, 600, 111, 1188, 376, 656, 698, 295, 1036, 1512, 1731, 1675, 730, 1050]. reductions [1445, 682, 1776, 853, 487, 1110, 1453]. Reeh [105, 1203]. reference [645]. Refined [22]. refinement [1752]. reflection [1543, 1898]. reflective [64]. reflectivity [307]. Reformulation [353, 369]. Regge [981, 1641]. regime [1097, 307]. region [527, 1244]. regions [1204, 1105, 1757, 1756]. Regular [1061, 1916, 1631]. Regularity [1434, 1214]. regularization [420, 693, 447, 761, 1289, 1067, 880]. regularized [175]. Regularizing [918]. Reissner [131]. related [526, 1704, 154, 758, 569, 984, 1191, 597, 509, 40, 1365, 1443, 830, 466, 537, 1547, 1685, 713, 1655, 1074]. Relating [106]. Relation [795, 1189, 136, 1331, 1216, 1779, 561, 255, 1124]. relations [182, 1212, 1570, 657, 603, 1640, 1353, 965, 1190]. Relative [349, 1934]. Relativistic [16, 1765, 726, 942, 1421, 1818, 104, 1666, 970, 4, 1580, 937, 1768, 237, 612, 1001, 1317, 219, 1876, 1881, 425, 1291, 543, 870, 1429, 983, 103, 496, 1353, 844, 108, 1911, 1456, 562, 584, 1834, 787, 1266, 1234, 222, 1235, 1255, 1432, 1006, 1715, 264]. Relativity [1757, 1105, 1756, 761, 1486, 512, 227, 899, 185, 624, 1164, 960, 442, 1204, 653, 301]. relaxation [339, 113]. remarkable [415]. Remarks [1186, 1187, 958, 879, 926, 83]. Remote [1937]. Renormalization [1352, 1534, 1071, 975, 1859, 1922, 1484, 488, 1942, 1588, 205]. Renormalized [340, 156]. replica [1086]. replication [1865]. Representation [471, 1309, 1844, 1533, 1140, 1951, 395, 630, 1331, 766, 1063, 1155, 371, 1450, 791, 747, 22, 1744, 382, 1408, 1631, 525, 303, 338]. Representations [1388, 680, 829, 801, 413, 1698, 1245, 278, 41, 626, 78, 1219, 1298, 1247, 1330, 448, 1189, 1417, 1837, 312, 1701, 137, 1708, 1286, 522, 71, 728, 286, 327, 380, 1496, 607, 581, 1118, 383, 1743, 1095, 967, 1867, 1305, 13, 208, 258, 456, 999, 489, 1229, 1339, 149, 388, 518, 1699]. reproducing [189]. repulsion [1431]. requirement [154]. reshaping [1078]. residue [603]. residues [1152, 1576]. resolution [807]. resolved [1759]. Resolvent [1315, 1433, 405, 1452, 1916]. Resonance [1084, 1088, 496, 646]. Resonances [245, 20, 471]. resonant [1943, 7, 293]. resources [1134]. respect [1297, 1671, 1645, 1533, 1647]. Response [739, 493]. rest [970]. rest-frame [970]. restricted [1084, 801]. restriction [1820]. restrictions [933]. result [807, 550]. Results [1232, 1617, 1022, 845, 1547]. resummation [497]. retarded [1176]. Return [229]. reversible [1527]. Reversing [990, 356]. review [1130, 1945]. revisited [1561, 88, 1884, 1789, 1150, 1340, 585]. revivals [110]. revolution [1915, 222]. revolutionizing [218]. Rham [1251]. ribbon [287, 821]. Riccati [1458, 1336]. Ricci [1047, 716, 1455, 1845, 1727]. Ricci-flat [1455, 716]. Richardson [115]. Riemann [895, 1106, 1435, 1574, 244, 1438, 1004, 965, 1727, 1783, 1523, 1270, 586, 1813, 1812]. Riemannian [1473, 1614, 620, 757, 1504, 1885, 1251, 1918, 1695, 416, 186, 1882, 111]. rigged [1309]. Right [1242]. rigid [1299, 339, 113, 594]. rigidity [898]. Rigorous [74, 1376, 356]. Rindler [1636, 1637]. Ring [1567, 851, 1880, 908, 1125]. Ring-shaped [1567]. ringed [882]. Riquier [1435]. Rivlin [541]. Robertson [1096, 957, 1186, 1837, 856]. Robust [1286, 755]. Rodrigues [803]. Rodrigues-type [803]. role [347, 1743, 1924, 1953]. root [988, 281, 1657, 774, 1026]. roots [1443, 985, 258, 1113]. Rose [120]. Rosen [51, 1843]. Rotating [514, 539, 425, 339, 113, 718, 1834]. rotation [540, 1666, 961, 1797]. rotational [970, 1666]. rotators [15]. Rouet [427, 698, 1157, 109]. Routh [212, 1444]. RTT} [657]. Ruijsenaars [476, 792, 889, 422, 294, 35, 1657, 768, 767, 1209, 256]. rule [1399, 1303, 914]. rules [838, 446, 579, 1118]. Runge [884].

S [1347, 1355, 785]. S-matrix [785]. S-wave [1347]. Sachs [368]. saddle [998]. saddle-point [998]. Salpeter [1243, 818, 1228]. Salpeter-type [1228]. sampling [92, 94]. Sasakian [727]. Satake [454]. Satija [1942]. satisfying [1665]. Savart [325, 546]. Sawyers [541]. Scalable [1128]. Scalar [836, 1237, 1294, 260, 1473, 1557, 620, 527, 927, 1633, 896, 897, 1516, 391, 604, 1093, 876, 864, 949, 1320]. Scalar-tensor [1237]. scalars [182]. Scale [1545, 1201, 1018, 1481, 695, 1663, 307]. Scaling [900, 613, 81, 973, 270, 255]. scatterers [743]. Scattering [1212, 1279, 1326, 571, 1195, 235, 775, 937, 1844, 16, 1068, 1463, 929, 1452, 742, 171, 1383, 1737, 175, 1565, 104, 1791, 590, 1766, 1855, 1456, 298, 1894, 1424, 38, 1448, 1713, 1949, 1006, 119, 1177, 618, 502]. scheme [811, 175, 844, 1522, 1109, 335, 1528]. schemes [313, 1353, 457]. Schild [762].

Schlesinger [194]. Schlieder [105, 1203]. Schmidt [463]. Schneider [889, 422, 476, 792, 361, 35, 1657]. Schouten [1778, 115]. Schrieffer [350, 1376]. Schrödinger [552, 1369, 216, 235, 150, 96, 236, 775, 1343, 1402, 1624, 1738, 1423, 1186, 1187, 1447, 1033, 376, 656, 1595, 1545, 1679, 483, 529, 530, 995, 1349, 1250, 178, 1261, 1192, 1601, 1038, 1879, 1658, 164, 925, 1194, 571, 702, 1831, 165, 166, 1480, 1751, 1816, 169, 1424, 1712, 1686, 1796, 848, 1015, 301, 502]. Schrödinger-type [1402, 376, 656]. Schwarz [135, 271]. Schwarzschild [316, 439, 958, 442]. Schwinger [1189, 1190, 52, 99, 426, 1165, 783]. screen [304]. screened [30, 920]. sea [391]. Searching [1638, 222]. second [16, 1671, 6, 207, 43, 1916, 1885, 411, 1026, 965, 549, 934, 62, 892, 1593]. second-order [207, 43, 1885, 62, 892]. secondary [1218]. Secrecy [1138]. section [1687]. sections [419]. sector [37]. secular [780]. seed [1345]. segments [1304]. Segur [1926]. Seiberg [341, 256, 1519, 501]. Seifert [1881]. selection [1118]. Self [1004, 93, 1445, 1738, 1511, 84, 601, 1589, 810, 1507, 485, 658, 853, 768, 849, 1539, 1775, 65, 335, 315, 623]. self-adjoint [1511, 485, 335]. self-adjointness [658, 768]. self-avoiding [84, 93]. self-consistent [315, 623]. Self-dual [1004, 1445, 810, 853, 849]. self-gravitating [601, 65]. self-graviting [1589]. self-polar [1507]. Self-testing [93]. Semi [18, 460, 614]. semi-circular [614]. Semi-classical [18]. semi-infinite [460]. Semiclassical [69, 912, 1601, 1149, 47, 1855, 1894, 1520, 1944, 708, 842, 1741, 975, 470]. Semiclassics [940]. semiconductors [235]. semidirect [1417]. semidiscrete [598]. semigroup [931]. semigroups [1297, 576, 1286, 1661, 1309]. semilinear [283]. semisimple [1263, 581]. Sen [1320]. separability [265, 174]. Separable [1801, 1690, 393, 1241, 954, 951]. separated [828]. separating [1686]. Separation [1065, 1167, 620, 1186, 1187, 177, 1079, 1733, 1492]. sequence [1730, 415, 1703]. sequences [385]. Sequential [816]. serial [743]. serial-structure [743]. series [41, 1117, 1862, 1877, 1312, 830, 163, 523, 147, 7, 293, 1255, 336]. set [1354, 1696, 1863, 1250, 1369, 1651, 1562, 862]. set-theoretic [1863]. sets [1079, 1307, 1203, 1039]. several [1488, 926, 1426]. shadow [295]. shallow [1842]. shape [888, 642, 54, 746, 917]. shaped [1567]. shapes [74]. Sharp [1581, 1888, 242]. Shear [983]. Shear-free [983]. sheets [1206, 1214]. shells [653, 492, 1293, 2, 493]. Sherrington [1244, 1086]. shift [1839, 742, 1608]. shift-invariant [1608]. shock [568, 373]. shocks [1841]. Short [922, 1705, 224, 1076, 1713, 1949]. short-range [1705, 1713, 1949]. Shouldn’t [1822, 1952]. Siggia [120]. Sigma [671, 328, 1430]. signals [329]. Signs [367]. Similarity [487, 730, 722, 1110, 1050]. Simons [1200, 810, 1589, 1263, 1004, 160, 582, 849, 467, 1518, 299, 707]. Simple [100, 651, 1391, 426, 1099, 198, 553, 147, 1655]. simulation [1188]. sine [381, 980, 651, 1446, 412]. sine-Gordon [381, 980, 651, 1446]. sine-Gordon/massive [980]. sine-Hilbert [412]. Single [851, 25, 1772, 743, 843, 701]. single-mode [743]. single-particle [701]. singly [48]. Singular [709, 37, 204, 1739, 1562, 1871, 164, 1759, 510, 1552, 1379, 1745]. singularities [480, 1060, 211, 964, 1707, 683]. Singularity [906, 1702, 1215, 1916, 1292, 371, 1357, 1815, 1749, 1515]. Sitter [1254, 1256, 1931, 1483, 345, 1264, 1, 916, 559, 736, 795, 564]. situation [225]. six [1163, 1057]. six-vertex [1057]. sixth [841]. size [1633, 813, 1913, 1854]. Skála [1703]. sketch [225]. Skyrme [777, 694, 1223, 841, 744, 563]. Skyrmed [1639]. skyrmions [977, 1464, 61]. Slater [1779]. SLE [1815]. SLE-type [1815]. slices [1915]. slowly [1834]. Small [775, 765, 236, 1616, 1131]. Small-energy [775, 236]. Smeared [519]. smooth [1910, 568, 1766, 364, 291]. solenoid [1742]. solenoids [1626]. solid [1898]. solid-on-solid [1898]. solitary [652, 334]. Soliton [720, 1772, 1688, 833, 1240, 980, 31, 929, 124, 925, 919, 1523, 387, 315]. Soliton-like [1240]. soliton-type [31]. soliton/particle [980]. Solitons [711, 1015, 1567, 1447, 1667, 896, 1168, 1326, 531, 1112, 795, 1465]. Solution [965, 1898, 1895, 937, 1877, 902, 66, 642, 994, 814, 1542, 1233, 34, 1725, 21, 1442, 1652, 163, 1166, 412, 1674, 350, 431, 1031, 1167, 1320, 387, 1903]. Solutions [612, 40, 1838, 404, 564, 341, 1873, 31, 1451, 720, 1001, 722, 1772, 1644, 1271, 574, 316, 1645, 481, 810, 715, 1896, 436, 1590, 1925, 897, 1688, 1861, 1642, 184, 483, 1099, 131, 841, 1355, 1863, 33, 530, 410, 269, 185, 1429, 1419, 1825, 817, 800, 1146, 1334, 1649, 116, 1651, 549, 1386, 179, 570, 1562, 381, 1110, 383, 1911, 1319, 1156, 572, 1336, 849, 1675, 1639, 1880, 1176, 1054, 1763, 772, 833, 1240, 563, 834]. solvability [947, 752]. Solvable [1528, 526, 1704, 722, 1560, 1769, 1921, 1299, 911, 44, 1191, 103, 451, 46, 466, 1019, 616, 1235, 1906]. solve [748]. Solving [1391, 1148, 1397, 1743, 702, 1410]. Some [838, 227, 817, 1283, 83, 845, 1547, 213, 1142, 1299, 1248, 1258, 1640, 758, 577, 55, 54, 1649, 943, 1675, 1897, 1492, 204]. source [788]. sourced [1557]. sources [11, 1212, 737, 798, 961, 864, 315, 623]. Sp [1300]. Space [1915, 736, 1724, 1032, 1561, 1623, 988, 692, 1204, 1105, 1757, 1756, 1557, 1088, 1422, 1932, 839, 1262, 1322, 927, 1806, 153, 741, 1933, 480, 1931, 1923, 1660, 342, 242, 317, 724, 1143, 1205, 409, 1047, 1220, 1798, 1604, 1605, 1634, 1801, 520, 1505, 1664, 1517, 1162, 1845, 1254, 1483, 345, 754, 1478, 604, 1860, 1264, 589, 1024, 1774, 1716, 1955, 513, 272, 1008, 1009, 797, 545, 1507, 1003, 578, 1302, 251, 924, 763, 916, 559, 1437, 1398, 382, 1508, 531, 749, 1766, 1194, 23, 1076, 609, 1636, 1637, 297, 353, 1759, 961, 1886, 1540, 1588]. space [1513, 333, 501, 469, 1067, 880, 1718, 1719, 471, 1203, 490, 1461, 443, 848, 303, 1555, 795, 564, 187]. Space-like [736]. spacelike [132]. spaces [1224, 1342, 189, 1530, 628, 1531, 1570, 342, 963, 1300, 1455, 930, 1771, 642, 1154, 1091, 1278, 326, 159, 1600, 474, 1080, 801, 1871, 1462, 72, 1182, 1577, 826, 1882, 322, 882, 1309, 736]. spacetimes [130, 602, 796]. spanned [767]. spanning [85]. spatial [1280, 1683, 1956, 357, 1838]. spatially [1204, 1105, 1757, 1756]. Special [1369, 1442, 1573, 1393, 1394, 1505, 394, 1250]. specifying [63]. spectra [1893, 57, 1712]. Spectral [405, 1705, 496, 1405, 254, 1491, 1257, 462, 1048, 1839, 1931, 88, 792, 1916, 1276, 1470, 916, 646, 1553, 976, 848, 1524, 935]. spectroscopy [21]. spectrum [15, 1088, 1343, 1738, 1349, 1261, 1527, 676, 164, 608, 414, 877, 1029, 615, 169, 468, 1020]. Spencer [1217]. Sphalerons [1470]. sphere [16, 972, 817, 172, 1, 1256, 103, 104, 747, 382, 1052, 117, 1783, 562, 891, 1384, 707, 1235, 1255, 1315, 437]. spheres [1493, 1342, 588, 941, 1948, 272, 624, 1164, 104, 928, 1954]. Spherical [1799, 527, 859, 1828, 1441, 1, 1256, 1399, 1303, 181, 566]. Spherically [841, 604, 1319, 240, 1590, 1925, 570, 1665, 1005, 1886, 1559, 404]. Spiked [871]. Spin [182, 1333, 1259, 888, 555, 1247, 1475, 765, 642, 1608, 1517, 100, 972, 1696, 1483, 1415, 1244, 1086, 699, 874, 591, 1294, 1630, 1378, 1912, 1741, 455, 554, 755, 470, 1610, 1760, 1207, 112, 1113]. spin- [642, 699, 1294]. spin-2 [1483, 112]. spin-glass [755]. spin-one [1517]. spinless [1884, 818, 108]. spinning [815, 961]. spinor [182, 1219, 762, 766, 369, 1808, 320, 525, 1807, 533]. spinor-like [525]. spinors [1538, 1075, 193, 1536, 144, 1807]. splitting [857, 55, 1662]. splittings [335]. Spontaneous [389, 1474, 887]. spontaneously [109]. spot [1880]. spring [1943]. Square [279, 491, 265, 281, 1495, 774, 1028, 257, 1712]. Square-integrable [279]. square-root [281, 774]. square-well [491, 257]. squares [1566]. squash [1854]. Squashed [1677]. squeezed [989, 166]. SRB [1827]. stabilities [57]. Stability [584, 1880, 1385, 1802, 866, 1208, 412]. Stabilization [1101]. Stable [938, 1931, 639, 505, 1796]. standard [573, 996, 353, 1835, 112]. Standard-Model [1835]. Star [1673, 1531, 1542, 590, 1166, 292]. Stark [468]. stars [1834, 1666]. State [295, 1623, 527, 1138, 1188, 8, 73, 567, 936, 1259, 1478, 1628, 699, 1722, 1141, 821, 1741, 1196, 260, 1067, 880, 847, 1582, 1069]. States [560, 4, 1580, 989, 639, 51, 1068, 1297, 78, 1347, 1933, 155, 1140, 1951, 5, 1280, 1810, 1190, 1837, 1098, 911, 576, 1608, 1477, 246, 589, 941, 1948, 558, 53, 843, 845, 1101, 1709, 1710, 393, 1683, 750, 1193, 1194, 1313, 1408, 166, 878, 1954, 259, 1875, 999, 1236, 470, 1554, 1796, 1723, 472, 1732, 1760]. Static [185, 1028, 1873, 130, 602, 928, 570, 1005, 1718, 1719, 1017]. Stationary [788, 418, 576, 1355, 1635, 1751]. statistic [423]. Statistical [406, 596, 200, 120, 1563, 1414, 567, 936, 230, 329, 1609, 86, 1747]. statistics [1680, 1905, 1095, 1287, 1784, 1179]. steepest [721]. steplike [96]. Stereographic [640]. Stewartson [1449, 179, 660]. Stieltjes [627]. stiff [1758]. Stochastic [1060, 1565, 1586, 354, 150, 552, 419, 421, 1878, 541, 558, 284, 484, 1905, 1817, 23, 787, 1646, 505]. stochastically [1018]. stochastics [1407]. Stokes [721, 1215, 362]. Stokes-flow [1215]. Stora [427, 698, 1157, 109]. straight [1098]. Stratified [1512, 358]. streaming [935]. stress [1040, 541, 1055]. stress-energy [1055]. String [1635, 685, 1920, 1317, 1318, 1660, 17, 674, 686, 665, 1638, 673, 670, 671]. stringlike [185]. Strings [676, 677, 158, 372, 670, 1158, 668]. stringy [683]. strip [1246]. stripping [1177]. strips [460]. strong [1117, 866, 1208, 596, 690, 842, 488, 1662]. strong-coupling [1117]. strongly [1143, 1840]. structural [1733]. Structure [150, 552, 630, 1043, 388, 1612, 692, 1105, 1756, 51, 890, 1648, 52, 897, 1604, 1605, 123, 743, 450, 1845, 312, 313, 1049, 35, 1734, 1556, 1602, 372, 121, 1485, 1076, 500, 1717, 946, 1296, 826, 65, 1236, 1491, 848, 619, 27, 1736, 337, 1199]. Structures [889, 1614, 722, 1322, 903, 964, 1591, 800, 195, 610, 14, 209, 127, 228, 1210, 1081, 1171, 511, 1090]. study [18, 1630, 1795, 1947]. Sturm [654, 1613, 374, 1893, 1031]. Sturmian [702]. sub [1614]. sub-Riemannian [1614]. subalgebra [1388]. subalgebras [528, 723, 802]. subgraphs [82]. subgroup [1065, 280]. Subharmonic [931]. submanifolds [392, 326]. Subsemigroups [1309]. subspace [1833, 767]. subspaces [1912, 945, 1855, 362]. sufficient [5, 1810, 785]. suitable [1562]. sum [444, 962, 1779, 1399, 1303, 907, 163, 821]. summability [1377]. summable [530, 1712]. summands [1899]. summation [1312, 449, 336]. Sums [859, 1289, 1228, 1899, 453, 1028, 1787, 1064, 1848, 1794]. Sundrum [681]. sunset [839]. Super [1042, 170, 1724, 655, 1730, 1272, 265, 710, 1330, 1595, 778, 1634, 99, 1771, 45, 1720, 81, 1156, 753, 338, 1247]. super- [1720]. super-Brownian [81]. super-disk [753]. Super-Grassmannian [1042]. super-integrable [778]. super-Jordanian [1724]. super-Poincaré [1247]. super-Schrödinger [1595]. super-separability [265]. super-Yangian [1595]. super-Yangians [1272, 1330]. superalgebra [140, 769, 967]. Superalgebras [823, 312, 313, 285, 48, 289, 832, 987]. superconducting [1158]. Superconductivity [1416, 460, 1802, 745]. superconductor [915, 944, 566]. Superconformal [1783, 417, 547, 473, 508, 680, 1685, 525]. supercovariant [42]. superenergy [1617]. Superfluidity [1721]. Supergeneralization [769]. supergroup [1506, 388]. supergroups [731]. Superheating [424, 1404]. Superintegrability [1227, 924, 947]. Superintegrable [1600, 532, 1746, 32, 565, 992, 180, 127, 752]. superlattice [1697]. superluminal [1912]. supermanifold [1714]. supermanifolds [908]. Superparticle [679]. superposition [269]. superpositions [612]. supersliver [1265]. supersonic [568]. Superspace [427, 250, 1025]. superstable [1739]. superstring [1265, 679]. Superstrings [668]. Supersymmetric [1730, 1532, 1714, 874, 508, 1428, 674, 681, 1100, 1852, 873, 1468, 1014, 255, 352, 820, 1425, 1773]. supersymmetry [858, 669, 220, 258, 1072, 170]. supertorus [67]. supported [104]. suppression [596]. surface [1802, 1257, 981, 244, 964, 369, 249, 1515, 1082, 1820]. surfaces [1583, 1409, 134, 542, 1281, 529, 318, 1915, 1063, 1155, 193, 1004, 1335, 124, 966, 582, 1195, 181, 385]. surfing [1206]. survey [224]. survival [1142, 561]. Suslov [1694]. SUSY [784]. Sutherland [422, 1348, 748, 1547, 1211]. Swendsen [90]. Swift [125]. swing [1943]. symbols [190, 281, 774, 251]. Symmetric [1, 1256, 1762, 1054, 1819, 805, 1297, 627, 240, 130, 1590, 1925, 796, 841, 604, 509, 797, 485, 139, 379, 570, 1319, 954, 1665, 1005, 961, 1886, 1559, 1718, 1719, 648, 1622, 1950, 404, 1113]. symmetric-positive [1622, 1950]. Symmetries [269, 1014, 397, 1559, 1719, 443, 789, 187, 1451, 1111, 516, 1403, 1482, 1642, 1457, 1051, 1591, 427, 428, 157, 1110, 1382, 1102, 1314, 1480, 127, 1239, 730, 1050, 174, 733, 362, 1887]. Symmetrized [484]. Symmetry [1097, 154, 574, 649, 1503, 517, 1770, 1817, 825, 381, 1812, 923, 613, 1187, 672, 1870, 1595, 98, 695, 1427, 1640, 1606, 642, 1263, 513, 1037, 746, 1649, 1496, 162, 719, 1691, 759, 1636, 877, 1029, 1102, 1336, 165, 166, 595, 417, 887, 261, 515, 1236, 1858, 111, 734, 735, 1325, 920, 1906, 705]. Symplectic [210, 913, 1105, 1756, 901, 1217, 1361, 1806, 1388, 903, 1318, 1356, 1591, 1752, 1380, 326, 1673, 786, 882]. symplectic-energy [1356]. syndrome [1900]. system [265, 1215, 459, 241, 574, 1200, 1362, 982, 897, 1574, 567, 936, 1035, 1479, 34, 1084, 1657, 1036, 1169, 966, 381, 1110, 1602, 599, 1549, 1674, 1071, 1487, 314, 431, 562, 1665, 1497, 1751, 826, 770, 129, 1520, 1944, 1491, 1622, 1950, 1524, 1903, 1173, 276]. Systematic [376]. Systems [595, 1445, 31, 988, 494, 923, 776, 1346, 1624, 555, 421, 375, 1746, 1048, 866, 1208, 991, 1931, 8, 73, 1769, 1921, 32, 852, 809, 1544, 279, 1476, 1678, 1143, 1482, 757, 1011, 765, 565, 1435, 1436, 911, 123, 779, 450, 758, 343, 1323, 1477, 1175, 794, 1049, 132, 268, 1034, 1374, 1627, 428, 1107, 942, 1421, 1818, 535, 1467, 1283, 1037, 1599, 1850, 1079, 1600, 1817, 452, 1739, 37, 306, 1522, 560, 80, 784, 953, 660, 1269, 701, 582, 126, 400, 954, 180, 1194, 1121, 583, 1485, 1855, 1924, 1953, 955, 584, 820, 1234]. systems [615, 510, 363, 355, 127, 1875, 760, 148, 356, 532, 892, 713, 771, 1917, 1270, 1492, 111, 752, 1668, 1015, 1736, 302, 511, 1760, 1174, 618]. syzygies [1727]. Szego [1892].

’t [719, 917]. table [1428]. tachyonic [670]. tachyons [666, 671]. tails [1910, 1038, 1366]. taken [1415]. Takesaki [1585]. tame [71]. Tau [1576]. Taub [1041]. TC [1172]. TD [893]. technique [1769, 1921, 1312, 873]. Teitelboim [1749]. Teller [639]. temperature [1243, 1292, 1244, 1416, 952]. temporal [1556]. Temporally [639, 1796]. Tensor [448, 1417, 331, 1316, 1040, 1698, 445, 1237, 895, 1106, 1047, 1640, 243, 905, 1617, 1332, 137, 1860, 1, 1256, 71, 606, 1619, 1727, 320, 260, 1055, 1559, 1887, 1719, 586]. tensors [1617, 509, 1882, 1077]. term [426, 658]. terms [1569, 805, 806, 1219, 1078, 827, 99, 780, 139, 1165, 783, 147, 1198]. test [1776, 1646, 756]. testing [93]. th [258]. th-roots [258]. their [324, 11, 51, 923, 1330, 792, 1363, 1060, 230, 1191, 737, 1, 1256, 1146, 560, 830, 595, 729, 1196, 1718, 1229, 1124, 563, 458, 518]. theorem [1422, 1932, 1387, 1361, 122, 1881, 910, 695, 1220, 368, 1427, 1120, 1281, 898, 851, 1083, 1488, 1154, 1244, 1441, 192, 331, 550, 1846, 1926, 455, 554, 647, 186, 1582, 733]. theorems [1613, 1260, 1203, 650]. theoretic [1714, 1863]. theoretical [1814, 153, 464, 1006]. theories [894, 1316, 389, 420, 903, 1096, 957, 432, 1051, 1263, 114, 1777, 427, 684, 544, 1927, 842, 1002, 157, 105, 250, 498, 954, 1487, 1076, 1685, 1430, 321, 1940, 260, 1112, 262, 26, 621]. Theory [1230, 837, 937, 11, 804, 691, 16, 1085, 613, 1117, 97, 1841, 929, 240, 154, 218, 1765, 1633, 546, 438, 1318, 1660, 98, 1564, 810, 60, 630, 1412, 17, 1435, 1516, 1428, 674, 243, 1457, 1061, 1643, 211, 1535, 868, 780, 9, 49, 19, 855, 1395, 870, 1707, 1869, 686, 689, 1446, 643, 1429, 665, 589, 1607, 1201, 102, 1585, 737, 1635, 248, 687, 104, 178, 284, 1003, 1202, 497, 1890, 1450, 498, 717, 1682, 659, 1042, 160, 783, 348, 349, 1835, 719, 310, 1697, 330, 1265, 212, 582, 126, 396, 1695, 351, 1731]. theory [500, 945, 1855, 1456, 467, 1638, 1044, 786, 1518, 255, 352, 256, 109, 87, 821, 615, 355, 307, 221, 222, 1728, 308, 664, 919, 223, 673, 670, 956, 532, 751, 975, 688, 1379, 1882, 1006, 1913, 1046, 949, 335, 671, 1703, 1788, 1376, 146, 1161, 666]. theory-differential [1869]. there [1822, 1952]. thermal [105, 1913]. Thermodynamic [1317, 1415, 1178, 704]. thermodynamical [1832]. Thermodynamics [538, 275, 1321, 1541, 433, 649, 121]. thermostatistics [1021]. Theta [1650, 767]. Thibon [1935]. thin [1279, 504]. thin-film [504]. third [1227, 1690, 1026, 322]. third-order [1227, 1690, 322]. Thirring [980]. Thompson [550]. Three [556, 1811, 167, 638, 533, 1844, 1422, 265, 79, 991, 1933, 8, 73, 1560, 1435, 642, 318, 1853, 1438, 282, 1035, 1084, 800, 1036, 746, 497, 1495, 559, 1398, 417, 213, 442, 1784, 1510, 1159, 1710]. three-body [265, 1560, 1035, 1084, 1495, 213]. three-dimensional [642, 1438, 282, 800, 746, 559, 442, 1510, 1159]. three-level [991]. Three-quark [556, 1811]. three-qutrit [1933]. three-wave [1398]. thresholds [1246, 1611]. Tiling [1577]. Time [1068, 715, 1681, 1542, 252, 1166, 1831, 165, 166, 864, 302, 865, 1561, 692, 1204, 1105, 1757, 1756, 239, 714, 1322, 927, 480, 910, 317, 765, 1220, 1798, 1517, 131, 1175, 1162, 439, 869, 1034, 1915, 1790, 272, 559, 738, 1437, 12, 739, 1838, 1840, 382, 1241, 173, 1766, 23, 561, 1636, 1637, 702, 1748, 1480, 202, 307, 1588, 300, 1045, 1956, 357, 795, 564]. time-delay [307]. Time-dependent [715, 1542, 1166, 165, 166, 865, 1175, 869, 1034, 1241, 173, 1480, 300]. Time-energy [1068]. time-harmonic [1766]. time-independent [702]. Time-optimal [302]. time-ordered [202]. time-periodic [131, 1840]. times [1557, 1923, 1205, 1047, 1664, 1162, 1845, 604, 1774, 1716, 1955, 513, 1008, 1009, 797, 545, 961, 1886, 1540, 1718, 1719, 1203, 443, 187]. Toda [1258, 176, 40]. Toda-like [1258]. together [503]. Tomita [218, 1585]. tomography [464, 1285]. top [1382]. Topics [1467, 569]. Topological [1136, 243, 1446, 1465, 837, 971, 1390, 1061, 1162, 1891, 378, 14, 209, 668]. Topologically [1672]. Topology [330, 210, 689, 3]. tops [1373]. tori [1039]. toroidal [1437, 413, 634, 275, 1010]. torsion [764, 373]. torus [1827, 909, 197, 1586, 633, 1829, 1553, 1236, 1229, 1069]. Total [1356]. trace [654, 1423, 857, 198, 553]. Trading [1134]. trajectories [883, 510]. transcendental [1064]. transfer [991, 535, 1350]. transform [1782, 1505, 1609, 551, 1509, 864, 1000]. Transformation [830, 96, 450, 267, 1489, 1398, 1358, 321, 1432, 1444, 1703, 29, 1732]. Transformations [70, 133, 1271, 1836, 649, 1681, 556, 1811, 1099, 1323, 995, 1154, 106, 194, 933, 784, 1430, 975, 1000, 623]. transforms [1346, 1481, 1277]. Transition [1659, 1493, 79, 851, 665]. transitions [21, 75]. transitive [1872]. translate [740]. Translation [756, 1229]. Translation-covariant [756]. transmission [1543]. transmutation [1803]. Transport [1327, 865, 1494, 743, 1056, 1125, 1797, 64]. transversality [1419]. trap [823]. trapped [538]. Traveling [1763]. treated [804, 805, 806]. treatment [651, 778, 369, 36, 1108]. tree [894, 1853, 1874]. tree-level [894]. trees [85]. Trefoil [733, 734, 735]. Treves [1846, 1926]. Tri [1490]. Tri-Hamiltonian [1490]. triangles [1399, 1303]. triangular [1687, 953]. tricritical [1020]. Trigonometric [1295, 1547]. Trions [812]. triple [444]. triplets [1196]. Tsallis [936, 567, 1934, 921]. Tsetlin [581]. tunneling [251]. Turaev [190]. turbulence [596]. turbulent [232]. turn [223]. Tutte [76]. twentieth [216]. Twist [220, 959]. Twisted [1115, 1330, 454]. Twisting [287, 35]. Twistor [1457, 1155, 1091, 1278]. twistors [39]. twists [1531, 802]. Two [992, 1885, 1489, 746, 348, 349, 359, 1729, 1732, 623, 1895, 804, 805, 806, 389, 419, 462, 1215, 1877, 1246, 991, 459, 241, 1474, 503, 1280, 1596, 1938, 565, 1435, 588, 1706, 642, 1016, 9, 49, 19, 1440, 994, 101, 77, 1626, 402, 269, 1264, 817, 699, 545, 1003, 1155, 924, 1284, 1165, 56, 380, 1849, 712, 1658, 1697, 1683, 400, 1071, 572, 609, 1637, 1150, 1340, 891, 1384, 1176, 1894, 1170, 1773, 1077, 1267, 26, 394, 1172, 1795, 1947, 1709]. Two- [746, 991]. two-center [804, 805, 806, 994]. two-component [1440, 402]. Two-dimensional [1885, 348, 349, 359, 1729, 462, 1246, 1474, 1596, 565, 642, 1016, 101, 77, 1264, 817, 699, 924, 1849, 1658, 1697, 572, 609, 1150, 1340, 891, 1384, 1170]. two-dimensional-Rindler [1637]. two-electron [1795, 1947]. two-form [389]. two-level [419, 459, 241, 1071]. two-particle [1706]. two-point [1003]. two-qubit [1895]. two-spheres [588]. two-surfaces [1155]. type [31, 16, 1388, 376, 656, 266, 1596, 1363, 681, 1323, 1664, 9, 49, 19, 1845, 1213, 360, 132, 995, 1753, 1107, 1805, 577, 545, 496, 192, 853, 514, 1815, 660, 1336, 985, 803, 1014, 826, 385, 48, 854, 1240, 29, 1402, 1423, 1488, 1770, 1228, 515, 1703]. type-D [515]. types [401, 715, 1732]. Tyutin [427, 815, 698, 1157, 109]. Tzetlin [1506]. Tzitzéica [904].

U [1455]. Uhlenbeck [398, 1661]. ultraviolet [1745]. Umbral [1879]. Unbraiding [1332]. Uncertainty [555, 738, 12, 739]. Underdetermined [375]. Underlying [182]. Understanding [964, 219]. Unification [1151, 1561, 277, 479, 675]. unified [1643, 606, 36, 1108]. uniform [1396, 146]. Unifying [1522]. unimodal [270]. Unimodular [512]. unique [1182]. Uniqueness [907, 1242, 1806, 1120, 1725, 350, 647]. unit [817, 1509]. unital [1147]. unitarity [894]. unitarization [1360]. Unitary [728, 1698, 1615, 448, 1939, 999, 1699]. unity [1026, 985]. universal [1864, 1039]. universality [206, 28]. universe [366, 1663, 544, 1854, 1167]. Universes [1592]. Unknown [1140, 1951]. unoriented [193]. unparametrized [725]. unsharp [1581, 242]. unstable [421, 232, 178]. Unsteady [1644, 364]. unstretchable [291]. unusual [1929]. Upper [1347, 843, 1904, 1892, 540, 1261, 1405]. upper-half [1261, 1405]. use [1795, 1947]. using [106, 499, 965, 1456, 1775, 1131].

V [401, 715]. vacua [1717]. Vacuum [6, 207, 1377, 715, 602, 696, 184, 814, 1355, 1265, 961, 670, 515]. vakonomic [126]. validity [534]. value [516, 6, 207, 486, 1066, 875, 1842, 1319, 1005]. valued [1613, 655, 1533, 1647, 1907, 1851, 1817]. values [1696, 817, 1081, 1082, 386, 1794]. Variable [620, 902, 658, 1080, 1883, 1492]. variables [888, 640, 1488, 1079, 861, 1094, 1065, 333, 1081, 1082, 1167, 472, 1524, 112]. variant [1070]. variation [1356, 1593]. Variational [702, 1715, 1529, 1841, 1488, 653, 1214, 288, 86]. variations [1607]. varieties [1276, 452]. various [984, 518]. Vassiliev [160]. Vavilov [1493]. vector [1316, 1613, 620, 764, 1218, 528, 325, 483, 1254, 345, 295, 1334, 578, 1302, 884, 68, 1044, 148, 770, 1324, 1103, 1554, 264]. vector-valued [1613]. vectorial [1911]. vectors [1623, 782, 289, 832]. velocities [1841]. velocity [418, 1154, 1053]. Verma [208, 13]. version [1056]. versus [1344, 1869, 544, 877, 1029, 1102, 1287, 1198]. Vertex [678, 149, 1753, 1057, 304]. vertices [201]. Vey [840]. via [603, 801, 487, 382, 91, 1912, 1350, 1453, 1184, 1732, 1611]. vicinity [339, 113, 608]. view [751]. viewpoint [1721, 1407]. VIII [599]. violating [1075]. violation [1824]. Virasoro [1569, 1720, 1602, 1637, 290]. Viro [190]. virtual [21]. viscoelastic [171]. Vlasov [1665, 1903]. Voisin [188]. Volkoff [323]. Volterra [1276]. volume [174]. vortex [477]. vortices [1460, 1910, 503, 1589, 1498, 1004, 707, 1461]. Vries [655, 1730, 710, 1525, 1334, 600, 1050, 271, 1240].

W. [1339]. Waals [314]. Walker [1096, 957, 856]. walks [84, 93]. wall [1290, 1057]. Wallach [1455]. walls [1215, 696, 683]. Wang [90]. Warped [1663, 478, 716, 1845, 684, 1540, 688]. water [1842]. Wave [1087, 628, 1383, 249, 1350, 110, 1177, 612, 1001, 1558, 1257, 1841, 910, 1427, 1861, 1436, 1640, 175, 211, 1625, 589, 283, 906, 853, 116, 1398, 1840, 1066, 875, 759, 978, 1894, 1763, 713, 864, 1103, 1203, 457, 1715, 334, 1622, 1950, 1185, 1240, 1320, 1347]. wave-packet [1894]. Wavefronts [1206]. waveguides [1383, 1098, 245, 878]. Wavelet [334, 1505]. Wavelets [191, 1679, 1023]. waves [11, 1212, 232, 652, 859, 1843, 737, 373, 1912, 364, 2, 119]. way [1900]. WDVV [352, 1211]. Weak [1568, 1419, 1628, 1197, 1404, 1517, 1826, 1562, 364]. weak- [1404]. weak-discontinuous [364]. weakening [356]. Weakly [832, 15, 1194, 1760]. wedge [344]. Weierstrass [574, 1412, 1063, 1454]. weight [285, 338, 149]. weighted [930, 1726]. Weingarten [385]. well [1819, 639, 491, 1598, 257, 1242]. well-posedness [1242]. Wess [1232, 1586]. wetting [79]. Weyl [449, 1806, 1190, 840, 1436, 1640, 184, 441, 1443, 1727, 1711, 561, 1336, 1077, 322, 1309]. Weyl-like [1077]. Wheeler [1641]. which [785]. white [971]. Whitham [1093]. Whittaker [1671]. Wichmann [1154]. wide [307]. wide-band [307]. Wiechert [1526]. Wightman [1940]. Wigner [554, 1224, 1342, 418, 1387, 641, 1680, 840, 1145, 1373, 1478, 579, 1284, 331, 1631, 455, 355, 636]. Wilson [448]. winding [1645]. within [1474, 567, 936, 1835]. without [1160, 540, 976, 1797]. Witt [1274]. Witten [341, 580, 1586, 256, 1519, 501]. WKB [1780, 1641]. world [682]. wormholes [1010]. Wouthuysen [1432]. wrapped [317]. Wronskian [1896]. Wu [1649, 1251, 1006]. WZW [676, 677].

XXZ [833, 1113].

Yanase [1478]. Yang [49, 900, 1005, 321, 1006, 404, 276, 1445, 1612, 1316, 389, 316, 1824, 674, 1331, 244, 9, 19, 1863, 1003, 250, 853, 559, 819, 1815, 570, 383, 749, 954, 1156, 255, 352, 1639, 1908, 795, 564]. Yangian [1595, 1648, 633]. Yangians [1272, 1330]. Yatsun [954]. Yau [188, 490]. Yetter [821]. York [569].

Zakharevich [1805]. Zakharov [274]. Zamastil [1703]. Zamolodchikov [1216]. Zanelli [1749]. Zeitlin [1118]. zero [506, 534, 698, 873, 845, 608]. zero-curvature [506]. zero-point [534]. zeros [374]. zeta [1089, 1700, 254]. Zetlin [449]. Zhang [1542, 1649]. Zheng [1793]. Zierau [47]. ZS [1524]. ZS-AKNS [1524]. Zumino [1586, 1232].

References

Higuchi:1987:STS

[1] Atsushi Higuchi. Symmetric tensor spherical harmonics on the N-sphere and their application to the de Sitter group SO(N, 1). Journal of Mathematical Physics, 28(7):1553–1566, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1256].

Wang:1992:GIP

[2] An Zhong Wang. Gravitational interaction of plane gravitational waves and matter shells. Journal of Mathematical Physics, 33(3):1065–1072, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comments [492, 493].

Dai:1994:IDL

[3] Xianzhe Dai and Daniel S. Freed. η-invariants and determinant lines. Topology and physics. Journal of Mathematical Physics, 35(10):5155– 5194, October 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [637].

Aldaya:1995:CCS

[4] V. Aldaya and J. Guerrero. Canonical coherent states for the relativistic harmonic oscillator. Journal of Mathematical Physics, 36(7):3191–3199, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1580].

Chadan:1997:SCE

[5] K. Chadan and R. Kobayashi. A sufficient condition for the existence of bound states in a potential. Journal of Mathematical Physics, 38(10): 4900–4908, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1810].

Branson:1998:VEV

[6] Thomas P. Branson, Peter B. Gilkey, and Dmitri V. Vassilevich. Vacuum expectation value asymptotics for second order differential operators on manifolds with boundary. Journal of Mathematical Physics, 39(2):1040– 1049, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [207].

Saenz:1998:LSA

[7] A. W. Sáenz. Lie-series approach to the evolution of resonant and nonresonant anharmonic oscillators in quantum mechanics. Journal of Mathematical Physics, 39(4):1887–1909, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [293].

Byrd:1998:DGA

[8] Mark Byrd. Differential geometry on SU (3 ) with applications to three state systems. Journal of Mathematical Physics, 39(11):6125–6136, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [73].

Fleischhack:1999:NTL

[9] Christian Fleischhack. A new type of loop independence and SU(N) quantum Yang–Mills theory in two dimensions. Journal of Mathematical Physics, 40(6):2584–2610, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p2584_s1. See erratum [49].

Ivashchuk:1999:HKM

[10] V. D. Ivashchuk, S.-W. Kim, and V. N. Melnikov. Hyperbolic Kac– Moody algebra from intersecting p-branes. Journal of Mathematical Physics, 40(8):4072–4083, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [835].

Ardavan:1999:MHD

[11] H. Ardavan. Method of handling the divergences in the radiation theory of sources that move faster than their waves. Journal of Mathematical Physics, 40(9):4331–4336, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [737].

Kudaka:1999:UPP

[12] Shoju Kudaka and Shuichi Matsumoto. Uncertainty principle for proper time and mass. Journal of Mathematical Physics, 40(3):1237–1245, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comments [738, 739].

Raghavan:1999:VBR

[13] K. N. Raghavan and P. Sankaran. On Verma bases for representations of sl(n, C). Journal of Mathematical Physics, 40(4):2190–2195, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [208].

Rainer:1999:CCS

[14] M. Rainer. Cones and causal structures on topological and differentiable manifolds. Journal of Mathematical Physics, 40(12):6589–6597, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [209].

Angelescu:2000:OPE

[15] Nicolae Angelescu, Robert A. Minlos, and Valentin A. Zagrebnov. The one-particle energy spectrum of weakly coupled quantum rotators. Journal of Mathematical Physics, 41(1):1–23, January 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avossevou:2000:RST

[16] G. Y. H. Avossevou and M. N. Hounkonnou. Relativistic scattering theory for a δ sphere plus a Coulomb interaction with boundary conditions of second type. Journal of Mathematical Physics, 41(1):24–39, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See note [1255].

Dimock:2000:LFS

[17] J. Dimock. Locality in free string field theory. Journal of Mathematical Physics, 41(1):40–61, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Faure:2000:SCS

[18] Frédéric Faure and Bernard Parisse. Semi-classical study of the origin of quantized Hall conductance in periodic potentials. Journal of Mathematical Physics, 41(1):62–75, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fleischhack:2000:NTL

[19] Christian Fleischhack. A new type of loop independence and SU (N ) quantum Yang–Mills theory in two dimensions. Journal of Mathematical Physics, 41(1):76–102, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hagedorn:2000:RB

[20] George A. Hagedorn and Bernhard Meller. Resonances in a box. Journal of Mathematical Physics, 41(1):103–117, January 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karpov:2000:FMV

[21] E. Karpov, I. Prigogine, T. Petrosky, and G. Pronko. Friedrichs model with virtual transitions. Exact solution and indirect spectroscopy. Journal of Mathematical Physics, 41(1):118–131, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Louko:2000:RAQ

[22] Jorma Louko and Carlo Rovelli. Refined algebraic quantization in the oscillator representation of SL(2 , R). Journal of Mathematical Physics, 41 (1):132–155, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendes:2000:GSC

[23] R. Vilela Mendes. Geometry, stochastic calculus, and quantum fields in a noncommutative space–time. Journal of Mathematical Physics, 41(1): 156–186, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pletyukhov:2000:GFF

[24] M. V. Pletyukhov and E. A. Tolkachev. Green’s function for the fivedimensional SU (2 ) MIC–Kepler problem. Journal of Mathematical Physics, 41(1):187–194, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Trueman:2000:SPI

[25] Colin Trueman and K. K. Wan. Single point interactions, quantization by parts, and boundary conditions. Journal of Mathematical Physics, 41 (1):195–205, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Verch:2000:ANE

[26] Rainer Verch. The averaged null energy condition for general quantum field theories in two dimensions. Journal of Mathematical Physics, 41 (1):206–217, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Walkup:2000:BPS

[27] John R. Walkup, Martin Dunn, and Deborah K. Watson. Branch-point structure and the energy level characterization of avoided crossings. Journal of Mathematical Physics, 41(1):218–239, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weaver:2000:UAE

[28] Nik Weaver. On the universality of almost every quantum logic gate. Journal of Mathematical Physics, 41(1):240–243, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [206].

Wittich:2000:TFK

[29] Olaf Wittich. A transformation of a Feynman–Kac formula for holomorphic families of type B. Journal of Mathematical Physics, 41(1):244–259, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruggen:2000:FES

[30] M. Brüggen and D. O. Gough. Free energy of a screened ion pair. Journal of Mathematical Physics, 41(1):260–283, January 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alber:2000:STS

[31] Mark S. Alber, Gregory G. Luther, and Charles A. Miller. On solitontype solutions of equations associated with N-component systems. Journal of Mathematical Physics, 41(1):284–316, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calzada:2000:CSH

[32] Juan A. Calzada, J. Negro, Mariano A. del Olmo, and Miguel A. Rodríguez. Contraction of superintegrable Hamiltonian systems. Journal of Mathematical Physics, 41(1):317–336, January 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geng:2000:AGS

[33] Xianguo Geng and H. H. Dai. Algebro-geometric solutions of (2 + 1)- dimensional coupled modified Kadomtsev–Petviashvili equations. Journal of Mathematical Physics, 41(1):337–348, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2000:ADS

[34] Boling Guo and Peicheng Zhu. Algebraic L^{2} decay for the solution to a class system of non-Newtonian fluid in R^{n}. Journal of Mathematical Physics, 41(1):349–356, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hou:2000:DTN

[35] Bo yu Hou and Wen-Li Yang. The dynamical twisting and nondynamical r-matrix structure of the elliptic Ruijsenaars–Schneider model. Journal of Mathematical Physics, 41(1):357–369, January 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karlovini:2000:UTC

[36] Max Karlovini and Kjell Rosquist. A unified treatment of cubic invariants at fixed and arbitrary energy. Journal of Mathematical Physics, 41(1): 370–384, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konopelchenko:2000:SSK

[37] Boris G. Konopelchenko, Luis Martínez Alonso, and Elena Medina. Singular sector of the Kadomtsev–Petviashvili hierarchy, ¯∂ operators of nonzero index, and associated integrable systems. Journal of Mathematical Physics, 41(1):385–413, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sabatier:2000:ESI

[38] Pierre C. Sabatier. Elbow scattering and inverse scattering applications to LKdV and KdV. Journal of Mathematical Physics, 41(1):414–436, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frauendiener:2000:LTC

[39] Jörg Frauendiener and George A. J. Sparling. Local twistors and the conformal field equations. Journal of Mathematical Physics, 41(1):437– 443, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivashchuk:2000:SIB

[40] V. D. Ivashchuk and Sung-Won Kim. Solutions with intersecting p-branes related to Toda chains. Journal of Mathematical Physics, 41(1):444–460, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basu:2000:CES

[41] Debabrata Basu, Subrata Bal, and K. V. Shajesh. The character of the exceptional series of representations of SU (1 , 1 ). Journal of Mathematical Physics, 41(1):461–467, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benhamou:2000:SOK

[42] N. Benhamou and T. Lhallabi. N = 2 supercovariant operators and Krichever–Novikov algebra. Journal of Mathematical Physics, 41(1):468– 479, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorodnitsyn:2000:LGC

[43] Vladimir Dorodnitsyn, Roman Kozlov, and Pavel Winternitz. Lie group classification of second-order ordinary difference equations. Journal of Mathematical Physics, 41(1):480–504, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fakhri:2000:MFA

[44] H. Fakhri and M. A. Jafarizadeh. Master function approach to quantum solvable models on SL(2, c) and SL(2, c)/GL(1, c) manifolds. Journal of Mathematical Physics, 41(1):505–546, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gould:2000:CIQ

[45] Mark D. Gould, Yao-Zhong Zhang, and Phillip S. Isaac. Casimir invariants from quasi-Hopf (super)algebras. Journal of Mathematical Physics, 41(1):547–568, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasovsky:2000:QSP

[46] I. V. Krasovsky. Quasiexactly solvable problems and the dual (q)-Hahn polynomials. Journal of Mathematical Physics, 41(1):569–577, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mehdi:2000:SLB

[47] Salah Mehdi. Semiclassical limit for the Binegar–Zierau quantization of the minimal nilpotent orbits of SO_{o}(2p, 2). Journal of Mathematical Physics, 41(1):578–601, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Su:2000:CFD

[48] Yucai Su. Classification of finite dimensional modules of singly atypical type over the Lie superalgebras sl(m/n). Journal of Mathematical Physics, 41(1):602–613, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fleischhack:2000:ENT

[49] Christian Fleischhack. Erratum: “A new type of loop independence and SU(N) quantum Yang–Mills theory in two dimensions”, [J. Math. Phys. 40, 2584 (1999)]. Journal of Mathematical Physics, 41(1):614, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [9].

Anastopoulos:2000:QFN

[50] C. Anastopoulos. Quantum fields in nonstatic background: A histories perspective. Journal of Mathematical Physics, 41(2):617–637, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arens:2000:CEP

[51] Richard Arens and V. S. Varadarajan. On the concept of Einstein– Podolsky–Rosen states and their structure. Journal of Mathematical Physics, 41(2):638–651, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daskaloyannis:2000:HAS

[52] C. Daskaloyannis, K. Kanakoglou, and I. Tsohantjis. Hopf algebraic structure of the parabosonic and parafermionic algebras and paraparticle generalization of the Jordan–Schwinger map. Journal of Mathematical Physics, 41(2):652–660, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiroshima:2000:GSM

[53] Fumio Hiroshima. Ground states of a model in nonrelativistic quantum electrodynamics. II. Journal of Mathematical Physics, 41(2):661–674, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarizadeh:2000:IDS

[54] M. A. Jafarizadeh, A. R. Esfandyari, and H. Panahi-Talemi. Isospectral deformation of some shape invariant potentials. Journal of Mathematical Physics, 41(2):675–700, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarizadeh:2000:CLS

[55] M. A. Jafarizadeh and S. Jalalzadeh. Calculation of the level splitting of energy and the decay rate of some one-dimensional potentials by instantons method. Journal of Mathematical Physics, 41(2):701–720, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kundu:2000:CQT

[56] Anjan Kundu. Construction of quasi-two- and higher-dimensional quantum integrable models. Journal of Mathematical Physics, 41(2):721–727, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2000:EEE

[57] Yanguang (Charles) Li. On 2D Euler equations. I. on the energy–Casimir stabilities and the spectra for linearized 2D Euler equations. Journal of Mathematical Physics, 41(2):728–758, February 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Osipov:2000:QIC

[58] Edward P. Osipov. Quantum interaction : ϕ_{4}^{4}:, the construction of quantum field defined as a bilinear form. Journal of Mathematical Physics, 41 (2):759–786, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prodan:2000:CEE

[59] E. Prodan. Cluster expansion for P(ϕ)_{∈}: explicit estimates. Journal of Mathematical Physics, 41(2):787–804, February 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cicogna:2000:AMT

[60] G. Cicogna and M. Santoprete. An approach to Mel’nikov theory in celestial mechanics. Journal of Mathematical Physics, 41(2):805–815, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kleihaus:2000:MS

[61] B. Kleihaus, D. H. Tchrakian, and F. Zimmerschied. Monopole Skyrmions. Journal of Mathematical Physics, 41(2):816–834, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mangiarotti:2000:GFS

[62] L. Mangiarotti and G. Sardanashvily. On the geodesic form of secondorder dynamic equations. Journal of Mathematical Physics, 41(2):835– 844, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marengo:2000:NPS

[63] Edwin A. Marengo and Richard W. Ziolkowski. A new procedure for specifying nonradiating current distributions and the fields they produce. Journal of Mathematical Physics, 41(2):845–866, February 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yuan:2000:CAN

[64] Guang wei Yuan. Convergence of P_{N} approximation for the neutron transport equation with reflective boundary condition. Journal of Mathematical Physics, 41(2):867–874, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

SanMiguel:2000:LPS

[65] A. San Miguel. Lie–Poisson structure for the homogeneous motion of self-gravitating compressible fluids. Journal of Mathematical Physics, 41 (2):875–881, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dadhich:2000:GDK

[66] Naresh Dadhich and L. K. Patel. Gravoelectric dual of the Kerr solution. Journal of Mathematical Physics, 41(2):882–890, February 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kachkachi:2000:PCS

[67] M. Kachkachi, O. Dafounansou, and M. Nazah. Polyakov conjecture on the supertorus. Journal of Mathematical Physics, 41(2):891–897, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kroon:2000:KVF

[68] Juan Antonio Valiente Kroon. On Killing vector fields and Newman– Penrose constants. Journal of Mathematical Physics, 41(2):898–923, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davids:2000:SLE

[69] Stefan Davids. Semiclassical limits of extended Racah coefficients. Journal of Mathematical Physics, 41(2):924–943, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doliwa:2000:TQL

[70] Adam Doliwa, Paolo Maria Santini, and Manuel Mañas. Transformations of quadrilateral lattices. Journal of Mathematical Physics, 41(2):944–990, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Howe:2000:MIT

[71] R. Michael Howe and Tuong Ton-That. Multiplicity, invariants and tensor product decompositions of tame representations of U(∞). Journal of Mathematical Physics, 41(2):991–1015, February 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nitta:2000:HFM

[72] Takashi Nitta and Tadashi Taniguchi. Hypergeometric functions and moduli spaces of null-correlation bundles. Journal of Mathematical Physics, 41(2):1016–1025, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Byrd:2000:EDG

[73] Mark Byrd. Erratum: “Differential geometry on SU(3) with applications to three state systems”, [J. Math. Phys. 39, 6125–6136 (1998)]. Journal of Mathematical Physics, 41(2):1026–1030, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [8].

Bodineau:2000:RPA

[74] T. Bodineau, D. Ioffe, and Y. Velenik. Rigorous probabilistic analysis of equilibrium crystal shapes. Journal of Mathematical Physics, 41(3): 1033–1098, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lyons:2000:PTN

[75] Russell Lyons. Phase transitions on nonamenable graphs. Journal of Mathematical Physics, 41(3):1099–1126, March 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Welsh:2000:PMT

[76] D. J. A. Welsh and C. Merino. The Potts model and the Tutte polynomial. Journal of Mathematical Physics, 41(3):1127–1152, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Georgii:2000:PNP

[77] Hans-Otto Georgii and Yasunari Higuchi. Percolation and number of phases in the two-dimensional Ising model. Journal of Mathematical Physics, 41(3):1153–1169, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biskup:2000:GSG

[78] M. Biskup, C. Borgs, J. T. Chayes, and R. Kotecký. Gibbs states of graphical representations of the Potts model with external fields. Journal of Mathematical Physics, 41(3):1170–1210, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolthausen:2000:AWT

[79] Erwin Bolthausen, Jean Dominique Deuschel, and Ofer Zeitouni. Absence of a wetting transition for a pinned harmonic crystal in dimensions three and larger. Journal of Mathematical Physics, 41(3):1211–1223, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lebowitz:2000:LDI

[80] Joel L. Lebowitz, Marco Lenci, and Herbert Spohn. Large deviations for ideal quantum systems. Journal of Mathematical Physics, 41(3):1224– 1243, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hara:2000:SLI

[81] Takashi Hara and Gordon Slade. The scaling limit of the incipient infinite cluster in high-dimensional percolation. II. integrated super-Brownian excursion. Journal of Mathematical Physics, 41(3):1244–1293, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benjamini:2000:EAB

[82] Itai Benjamini, Olle Häggström, and Oded Schramm. On the effect of adding ϵ-Bernoulli percolation to everywhere percolating subgraphs of Z^{d}. Journal of Mathematical Physics, 41(3):1294–1297, March 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kesten:2000:SRP

[83] Harry Kesten and Zhonggen Su. Some remarks on AB-percolation in high dimensions. Journal of Mathematical Physics, 41(3):1298–1320, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borgs:2000:ASA

[84] Christian Borgs, Jennifer Chayes, Christopher King, and Neal Madras. Anisotropic self-avoiding walks. Journal of Mathematical Physics, 41(3): 1321–1337, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kenyon:2000:LRP

[85] Richard Kenyon. Long-range properties of spanning trees. Journal of Mathematical Physics, 41(3):1338–1363, March 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shlosman:2000:GVP

[86] Senya Shlosman. Geometric variational problems of statistical mechanics and of combinatorics. Journal of Mathematical Physics, 41(3):1364–1370, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pemantle:2000:TTN

[87] Robin Pemantle. Towards a theory of negative dependence. Journal of Mathematical Physics, 41(3):1371–1390, March 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cancrini:2000:SGK

[88] N. Cancrini and F. Martinelli. On the spectral gap of Kawasaki dynamics under a mixing condition revisited. Journal of Mathematical Physics, 41 (3):1391–1423, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

denHollander:2000:MNC

[89] F. den Hollander, E. Olivieri, and E. Scoppola. Metastability and nucleation for conservative dynamics. Journal of Mathematical Physics, 41 (3):1424–1498, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cooper:2000:MPS

[90] Colin Cooper, Martin E. Dyer, Alan M. Frieze, and Rachel Rue. Mixing properties of the Swendsen–Wang process on the complete graph and narrow grids. Journal of Mathematical Physics, 41(3):1499–1527, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maes:2000:DEP

[91] Christian Maes, Frank Redig, and Annelies Van Moffaert. On the definition of entropy production, via examples. Journal of Mathematical Physics, 41(3):1528–1554, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vigoda:2000:IBS

[92] Eric Vigoda. Improved bounds for sampling colorings. Journal of Mathematical Physics, 41(3):1555–1569, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Randall:2000:STA

[93] Dana Randall and Alistair Sinclair. Self-testing algorithms for selfavoiding walks. Journal of Mathematical Physics, 41(3):1570–1584, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandenBerg:2000:RSM

[94] J. van den Berg and R. Brouwer. Random sampling for the monomer– dimer model on a lattice. Journal of Mathematical Physics, 41(3):1585– 1597, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Randall:2000:AGD

[95] Dana Randall and Prasad Tetali. Analyzing Glauber dynamics by comparison of Markov chains. Journal of Mathematical Physics, 41(3):1598– 1615, March 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:2000:DTS

[96] Tuncay Aktosun. Darboux transformation for the Schrödinger equation with steplike potentials. Journal of Mathematical Physics, 41(4):1619– 1631, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bibikov:2000:FTS

[97] P. N. Bibikov. Field theory in su_{q} (2). Journal of Mathematical Physics, 41(4):1632–1646, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chaichian:2000:QFT

[98] M. Chaichian, A. Demichev, and P. Pres̆najder. Quantum field theory on the noncommutative plane with E_{q}(2) symmetry. Journal of Mathematical Physics, 41(4):1647–1671, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ekstrand:2000:NPS

[99] C. Ekstrand. Neutral particles and super Schwinger terms. Journal of Mathematical Physics, 41(4):1672–1680, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freidel:2000:SSN

[100] L. Freidel and K. Krasnov. Simple spin networks as Feynman graphs. Journal of Mathematical Physics, 41(4):1681–1690, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fry:2000:CLT

[101] M. P. Fry. Chiral limit of the two-dimensional fermionic determinant in a general magnetic field. Journal of Mathematical Physics, 41(4):1691– 1710, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Halvorson:2000:GBC

[102] Hans Halvorson and Rob Clifton. Generic Bell correlation between arbitrary local algebras in quantum field theory. Journal of Mathematical Physics, 41(4):1711–1717, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:2000:ESM

[103] M. N. Hounkonnou and G. Y. H. Avossevou. Exactly solvable models of δ′-sphere interactions in relativistic quantum mechanics. Journal of Mathematical Physics, 41(4):1718–1734, April 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See note [1255].

Hounkonnou:2000:RST

[104] M. N. Hounkonnou and G. Y. H. Avossevou. Relativistic scattering theory for finitely many δ-sphere interactions supported by concentric spheres. Journal of Mathematical Physics, 41(4):1735–1744, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See note [1255].

Jakel:2000:RSP

[105] Christian D. Jäkel. The Reeh–Schlieder property for thermal field theories. Journal of Mathematical Physics, 41(4):1745–1754, April 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joglekar:2000:RGF

[106] Satish D. Joglekar and A. Misra. Relating Green’s functions in axial and Lorentz gauges using finite field-dependent BRS transformations. Journal of Mathematical Physics, 41(4):1755–1767, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khrennikov:2000:NKP

[107] Andrei Khrennikov. Non-Kolmogorov probability models and modified Bell’s inequality. Journal of Mathematical Physics, 41(4):1768–1777, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lucha:2000:ODS

[108] Wolfgang Lucha and Franz F. Schöberl. The one-dimensional spinless relativistic Coulomb problem. Journal of Mathematical Physics, 41(4): 1778–1787, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okumura:2000:BRS

[109] Yoshitaka Okumura. Becchi–Rouet–Stora–Tyutin invariant formulation of spontaneously broken gauge theory in a generalized differential geometry. Journal of Mathematical Physics, 41(4):1788–1800, April 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Robinett:2000:WPR

[110] R. W. Robinett. Wave packet revivals and quasirevivals in onedimensional power law potentials. Journal of Mathematical Physics, 41 (4):1801–1813, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanimura:2000:RQS

[111] Shogo Tanimura and Toshihiro Iwai. Reduction of quantum systems on Riemannian manifolds with symmetry and application to molecular mechanics. Journal of Mathematical Physics, 41(4):1814–1842, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deFreitas:2000:ELS

[112] Luciane R. de Freitas. Exceptional Lagrangians for spin-2 field: Standard variables. Journal of Mathematical Physics, 41(4):1843–1853, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Efroimsky:2000:PFR

[113] Michael Efroimsky. Precession of a freely rotating rigid body. Inelastic relaxation in the vicinity of poles. Journal of Mathematical Physics, 41 (4):1854–1888, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [339].

Ferraris:2000:DLF

[114] M. Ferraris, M. Francaviglia, and M. Raiteri. Dual Lagrangian field theories. Journal of Mathematical Physics, 41(4):1889–1915, April 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Gozzi:2000:NLS

[115] E. Gozzi and D. Mauro. A new look at the Schouten–Nijenhuis, Frölicher–Nijenhuis, and Nijenhuis–Richardson brackets. Journal of Mathematical Physics, 41(4):1916–1933, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiselev:2000:HLS

[116] Aleksei P. Kiselev and Maria V. Perel. Highly localized solutions of the wave equation. Journal of Mathematical Physics, 41(4):1934–1955, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mkrtchian:2000:DEM

[117] V. E. Mkrtchian and R. v. Baltz. Dynamical electromagnetic modes for an expanding sphere. Journal of Mathematical Physics, 41(4):1956– 1960, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Neshveyev:2000:KPQ

[118] S. V. Neshveyev. On the K-property of quantized Arnold cat maps. Journal of Mathematical Physics, 41(4):1961–1965, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2000:BIS

[119] Jenn-Nan Wang. A bistatic inverse scattering problem for electromagnetic waves. Journal of Mathematical Physics, 41(4):1966–1978, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersen:2000:FGM

[120] Hans C. Andersen. Functional and graphical methods for classical statistical dynamics. I. A formulation of the Martin–Siggia–Rose method. Journal of Mathematical Physics, 41(4):1979–2020, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matolcsi:2000:MST

[121] T. Matolcsi. On the mathematical structure of thermodynamics. Journal of Mathematical Physics, 41(4):2021–2042, April 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bogoyavlenskij:2000:CPT

[122] Oleg I. Bogoyavlenskij. Counterexamples to Parker’s theorem. Journal of Mathematical Physics, 41(4):2043–2057, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Engui:2000:HNE

[123] Fan Engui and Zhang Hongqing. A hierarchy of nonlinear evolution equations, its bi-Hamiltonian structure, and finite-dimensional integrable systems. Journal of Mathematical Physics, 41(4):2058–2065, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2000:MSG

[124] Yishen Li and Chunli Chen. The motion of surfaces in geodesic coordinates and 2 + 1-dimensional breaking soliton equation. Journal of Mathematical Physics, 41(4):2066–2076, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2000:ADD

[125] Guoguang Lin, Hongjun Gao, Jinqiao Duan, and Vincent J. Ervin. Asymptotic dynamical difference between the nonlocal and local Swift– Hohenberg models. Journal of Mathematical Physics, 41(4):2077–2089, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez:2000:GTC

[126] Sonia Martínez, Jorge Cortés, and Manuel de León. The geometrical theory of constraints applied to the dynamics of vakonomic mechanical systems: The vakonomic bracket. Journal of Mathematical Physics, 41 (4):2090–2120, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:2000:DSB

[127] Manuel F. Rañada. Dynamical symmetries, bi-Hamiltonian structures, and superintegrable n = 2 systems. Journal of Mathematical Physics, 41 (4):2121–2134, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seredynska:2000:NHE

[128] M. Seredyńska and A. Hanyga. Nonlinear Hamiltonian equations with fractional damping. Journal of Mathematical Physics, 41(4):2135–2156, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2000:FDI

[129] Qi-Yan Shi and Si-Ming Zhu. A finite-dimensional integrable system and the r-matrix method. Journal of Mathematical Physics, 41(4):2157– 2166, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bokhari:2000:CCC

[130] Ashfaque H. Bokhari, Abdul R. Kashif, and Asghar Qadir. Classification of curvature collineations of plane symmetric static spacetimes. Journal of Mathematical Physics, 41(4):2167–2172, April 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2000:NET

[131] Felix Finster, Joel Smoller, and Shing-Tung Yau. Non-existence of timeperiodic solutions of the Dirac equation in a Reissner–Nordström black hole background. Journal of Mathematical Physics, 41(4):2173–2194, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Friedrich:2000:BTS

[132] Helmut Friedrich and János Kánnár. Bondi-type systems near spacelike infinity and the calculation of the Newman–Penrose constants. Journal of Mathematical Physics, 41(4):2195–2232, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bandelloni:2000:ACT

[133] G. Bandelloni and S. Lazzarini. W -algebras from canonical transformations. Journal of Mathematical Physics, 41(4):2233–2250, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ceyhan:2000:DSA

[134] Ö. Ceyhan, A. S. Fokas, and M. Gürses. Deformations of surfaces associated with integrable Gauss–Mainardi–Codazzi equations. Journal of Mathematical Physics, 41(4):2251–2270, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2000:ENS

[135] Shun-Jen Cheng, Victor G. Kac, and Minoru Wakimoto. Extensions of Neveu–Schwarz conformal modules. Journal of Mathematical Physics, 41 (4):2271–2294, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cutler:2000:RBA

[136] R. W. Cutler III and T. W. Kephart. A relation between anomaly coefficients and the group G_{2}. Journal of Mathematical Physics, 41(4): 2295–2298, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gladush:2000:GCO

[137] V. D. Gladush and R. A. Konoplya. The generalized Casimir operator and tensor representations of groups. Journal of Mathematical Physics, 41(4):2299–2309, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jakobsen:2000:CQM

[138] Hans Plesner Jakobsen and Hechun Zhang. A class of quadratic matrix algebras arising from the quantized enveloping algebra U_{⨿}(A_{∈\−∞}). Journal of Mathematical Physics, 41(4):2310–2336, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jorgensen:2000:DTS

[139] Palle E. T. Jorgensen. Off-diagonal terms in symmetric operators. Journal of Mathematical Physics, 41(4):2337–2349, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Juszczak:2000:CML

[140] Cezary Juszczak. Classical r-matrices for the osp(2 /2 ) Lie superalgebra. Journal of Mathematical Physics, 41(4):2350–2359, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kempf:2000:BHB

[141] Achim Kempf. Black holes, bandwidths and Beethoven. Journal of Mathematical Physics, 41(4):2360–2374, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kowalski:2000:GBD

[142] Kenneth L. Kowalski. Generalized binomial distributions. Journal of Mathematical Physics, 41(4):2375–2382, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lance:2000:CBC

[143] E. C. Lance and A. Paolucci. Conjugation in braided C^{∗}-categories and orthogonal quantum groups. Journal of Mathematical Physics, 41(4):

2383–2394, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moroianu:2000:PSH

[144] Andrei Moroianu and Uwe Semmelmann. Parallel spinors and holonomy groups. Journal of Mathematical Physics, 41(4):2395–2402, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parashar:2000:CQG

[145] Deepak Parashar and Roger J. McDermott. Contraction of the G_{r,s} quantum group to its nonstandard analog and corresponding colored quantum groups. Journal of Mathematical Physics, 41(4):2403–2416, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parry:2000:PDD

[146] Michelle Parry and Ephraim Fischbach. Probability distribution of distance in a uniform ellipsoid: Theory and applications to physics. Journal of Mathematical Physics, 41(4):2417–2433, April 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reinsch:2000:SET

[147] Matthias W. Reinsch. A simple expression for the terms in the Baker– Campbell–Hausdorff series. Journal of Mathematical Physics, 41(4): 2434–2442, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schork:2000:GSB

[148] Matthias Schork. Generalized bc-systems based on Hermitian vector bundles. Journal of Mathematical Physics, 41(4):2443–2459, April 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2000:VOH

[149] Wen-Li Yang and Yao-Zhong Zhang. Vertex operators of U_{q}[gl(N̂|N)] and highest weight representations of U_{q}[gl(ˆ2|2)]. Journal of Mathematical Physics, 41(4):2460–2481, April 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:2000:SPH

[150] Stephen L. Adler and Lawrence P. Horwitz. Structure and properties of Hughston’s stochastic extension of the Schrödinger equation. Journal of Mathematical Physics, 41(5):2485–2499, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [552].

Beltrametti:2000:ECM

[151] E. G. Beltrametti, S. Bugajski, and V. S. Varadarajan. Extensions of convexity models. Journal of Mathematical Physics, 41(5):2500–2514, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bjerrum-Bohr:2000:ENQ

[152] Niels Emil Jannik Bjerrum-Bohr. 1/N expansions in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 41(5):2515–2536, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bojowald:2000:GTQ

[153] Martin Bojowald and Thomas Strobl. Group theoretical quantization and the example of a phase space S^{1} × R^{+}. Journal of Mathematical Physics, 41(5):2537–2567, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bokanowski:2000:SRD

[154] Olivier Bokanowski. Symmetry requirement for a deformation operator related to density functional theory. Journal of Mathematical Physics, 41 (5):2568–2585, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brody:2000:ICQ

[155] Dorje C. Brody and Lane P. Hughston. Information content for quantum states. Journal of Mathematical Physics, 41(5):2586–2592, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2000:RHP

[156] Ariel Fernández and Andrés Colubri. Renormalized Hamiltonian for a peptide chain: Digitalizing the protein folding problem. Journal of Mathematical Physics, 41(5):2593–2603, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [340].

Ivanova:2000:DSH

[157] T. A. Ivanova and A. D. Popov. Dressing symmetries of holomorphic BF theories. Journal of Mathematical Physics, 41(5):2604–2615, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2000:BPC

[158] Ying Jiang and Yishi Duan. The branch process of cosmic strings. Journal of Mathematical Physics, 41(5):2616–2628, May 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:2000:COD

[159] E. G. Kalnins, W. Miller, Jr., and G. S. Pogosyan. Coulomb-oscillator duality in spaces of constant curvature. Journal of Mathematical Physics, 41(5):2629–2657, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Labastida:2000:CFV

[160] J. M. F. Labastida and Esther Pérez. Combinatorial formulas for Vassiliev invariants from Chern–Simons gauge theory. Journal of Mathematical Physics, 41(5):2658–2699, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leung:2000:PIM

[161] C. N. Leung and J. A. Murakowski. Padé interpolation: Methodology and application to quarkonium. Journal of Mathematical Physics, 41 (5):2700–2705, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2000:DSA

[162] Hui Li and Dimitri Kusnezov. Dynamical symmetry approach to periodic Hamiltonians. Journal of Mathematical Physics, 41(5):2706–2722, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lin:2000:PIS

[163] De-Hone Lin. Path integral solution by sum over perturbation series. Journal of Mathematical Physics, 41(5):2723–2731, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mantoiu:2000:ASS

[164] Marius Măntoiu and Serge Richard. Absence of singular spectrum for Schrödinger operators with anisotropic potentials and magnetic fields. Journal of Mathematical Physics, 41(5):2732–2740, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nieto:2000:TDSa

[165] Michael Martin Nieto and D. Rodney Truax. Time-dependent Schrödinger equations having isomorphic symmetry algebras. I. classes of interrelated equations. Journal of Mathematical Physics, 41(5):2741– 2752, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nieto:2000:TDSb

[166] Michael Martin Nieto and D. Rodney Truax. Time-dependent Schrödinger equations having isomorphic symmetry algebras. II. symmetry algebras, coherent and squeezed states. Journal of Mathematical Physics, 41(5):2753–2767, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohlsson:2000:TFN

[167] Tommy Ohlsson and Håkan Snellman. Three flavor neutrino oscillations in matter. Journal of Mathematical Physics, 41(5):2768–2788, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [638].

Presnajder:2000:OCA

[168] P. Pres̆najder. The origin of chiral anomaly and the noncommutative geometry. Journal of Mathematical Physics, 41(5):2789–2804, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pule:2000:SMS

[169] J. V. Pulé and M. Scrowston. The spectrum of a magnetic Schrödinger operator with randomly located delta impurities. Journal of Mathematical Physics, 41(5):2805–2825, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wyllard:2000:SCM

[170] Niclas Wyllard. (super)conformal many-body quantum mechanics with extended supersymmetry. Journal of Mathematical Physics, 41(5):2826– 2838, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2000:ISI

[171] Chang jun Cheng and Xian yao Chen. Inverse scattering for inhomogeneous viscoelastic media. Journal of Mathematical Physics, 41(5):2839– 2850, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurcan:2000:RMH

[172] Ö. D. Gürcan, V. V. Mirnov, and D. Üçer. Radial motion of highly conducting sphere in magnetic field. Journal of Mathematical Physics, 41(5):2851–2857, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mangiarotti:2000:CHT

[173] L. Mangiarotti and G. Sardanashvily. Constraints in Hamiltonian timedependent mechanics. Journal of Mathematical Physics, 41(5):2858– 2876, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarlet:2000:ASS

[174] W. Sarlet and A. Ramos. Adjoint symmetries, separability, and volume forms. Journal of Mathematical Physics, 41(5):2877–2888, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dye:2000:ISS

[175] J. M. Dye and A. Parker. An inverse scattering scheme for the regularized long-wave equation. Journal of Mathematical Physics, 41(5):2889–2904, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Faybusovich:2000:ETO

[176] Leonid Faybusovich and Michael Gekhtman. Elementary Toda orbits and integrable lattices. Journal of Mathematical Physics, 41(5):2905– 2921, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gascon:2000:SBN

[177] F. G. Gascón, D. Peralta Salas, and J. Ruiz Sancho. A separation bound for non-Hamiltonian differential equations with proper first integrals. Journal of Mathematical Physics, 41(5):2922–2930, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2000:DTP

[178] Nian-Ning Huang, Sien Chi, B. L. Lou, and Xiang-Jun Chen. A direct theory for the perturbed unstable nonlinear Schrödinger equation. Journal of Mathematical Physics, 41(5):2931–2942, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2000:EBS

[179] Yongsheng Li, Boling Guo, and Murong Jiang. Existence and blowup of solutions to degenerate Davey–Stewartson equations. Journal of Mathematical Physics, 41(5):2943–2956, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McSween:2000:ISH

[180] Eric McSween and Pavel Winternitz. Integrable and superintegrable Hamiltonian systems in magnetic fields. Journal of Mathematical Physics, 41(5):2957–2967, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reyes:2000:CLC

[181] Enrique G. Reyes. Conservation laws and Calapso–Guichard deformations of equations describing pseudo-spherical surfaces. Journal of Mathematical Physics, 41(5):2968–2989, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersson:2000:SCL

[182] F. Andersson and S. B. Edgar. Spin coefficients as Lanczos scalars: Underlying spinor relations. Journal of Mathematical Physics, 41(5): 2990–3001, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chauvineau:2000:GGN

[183] B. Chauvineau. A generalization of the geodesic notion. Journal of Mathematical Physics, 41(5):3002–3006, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dzhunushaliev:2000:NVS

[184] V. Dzhunushaliev and H.-J. Schmidt. New vacuum solutions of conformal Weyl gravity. Journal of Mathematical Physics, 41(5):3007–3015, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gravel:2000:SSS

[185] Pierre Gravel. Static stringlike solutions in five-dimensional relativity. II. Journal of Mathematical Physics, 41(5):3016–3027, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scipioni:2000:EOT

[186] R. Scipioni. On the extension of the Obukhov theorem in non-Riemannian gravity. I. Journal of Mathematical Physics, 41(5):3028– 3034, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vonderGonna:2000:SAF

[187] U. von der Gönna and A. Pravdová. Symmetries of asymptotically flat axisymmetric space–times with null dust. Journal of Mathematical Physics, 41(5):3035–3044, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abe:2000:CYM

[188] Mitsuko Abe and Masamichi Sato. Calabi–Yau manifolds by Borcea– Voisin method. Journal of Mathematical Physics, 41(5):3045–3056, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Askour:2000:EFR

[189] Nour Eddine Askour, Ahmed Intissar, and Zouhaïr Mouayn. Explicit formulas for reproducing kernels of generalized Bargmann spaces on C^{n}. Journal of Mathematical Physics, 41(5):3057–3067, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carbone:2000:TVI

[190] Gaspare Carbone. Turaev–Viro invariant and 3 nj symbols. Journal of Mathematical Physics, 41(5):3068–3085, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dai:2000:WOP

[191] Dao-Qing Dai. Wavelets and orthogonal polynomials based on harmonic oscillator eigenstates. Journal of Mathematical Physics, 41(5):3086–3103, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Johnson:2000:DTC

[192] G. W. Johnson and J. G. Kim. A dominated-type convergence theorem for the Feynman integral. Journal of Mathematical Physics, 41(5):3104– 3112, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalogeropoulos:2000:MSU

[193] Nikolaos Kalogeropoulos. Majorana spinors on unoriented surfaces. Journal of Mathematical Physics, 41(5):3113–3124, May 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Korotkin:2000:STE

[194] D. Korotkin, N. Manojlović, and H. Samtleben. Schlesinger transformations for elliptic isomonodromic deformations. Journal of Mathematical Physics, 41(5):3125–3141, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kosinski:2000:CDS

[195] Piotr Kosiński and Paweł Maślanka. The contraction of SU_{μ}(2) and its differential structures to E_{κ}(2). Journal of Mathematical Physics, 41 (5):3142–3159, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kraenkel:2000:IPC

[196] R. A. Kraenkel, M. Senthilvelan, and A. I. Zenchuk. On the integrable perturbations of the Camassa–Holm equation. Journal of Mathematical Physics, 41(5):3160–3169, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Laurence:2000:AMP

[197] Peter Laurence and Edward Stredulinsky. Asymptotic Massey products, induced currents and Borromean torus links. Journal of Mathematical Physics, 41(5):3170–3191, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Macfarlane:2000:CET

[198] A. J. Macfarlane and Hendryk Pfeiffer. On characteristic equations, trace identities and Casimir operators of simple Lie algebras. Journal of Mathematical Physics, 41(5):3192–3225, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [553].

Mao:2000:CAP

[199] Jian min Mao. A condition for an area-preserving mapping to be in the Engel’s form. Journal of Mathematical Physics, 41(5):3226–3232, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mehlig:2000:SPE

[200] B. Mehlig and J. T. Chalker. Statistical properties of eigenvectors in non-Hermitian Gaussian random matrix ensembles. Journal of Mathematical Physics, 41(5):3233–3256, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Minguzzi:2000:OPM

[201] Ettore Minguzzi. Orthogonal polynomial method and odd vertices in matrix models. Journal of Mathematical Physics, 41(5):3257–3267, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oteo:2000:TOP

[202] J. A. Oteo and J. Ros. From time-ordered products to Magnus expansion. Journal of Mathematical Physics, 41(5):3268–3277, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prato:2000:FLO

[203] Domingo Prato and Constantino Tsallis. Functions of linear operators: Parameter differentiation. Journal of Mathematical Physics, 41(5):3278– 3282, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhevandrov:2000:SBP

[204] P. Zhevandrov and E. Alcántar. Singular boundary perturbations for some eigenvalue problems. Journal of Mathematical Physics, 41(5):3283– 3289, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ziane:2000:CRG

[205] M. Ziane. On a certain renormalization group method. Journal of Mathematical Physics, 41(5):3290–3299, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weaver:2000:CUA

[206] Nik Weaver. Comment on “On the universality of almost every quantum logic gate”, [J. Math. Phys. 41, 240 (2000)]. Journal of Mathematical Physics, 41(5):3300, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [28].

Branson:2000:EVE

[207] Thomas P. Branson, Peter B. Gilkey, and Dmitri V. Vassilevich. Erratum: “Vacuum expectation value asymptotics for second-order differential operators on manifolds with boundary”, [J. Math. Phys. 39, 1040 (1998)]. Journal of Mathematical Physics, 41(5):3301, May 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [6].

Raghavan:2000:EVB

[208] K. N. Raghavan and P. Sankaran. Erratum: “On Verma bases for representations of sl(n, C)”, [J. Math. Phys. 40, 2190 (1999)]. Journal of Mathematical Physics, 41(5):3302, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [13].

Rainer:2000:ECC

[209] M. Rainer. Erratum: “Cones and causal structures on topological and differentiable manifolds”, [J. Math. Phys. 40, 6589 (1999)]. Journal of Mathematical Physics, 41(5):3303, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [14].

Arnold:2000:SGT

[210] V. I. Arnold. Symplectic geometry and topology. Journal of Mathematical Physics, 41(6):3307–3343, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ehlers:2000:TCW

[211] Jürgen Ehlers and Ezra T. Newman. The theory of caustics and wave front singularities with physical applications. Journal of Mathematical Physics, 41(6):3344–3378, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marsden:2000:RTL

[212] Jerrold E. Marsden, Tudor S. Ratiu, and Jürgen Scheurle. Reduction theory and the Lagrange–Routh equations. Journal of Mathematical Physics, 41(6):3379–3429, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Posch:2000:SAC

[213] H. A. Posch and W. Thirring. Some aspects of the classical three-body problem that are close or foreign to physical intuition. Journal of Mathematical Physics, 41(6):3430–3447, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hunziker:2000:QBP

[214] W. Hunziker and I. M. Sigal. The quantum N-body problem. Journal of Mathematical Physics, 41(6):3448–3510, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rokhlenko:2000:IMA

[215] Alexander Rokhlenko and Joel L. Lebowitz. Ionization of a model atom by perturbations of the potential. Journal of Mathematical Physics, 41 (6):3511–3522, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simon:2000:SOT

[216] Barry Simon. Schrödinger operators in the twentieth century. Journal of Mathematical Physics, 41(6):3523–3555, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Streater:2000:CQP

[217] R. F. Streater. Classical and quantum probability. Journal of Mathematical Physics, 41(6):3556–3603, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borchers:2000:RQF

[218] H. J. Borchers. On revolutionizing quantum field theory with Tomita’s modular theory. Journal of Mathematical Physics, 41(6):3604–3673, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buchholz:2000:QUR

[219] Detlev Buchholz and Rudolf Haag. The quest for understanding in relativistic quantum physics. Journal of Mathematical Physics, 41(6):3674– 3697, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grandjean:2000:TFB

[220] Olivier Grandjean and Arthur Jaffe. Twist fields and broken supersymmetry. Journal of Mathematical Physics, 41(6):3698–3763, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rivasseau:2000:CFT

[221] V. Rivasseau. Constructive field theory and applications: Perspectives and open problems. Journal of Mathematical Physics, 41(6):3764–3775, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rovelli:2000:CIR

[222] Carlo Rovelli. The century of the incomplete revolution: Searching for general relativistic quantum field theory. Journal of Mathematical Physics, 41(6):3776–3800, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schroer:2000:PPQ

[223] Bert Schroer. Particle physics and quantum field theory at the turn of the century: Old principles with new concepts. Journal of Mathematical Physics, 41(6):3801–3831, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Connes:2000:SSN

[224] Alain Connes. A short survey of noncommutative geometry. Journal of Mathematical Physics, 41(6):3832–3866, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kastler:2000:NGF

[225] Daniel Kastler. Noncommutative geometry and fundamental physical interactions: The Lagrangian level — Historical sketch and description of the present situation. Journal of Mathematical Physics, 41(6):3867– 3891, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:2000:QGN

[226] Shahn Majid. Quantum groups and noncommutative geometry. Journal of Mathematical Physics, 41(6):3892–3942, June 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2000:SRP

[227] Felix Finster, Joel Smoller, and Shing-Tung Yau. Some recent progress in classical general relativity. Journal of Mathematical Physics, 41(6):3943– 3963, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Regge:2000:DSG

[228] Tullio Regge and Ruth M. Williams. Discrete structures in gravity. Journal of Mathematical Physics, 41(6):3964–3984, June 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bach:2000:RE

[229] Volker Bach, Jürg Fröhlich, and Israel Michael Sigal. Return to equilibrium. Journal of Mathematical Physics, 41(6):3985–4060, June 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gallavotti:2000:NSF

[230] Giovanni Gallavotti. Nonequilibrium in statistical and fluid mechanics. Ensembles and their equivalence. Entropy driven intermittency. Journal of Mathematical Physics, 41(6):4061–4081, June 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sabatier:2000:PFI

[231] Pierre C. Sabatier. Past and future of inverse problems. Journal of Mathematical Physics, 41(6):4082–4124, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cai:2000:CTB

[232] David Cai and David W. McLaughlin. Chaotic and turbulent behavior of unstable one-dimensional nonlinear dispersive waves. Journal of Mathematical Physics, 41(6):4125–4153, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cartier:2000:FI

[233] Pierre Cartier and Cécile DeWitt-Morette. Functional integration. Journal of Mathematical Physics, 41(6):4154–4187, June 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fokas:2000:ILN

[234] A. S. Fokas. On the integrability of linear and nonlinear partial differential equations. Journal of Mathematical Physics, 41(6):4188–4237, June 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abdallah:2000:MSP

[235] Naoufel Ben Abdallah. On a multidimensional Schrödinger–Poisson scattering model for semiconductors. Journal of Mathematical Physics, 41 (7):4241–4261, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:2000:FSE

[236] Tuncay Aktosun. Factorization and small-energy asymptotics for the radial Schrödinger equation. Journal of Mathematical Physics, 41(7):

4262–4270, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arai:2000:PFH

[237] Asao Arai. A particle-field Hamiltonian in relativistic quantum electrodynamics. Journal of Mathematical Physics, 41(7):4271–4283, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baker:2000:CEB

[238] Linda Baker and David Fairlie. Companion equations for branes. Journal of Mathematical Physics, 41(7):4284–4292, July 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bechouche:2000:TEM

[239] P. Bechouche, J. Nieto, E. Ruiz Arriola, and J. Soler. On the time evolution of the mean-field polaron. Journal of Mathematical Physics, 41(7):4293–4312, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bojowald:2000:ATS

[240] Martin Bojowald. Abelian BF -theory and spherically symmetric electromagnetism. Journal of Mathematical Physics, 41(7):4313–4329, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyd:2000:PAD

[241] John K. Boyd. Probability amplitude dynamics for a two-level system. Journal of Mathematical Physics, 41(7):4330–4364, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [459].

Cattaneo:2000:USQ

[242] G. Cattaneo and G. Nisticò. From unsharp to sharp quantum observables: The general Hilbert space case. Journal of Mathematical Physics, 41(7):4365–4378, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duan:2000:TTC

[243] Yishi Duan, Libin Fu, and Guang Jia. Topological tensor current of p̃- branes in the ϕ-mapping theory. Journal of Mathematical Physics, 41 (7):4379–4386, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fine:2000:EYM

[244] Dana Fine. Energy in Yang–Mills on a Riemann surface. Journal of Mathematical Physics, 41(7):4387–4390, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frolov:2000:RLC

[245] S. V. Frolov and I. Yu. Popov. Resonances for laterally coupled quantum waveguides. Journal of Mathematical Physics, 41(7):4391–4405, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujii:2000:NCS

[246] Kazuyuki Fujii. Note on coherent states and adiabatic connections, curvatures. Journal of Mathematical Physics, 41(7):4406–4412, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Furutsu:2000:ACB

[247] H. Furutsu and T. Kojima. The U_{q}(ŝl_{n}) analogue of the XXZ chain with a boundary. Journal of Mathematical Physics, 41(7):4413–4436, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Horan:2000:ADQ

[248] Robin Horan, Martin Lavelle, and David McMullan. Asymptotic dynamics in quantum field theory. Journal of Mathematical Physics, 41 (7):4437–4451, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jaksic:2000:WOS

[249] Vojkan Jaks̆ić and Stanislav Molchanov. Wave operators for the surface Maryland model. Journal of Mathematical Physics, 41(7):4452–4463, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joglekar:2000:SFA

[250] Satish D. Joglekar and Bhabani Prasad Mandal. A superspace formulation of an “asymptotic” OSp(3, 1|2) invariance of Yang–Mills theories. Journal of Mathematical Physics, 41(7):4464–4477, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jung:2000:PST

[251] Klaus Jung. Phase space tunneling for operators with symbols in a Gevrey class. Journal of Mathematical Physics, 41(7):4478–4496, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lam:2000:TOE

[252] C. S. Lam. Time ordering, energy ordering, and factorization. Journal of Mathematical Physics, 41(7):4497–4507, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Malyshev:2000:QG

[253] V. A. Malyshev. Quantum grammars. Journal of Mathematical Physics, 41(7):4508–4520, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nesterenko:2000:SZF

[254] V. V. Nesterenko and I. G. Pirozhenko. Spectral zeta functions for a cylinder and a circle. Journal of Mathematical Physics, 41(7):4521–4531, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:2000:DDS

[255] Yūji Ohta. Direct derivation of scaling relation of prepotential in N = 2 supersymmetric G_{2} Yang–Mills theory. Journal of Mathematical Physics, 41(7):4532–4540, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:2000:ICP

[256] Yūji Ohta. Instanton correction of prepotential in Ruijsenaars model associated with N = 2 SU (2 ) Seiberg–Witten theory. Journal of Mathematical Physics, 41(7):4541–4550, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paul:2000:ELF

[257] Prabasaj Paul and Daniel Nkemzi. On the energy levels of a finite squarewell potential. Journal of Mathematical Physics, 41(7):4551–4555, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [491].

RauschdeTraubenberg:2000:FST

[258] M. Rausch de Traubenberg and M. J. Slupinski. Fractional supersymmetry and F th-roots of representations. Journal of Mathematical Physics, 41(7):4556–4571, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reitz:2000:NEB

[259] John R. Reitz and Frederick J. Mayer. New electromagnetic bound states. Journal of Mathematical Physics, 41(7):4572–4581, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salehi:2000:STT

[260] H. Salehi, Y. Bisabr, and H. Ghafarnejad. Scalar tensor theories and Hadamard state condition. Journal of Mathematical Physics, 41(7):4582– 4589, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scolarici:2000:QSG

[261] Giuseppe Scolarici and Luigi Solombrino. Quaternionic symmetry groups and particle multiplets. Journal of Mathematical Physics, 41(7):4590– 4603, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sniady:2000:IFT

[262] Piotr Şniady and Marcin Zygmunt. The U(1)-invariant field theories with normal field operators. Journal of Mathematical Physics, 41(7):4604– 4606, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aaberge:2000:CQL

[263] Terje Aaberge. Classical and quantal Lorentz covariant models of the electromagnetic radiation field. Journal of Mathematical Physics, 41(7): 4607–4621, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Woodside:2000:CFV

[264] Dale A. Woodside. Classical four-vector fields in the relativistic longitudinal gauge. Journal of Mathematical Physics, 41(7):4622–4653, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benenti:2000:SST

[265] S. Benenti, C. Chanu, and G. Rastelli. The super-separability of the three-body inverse-square Calogero system. Journal of Mathematical Physics, 41(7):4654–4678, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caseiro:2000:ALD

[266] R. Caseiro, J.-P. Françoise, and R. Sasaki. Algebraic linearization of dynamics of Calogero type for any Coxeter group. Journal of Mathematical Physics, 41(7):4679–4686, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2000:DTD

[267] Jining Gao. Darboux transformation for differential–difference principal chiral equation and its continuous limits. Journal of Mathematical Physics, 41(7):4687–4694, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gawedzki:2000:HSL

[268] Krzysztof Gawedzki and Pascal Tran-Ngoc-Bich. Hitchin systems at low genera. Journal of Mathematical Physics, 41(7):4695–4712, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gordoa:2000:SES

[269] Pilar R. Gordoa. Symmetries, exact solutions, and nonlinear superposition formulas for two integrable partial differential equations. Journal of Mathematical Physics, 41(7):4713–4731, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mestel:2000:ASP

[270] B. D. Mestel and A. H. Osbaldestin. Asymptotics of scaling parameters for period-doubling in unimodal maps with asymmetric critical points. Journal of Mathematical Physics, 41(7):4732–4746, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Toda:2000:ISK

[271] Kouichi Toda and Song-Ju Yu. The investigation into the Schwarz– Korteweg–de Vries equation and the Schwarz derivative in (2+1) dimensions. Journal of Mathematical Physics, 41(7):4747–4751, July 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Harko:2000:ACF

[272] T. Harko and M. K. Mak. Anisotropic charged fluid spheres in D space– time dimensions. Journal of Mathematical Physics, 41(7):4752–4764, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrero:2000:KMO

[273] Alicia Herrero and Juan Antonio Morales. A kinematic method to obtain conformal factors. Journal of Mathematical Physics, 41(7):4765– 4776, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manojlovic:2000:BZF

[274] Nenad Manojlović and Aleksandar Miković. Belinskii–Zakharov formulation for Bianchi models and Painlevé III equation. Journal of Mathematical Physics, 41(7):4777–4782, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peca:2000:TTB

[275] Claudia S. Peça and José P. S. Lemos. Thermodynamics of toroidal black holes. Journal of Mathematical Physics, 41(7):4783–4789, July 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zotov:2000:DSA

[276] M. Yu. Zotov. Dynamical system analysis for the Einstein–Yang–Mills equations. Journal of Mathematical Physics, 41(7):4790–4807, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anderson:2000:UMBa

[277] Gregory W. Anderson and Tomás̆ Blažek. E_{6} unification model building. I. Clebsch–Gordan coefficients of 27 ⊗ \overline{27}. Journal of Mathematical Physics, 41(7):4808–4816, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arratia:2000:IRQ

[278] Oscar Arratia and Mariano A. del Olmo. Induced representations of quantum (1 + 1) Galilei algebras. Journal of Mathematical Physics, 41 (7):4817–4832, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassinelli:2000:SII

[279] G. Cassinelli, E. De Vito, and A. Levrero. Square-integrable imprimitivity systems. Journal of Mathematical Physics, 41(7):4833–4859, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGuise:2000:GME

[280] H. de Guise, J. Patera, and R. T. Sharp. SL(3, C) generator matrix elements in a Pauli subgroup basis. Journal of Mathematical Physics, 41 (7):4860–4880, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fishman:2000:ECS

[281] Louis Fishman, Maarten V. de Hoop, and Mattheus J. N. van Stralen. Exact constructions of square-root Helmholtz operator symbols: The focusing quadratic profile. Journal of Mathematical Physics, 41(7):4881– 4938, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [774].

Gomez:2000:CTD

[282] Xavier Gomez. Classification of three-dimensional Lie bialgebras. Journal of Mathematical Physics, 41(7):4939–4956, July 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2000:DGA

[283] Yu Huang, Zhao Yi, and Zhaoyang Yin. On the dimension of the global attractor for a damped semilinear wave equation with critical exponent. Journal of Mathematical Physics, 41(7):4957–4966, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hudson:2000:ATP

[284] R. L. Hudson and S. Pulmannová. Algebraic theory of product integrals in quantum stochastic calculus. Journal of Mathematical Physics, 41 (7):4967–4980, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2000:CKM

[285] Jeong-Ah Kim and Dong-Uy Shin. Construction of Kac–Moody superalgebras as minimal graded Lie superalgebras and weight multiplicities for Kac–Moody superalgebras. Journal of Mathematical Physics, 41(7): 4981–5001, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

King:2000:ABFa

[286] R. C. King and B. G. Wybourne. Analogies between finite-dimensional irreducible representations of SO(2n) and infinite-dimensional irreducible representations of Sp(2n, R). I. characters and products. Journal of Mathematical Physics, 41(7):5002–5019, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Links:2000:TIL

[287] J. R. Links, M. D. Gould, and Y.-Z. Zhang. Twisting invariance of link polynomials derived from ribbon quasi-Hopf algebras. Journal of Mathematical Physics, 41(7):5020–5032, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:2000:AMV

[288] Xing-Bin Pan and Yuanwei Qi. Asymptotics of minimizers of variational problems involving curl functional. Journal of Mathematical Physics, 41 (7):5033–5063, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Su:2000:PVK

[289] Yucai Su, J. W. B. Hughes, and R. C. King. Primitive vectors of Kacmodules of the Lie superalgebras sl(m/n). Journal of Mathematical Physics, 41(7):5064–5087, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Twarock:2000:AVA

[290] R. Twarock. Aperiodic Virasoro algebra. Journal of Mathematical Physics, 41(7):5088–5106, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Venkataramani:2000:LSC

[291] S. C. Venkataramani, T. A. Witten, E. M. Kramer, and R. P. Geroch. Limitations on the smooth confinement of an unstretchable manifold. Journal of Mathematical Physics, 41(7):5107–5128, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zachos:2000:GES

[292] Cosmas Zachos. Geometrical evaluation of star products. Journal of Mathematical Physics, 41(7):5129–5134, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saenz:2000:ELS

[293] A. W. Sáenz. Erratum: “Lie-series approach to the evolution of resonant and nonresonant anharmonic oscillators in quantum mechanics”, [J. Math. Phys. 39, 1887 (1998)]. Journal of Mathematical Physics, 41 (7):5135, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [7].

Chalykh:2000:BQR

[294] Oleg A. Chalykh. Bispectrality for the quantum Ruijsenaars model and its integrable deformation. Journal of Mathematical Physics, 41(8):5139– 5167, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heller:2000:SVR

[295] Michael Heller, Wiesław Sasin, and Zdzisław Odrzygózdz. State vector reduction as a shadow of a noncommutative dynamics. Journal of Mathematical Physics, 41(8):5168–5179, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krausz:2000:EEO

[296] Nurit Krausz and M. S. Marinov. Exact evolution operator on noncompact group manifolds. Journal of Mathematical Physics, 41(8):5180– 5208, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakamura:2000:PSM

[297] Toru Nakamura. Path space measure for the 3 + 1-dimensional Dirac equation in momentum space. Journal of Mathematical Physics, 41(8): 5209–5222, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nicoleau:2000:ISP

[298] François Nicoleau. An inverse scattering problem with the Aharonov– Bohm effect. Journal of Mathematical Physics, 41(8):5223–5237, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okumura:2000:GCS

[299] Yoshitaka Okumura. Generalized Chern–Simons form and descent equation. Journal of Mathematical Physics, 41(8):5238–5244, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sardanashvily:2000:CQM

[300] G. Sardanashvily. Classical and quantum mechanics with time-dependent parameters. Journal of Mathematical Physics, 41(8):5245–5255, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vuille:2000:SEG

[301] Chris Vuille. Schrödinger’s equation and General Relativity. Journal of Mathematical Physics, 41(8):5256–5261, August 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weaver:2000:TOC

[302] Nik Weaver. Time-optimal control of finite quantum systems. Journal of Mathematical Physics, 41(8):5262–5269, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wichnoski:2000:RSM

[303] Bruno Wichnoski. A representation space for minimal coupling. Journal of Mathematical Physics, 41(8):5270–5276, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2000:VOS

[304] Yao-Zhong Zhang and Mark D. Gould. U_{q}[sl(ˆ2|)] vertex operators, screen currents, and correlation functions at an arbitrary level. Journal of Mathematical Physics, 41(8):5277–5291, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2000:GIC

[305] Antonio Fernández, Pedro L. García, and J. Muñoz Masqué. Gaugeinvariant covariant Hamiltonians. Journal of Mathematical Physics, 41 (8):5292–5303, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krupkova:2000:HOM

[306] Olga Krupková. Higher-order mechanical systems with constraints. Journal of Mathematical Physics, 41(8):5304–5324, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rebollo-Neira:2000:RJT

[307] L. Rebollo-Neira, A. Plastino, and J. Fernandez-Rubio. Reconstruction of the joint time-delay Doppler-scale reflectivity density in the wideband regime. A frame theory based approach. Journal of Mathematical Physics, 41(8):5325–5341, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saenz:2000:DAC

[308] A. W. Sáenz. Dynamics of axial channeling in quasicrystals: An averaging-theory approach. Journal of Mathematical Physics, 41(8): 5342–5368, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [664].

Soloviev:2000:BPB

[309] Vladimir O. Soloviev. Bering’s proposal for boundary contribution to the Poisson bracket. Journal of Mathematical Physics, 41(8):5369–5380, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Madureira:2000:MLN

[310] Justino R. Madureira. The Markovian limit in a nonlinear quantum kinetic theory. Journal of Mathematical Physics, 41(8):5381–5390, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2000:MMB

[311] Jen-Hsu Chang and Ming-Hsien Tu. On the Miura Map between the dispersionless KP and dispersionless modified KP hierarchies. Journal of Mathematical Physics, 41(8):5391–5406, August 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forger:2000:LSMa

[312] Michael Forger and Sebastian Sachse. Lie superalgebras and the multiplet structure of the genetic code. I. codon representations. Journal of Mathematical Physics, 41(8):5407–5422, August 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forger:2000:LSMb

[313] Michael Forger and Sebastian Sachse. Lie superalgebras and the multiplet structure of the genetic code. II. branching schemes. Journal of Mathematical Physics, 41(8):5423–5444, August 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mondejar:2000:NGV

[314] Félix Mondéjar and Sebastián Ferrer. On the nonintegrability of the generalized van der Waals Hamiltonian system. Journal of Mathematical Physics, 41(8):5445–5452, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:2000:ISH

[315] Yunbo Zeng, Wen-Xiu Ma, and Runliang Lin. Integration of the soliton hierarchy with self-consistent sources. Journal of Mathematical Physics, 41(8):5453–5489, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:2000:FEA

[316] Y. Brihaye, A. Chakrabarti, and D. H. Tchrakian. Finite energy–action solutions of p_{1} Yang–Mills equations on p_{2} Schwarzschild and deSitter backgrounds in dimensions d ≥ 4. Journal of Mathematical Physics, 41 (8):5490–5509, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chamblin:2000:EMF

[317] Andrew Chamblin. Existence of Majorana fermions for M-branes wrapped in space and time. Journal of Mathematical Physics, 41(8): 5510–5516, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forni:2000:NSF

[318] Diego M. Forni, Mirta Iriondo, and Carlos N. Kozameh. Null surfaces formulation in three dimensions. Journal of Mathematical Physics, 41(8): 5517–5534, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:2000:IPE

[319] Simonetta Frittelli and Roberto Gomez. Ill-posedness in the Einstein equations. Journal of Mathematical Physics, 41(8):5535–5549, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pelykh:2000:EST

[320] Volodymyr Pelykh. Equivalence of the spinor and tensor methods in the positive energy problem. Journal of Mathematical Physics, 41(8): 5550–5556, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pons:2000:GTE

[321] J. M. Pons, D. C. Salisbury, and L. C. Shepley. Gauge transformation in Einstein–Yang–Mills theories. Journal of Mathematical Physics, 41(8):

5557–5571, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tod:2000:EWS

[322] K. P. Tod. Einstein–Weyl spaces and third-order differential equations. Journal of Mathematical Physics, 41(8):5572–5581, August 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Winter:2000:ACO

[323] Dale Winter. Analysis of the cosmological Oppenheimer–Volkoff equations. Journal of Mathematical Physics, 41(8):5582–5597, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abramov:2000:EDH

[324] V. Abramov and R. Kerner. Exterior differentials of higher order and their covariant generalization. Journal of Mathematical Physics, 41(8): 5598–5614, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cantarella:2000:IPH

[325] Jason Cantarella, Dennis DeTurck, Herman Gluck, and Mikhail Teytel. Isoperimetric problems for the helicity of vector fields and the Biot– Savart and curl operators. Journal of Mathematical Physics, 41(8):5615– 5641, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Janeczko:2000:LSP

[326] S. Janeczko. Lagrangian submanifolds in product symplectic spaces. Journal of Mathematical Physics, 41(8):5642–5655, August 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

King:2000:ABFb

[327] R. C. King and B. G. Wybourne. Analogies between finite-dimensional irreducible representations of SO(2n) and infinite-dimensional irreducible representations of Sp(2n, R). II. plethysms. Journal of Mathematical Physics, 41(8):5656–5690, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liebling:2000:CPN

[328] Steven L. Liebling, Eric W. Hirschmann, and James Isenberg. Critical phenomena in nonlinear sigma models. Journal of Mathematical Physics, 41(8):5691–5700, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mandelkern:2000:SAG

[329] Mark Mandelkern and Jonas Schultz. The statistical analysis of Gaussian and Poisson signals near physical boundaries. Journal of Mathematical Physics, 41(8):5701–5709, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manjavidze:2000:TPT

[330] J. Manjavidze. Topology and perturbation theory. Journal of Mathematical Physics, 41(8):5710–5734, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marotta:2000:TOW

[331] V. Marotta and A. Sciarrino. Tensor operators and Wigner–Eckart theorem for U_{q→0}(sl(2)). Journal of Mathematical Physics, 41(8):5735–5744, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Otway:2000:NHM

[332] Thomas H. Otway. Nonlinear Hodge maps. Journal of Mathematical Physics, 41(8):5745–5766, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saitoh:2000:DPR

[333] Naoko Saitoh and Hiroaki Yoshida. q-deformed Poisson random variables on q-Fock space. Journal of Mathematical Physics, 41(8):5767–5772, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tian:2000:WAI

[334] Lixin Tian. Wavelet approximate inertial manifold in nonlinear solitary wave equation. Journal of Mathematical Physics, 41(8):5773–5792, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsaune:2000:GSS

[335] A. Ya. Tsaune and M. P. Dyachenko. General scheme of splittings of degenerate eigensubspaces and eigenelements of self-adjoint operators in high orders of perturbation theory. Journal of Mathematical Physics, 41 (8):5793–5813, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Walkup:2000:LOS

[336] J. R. Walkup, M. Dunn, and D. K. Watson. Local optimization of the summation of divergent power series. Journal of Mathematical Physics, 41(8):5814–5831, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2000:LPS

[337] Yongtang Wu and Dianlou Du. On the Lie–Poisson structure of the nonlinearized discrete eigenvalue problem. Journal of Mathematical Physics, 41(8):5832–5848, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2000:LOH

[338] Wen-Li Yang and Yao-Zhong Zhang. Level-one highest weight representation of U_{q}[sl(N̂ |1 )] and Bosonization of the multicomponent Super t–J model. Journal of Mathematical Physics, 41(8):5849–5869, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Efroimsky:2000:EPF

[339] Michael Efroimsky. Erratum: “Precession of a freely rotating rigid body. Inelastic relaxation in the vicinity of poles”, [J. Math. Phys. 41, 1854 (2000)]. Journal of Mathematical Physics, 41(8):5870, August 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [113].

Fernandez:2000:ERH

[340] Ariel Fernández and Andrés Colubri. Erratum: “Renormalized Hamiltonian for a peptide chain: Digitalizing the protein folding problem”, [J. Math. Phys. 41, 2593 (2000)]. Journal of Mathematical Physics, 41 (8):5871, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [156].

Adam:2000:NSS

[341] Christoph Adam, Bruno Muratori, and Charles Nash. Non- L^{2} solutions to the Seiberg–Witten equations. Journal of Mathematical Physics, 41 (9):5875–5882, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassinelli:2000:PSO

[342] Gianni Cassinelli, Ernesto De Vito, Pekka Lahti, and Alberto Levrero. Phase space observables and isotypic spaces. Journal of Mathematical Physics, 41(9):5883–5896, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feldmann:2000:CHO

[343] Tova Feldmann, M. Cohen, and B. L. Burrows. Coupled harmonic oscillator systems: An elementary algebraic decoupling approach. Journal of Mathematical Physics, 41(9):5897–5909, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaier:2000:GMA

[344] Johanna Gaier and Jakob Yngvason. Geometric modular action, wedge duality, and Lorentz covariance are equivalent for generalized free fields. Journal of Mathematical Physics, 41(9):5910–5919, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gazeau:2000:MVF

[345] J-P. Gazeau and M. V. Takook. “massive” vector field in de Sitter space. Journal of Mathematical Physics, 41(9):5920–5933, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1254].

Khrennikov:2000:PCI

[346] Andrei Khrennikov. A perturbation of CHSH inequality induced by fluctuations of ensemble distributions. Journal of Mathematical Physics, 41 (9):5934–5944, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kupsch:2000:RID

[347] Joachim Kupsch. The role of infrared divergence for decoherence. Journal of Mathematical Physics, 41(9):5945–5953, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

LeGuennec:2000:TDTa

[348] Patrick Le Guennec. Two-dimensional theory of chirality. I. absolute chirality. Journal of Mathematical Physics, 41(9):5954–5985, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

LeGuennec:2000:TDTb

[349] Patrick Le Guennec. Two-dimensional theory of chirality. II. relative chirality and the chirality of complex fields. Journal of Mathematical Physics, 41(9):5986–6006, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McLeod:2000:UAP

[350] J. Bryce McLeod and Yisong Yang. The uniqueness and approximation of a positive solution of the Bardeen–Cooper–Schrieffer gap equation. Journal of Mathematical Physics, 41(9):6007–6025, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muljarov:2000:HTA

[351] E. A. Muljarov, A. L. Yablonskii, S. G. Tikhodeev, A. E. Bulatov, and Joseph L. Birman. Hyperspherical theory of anisotropic exciton. Journal of Mathematical Physics, 41(9):6026–6041, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:2000:DWE

[352] Yūji Ohta. Dual WDVV equations in N = 2 supersymmetric Yang–Mills theory. Journal of Mathematical Physics, 41(9):6042–6047, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okumura:2000:RSM

[353] Yoshitaka Okumura and Kohei Mori. Reformulation of the standard model in a generalized differential geometry on the discrete space M_{4} × X_{2}. Journal of Mathematical Physics, 41(9):6048–6059, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pavon:2000:SMF

[354] Michele Pavon. Stochastic mechanics and the Feynman integral. Journal of Mathematical Physics, 41(9):6060–6078, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramacher:2000:MLE

[355] Pablo Ramacher. Modular localization of elementary systems in the theory of Wigner. Journal of Mathematical Physics, 41(9):6079–6089, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schwartz:2000:ECW

[356] Moshe Schwartz, Avishay Efrat, and James L. Monroe. Effect of couplings weakening and reversing in ferromagnetic Ising systems — Rigorous inequalities. Journal of Mathematical Physics, 41(9):6090–6092, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VonZuben:2000:QTS

[357] Francis S. G. Von Zuben. Quantum time and spatial localization: An analysis of the Hegerfeldt paradox. Journal of Mathematical Physics, 41 (9):6093–6115, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1956].

BoutetdeMonvel:2000:FDI

[358] Anne Boutet de Monvel and Dmitry Shepelsky. A frequency–domain inverse problem for a dispersive stratified chiral medium. Journal of Mathematical Physics, 41(9):6116–6129, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Margetis:2000:TDH

[359] Dionisios Margetis and George Fikioris. Two-dimensional, highly directive currents on large circular loops. Journal of Mathematical Physics, 41(9):6130–6172, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1729].

Foursov:2000:CCI

[360] Mikhail V. Foursov. Classification of certain integrable coupled potential KdV and modified KdV-type equations. Journal of Mathematical Physics, 41(9):6173–6185, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giraldo:2000:NPG

[361] Luis Giraldo and Francisco G. Gascón. New proof and generalizations of the Demidowitsch–Schneider criterion. Journal of Mathematical Physics, 41(9):6186–6192, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nunez:2000:ISP

[362] Manuel Núñez. Invariant subspaces of the periodic Navier–Stokes and magnetohydrodynamics equations: Symmetries and inverse cascades. Journal of Mathematical Physics, 41(9):6193–6197, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Radak:2000:MIN

[363] Branislav Radak. Moment invariants for nonlinear Hamiltonian systems. Journal of Mathematical Physics, 41(9):6198–6232, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sano:2000:GSW

[364] Yukio Sano and Isamu Miyamoto. Generalized smooth and weakdiscontinuous unsteady waves. Journal of Mathematical Physics, 41(9): 6233–6247, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zenchuk:2000:MHD

[365] A. I. Zenchuk. Multidimensional hierarchies of (1 + 1)-dimensional integrable partial differential equations. nonsymmetric ¯∂-dressing. Journal of Mathematical Physics, 41(9):6248–6276, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Billyard:2000:CBE

[366] Andrew P. Billyard, Alan A. Coley, and James E. Lidsey. Cyclical behavior in early universe cosmologies. Journal of Mathematical Physics, 41(9):6277–6283, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bozza:2000:SCB

[367] V. Bozza. Signs of the cusps in binary lenses. Journal of Mathematical Physics, 41(9):6284–6295, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dain:2000:GST

[368] Sergio Dain and Osvaldo M. Moreschi. The Goldberg–Sachs theorem in linearized gravity. Journal of Mathematical Physics, 41(9):6296–6299, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:2000:SRN

[369] Simonetta Frittelli, Lionel Mason, and Ezra T. Newman. A spinor reformulation of the null surface treatment of GR. Journal of Mathematical Physics, 41(9):6300–6317, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Itin:2000:CIC

[370] Yakov Itin and Shmuel Kaniel. On a class of invariant coframe operators with application to gravity. Journal of Mathematical Physics, 41(9): 6318–6340, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivashchuk:2000:BRM

[371] V. D. Ivashchuk and V. N. Melnikov. Billiard representation for multidimensional cosmology with intersecting p-branes near the singularity. Journal of Mathematical Physics, 41(9):6341–6363, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Low:2000:SCS

[372] Robert J. Low. The structure of cosmic strings. Journal of Mathematical Physics, 41(9):6364–6368, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Macias:2000:CCW

[373] Alfredo Macías, Claus Lämmerzahl, and Alberto García. A class of colliding waves in metric-affine gravity, nonmetricity and torsion shock waves. Journal of Mathematical Physics, 41(9):6369–6380, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:2000:CIZ

[374] Carl M. Bender, Stefan Boettcher, and Van M. Savage. Conjecture on the interlacing of zeros in complex Sturm–Liouville problems. Journal of Mathematical Physics, 41(9):6381–6387, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:2000:USP

[375] Carl M. Bender, Gerald V. Dunne, and Lawrence R. Mead. Underdetermined systems of partial differential equations. Journal of Mathematical Physics, 41(9):6388–6398, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2000:NST

[376] F. Calogero, A. Degasperis, and Ji Xiaoda. Nonlinear Schrödinger-type equations from multiscale reduction of PDEs. I. Systematic derivation. Journal of Mathematical Physics, 41(9):6399–6443, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferrari:2000:BFP

[377] Franco Ferrari and Jan Sobczyk. Bosonic field propagators on algebraic curves. Journal of Mathematical Physics, 41(9):6444–6462, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2000:NTA

[378] Ying Jiang and Yishi Duan. A new topological aspect of the arbitrary dimensional topological defects. Journal of Mathematical Physics, 41 (9):6463–6476, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuniba:2000:PDC

[379] Atsuo Kuniba, Kailash C. Misra, Masato Okado, Taichiro Takagi, and Jun Uchiyama. Paths, Demazure crystals, and symmetric functions. Journal of Mathematical Physics, 41(9):6477–6486, September 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ky:2000:IRT

[380] Nguyen Anh Ky. Induced representations of the two parametric quantum deformation U_{pq}[gl(2/2)]. Journal of Mathematical Physics, 41(9):6487– 6508, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:2000:SAE

[381] Sen yue Lou. Symmetry analysis and exact solutions of the 2 + 1 dimensional sine-Gordon system. Journal of Mathematical Physics, 41(9): 6509–6524, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Low:2000:CCR

[382] Robert J. Low. Continuous causal representation of space–time via sphere orderings. Journal of Mathematical Physics, 41(9):6525–6528, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luan:2000:CSY

[383] Pi-Gang Luan, H. C. Lee, and R. B. Zhang. Colored solutions of the Yang–Baxter equation from representations of U_{q} gl (2). Journal of Mathematical Physics, 41(9):6529–6543, September 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:2000:CGC

[384] D. J. Rowe and C. Bahri. Clebsch–Gordan coefficients of SU (3 ) in SU (2 ) and SO(3 ) bases. Journal of Mathematical Physics, 41(9):6544–6565, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schief:2000:LDT

[385] W. K. Schief. On Laplace–Darboux-type sequences of generalized Weingarten surfaces. Journal of Mathematical Physics, 41(9):6566–6599, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanAssche:2000:FGP

[386] W. Van Assche, R. J. Yáñez, R. González-Férez, and Jesús S. Dehesa. Functionals of Gegenbauer polynomials and D-dimensional hydrogenic momentum expectation values. Journal of Mathematical Physics, 41(9): 6600–6613, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2000:CEL

[387] Jianke Yang. Complete eigenfunctions of linearized integrable equations expanded around a soliton solution. Journal of Mathematical Physics, 41(9):6614–6638, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2000:SRQ

[388] R. B. Zhang. Structure and representations on the quantum supergroup OSP_{q} (2—2 n ). Journal of Mathematical Physics, 41(9):6639–6656, September 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barcelos-Neto:2000:SCT

[389] J. Barcelos-Neto. Spontaneous compactification of D = 5 two-form gauge fields and the obtainment of Maxwell and Yang–Mills theories. Journal of Mathematical Physics, 41(10):6661–6670, October 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DePietri:2000:FDG

[390] Roberto De Pietri and Carlo Petronio. Feynman diagrams of generalized matrix models and the associated manifolds in dimension four. Journal of Mathematical Physics, 41(10):6671–6688, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:2000:LCE

[391] Felix Finster. Light-cone expansion of the Dirac sea in the presence of chiral and scalar potentials. Journal of Mathematical Physics, 41(10): 6689–6746, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerrero:2000:IMP

[392] J. Guerrero and V. Aldaya. Invariant measures on polarized submanifolds in group quantization. Journal of Mathematical Physics, 41(10):6747– 6765, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lockhart:2000:OEL

[393] Robert Lockhart. Optimal ensemble length of mixed separable states. Journal of Mathematical Physics, 41(10):6766–6771, October 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vollbrecht:2000:WTQ

[394] K. G. H. Vollbrecht and R. F. Werner. Why two qubits are special. Journal of Mathematical Physics, 41(10):6772–6782, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clegg:2000:RIL

[395] J. R. Clegg, P. K. Browning, P. Laurence, B. J. I. Bromage, and E. Stredulinsky. On the representation of inhomogeneous linear forcefree fields. Journal of Mathematical Physics, 41(10):6783–6807, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McLean:2000:CFT

[396] M. McLean and L. K. Norris. Covariant field theory on frame bundles of fibered manifolds. Journal of Mathematical Physics, 41(10):6808–6823, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saller:2000:SCC

[397] Dirk Saller and Raffaele Vitolo. Symmetries in covariant classical mechanics. Journal of Mathematical Physics, 41(10):6824–6842, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garbaczewski:2000:OUC

[398] Piotr Garbaczewski and Robert Olkiewicz. Ornstein–Uhlenbeck–Cauchy process. Journal of Mathematical Physics, 41(10):6843–6860, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyazawa:2000:GLE

[399] Toru Miyazawa. Generalized low-energy expansion formula for Green’s function of the Fokker–Planck equation. Journal of Mathematical Physics, 41(10):6861–6878, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McLenaghan:2000:CLI

[400] Raymond G. McLenaghan and Roman G. Smirnov. A class of Liouvilleintegrable Hamiltonian systems with two degrees of freedom. Journal of Mathematical Physics, 41(10):6879–6889, October 2000. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banach:2000:ECP

[401] Zbigniew Banach and Slawomir Piekarski. Equivalence classes of perturbations in cosmologies of Bianchi types I and V: Propagation and constraint equations. Journal of Mathematical Physics, 41(10):6890–6905, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goliath:2000:ITC

[402] Martin Goliath and Ulf S. Nilsson. Isotropization of two-component fluids. Journal of Mathematical Physics, 41(10):6906–6917, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mustafa:2000:AHM

[403] M. T. Mustafa. Applications of harmonic morphisms to gravity. Journal of Mathematical Physics, 41(10):6918–6929, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wasserman:2000:SSS

[404] Arthur G. Wasserman. Solutions of the spherically symmetric SU (2 ) Einstein–Yang–Mills equations defined in the far field. Journal of Mathematical Physics, 41(10):6930–6936, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Askour:2000:SDR

[405] N. Askour and Z. Mouayn. Spectral decomposition and resolvent kernel for a magnetic Laplacian in C^{n}. Journal of Mathematical Physics, 41(10): 6937–6943, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bombelli:2000:SLG

[406] Luca Bombelli. Statistical Lorentzian geometry and the closeness of Lorentzian manifolds. Journal of Mathematical Physics, 41(10):6944– 6958, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borowiec:2000:CPA

[407] Andrzej Borowiec and Władysław Marcinek. On crossed product of algebras. Journal of Mathematical Physics, 41(10):6959–6975, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:2000:DG

[408] Salih Çelik. Differential geometry of GL_{p,q}(1|1). Journal of Mathematical Physics, 41(10):6976–6994, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Conradt:2000:MSC

[409] Oliver Conradt. Mechanics in space and counterspace. Journal of Mathematical Physics, 41(10):6995–7028, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goenner:2000:ESG

[410] Hubert Goenner and Peter Havas. Exact solutions of the generalized Lane–Emden equation. Journal of Mathematical Physics, 41(10):7029– 7042, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grosse:2000:SQQ

[411] Harald Grosse and Karl-Georg Schlesinger. On second quantization of quantum groups. Journal of Mathematical Physics, 41(10):7043–7060, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsuno:2000:LSA

[412] Yoshimasa Matsuno. Linear stability analysis for the multikink solution of the sine-Hilbert equation. Journal of Mathematical Physics, 41(10): 7061–7078, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miki:2000:RQT

[413] Kei Miki. Representations of quantum toroidal algebra U_{q}(sl_{n+1 ,tor}) (n ≥ 2). Journal of Mathematical Physics, 41(10):7079–7098, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morame:2000:ACS

[414] Abderemane Morame. The absolute continuity of the spectrum of Maxwell operator in a periodic media. Journal of Mathematical Physics, 41(10):7099–7108, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morgan:2000:RSD

[415] Dennis R. Morgan. A remarkable sequence derived from Euler products. Journal of Mathematical Physics, 41(10):7109–7121, October 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parfionov:2000:CDP

[416] G. N. Parfionov and Roman R. Zapatrin. Connes duality in pseudo-Riemannian geometry. Journal of Mathematical Physics, 41(10):7122– 7128, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:2000:SST

[417] Jeong-Hyuck Park. Superconformal symmetry in three dimensions. Journal of Mathematical Physics, 41(10):7129–7161, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnold:2000:DVS

[418] Anton Arnold, Horst Lange, and Paul F. Zweifel. A discrete-velocity, stationary Wigner equation. Journal of Mathematical Physics, 41(11): 7167–7180, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barchielli:2000:QSM

[419] Alberto Barchielli and Giancarlo Lupieri. Quantum stochastic models of two-level atoms and electromagnetic cross sections. Journal of Mathematical Physics, 41(11):7181–7205, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basu:2000:NRN

[420] Anirban Basu and Satish D. Joglekar. Nonlocal regularization for non-Abelian gauge theories for arbitrary gauge parameter. Journal of Mathematical Physics, 41(11):7206–7219, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belavkin:2000:QSD

[421] V. P. Belavkin and P. Staszewski. Quantum stochastic differential equation for unstable systems. Journal of Mathematical Physics, 41(11):7220– 7233, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergere:2000:CPE

[422] M. C. Bergère. “composite particles” and the eigenstates of Calogero– Sutherland and Ruijsenaars–Schneider. Journal of Mathematical Physics, 41(11):7234–7251, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergere:2000:FS

[423] M. C. Bergère. Fractional statistic. Journal of Mathematical Physics, 41(11):7252–7262, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolley:2000:SFG

[424] Catherine Bolley, Françoise Foucher, and Bernard Helffer. Superheating field for the Ginzburg–Landau equations in the case of a large bounded interval. Journal of Mathematical Physics, 41(11):7263–7289, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cotaescu:2000:GMD

[425] Ion I. Cotăescu and Ion I. Cotăescu, Jr. Geometric models of (d + 1)- dimensional relativistic rotating oscillators. Journal of Mathematical Physics, 41(11):7290–7293, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ekstrand:2000:SAD

[426] C. Ekstrand. A simple algebraic derivation of the covariant anomaly and Schwinger term. Journal of Mathematical Physics, 41(11):7294–7303, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geyer:2000:SFG

[427] B. Geyer and D. Mülsch. Superspace formulation of general massive gauge theories and geometric interpretation of mass-dependent Becchi– Rouet–Stora–Tyutin symmetries. Journal of Mathematical Physics, 41 (11):7304–7332, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gracia:2000:CNS

[428] Xavier Gràcia and Josep M. Pons. Canonical Noether symmetries and commutativity properties for gauge systems. Journal of Mathematical Physics, 41(11):7333–7351, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lahti:2000:CCP

[429] Pekka Lahti and Juha-Pekka Pellonpää. Characterizations of the canonical phase observable. Journal of Mathematical Physics, 41(11):7352– 7381, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levay:2000:KP

[430] Péter Lévay. On the SU(2) Kepler problem. Journal of Mathematical Physics, 41(11):7382–7385, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Musso:2000:ESQ

[431] Fabio Musso and Orlando Ragnisco. Exact solution of the quantum Calogero–Gaudin system and of its q deformation. Journal of Mathematical Physics, 41(11):7386–7401, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Echeverria-Enriquez:2000:GMH

[432] Arturo Echeverria-Enríquez, Miguel C. Muñoz-Lecanda, and Narciso Román-Roy. Geometry of multisymplectic Hamiltonian first-order field theories. Journal of Mathematical Physics, 41(11):7402–7444, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2000:KFE

[433] M. Chen. Kinetic foundation of extended irreversible thermodynamics. Journal of Mathematical Physics, 41(11):7445–7457, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Semkat:2000:KBE

[434] D. Semkat, D. Kremp, and M. Bonitz. Kadanoff–Baym equations and non-Markovian Boltzmann equation in generalized T -matrix approximation. Journal of Mathematical Physics, 41(11):7458–7467, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bering:2000:PEP

[435] K. Bering. Putting an edge to the Poisson bracket. Journal of Mathematical Physics, 41(11):7468–7500, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dai:2000:DMK

[436] H. H. Dai and Xianguo Geng. On the decomposition of the modified Kadomtsev–Petviashvili equation and explicit solutions. Journal of Mathematical Physics, 41(11):7501–7509, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Topalov:2000:FMG

[437] Petar Topalov. Families of metrics geodesically equivalent to the analogs of the Poisson sphere. Journal of Mathematical Physics, 41(11):7510– 7520, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cartas-Fuentevilla:2000:AOP

[438] R. Cartas-Fuentevilla. Adjoint operators and perturbation theory of black holes. Journal of Mathematical Physics, 41(11):7521–7528, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Friedman:2000:SPT

[439] John L. Friedman and Michael S. Morris. Schwarzschild perturbations die in time. Journal of Mathematical Physics, 41(11):7529–7534, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gariel:2000:LM

[440] J. Gariel, G. Marcilhacy, and N. O. Santos. On the Lewis metric. Journal of Mathematical Physics, 41(11):7535–7543, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:2000:MMM

[441] L. Herrera and J. L. Hernández Pastora. Measuring multipole moments of Weyl metrics by means of gyroscopes. Journal of Mathematical Physics, 41(11):7544–7555, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sen:2000:KSB

[442] Dipak K. Sen. Kerr and Schwarzschild black-holes in three-dimensional relativity. Journal of Mathematical Physics, 41(11):7556–7572, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsamparlis:2000:SBS

[443] Michael Tsamparlis and Pantelis S. Apostolopoulos. Symmetries of Bianchi I space–times. Journal of Mathematical Physics, 41(11):7573– 7588, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alisauskas:2000:TSF

[444] Sigitas Alis̆auskas. The triple sum formulas for 9j coefficients of SU(2) and u_{q}(2). Journal of Mathematical Physics, 41(11):7589–7610, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Begin:2000:GFMa

[445] L. Bégin, C. Cummins, and P. Mathieu. Generating-function method for tensor products. Journal of Mathematical Physics, 41(11):7611–7639, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Begin:2000:GFMb

[446] L. Bégin, C. Cummins, and P. Mathieu. Generating-function method for fusion rules. Journal of Mathematical Physics, 41(11):7640–7674, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blanchet:2000:HR

[447] Luc Blanchet and Guillaume Faye. Hadamard regularization. Journal of Mathematical Physics, 41(11):7675–7714, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buffenoir:2000:TPP

[448] E. Buffenoir and Ph. Roche. Tensor product of principal unitary representations of quantum Lorentz group and Askey–Wilson polynomials. Journal of Mathematical Physics, 41(11):7715–7751, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrev:2000:NGW

[449] V. K. Dobrev, A. D. Mitov, and P. Truini. Normalized U_{q}(sl (3)) Gel’fand–(Weyl)–Zetlin basis and new summation formulas for qhypergeometric functions. Journal of Mathematical Physics, 41(11): 7752–7768, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2000:IES

[450] Engui Fan. Integrable evolution systems based on Gerdjikov–Ivanov equations, bi-Hamiltonian structure, finite-dimensional integrable systems and N-fold Darboux transformation. Journal of Mathematical Physics, 41(11):7769–7782, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarizadeh:2000:GMF

[451] M. A. Jafarizadeh and S. J. Akhtarshenas. Generalized master function approach to one-dimensional quasiexactly solvable models. Journal of Mathematical Physics, 41(11):7783–7796, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kjiri:2000:GHS

[452] M. Kjiri. The G-generalized Hitchin systems and Prym varieties. Journal of Mathematical Physics, 41(11):7797–7807, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McPhedran:2000:LSG

[453] R. C. McPhedran, N. A. Nicorovici, L. C. Botten, and K. A. Grubits. Lattice sums for gratings and arrays. Journal of Mathematical Physics, 41(11):7808–7816, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pati:2000:SDI

[454] K. C. Pati and B. Das. Satake diagrams and Iwasawa decomposition of twisted Kac–Moody algebras. Journal of Mathematical Physics, 41(11): 7817–7831, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Poschadel:2000:NFS

[455] Norbert Poschadel. A note on families of spin observables and a generalization of Wigner’s theorem in C^{2}. Journal of Mathematical Physics, 41(11):7832–7838, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [554].

Ruan:2000:IRN

[456] Dong Ruan, Linhong Chen, and Wei Ruan. Indecomposable representations of the nonlinear Lie algebras. Journal of Mathematical Physics, 41(11):7839–7853, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:2000:CMS

[457] Ya Juan Sun and Meng Zhao Qin. Construction of multisymplectic schemes of any finite order for modified wave equations. Journal of Mathematical Physics, 41(11):7854–7868, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yanovski:2000:LBL

[458] A. B. Yanovski. Linear bundles of Lie algebras and their applications. Journal of Mathematical Physics, 41(11):7869–7882, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyd:2000:EPA

[459] John K. Boyd. Erratum: “Probability amplitude dynamics for a twolevel system”, [J. Math. Phys. 41, 4330 (2000)]. Journal of Mathematical Physics, 41(11):7883, November 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [241].

Almog:2000:OSS

[460] Y. Almog. The onset of superconductivity in semi-infinite strips. Journal of Mathematical Physics, 41(12):7889–7905, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Angulo:2000:IAD

[461] J. C. Angulo, E. Romera, and J. S. Dehesa. Inverse atomic densities and inequalities among density functionals. Journal of Mathematical Physics, 41(12):7906–7917, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Besch:2000:ESG

[462] Alexander Besch. Eigenvalues in spectral gaps of the two-dimensional Pauli operator. Journal of Mathematical Physics, 41(12):7918–7931, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carteret:2000:MGS

[463] H. A. Carteret, A. Higuchi, and A. Sudbery. Multipartite generalization of the Schmidt decomposition. Journal of Mathematical Physics, 41(12): 7932–7939, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassinelli:2000:GTQ

[464] G. Cassinelli, G. M. D’Ariano, E. De Vito, and A. Levrero. Group theoretical quantum tomography. Journal of Mathematical Physics, 41 (12):7940–7951, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Froman:2000:PIF

[465] Per Olof Fröman. Phase-integral formulas for quantal matrix elements. Journal of Mathematical Physics, 41(12):7952–7963, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Negro:2000:CHE

[466] J. Negro, L. M. Nieto, and O. Rosas-Ortiz. Confluent hypergeometric equations and related solvable potentials in quantum mechanics. Journal of Mathematical Physics, 41(12):7964–7996, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nieto:2000:MTC

[467] J. A. Nieto and M. C. Marín. Matroid theory and Chern–Simons. Journal of Mathematical Physics, 41(12):7997–8005, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sahbani:2000:ACS

[468] Jaouad Sahbani. On the absolutely continuous spectrum of Stark Hamiltonians. Journal of Mathematical Physics, 41(12):8006–8015, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salasnich:2000:IQG

[469] Luca Salasnich. Ideal quantum gases in D-dimensional space and powerlaw potentials. Journal of Mathematical Physics, 41(12):8016–8024, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stone:2000:SPS

[470] Michael Stone, Kee-Su Park, and Anupam Garg. The semiclassical propagator for spin coherent states. Journal of Mathematical Physics, 41(12): 8025–8049, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strauss:2000:RQM

[471] Y. Strauss, L. P. Horwitz, and E. Eisenberg. Representation of quantum mechanical resonances in the Lax–Phillips Hilbert space. Journal of Mathematical Physics, 41(12):8050–8071, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Watson:2000:GAC

[472] Glenn Watson and John R. Klauder. Generalized affine coherent states: A natural framework for the quantization of metric-like variables. Journal of Mathematical Physics, 41(12):8072–8082, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deotto:2000:NSM

[473] E. Deotto, G. Furlan, and E. Gozzi. A new superconformal mechanics. Journal of Mathematical Physics, 41(12):8083–8107, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Keane:2000:CKP

[474] Aidan J. Keane, Richard K. Barrett, and John F. L. Simmons. The classical Kepler problem and geodesic motion on spaces of constant curvature. Journal of Mathematical Physics, 41(12):8108–8116, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:2000:PAA

[475] Jen-Hsu Chang and Ming-Hsien Tu. Poisson algebras associated with constrained dispersionless modified Kadomtsev–Petviashvili hierarchies. Journal of Mathematical Physics, 41(12):8117–8131, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2000:IRS

[476] Kai Chen, Bo yu Hou, and Wen-Li Yang. Integrability of the C_{n} and BC_{n} Ruijsenaars–Schneider models. Journal of Mathematical Physics, 41(12):8132–8147, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shashikanth:2000:GPC

[477] B. N. Shashikanth and P. K. Newton. Geometric phases for corotating elliptical vortex patches. Journal of Mathematical Physics, 41(12):8148– 8162, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:2000:MWP

[478] Jaedong Choi. Multiply warped products with nonsmooth metrics. Journal of Mathematical Physics, 41(12):8163–8169, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anderson:2000:UMBb

[479] Gregory W. Anderson and Tomás Blazek. E_{6} unification model building II. Clebsch–Gordan coefficients of 78 ⊗ 78. Journal of Mathematical Physics, 41(12):8170–8189, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bromberg:2000:HST

[480] Shirley Bromberg and Alberto Medina. A homogeneous space–time model with singularities. Journal of Mathematical Physics, 41(12):8190– 8195, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2000:ASA

[481] Jin-Quan Chen, Peng-Dong Fan, Luke McAven, and Philip Butler. Algebraic solutions for all dihedral groups. Journal of Mathematical Physics, 41(12):8196–8222, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eidus:2000:PLA

[482] D. Eidus. A parametric limiting absorption problem with degeneration. Journal of Mathematical Physics, 41(12):8223–8235, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eilbeck:2000:QPS

[483] J. C. Eilbeck, V. Z. Enolskii, and N. A. Kostov. Quasiperiodic and periodic solutions for vector nonlinear Schrödinger equations. Journal of Mathematical Physics, 41(12):8236–8248, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hudson:2000:SDQ

[484] R. L. Hudson and S. Pulmannová. Symmetrized double quantum stochastic product integrals. Journal of Mathematical Physics, 41(12):8249– 8262, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jorgensen:2000:CSA

[485] Palle E. T. Jorgensen and Steen Pedersen. Commuting self-adjoint extensions of symmetric operators defined from the partial derivatives. Journal of Mathematical Physics, 41(12):8263–8278, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kwek:2000:IBV

[486] Keng-Huat Kwek, Hongjun Gao, Weinian Zhang, and Chaochun Qu. An initial boundary value problem of Camassa–Holm equation. Journal of Mathematical Physics, 41(12):8279–8285, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:2000:SCS

[487] Sen yue Lou, Xiao yan Tang, and Ji Lin. Similarity and conditional similarity reductions of a (2 + 1)-dimensional KdV equation via a direct method. Journal of Mathematical Physics, 41(12):8286–8303, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mestel:2000:GMR

[488] B. D. Mestel, A. H. Osbaldestin, and B. Winn. Golden mean renormalization for the Harper equation: The strong coupling fixed point. Journal of Mathematical Physics, 41(12):8304–8330, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scolarici:2000:CPQ

[489] G. Scolarici and L. Solombrino. Central projective quaternionic representations. Journal of Mathematical Physics, 41(12):8331–8338, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sugiyama:2000:KPM

[490] Katsuyuki Sugiyama. Kähler potential of moduli space of Calabi–Yau d-fold embedded in CP^{d} + 1. Journal of Mathematical Physics, 41(12): 8339–8348, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aronstein:2000:CEL

[491] David L. Aronstein and C. R. Stroud, Jr. Comment on “On the energy levels of a finite square-well potential”, [J. Math. Phys. 41, 4551 (2000)]. Journal of Mathematical Physics, 41(12):8349–8350, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [257].

Halilsoy:2000:CCN

[492] M. Halilsoy. Comment on “Colliding null shells of matter”, [J. Math. Phys. 33, 1065 (1992)]. Journal of Mathematical Physics, 41(12):8351– 8353, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2, 493].

Wang:2000:RCC

[493] Anzhong Wang. Response to “Comment on ‘Colliding null shells of matter’ ”, [J. Math. Phys. 41, 8351 (2000)]. Journal of Mathematical Physics, 41(12):8354–8355, December 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2, 492].

Avron:2001:FIP

[494] J. E. Avron and L. Sadun. Fredholm indices and the phase diagram of quantum Hall systems. Journal of Mathematical Physics, 42(1):1–14, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyd:2001:ODC

[495] John K. Boyd. One-dimensional crystal with a complex periodic potential. Journal of Mathematical Physics, 42(1):15–29, January 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:2001:SRP

[496] M. N. Hounkonnou and G. Y. H. Avossevou. Spectral and resonance properties of δ- and δ′-type interactions in relativistic quantum mechanics. Journal of Mathematical Physics, 42(1):30–51, January 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See note [1255].

Jasch:2001:FCR

[497] Florian Jasch and Hagen Kleinert. Fast-convergent resummation algorithm and critical exponents of ϕ^{4}-theory in three dimensions. Journal of Mathematical Physics, 42(1):52–73, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kahler:2001:MTR

[498] Ralf Kähler and Hans-Werner Wiesbrock. Modular theory and the reconstruction of four-dimensional quantum field theories. Journal of Mathematical Physics, 42(1):74–86, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

King:2001:CQC

[499] Christopher King and Mary Beth Ruskai. Capacity of quantum channels using product measurements. Journal of Mathematical Physics, 42(1): 87–98, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nagamachi:2001:HQF

[500] S. Nagamachi and E. Brüning. Hyperfunction quantum field theory: Analytic structure, modular aspects, and local observable algebras. Journal of Mathematical Physics, 42(1):99–129, January 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sako:2001:ENI

[501] A. Sako and T. Sasaki. Euler number of instanton moduli space and Seiberg–Witten invariants. Journal of Mathematical Physics, 42(1):130– 157, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanderMee:2001:ISS

[502] Cornelis van der Mee and Vjacheslav Pivovarchik. Inverse scattering for a Schrödinger equation with energy dependent potential. Journal of Mathematical Physics, 42(1):158–181, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burzlaff:2001:EDP

[503] J. Burzlaff and E. Kellegher. Expansion in the distance parameter for two vortices close together. Journal of Mathematical Physics, 42(1):182–191, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jin:2001:EEK

[504] Weimin Jin and Peter Sternberg. Energy estimates for the von Kármán model of thin-film blistering. Journal of Mathematical Physics, 42(1): 192–199, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schertzer:2001:FFP

[505] D. Schertzer, M. Larchevêque, J. Duan, V. V. Yanovsky, and S. Lovejoy. Fractional Fokker–Planck equation for nonlinear stochastic differential equations driven by non-Gaussian Lévy stable noises. Journal of Mathematical Physics, 42(1):200–212, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adams:2001:HZC

[506] Malcolm R. Adams and Sasa Kresic-Juric. Hamiltonians and zerocurvature equations for integrable partial differential equations. Journal of Mathematical Physics, 42(1):213–224, January 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaszak:2001:LAA

[507] Maciej Blaszak and Andrzej Szum. Lie algebraic approach to the construction of (2 + 1)-dimensional lattice-field and field integrable Hamiltonian equations. Journal of Mathematical Physics, 42(1):225–259, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Devchand:2001:SCH

[508] Chandrashekar Devchand and Jeremy Schiff. The supersymmetric Camassa–Holm equation and geodesic flow on the superconformal group. Journal of Mathematical Physics, 42(1):260–273, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hatayama:2001:ARC

[509] G. Hatayama, K. Hikami, R. Inoue, A. Kuniba, T. Takagi, and T. Tokihiro. The A_{M}^{(1)} automata related to crystals of symmetric tensors. Journal of Mathematical Physics, 42(1):274–308, January 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pugliese:2001:DTL

[510] F. Pugliese and A. M. Vinogradov. Discontinuous trajectories of Lagrangian systems with singular hypersurface. Journal of Mathematical Physics, 42(1):309–329, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yan:2001:HGA

[511] Zhen ya Yan and Hong qing Zhang. A hierarchy of generalized AKNS equations, N-Hamiltonian structures and finite-dimensional involutive systems and integrable systems. Journal of Mathematical Physics, 42 (1):330–339, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkelstein:2001:URC

[512] David R. Finkelstein, Andrei A. Galiautdinov, and James E. Baugh. Unimodular relativity and cosmological constant. Journal of Mathematical Physics, 42(1):340–346, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2001:PSN

[513] G. S. Hall and S. Khan. Projective symmetry in null Einstein–Maxwell space–times. Journal of Mathematical Physics, 42(1):347–354, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasinski:2001:RBT

[514] Andrzej Krasinski. Rotating Bianchi type V dust models generalizing the k = −1 Friedmann model. Journal of Mathematical Physics, 42(1): 355–367, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silva-Ortigoza:2001:MAO

[515] Gilberto Silva-Ortigoza. The method of adjoint operators and a symmetry operator for the Maxwell equations in type-D vacuum backgrounds. Journal of Mathematical Physics, 42(1):368–375, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Basarab-Horwath:2001:IVP

[516] P. Basarab-Horwath and R. Z. Zhdanov. Initial-value problems for evolutionary partial differential equations and higher-order conditional symmetries. Journal of Mathematical Physics, 42(1):376–389, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Craddock:2001:SGM

[517] M. J. Craddock and A. H. Dooley. Symmetry group methods for heat kernels. Journal of Mathematical Physics, 42(1):390–418, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deAzcarraga:2001:ODR

[518] J. A. de Azcárraga and A. J. Macfarlane. Optimally defined Racah– Casimir operators for su(n) and their eigenvalues for various classes of representations. Journal of Mathematical Physics, 42(1):419–433, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dowker:2001:SHK

[519] J. S. Dowker and Klaus Kirsten. Smeared heat-kernel coefficients on the ball and generalized cone. Journal of Mathematical Physics, 42(1):434– 452, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v42/i1/p434_s1.

Fairlie:2001:CBE

[520] D. B. Fairlie and A. N. Leznov. The complex Bateman equation in a space of arbitrary dimension. Journal of Mathematical Physics, 42(1):453– 462, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v42/i1/p453_s1.

Gao:2001:DCQ

[521] Ya-Jun Gao and Yuan-Xing Gui. Differential calculuses on the quantized braided groups. Journal of Mathematical Physics, 42(1):463–471, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v42/i1/p463_s1.

Havlicek:2001:CIF

[522] M. Havlícek and S. Posta. On the classification of irreducible finitedimensional representations of U′_{q}(so_{3}) algebra. Journal of Mathematical Physics, 42(1):472–500, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v42/i1/p472_s1.

Moan:2001:CEL

[523] P. C. Moan and J. A. Oteo. Convergence of the exponential Lie series. Journal of Mathematical Physics, 42(1):501–508, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v42/i1/p501_s1.

Montani:2001:NCQ

[524] H. Montani and R. Trinchero. Nonstandard comodules for quantum matrix bialgebras. Journal of Mathematical Physics, 42(1):509–525, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v42/i1/p509_s1.

Poletaeva:2001:SLR

[525] Elena Poletaeva. A spinor-like representation of the contact superconformal algebra K′(4). Journal of Mathematical Physics, 42(1):526–540, January 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baran:2001:SPR

[526] S. A. Baran, A. Dalgic, and G. A. Kerimov. On solvable potentials related to SO(2, 2). Journal of Mathematical Physics, 42(2):543–561, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

BezerradeMello:2001:GSE

[527] E. R. Bezerra de Mello, V. B. Bezerra, and N. R. Khusnutdinov. Ground state energy of massive scalar field inside a spherical region in the global monopole background. Journal of Mathematical Physics, 42(2):562–581, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boniver:2001:MSV

[528] F. Boniver and P. Mathonet. Maximal subalgebras of vector fields for equivariant quantizations. Journal of Mathematical Physics, 42(2):582– 589, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferapontov:2001:ISO

[529] E. V. Ferapontov and A. P. Veselov. Integrable Schrödinger operators with magnetic fields: Factorization method on curved surfaces. Journal of Mathematical Physics, 42(2):590–607, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giller:2001:BSS

[530] Stefan Giller and Piotr Milczarski. Borel summable solutions to onedimensional Schrödinger equation. Journal of Mathematical Physics, 42 (2):608–640, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manjavidze:2001:QSC

[531] J. Manjavidze and A. Sissakian. Quantization of solitons in coset space. Journal of Mathematical Physics, 42(2):641–658, February 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sheftel:2001:SSQ

[532] M. B. Sheftel, P. Tempesta, and P. Winternitz. Superintegrable systems in quantum mechanics and classical Lie theory. Journal of Mathematical Physics, 42(2):659–673, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wightman:2001:TFD

[533] A. S. Wightman. Three formulas for the off-diagonal density matrix of a Dirac spinor, with an application. Journal of Mathematical Physics, 42 (2):674–685, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Battezzati:2001:EGV

[534] M. Battezzati. An expansion of general validity for the diffusive parameters of a charged particle in a zero-point field. Journal of Mathematical Physics, 42(2):686–697, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Healy:2001:CLT

[535] Liam M. Healy. Computation of Lie transfer maps for perturbed Hamiltonian systems. Journal of Mathematical Physics, 42(2):698–712, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Margetis:2001:ECF

[536] Dionisios Margetis and Tai Tsun Wu. Exactly calculable field components of electric dipoles in planar boundary. Journal of Mathematical Physics, 42(2):713–745, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nucci:2001:HKR

[537] M. C. Nucci and P. G. L. Leach. The harmony in the Kepler and related problems. Journal of Mathematical Physics, 42(2):746–764, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gu:2001:TTI

[538] Shi-Jian Gu, You-Quan Li, and Zu-Jian Ying. Thermodynamics of trapped interacting bosons in one dimension. Journal of Mathematical Physics, 42(2):765–772, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Constantin:2001:LBI

[539] Peter Constantin, Chris Hallstrom, and Vachtang Poutkaradze. Logarithmic bounds for infinite Prandtl number rotating convection. Journal of Mathematical Physics, 42(2):773–783, February 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doering:2001:UBI

[540] Charles R. Doering and Peter Constantin. On upper bounds for infinite Prandtl number convection with or without rotation. Journal of Mathematical Physics, 42(2):784–795, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feigl:2001:ECR

[541] Kathleen Feigl and Hans Christian Öttinger. The equivalence of the class of Rivlin–Sawyers equations and a class of stochastic models for polymer stress. Journal of Mathematical Physics, 42(2):796–817, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Claudel:2001:GPS

[542] Clarissa-Marie Claudel, K. S. Virbhadra, and G. F. R. Ellis. The geometry of photon surfaces. Journal of Mathematical Physics, 42(2):818–838, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez-Jambrina:2001:HOR

[543] L. Fernández-Jambrina and C. Hoenselaers. High order relativistic corrections to Keplerian motion. Journal of Mathematical Physics, 42(2):

839–855, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golda:2001:AEU

[544] Zdzislaw A. Golda and Andrzej Woszczyna. Acoustics of early universe — Lifshitz versus gauge-invariant theories. Journal of Mathematical Physics, 42(2):856–862, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hodgkinson:2001:PTE

[545] D. E. Hodgkinson. Petrov type D Einstein space–times of embedding class two. Journal of Mathematical Physics, 42(2):863–875, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cantarella:2001:BSO

[546] Jason Cantarella, Dennis DeTurck, and Herman Gluck. The Biot–Savart operator for application to knot theory, fluid dynamics, and plasma physics. Journal of Mathematical Physics, 42(2):876–905, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:2001:FCM

[547] Shun-Jen Cheng and Ngau Lam. Finite conformal modules over N = 2, 3, 4 superconformal algebras. Journal of Mathematical Physics, 42(2): 906–933, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2001:BCQ

[548] Ya-Jun Gao and Yuan-Xing Gui. Braided coaddition on the quantized braided matrices. Journal of Mathematical Physics, 42(2):934–944, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leiva:2001:BSS

[549] Hugo Leiva. Bounded solutions of a second order evolution equation and applications. Journal of Mathematical Physics, 42(2):945–952, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morales:2001:MPR

[550] Daniel A. Morales. A mathematical proof of a result derived from the application of the Thompson–Lampard theorem of electrostatics. Journal of Mathematical Physics, 42(2):953–965, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morales:2001:GDT

[551] J. Morales, J. J. Peña, and J. L. López-Bonilla. Generalization of the Darboux transform. Journal of Mathematical Physics, 42(2):966–975, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:2001:ESP

[552] Stephen L. Adler and Lawrence P. Horwitz. Erratum: “Structure and properties of Hughston’s stochastic extension of the Schrödinger equation”, [J. Math. Phys. 41, 2485 (2000)]. Journal of Mathematical Physics, 42(2):976, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [150].

Macfarlane:2001:ECE

[553] A. J. Macfarlane and Hendryk Pfeiffer. Erratum: “On characteristic equations, trace identities and Casimir operators of simple Lie algebras”, [J. Math. Phys. 41, 3192–3225 (2000)]. Journal of Mathematical Physics, 42(2):977, February 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [198].

Poschadel:2001:ENF

[554] Norbert Poschadel. Erratum: “A note on families of spin observables and a generalization of Wigner’s theorem in C^{∈}”, [J. Math. Phys. 41, 7832 (2000)]. Journal of Mathematical Physics, 42(2):978, February 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [455].

BarroseSa:2001:USS

[555] Nuno Barros e Sá. Uncertainty for spin systems. Journal of Mathematical Physics, 42(3):981–990, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dmitrasinovic:2001:TQE

[556] V. Dmitrasinovic. Three-quark exchange operators, crossing matrices and Fierz transformations in SU(2) and SU(3). Journal of Mathematical Physics, 42(3):991–995, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1811].

Fendrik:2001:TAF

[557] A. J. Fendrik and M. J. Sánchez. Towards an analytical formula for the eigenvalues of the Aharonov–Bohm annular billiard. Journal of Mathematical Physics, 42(3):996–1005, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hertfelder:2001:CAQ

[558] C. Hertfelder and B. Kümmerer. Classical analogs of quasifree quantum stochastic processes given by stochastic states of the quantized electromagnetic field. Journal of Mathematical Physics, 42(3):1006–1017, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kotecha:2001:IYM

[559] V. Kotecha and R. S. Ward. Integrable Yang–Mills–Higgs equations in three-dimensional de Sitter space–time. Journal of Mathematical Physics, 42(3):1018–1025, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kupsch:2001:SQS

[560] Joachim Kupsch, Oleg G. Smolyanov, and Nadejda A. Sidorova. States of quantum systems and their liftings. Journal of Mathematical Physics, 42 (3):1026–1037, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyamoto:2001:GWR

[561] Manabu Miyamoto. A generalized Weyl relation approach to the time operator and its connection to the survival probability. Journal of Mathematical Physics, 42(3):1038–1052, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nouicer:2001:PIR

[562] Kh. Nouicer and L. Chetouani. Path integral for relativistic Aharonov– Bohm–Coulomb system on the pseudo-sphere. Journal of Mathematical Physics, 42(3):1053–1065, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wospakrik:2001:ASM

[563] H. J. Wospakrik and W. J. Zakrzewski. Alternative SU(N) Skyrme models and their solutions. Journal of Mathematical Physics, 42(3): 1066–1084, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2001:SYM

[564] Zixiang Zhou. Solutions of the Yang–Mills–Higgs equations in 2 + 1dimensional anti-de Sitter space–time. Journal of Mathematical Physics, 42(3):1085–1099, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daskaloyannis:2001:QPA

[565] C. Daskaloyannis. Quadratic Poisson algebras of two-dimensional classical superintegrable systems and quadratic associative algebras of quantum superintegrable systems. Journal of Mathematical Physics, 42(3): 1100–1119, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wakano:2001:AER

[566] Masami Wakano and John A. Wheeler. Asymptotics of electromagnetic radiation off a spherical superconductor. Journal of Mathematical Physics, 42(3):1120–1147, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fa:2001:EES

[567] Kwok Sau Fa and E. K. Lenzi. Exact equation of state for 2-dimensional gravitating system within Tsallis statistical mechanics. Journal of Mathematical Physics, 42(3):1148–1153, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [936].

Chen:2001:HDS

[568] Shuxing Chen. How does a shock in supersonic flow grow out of smooth data? Journal of Mathematical Physics, 42(3):1154–1172, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fatibene:2001:NCB

[569] L. Fatibene, M. Ferraris, M. Francaviglia, and M. Raiteri. Nöther charges, Brown–York quasilocal energy, and related topics. Journal of Mathematical Physics, 42(3):1173–1195, March 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Linden:2001:FFB

[570] Alexander N. Linden. Far field behavior of noncompact static spherically symmetric solutions of Einstein SU(2) Yang Mills equations. Journal of Mathematical Physics, 42(3):1196–1201, March 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melnikov:2001:SGO

[571] Yu. Melnikov and B. Pavlov. Scattering on graphs and one-dimensional approximations to N-dimensional Schrödinger operators. Journal of Mathematical Physics, 42(3):1202–1228, March 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moayedi:2001:FES

[572] S. K. Moayedi and F. Darabi. Families of exact solutions of a twodimensional gravity model minimally coupled to electrodynamics. Journal of Mathematical Physics, 42(3):1229–1235, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aneva:2001:CSN

[573] B. L. Aneva, D. Arnaudon, A. Chakrabarti, V. K. Dobrev, and S. G. Mihov. On combined standard–nonstandard or hybrid (q, h)-deformations. Journal of Mathematical Physics, 42(3):1236–1249, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bracken:2001:SPE

[574] P. Bracken and A. M. Grundland. Symmetry properties and explicit solutions of the generalized Weierstrass system. Journal of Mathematical Physics, 42(3):1250–1282, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:2001:DG

[575] Salih Çelik. Differential geometry of GL_{h}(1|1). Journal of Mathematical Physics, 42(3):1283–1295, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fagnola:2001:ESS

[576] Franco Fagnola and Rolando Rebolledo. On the existence of stationary states for quantum dynamical semigroups. Journal of Mathematical Physics, 42(3):1296–1308, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurses:2001:ROS

[577] M. Gürses and K. Zheltukhin. Recursion operators of some equations of hydrodynamic type. Journal of Mathematical Physics, 42(3):1309–1325, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iorgov:2001:LOH

[578] N. Z. Iorgov and A. U. Klimyk. The q-Laplace operator and q-harmonic polynomials on the quantum vector space. Journal of Mathematical Physics, 42(3):1326–1345, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joyce:2001:DPR

[579] William P. Joyce. Diagram projection rules for recoupling diagrams in the Racah–Wigner category. Journal of Mathematical Physics, 42(3): 1346–1363, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leandre:2001:QLG

[580] Rémi Léandre. Quotient of a loop group and Witten genus. Journal of Mathematical Physics, 42(3):1364–1383, March 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leznov:2001:GTR

[581] A. N. Leznov. To the Gel’fand–Tsetlin realization of irreducible representations of classical semisimple algebras. Journal of Mathematical Physics, 42(3):1384–1396, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martina:2001:DSI

[582] L. Martina, Kur. Myrzakul, R. Myrzakulov, and G. Soliani. Deformation of surfaces, integrable systems, and Chern–Simons theory. Journal of Mathematical Physics, 42(3):1397–1417, March 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Milewski:2001:HFA

[583] Jan Milewski and Tadeusz Lulek. Harmonic functions and angular momentum for anyon systems. Journal of Mathematical Physics, 42(3): 1418–1425, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ortiz:2001:SNH

[584] Omar E. Ortiz. Stability of nonconservative hyperbolic systems and relativistic dissipative fluids. Journal of Mathematical Physics, 42(3): 1426–1442, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takasaki:2001:PCC

[585] Kanehisa Takasaki. Painlevé–Calogero correspondence revisited. Journal of Mathematical Physics, 42(3):1443–1473, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zakhary:2001:PAC

[586] E. Zakhary and J. Carminati. On the problem of algebraic completeness for the invariants of the Riemann tensor: I. Journal of Mathematical Physics, 42(3):1474–1485, March 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baugh:2001:CAQ

[587] James Baugh, David Ritz Finkelstein, Andrei Galiautdinov, and Heinrich Saller. Clifford algebra as quantum language. Journal of Mathematical Physics, 42(4):1489–1500, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dowker:2001:MFF

[588] J. S. Dowker. Magnetic fields and factored two-spheres. Journal of Mathematical Physics, 42(4):1501–1532, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gufan:2001:GIT

[589] Yu. M. Gufan, Al. V. Popov, G. Sartori, V. Talamini, G. Valente, and E. B. Vinberg. Geometric invariant theory approach to the determination of ground states of D-wave condensates in isotropic space. Journal of Mathematical Physics, 42(4):1533–1562, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kostrykin:2001:GSP

[590] V. Kostrykin and R. Schrader. The generalized star product and the factorization of scattering matrices on graphs. Journal of Mathematical Physics, 42(4):1563–1598, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marques:2001:CCS

[591] G. C. Marques and D. Spehler. Chirality in the context of spin (3/2) particles. Journal of Mathematical Physics, 42(4):1599–1611, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garnier:2001:HFAa

[592] Josselin Garnier. High-frequency asymptotics for Maxwell’s equations in anisotropic media Part I: Linear geometric and diffractive optics. Journal of Mathematical Physics, 42(4):1612–1635, April 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garnier:2001:HFAb

[593] Josselin Garnier. High-frequency asymptotics for Maxwell’s equations in anisotropic media. Part II: Nonlinear propagation and frequency conversion. Journal of Mathematical Physics, 42(4):1636–1654, April 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lawton:2001:CIO

[594] Wayne Lawton and Lyle Noakes. Computing the inertia operator of a rigid body. Journal of Mathematical Physics, 42(4):1655–1665, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nikitin:2001:SRD

[595] A. G. Nikitin and R. J. Wiltshire. Systems of reaction diffusion equations and their symmetry properties. Journal of Mathematical Physics, 42(4): 1666–1688, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Botelho:2001:STS

[596] Luiz C. L. Botelho. Statistical turbulence suppression by a strong magnetic field: A Feynman path integral argument. Journal of Mathematical Physics, 42(4):1689–1696, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gordoa:2001:NIE

[597] Pilar R. Gordoa and Andrew Pickering. New integrable equations of fourth order and higher degree related to Cosgrove’s equation. Journal of Mathematical Physics, 42(4):1697–1707, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kozlov:2001:CLS

[598] Roman Kozlov. Conservation laws of semidiscrete Hamiltonian equations. Journal of Mathematical Physics, 42(4):1708–1727, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maciejewski:2001:NBV

[599] Andrzej J. Maciejewski, Jean-Marie Strelcyn, and Marek Szydlowski. Nonintegrability of Bianchi VIII Hamiltonian system. Journal of Mathematical Physics, 42(4):1728–1743, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsuno:2001:RDC

[600] Yoshimasa Matsuno. Reduction of dispersionless coupled Korteweg–de Vries equations to the Euler–Darboux equation. Journal of Mathematical Physics, 42(4):1744–1760, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carrero:2001:LAD

[601] J. Carrero. Lie aspects of the dynamics of self-gravitating ellipsoids in n dimensions. Journal of Mathematical Physics, 42(4):1761–1778, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chrusciel:2001:TCS

[602] Piotr T. Chrusciel and Walter Simon. Towards the classification of static vacuum spacetimes with negative cosmological constant. Journal of Mathematical Physics, 42(4):1779–1817, April 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dalal:2001:MRG

[603] Neal Dalal and Jeffrey M. Rabin. Magnification relations in gravitational lensing via multidimensional residue integrals. Journal of Mathematical Physics, 42(4):1818–1836, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goenner:2001:SSS

[604] Hubert Goenner and Peter Havas. Spherically symmetric space–times with constant curvature scalar. Journal of Mathematical Physics, 42(4): 1837–1859, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garoni:2001:CAE

[605] T. M. Garoni, N. E. Frankel, and M. L. Glasser. Complete asymptotic expansions of the Fermi–Dirac integrals F_{√}(η) = ∞/−(√ +

∞) ∫_{′}^{∞}[ϵ^{√}/(∞ + ⌉^{ϵ−η})] ⌈ϵ. Journal of Mathematical Physics, 42(4):1860– 1868, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hurley:2001:UFL

[606] D. J. Hurley and M. A. Vandyck. A unified framework for Lie and covariant differentiation (with application to tensor fields). Journal of Mathematical Physics, 42(4):1869–1886, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lesimple:2001:DMR

[607] Marc Lesimple and Georges Pinczon. Deformations of the metaplectic representations. Journal of Mathematical Physics, 42(4):1887–1899, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melnikov:2001:CSI

[608] Yu. Melnikov. On the concentration of the spectrum of integral-difference collision operator with Gaussian equilibrium distribution function in a vicinity of zero. Journal of Mathematical Physics, 42(4):1900–1906, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molnar:2001:OAH

[609] Lajos Molnár and Zsolt Páles. ⊥-order automorphisms of Hilbert space effect algebras: The two-dimensional case. Journal of Mathematical Physics, 42(4):1907–1912, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Przybysz:2001:OCE

[610] Radoslaw Przybysz. On one class of exact Poisson structures. Journal of Mathematical Physics, 42(4):1913–1920, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thorsley:2001:AEH

[611] Michael D. Thorsley and Marita C. Chidichimo. An asymptotic expansion for the hypergeometric function _{2}F_{1}(a, b; c; x). Journal of Mathematical Physics, 42(4):1921–1930, April 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagrov:2001:SRW

[612] V. G. Bagrov, D. M. Gitman, and V. B. Tlyachev. Solutions of relativistic wave equations in superpositions of Aharonov–Bohm, magnetic, and electric fields. Journal of Mathematical Physics, 42(5):1933–1959, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:2001:DSL

[613] Carl M. Bender, Stefan Boettcher, H. F. Jones, and Peter N. Meisinger. Double-scaling limit of a broken symmetry quantum field theory in dimensions D < 2. Journal of Mathematical Physics, 42(5):1960–1973, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nesterenko:2001:CES

[614] V. V. Nesterenko, G. Lambiase, and G. Scarpetta. Casimir energy of a semi-circular infinite cylinder. Journal of Mathematical Physics, 42 (5):1974–1986, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Povzner:2001:PTE

[615] A. Ya. Povzner and Yu. A. Kukharenko. Perturbation theory and the energy spectrum of normal Fermi systems. Journal of Mathematical Physics, 42(5):1987–1995, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roychoudhury:2001:CAC

[616] Rajkumar Roychoudhury, Pinaki Roy, Miloslav Znojil, and Géza Lévai. Comprehensive analysis of conditionally exactly solvable models. Journal of Mathematical Physics, 42(5):1996–2007, May 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strahov:2001:BPC

[617] E. Strahov. Berry’s phase for compact Lie groups. Journal of Mathematical Physics, 42(5):2008–2022, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanWinter:2001:CIP

[618] Clasine van Winter. Chaotic and irreversible properties of quantum scattering systems. Journal of Mathematical Physics, 42(5):2023–2054, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Viazminsky:2001:MSE

[619] C. P. Viazminsky and James P. Vary. On mathematical structure of effective observables. Journal of Mathematical Physics, 42(5):2055–2064, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benenti:2001:VSN

[620] S. Benenti, C. Chanu, and G. Rastelli. Variable separation for natural Hamiltonians with scalar and vector potentials on Riemannian manifolds. Journal of Mathematical Physics, 42(5):2065–2091, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:2001:CML

[621] Manuel de León, Eugenio Merino, and Modesto Salgado. k-cosymplectic manifolds and Lagrangian field theories. Journal of Mathematical Physics, 42(5):2092–2104, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2001:MMB

[622] Q. P. Liu. Miura map between lattice Kadomtsev–Petviashvili and its modification is canonical. Journal of Mathematical Physics, 42(5):2105– 2112, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:2001:TBD

[623] Yunbo Zeng, Wen-Xiu Ma, and Yijun Shao. Two binary Darboux transformations for the KdV hierarchy with self-consistent sources. Journal of Mathematical Physics, 42(5):2113–2128, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:2001:CFA

[624] L. Herrera, A. Di Prisco, J. Ospino, and E. Fuenmayor. Conformally flat anisotropic spheres in general relativity. Journal of Mathematical Physics, 42(5):2129–2143, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alekseev:2001:CCA

[625] Anton Alekseev, Volker Schomerus, and Thomas Strobl. Closed constraint algebras and path integrals for loop group actions. Journal of Mathematical Physics, 42(5):2144–2155, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beggs:2001:BRD

[626] E. J. Beggs and P. R. Johnson. The braiding for representations of qdeformed affine sl_{2}. Journal of Mathematical Physics, 42(5):2156–2166, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:2001:NEP

[627] Carl M. Bender and Ernst Joachim Weniger. Numerical evidence that the perturbation expansion for a non-Hermitian PT -symmetric Hamiltonian is Stieltjes. Journal of Mathematical Physics, 42(5):2167–2183, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertini:2001:WEP

[628] Massimo Bertini, Diego Noja, and Andrea Posilicano. Wave equations with point interactions in finite energy spaces. Journal of Mathematical Physics, 42(5):2184–2202, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cottrill-Shepherd:2001:FDF

[629] Kathleen Cottrill-Shepherd and Mark Naber. Fractional differential forms. Journal of Mathematical Physics, 42(5):2203–2212, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dai:2001:SRT

[630] Jian Dai and Xing-Chang Song. Structure and representation theory for the double group of the four-dimensional cubic group. Journal of Mathematical Physics, 42(5):2213–2235, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeLeo:2001:QDO

[631] Stefano De Leo and Gisele Ducati. Quaternionic differential operators. Journal of Mathematical Physics, 42(5):2236–2265, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ghanam:2001:HLA

[632] Ryad Ghanam and G. Thompson. The holonomy Lie algebras of neutral metrics in dimension four. Journal of Mathematical Physics, 42(5):2266– 2284, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lipinski:2001:YCC

[633] Z. Lipinski. Yangian construction of a centrally extended Lie algebra of a diffeomorphism group of a torus. Journal of Mathematical Physics, 42 (5):2285–2292, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miki:2001:QTA

[634] Kei Miki. Quantum toroidal algebra U_{q}(sl_{2,tor}) and R matrices. Journal of Mathematical Physics, 42(5):2293–2308, May 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Romera:2001:HEM

[635] E. Romera, J. C. Angulo, and J. S. Dehesa. The Hausdorff entropic moment problem. Journal of Mathematical Physics, 42(5):2309–2314, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1400].

Rowe:2001:ALW

[636] D. J. Rowe, H. de Guise, and B. C. Sanders. Asymptotic limits of SU(2) and SU(3) Wigner functions. Journal of Mathematical Physics, 42(5): 2315–2342, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dai:2001:EID

[637] Xianzhe Dai and Daniel S. Freed. Erratum: “η-invariants and determinant lines”, [J. Math. Phys. 35, 5155 (1994)]. Journal of Mathematical Physics, 42(5):2343–2344, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [3].

Ohlsson:2001:ETF

[638] Tommy Ohlsson and Håkan Snellman. Erratum: “Three flavor neutrino oscillations in matter”, [J. Math. Phys. 41, 2768 (2000)]. Journal of Mathematical Physics, 42(5):2345, May 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [167].

Antoine:2001:TSC

[639] J.-P. Antoine, J.-P. Gazeau, P. Monceau, J. R. Klauder, and K. A. Penson. Temporally stable coherent states for infinite well and Pöschl– Teller potentials. Journal of Mathematical Physics, 42(6):2349–2387, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arik:2001:SVM

[640] M. Arik and A. S. Arikan. Stereographic variables for multidimensional quantum group covariant q-oscillators. Journal of Mathematical Physics, 42(6):2388–2395, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Curtright:2001:GAW

[641] Thomas Curtright, Tsuneo Uematsu, and Cosmas Zachos. Generating all Wigner functions. Journal of Mathematical Physics, 42(6):2396–2415, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fakhri:2001:ESD

[642] H. Fakhri and N. Abbasi. Exact solution of the Dirac equation for a spin-(1/2) charged particle in two-dimensional and three-dimensional Euclidean spaces with shape invariance symmetry. Journal of Mathematical Physics, 42(6):2416–2437, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gruber:2001:NBT

[643] Michael J. Gruber. Noncommutative Bloch theory. Journal of Mathematical Physics, 42(6):2438–2465, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2001:CBA

[644] Guang-Liang Li, Rui-Hong Yue, and Kang-Jie Shi. Coordinate Bethe ansatz for the one-dimensional SU(n) Hubbard model with open boundary conditions. Journal of Mathematical Physics, 42(6):2466–2476, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazzucchi:2001:ODQ

[645] S. Mazzucchi. On the observables describing a quantum reference frame. Journal of Mathematical Physics, 42(6):2477–2489, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Motovilov:2001:PLS

[646] A. K. Motovilov, W. Sandhas, and V. B. Belyaev. Perturbation of a lattice spectral band by a nearby resonance. Journal of Mathematical Physics, 42(6):2490–2506, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudolph:2001:UTE

[647] Oliver Rudolph. A uniqueness theorem for entanglement measures. Journal of Mathematical Physics, 42(6):2507–2512, June 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shin:2001:ESO

[648] K. C. Shin. On the eigenproblems of PT -symmetric oscillators. Journal of Mathematical Physics, 42(6):2513–2530, June 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2001:STE

[649] M. Chen. Symmetry transformations in extended irreversible thermodynamics. Journal of Mathematical Physics, 42(6):2531–2539, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shiino:2001:FEB

[650] Masatoshi Shiino. Free energies based on generalized entropies and Htheorems for nonlinear Fokker–Planck equations. Journal of Mathematical Physics, 42(6):2540–2553, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2001:STS

[651] Shi-Rong Chen and Nian-Ning Huang. A simple treatment of the sine-Gordon equation. Journal of Mathematical Physics, 42(6):2554–2566, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dye:2001:BFO

[652] J. M. Dye and A. Parker. On bidirectional fifth-order nonlinear evolution equations, Lax pairs, and directionally dependent solitary waves. Journal of Mathematical Physics, 42(6):2567–2589, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gladush:2001:VPD

[653] Valentin D. Gladush. On the variational principle for dust shells in General Relativity. Journal of Mathematical Physics, 42(6):2590–2610, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adiguzelov:2001:TFS

[654] Ehliman Adigüzelov, Hamit Avci, and Erdal Gül. The trace formula for Sturm–Liouville operator with operator coefficient. Journal of Mathematical Physics, 42(6):2611–2624, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andrea:2001:OVE

[655] S. Andrea, A. Restuccia, and A. Sotomayor. An operator valued extension of the super Korteweg–de Vries equations. Journal of Mathematical Physics, 42(6):2625–2634, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2001:NST

[656] F. Calogero, A. Degasperis, and Ji Xiaoda. Nonlinear Schrödinger-type equations from multiscale reduction of PDEs. II. Necessary conditions of integrability for real PDEs. Journal of Mathematical Physics, 42(6): 2635–2652, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakrabarti:2001:RRM

[657] A. Chakrabarti. RTT relations, a modified braid equation and noncommutative planes. Journal of Mathematical Physics, 42(6):2653–2666, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalf:2001:ESA

[658] Hubert Kalf and Osanobu Yamada. Essential self-adjointness of ndimensional Dirac operators with a variable mass term. Journal of Mathematical Physics, 42(6):2667–2676, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2001:MTC

[659] Eunsang Kim, Hoil Kim, and Chang-Yeong Lee. Matrix theory compactification on noncommutative T^{4}/Z_{2}. Journal of Mathematical Physics, 42 (6):2677–2688, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maccari:2001:CID

[660] Attilio Maccari. A class of integrable Davey–Stewartson type systems. Journal of Mathematical Physics, 42(6):2689–2700, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maciejewski:2001:EEL

[661] Andrzej J. Maciejewski, Sasho Ivanov Popov, and Jean-Marie Strelcyn. The Euler equations on Lie algebra so(4): An elementary approach to integrability condition. Journal of Mathematical Physics, 42(6):2701– 2717, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruan:2001:BDR

[662] Dong Ruan, Yufeng Jia, and Wei Ruan. Boson and differential realizations of polynomial angular momentum algebra. Journal of Mathematical Physics, 42(6):2718–2724, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shibukawa:2001:CO

[663] Youichi Shibukawa. Classification of R-operators. Journal of Mathematical Physics, 42(6):2725–2745, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saenz:2001:EDA

[664] A. W. Sáenz. Erratum: “Dynamics of axial channeling in quasicrystals: An averaging-theory approach”, [J. Math. Phys. 41, 5342 (2000)]. Journal of Mathematical Physics, 42(6):2746, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [308].

Gubser:2001:HTN

[665] S. S. Gubser, S. Gukov, I. R. Klebanov, M. Rangamani, and E. Witten. The Hagedorn transition in noncommutative open string theory. Journal of Mathematical Physics, 42(7):2749–2764, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harvey:2001:NTK

[666] Jeffrey A. Harvey and Gregory Moore. Noncommutative tachyons and K-theory. Journal of Mathematical Physics, 42(7):2765–2780, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Myers:2001:NAD

[667] Robert Myers. Non-Abelian D-branes and noncommutative geometry. Journal of Mathematical Physics, 42(7):2781–2797, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vafa:2001:STS

[668] Cumrun Vafa. Superstrings and topological strings at large N. Journal of Mathematical Physics, 42(7):2798–2817, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Douglas:2001:BCS

[669] Michael R. Douglas. D-branes, categories and N = 1 supersymmetry. Journal of Mathematical Physics, 42(7):2818–2843, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sen:2001:FSO

[670] Ashoke Sen. Fundamental strings in open string theory at the tachyonic vacuum. Journal of Mathematical Physics, 42(7):2844–2853, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tseytlin:2001:SMA

[671] A. A. Tseytlin. Sigma model approach to string theory effective actions with tachyons. Journal of Mathematical Physics, 42(7):2854–2871, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergshoeff:2001:NAB

[672] E. A. Bergshoeff, M. de Roo, and A. Sevrin. Non-Abelian Born–Infeld and kappa-symmetry. Journal of Mathematical Physics, 42(7):2872–2888, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schwarz:2001:FST

[673] John H. Schwarz. The future of string theory. Journal of Mathematical Physics, 42(7):2889–2895, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Drukker:2001:EPS

[674] Nadav Drukker and David J. Gross. An exact prediction of N = 4 supersymmetric Yang–Mills theory for string theory. Journal of Mathematical Physics, 42(7):2896–2914, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frampton:2001:FAC

[675] Paul H. Frampton. Finite N AdS/CFT correspondence for Abelian and non-Abelian orbifolds, and gauge coupling unification. Journal of Mathematical Physics, 42(7):2915–2928, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maldacena:2001:SWMa

[676] Juan Maldacena and Hirosi Ooguri. Strings in A d S_{3} and the SL(2, R) WZW model. I: The spectrum. Journal of Mathematical Physics, 42 (7):2929–2960, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maldacena:2001:SWMb

[677] Juan Maldacena, Hirosi Ooguri, and John Son. Strings in AdS_{3} and the SL(2, R) WZW model. II: Euclidean black hole. Journal of Mathematical Physics, 42(7):2961–2977, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dolan:2001:VOA

[678] L. Dolan. Vertex operators for AdS3 with Ramond background. Journal of Mathematical Physics, 42(7):2978–2986, July 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Metsaev:2001:SSR

[679] R. R. Metsaev and A. A. Tseytlin. Superparticle and superstring in AdS_{3} × S^{3} Ramond–Ramond background in the light-cone gauge. Journal of Mathematical Physics, 42(7):2987–3014, July 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferrara:2001:RSA

[680] Sergio Ferrara and Emery Sokatchev. Representations of superconformal algebras in the AdS_{7/4}/CFT_{6/3} correspondence. Journal of Mathematical Physics, 42(7):3015–3026, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duff:2001:STI

[681] M. J. Duff, James T. Liu, and K. S. Stelle. A supersymmetric type IIB Randall–Sundrum realization. Journal of Mathematical Physics, 42 (7):3027–3047, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cvetic:2001:BWK

[682] M. Cvetic, H. Lü, and C. N. Pope. Brane-world Kaluza–Klein reductions and branes on the brane. Journal of Mathematical Physics, 42(7):3048– 3070, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kallosh:2001:ESS

[683] Renata Kallosh, Thomas Mohaupt, and Marina Shmakova. Excision of singularities by stringy domain walls. Journal of Mathematical Physics, 42(7):3071–3081, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giddings:2001:ETB

[684] Steven B. Giddings and Emanuel Katz. Effective theories and black hole production in warped compactifications. Journal of Mathematical Physics, 42(7):3082–3102, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aldazabal:2001:CSC

[685] G. Aldazabal, S. Franco, L. E. Ibáñez, R. Rabadán, and A. M. Uranga. D = 4 chiral string compactifications from intersecting branes. Journal of Mathematical Physics, 42(7):3103–3126, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gomis:2001:NCS

[686] Jaume Gomis and Hirosi Ooguri. Nonrelativistic closed string theory. Journal of Mathematical Physics, 42(7):3127–3151, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Horowitz:2001:BMT

[687] Gary T. Horowitz and Leonard Susskind. Bosonic M theory. Journal of Mathematical Physics, 42(7):3152–3160, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silverstein:2001:EOW

[688] Eva Silverstein. Extended objects from warped compactifications of M theory. Journal of Mathematical Physics, 42(7):3161–3170, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Greene:2001:DTC

[689] Brian R. Greene, Koenraad Schalm, and Gary Shiu. Dynamical topology change in M theory. Journal of Mathematical Physics, 42(7):3171–3187, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gibbons:2001:MSC

[690] G. W. Gibbons and P. C. West. The metric and strong coupling limit of the M5-brane. Journal of Mathematical Physics, 42(7):3188–3208, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atiyah:2001:TFL

[691] Michael Atiyah, Juan Maldacena, and Cumrun Vafa. An M-theory flop as a large N duality. Journal of Mathematical Physics, 42(7):3209–3220, July 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anastopoulos:2001:CTH

[692] Charis Anastopoulos. Continuous-time histories: Observables, probabilities, phase space structure and the classical limit. Journal of Mathematical Physics, 42(8):3225–3259, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beneventano:2001:EAT

[693] C. G. Beneventano and E. M. Santangelo. Effective action for QED_{4} through ζ function regularization. Journal of Mathematical Physics, 42 (8):3260–3269, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:2001:MDG

[694] Y. Brihaye, B. Hartmann, and D. H. Tchrakian. Monopoles and dyons in SO(3) gauged Skyrme models. Journal of Mathematical Physics, 42(8): 3270–3281, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See corrigendum [1223].

Chodos:2001:GGT

[695] A. Chodos and G. Gallatin. Generalized Goldstone theorem: Automatic imposition of the Higgs mechanism and application to scale and conformal symmetry breaking. Journal of Mathematical Physics, 42(8):3282– 3291, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dias:2001:FVD

[696] Óscar J. C. Dias and José P. S. Lemos. False vacuum decay: Effective one-loop action for pair creation of domain walls. Journal of Mathematical Physics, 42(8):3292–3298, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkelstein:2001:C

[697] David R. Finkelstein and Andrei A. Galiautdinov. Cliffordons. Journal of Mathematical Physics, 42(8):3299–3314, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grigoriev:2001:BRS

[698] M. A. Grigoriev, A. M. Semikhatov, and I. Yu. Tipunin. Becchi–Rouet– Stora–Tyutin formalism and zero locus reduction. Journal of Mathematical Physics, 42(8):3315–3333, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hirokawa:2001:GSS

[699] Masao Hirokawa and Osamu Ogurisu. Ground state of a spin-1/2 charged particle in a two-dimensional magnetic field. Journal of Mathematical Physics, 42(8):3334–3343, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuru:2001:IIP

[700] S. Kuru, A. Tegmen, and A. Verçin. Intertwined isospectral potentials in an arbitrary dimension. Journal of Mathematical Physics, 42(8):3344– 3360, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

March:2001:FDL

[701] N. H. March and L. M. Nieto. Functional differentiation, line integration, and departures from homogeneity of the single-particle kinetic energy functional for one-dimensional systems of N fermions. Journal of Mathematical Physics, 42(8):3361–3371, August 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:2001:VSA

[702] Ali Mostafazadeh. Variational Sturmian approximation: A nonperturbative method of solving time-independent Schrödinger equation. Journal of Mathematical Physics, 42(8):3372–3389, August 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prodan:2001:HAF

[703] E. Prodan and P. Nordlander. Hartree approximation I: The fixed point approach. Journal of Mathematical Physics, 42(8):3390–3406, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prodan:2001:HAIa

[704] E. Prodan and P. Nordlander. Hartree approximation II: The thermodynamic limit. Journal of Mathematical Physics, 42(8):3407–3423, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prodan:2001:HAIb

[705] E. Prodan and P. Nordlander. Hartree approximation III: Symmetry breaking. Journal of Mathematical Physics, 42(8):3424–3438, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Richter:2001:HAQ

[706] T. Richter and H. Schulz-Baldes. Homotopy arguments for quantized Hall conductivity. Journal of Mathematical Physics, 42(8):3439–3444, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Romao:2001:QCS

[707] Nuno M. Romão. Quantum Chern–Simons vortices on a sphere. Journal of Mathematical Physics, 42(8):3445–3469, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thorsley:2001:SCD

[708] Michael D. Thorsley and Marita C. Chidichimo. Semiclassical Coulomb differential excitation function: Asymptotic expansions. Journal of Mathematical Physics, 42(8):3470–3496, August 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garajeu:2001:SNE

[709] Daniela Garajeu and Annamaria Kiss. Singular and nonsingular eigenvectors for the Gaudin model. Journal of Mathematical Physics, 42(8): 3497–3516, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bourque:2001:PAS

[710] S. Bourque and P. Mathieu. The Painlevé analysis for N = 2 super Korteweg–de Vries equations. Journal of Mathematical Physics, 42(8): 3517–3539, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kasman:2001:SAI

[711] Alex Kasman and Michael Gekhtman. Solitons and almost-intertwining matrices. Journal of Mathematical Physics, 42(8):3540–3551, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2001:LPT

[712] Yanguang (Charles) Li. A Lax pair for the two dimensional Euler equation. Journal of Mathematical Physics, 42(8):3552–3553, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shi:2001:FDI

[713] Qi-Yan Shi. Finite-dimensional integrable systems related to the nwave interaction equations. Journal of Mathematical Physics, 42(8): 3554–3564, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bengtsson:2001:GCT

[714] Ingemar Bengtsson and Johan Brännlund. 2 + 1 gravity, chaos and time machines. Journal of Mathematical Physics, 42(8):3565–3579, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Christodoulakis:2001:TDA

[715] T. Christodoulakis, G. Kofinas, E. Korfiatis, G. O. Papadopoulos, and A. Paschos. Time-dependent automorphism inducing diffeomorphisms in vacuum Bianchi cosmologies and the complete closed form solutions for types II and V. Journal of Mathematical Physics, 42(8):3580–3608, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dancer:2001:RFW

[716] Andrew Dancer and McKenzie Y. Wang. Ricci-flat warped products and Painlevé analysis. Journal of Mathematical Physics, 42(8):3609–3614, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khelili:2001:NGT

[717] F. Khelili, J. Mimouni, and N. Mebarki. Nonsymmetric gravitation theory in noncommutative geometry. Journal of Mathematical Physics, 42 (8):3615–3627, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasinski:2001:FLR

[718] Andrzej Krasinski. Friedmann limits of rotating hypersurface– homogeneous dust models. Journal of Mathematical Physics, 42(8): 3628–3664, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2001:CHE

[719] X. F. Liu and C. P. Sun. Consequences of ’t Hooft’s equivalence class theory and symmetry by coarse graining. Journal of Mathematical Physics, 42(8):3665–3672, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anders:2001:SAN

[720] Igor Anders and Anne Boutet de Monvel. Soliton asymptotics of nondecaying solutions of the modified Kadomtsev–Petviashvili–I equation. Journal of Mathematical Physics, 42(8):3673–3690, August 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aoki:2001:ESD

[721] Takashi Aoki, Takahiro Kawai, and Yoshitsugu Takei. On the exact steepest descent method: A new method for the description of Stokes curves. Journal of Mathematical Physics, 42(8):3691–3713, August 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barco:2001:ASS

[722] M. A. Barco. An application of solvable structures to classical and nonclassical similarity solutions. Journal of Mathematical Physics, 42(8): 3714–3734, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyallian:2001:CLS

[723] Carina Boyallian and Jose I. Liberati. Classical Lie subalgebras of the Lie algebra of matrix differential operators on the circle. Journal of Mathematical Physics, 42(8):3735–3753, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chistyakov:2001:CRE

[724] G. Chistyakov, Yu. Lyubarskii, and L. Pastur. On completeness of random exponentials in the Bargmann–Fock space. Journal of Mathematical Physics, 42(8):3754–3768, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durhuus:2001:DAI

[725] Bergfinnur Durhuus and Thordur Jonsson. Discrete approximations to integrals over unparametrized paths. Journal of Mathematical Physics, 42(8):3769–3788, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geroch:2001:RLF

[726] Robert Geroch, G. Nagy, and O. Reula. Relativistic Lagrange formulation. Journal of Mathematical Physics, 42(8):3789–3808, August 2001.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grantcharov:2001:RSM

[727] Gueo Grantcharov and Liviu Ornea. Reduction of Sasakian manifolds. Journal of Mathematical Physics, 42(8):3809–3816, August 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jakobsen:2001:UIR

[728] H. P. Jakobsen and C.-W. H. Lee. Unitary irreducible representations of a Lie algebra for matrix chain models. Journal of Mathematical Physics, 42(8):3817–3838, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Patera:2001:FGT

[729] Jirí Patera, Edita Pelantová, and Milena Svobodová. Fine gradings of o(5, C), sp(4, C) and of their real forms. Journal of Mathematical Physics, 42(8):3839–3853, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sahadevan:2001:SRG

[730] R. Sahadevan and S. Khousalya. Similarity reduction, generalized symmetries and integrability of Belov–Chaltikian and Blaszak–Marciniak lattice equations. Journal of Mathematical Physics, 42(8):3854–3870, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scheunert:2001:IIC

[731] M. Scheunert and R. B. Zhang. Invariant integration on classical and quantum Lie supergroups. Journal of Mathematical Physics, 42(8):3871– 3897, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Topalov:2001:GHI

[732] Peter Topalov. Geodesic hierarchies and involutivity. Journal of Mathematical Physics, 42(8):3898–3914, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wills-Toro:2001:TSC

[733] L. A. Wills-Toro. Trefoil symmetries I. Clover extensions beyond Coleman–Mandula theorem. Journal of Mathematical Physics, 42(8):

3915–3934, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wills-Toro:2001:TSIa

[734] L. A. Wills-Toro, L. A. Sánchez, J. M. Osorio, and D. E. Jaramillo. Trefoil symmetry II. Another clover extension. Journal of Mathematical Physics, 42(8):3935–3946, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wills-Toro:2001:TSIb

[735] L. A. Wills-Toro. Trefoil symmetry III. The full clover extension. Journal of Mathematical Physics, 42(8):3947–3964, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ximin:2001:SLH

[736] Liu Ximin. Space-like hypersurfaces in the de Sitter spaces. Journal of Mathematical Physics, 42(8):3965–3972, August 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hannay:2001:CMH

[737] J. H. Hannay. Comment on “Method of handling the divergences in the radiation theory of sources that move faster than their waves”, [J. Math. Phys. 40, 4331 (1999)]. Journal of Mathematical Physics, 42(8): 3973–3974, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [11].

Krikorian:2001:CUP

[738] R. A. Krikorian. Comment on “Uncertainty principle for proper time and mass”, [J. Math. Phys 40, 1237 (1999)]. Journal of Mathematical Physics, 42(8):3975–3977, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [12, 739].

Kudaka:2001:RCU

[739] Shoju Kudaka and Shuichi Matsumoto. Response to “Comment on ‘Uncertainty principle for proper time and mass’ ”, [J. Math. Phys. 42, 3975 (2001)]. Journal of Mathematical Physics, 42(8):3978–3979, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [738].

Bergeron:2001:CQM

[740] H. Bergeron. From classical to quantum mechanics: “How to translate physical ideas into mathematical language”. Journal of Mathematical Physics, 42(9):3983–4019, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bolivar:2001:CLF

[741] A. O. Bolivar. Classical limit of fermions in phase space. Journal of Mathematical Physics, 42(9):4020–4030, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chadan:2001:ADP

[742] Khosrow Chadan, Reido Kobayashi, and Takao Kobayashi. The absolute definition of the phase-shift in potential scattering. Journal of Mathematical Physics, 42(9):4031–4049, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2001:SMQ

[743] P. Exner, M. Tater, and D. Vaněk. A single-mode quantum transport in serial-structure geometric scatterers. Journal of Mathematical Physics, 42(9):4050–4078, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Houghton:2001:FSM

[744] Conor J. Houghton and Steffen Krusch. Folding in the Skyrme model. Journal of Mathematical Physics, 42(9):4079–4100, September 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jadallah:2001:OSD

[745] Hala T. Jadallah. The onset of superconductivity in a domain with a corner. Journal of Mathematical Physics, 42(9):4101–4121, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarizadeh:2001:TTD

[746] M. A. Jafarizadeh, H. Panahi-Talemi, and E. Faizi. Two- and threedimensional Hamiltonians with shape invariance symmetry. Journal of Mathematical Physics, 42(9):4122–4137, September 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kowalski:2001:BRQ

[747] K. Kowalski and J. Rembieliński. The Bargmann representation for the quantum mechanics on a sphere. Journal of Mathematical Physics, 42 (9):4138–4147, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langmann:2001:ASQ

[748] Edwin Langmann. Algorithms to solve the (quantum) Sutherland model. Journal of Mathematical Physics, 42(9):4148–4157, September 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manjavidze:2001:YMF

[749] J. Manjavidze and A. Sissakian. Yang–Mills field quantization in the factor space. Journal of Mathematical Physics, 42(9):4158–4180, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mathur:2001:CSS

[750] Manu Mathur and Diptiman Sen. Coherent states for SU(3). Journal of Mathematical Physics, 42(9):4181–4196, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shvedov:2001:LTA

[751] O. Yu. Shvedov. Large-N theory from the axiomatic point of view. Journal of Mathematical Physics, 42(9):4197–4247, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tempesta:2001:ESS

[752] Piergiulio Tempesta, Alexander V. Turbiner, and Pavel Winternitz. Exact solvability of superintegrable systems. Journal of Mathematical Physics, 42(9):4248–4257, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Turgut:2001:GQS

[753] Osman Teoman Turgut. Geometric quantization on the super-disk. Journal of Mathematical Physics, 42(9):4258–4271, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giachetta:2001:CIO

[754] G. Giachetta, L. Mangiarotti, and G. Sardanashvily. Cohomology of the infinite-order jet space and the inverse problem. Journal of Mathematical Physics, 42(9):4272–4282, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roy:2001:RAP

[755] Tapon Roy. A robust approach to protein foldability measures based on spin-glass models. Journal of Mathematical Physics, 42(9):4283–4290, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vacchini:2001:TCM

[756] Bassano Vacchini. Translation-covariant Markovian master equation for a test particle in a quantum fluid. Journal of Mathematical Physics, 42 (9):4291–4312, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crampin:2001:CNL

[757] M. Crampin and W. Sarlet. A class of nonconservative Lagrangian systems on Riemannian manifolds. Journal of Mathematical Physics, 42 (9):4313–4326, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2001:FCI

[758] Engui Fan. A family of completely integrable multi-Hamiltonian systems explicitly related to some celebrated equations. Journal of Mathematical Physics, 42(9):4327–4344, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2001:BSC

[759] Wen-Xiu Ma and Zixiang Zhou. Binary symmetry constraints of N -wave interaction equations in 1 + 1 and 2 + 1 dimensions. Journal of Mathematical Physics, 42(9):4345–4382, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saavedra:2001:DDS

[760] Joel Saavedra, Ricardo Troncoso, and Jorge Zanelli. Degenerate dynamical systems. Journal of Mathematical Physics, 42(9):4383–4390, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blanchet:2001:LRP

[761] Luc Blanchet and Guillaume Faye. Lorentzian regularization and the problem of point-like particles in general relativity. Journal of Mathematical Physics, 42(9):4391–4418, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cadavid:2001:NSK

[762] A. C. Cadavid and R. J. Finkelstein. Nonlinear spinor in a Kerr–Schild background. Journal of Mathematical Physics, 42(9):4419–4439, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klauder:2001:NQG

[763] John R. Klauder. Noncanonical quantization of gravity. II. Constraints and the physical Hilbert space. Journal of Mathematical Physics, 42(9): 4440–4465, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berthomieu:2001:ATA

[764] A. Berthomieu. Analytic torsion of all vector bundles over an elliptic curve. Journal of Mathematical Physics, 42(9):4466–4487, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DAlessandro:2001:STC

[765] Domenico D’Alessandro. Small time controllability of systems on compact Lie groups and spin angular momentum. Journal of Mathematical Physics, 42(9):4488–4496, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fredsted:2001:ESR

[766] John Fredsted. Exponentiation of the spinor representation of the Lorentz group. Journal of Mathematical Physics, 42(9):4497–4502, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Komori:2001:RCD

[767] Yasushi Komori. Ruijsenaars’ commuting difference operators and invariant subspace spanned by theta functions. Journal of Mathematical Physics, 42(9):4503–4522, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Komori:2001:ESA

[768] Yasushi Komori. Essential self-adjointness of the elliptic Ruijsenaars models. Journal of Mathematical Physics, 42(9):4523–4553, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okubo:2001:SDK

[769] Susumu Okubo. Supergeneralization of Duffin–Kemmer–Petiau Algebra and Lie superalgebra osp(N, M). Journal of Mathematical Physics, 42 (9):4554–4562, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schork:2001:CFV

[770] Matthias Schork. On the correlation functions of the vector bundle generalization of the bc-system. Journal of Mathematical Physics, 42(9): 4563–4569, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skrypnyk:2001:QLA

[771] T. Skrypnyk. Quasigraded Lie algebras on hyperelliptic curves and classical integrable systems. Journal of Mathematical Physics, 42(9):4570– 4581, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suarez:2001:LAA

[772] Rodolfo Suárez and Leonardo Sáenz. Lie algebras associated with the exponential solutions of nonautonomous linear differential equations. Journal of Mathematical Physics, 42(9):4582–4605, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szmytkowski:2001:DDE

[773] Radosław Szmytkowski. Discontinuities in Dirac eigenfunction expansions. Journal of Mathematical Physics, 42(9):4606–4617, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jordan:2001:CEC

[774] P. M. Jordan. Comment on “Exact constructions of square-root Helmholtz operator symbols: The focusing quadratic profile”, [J. Math. Phys. 41, 4881 (2000)]. Journal of Mathematical Physics, 42(9):4618– 4623, September 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [281].

Aktosun:2001:SEA

[775] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. Small-energy asymptotics of the scattering matrix for the matrix Schrödinger equation on the line. Journal of Mathematical Physics, 42(10):4627–4652, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barker:2001:CQS

[776] Laurence Barker. Continuum quantum systems as limits of discrete quantum systems. III. Operators. Journal of Mathematical Physics, 42(10): 4653–4668, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:2001:IFD

[777] Y. Brihaye, V. Paturyan, B. M. A. G. Piette, and D. H. Tchrakian. Instantons in four-dimensional gauged O(5) Skyrme models. Journal of Mathematical Physics, 42(10):4669–4683, October 2001. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chetouani:2001:ATS

[778] L. Chetouani, L. Guechi, and T. F. Hammann. Algebraic treatment of super-integrable potentials. Journal of Mathematical Physics, 42(10): 4684–4706, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2001:ACM

[779] P. Exner and A. Joye. Avoided crossings in mesoscopic systems: Electron propagation on a nonuniform magnetic cylinder. Journal of Mathematical Physics, 42(10):4707–4738, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:2001:PTF

[780] Francisco M. Fernández. Perturbation theory free from secular terms for the equations of motion of anharmonic oscillators. Journal of Mathematical Physics, 42(10):4739–4748, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kozlov:2001:DDD

[781] G. A. Kozlov. The disorder deviation in the deconfined phase. Journal of Mathematical Physics, 42(10):4749–4756, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kulish:2001:COB

[782] P. P. Kulish and N. Manojlović. Creation operators and Bethe vectors of the osp(1|2) Gaudin model. Journal of Mathematical Physics, 42(10): 4757–4778, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langmann:2001:AST

[783] E. Langmann, J. Mickelsson, and S. Rydh. Anomalies and Schwinger terms in NCG field theory models. Journal of Mathematical Physics, 42(10):4779–4801, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leung:2001:STQ

[784] P. T. Leung, Alec Maassen van den Brink, W. M. Suen, C. W. Wong, and K. Young. SUSY transformations for quasinormal modes of open systems. Journal of Mathematical Physics, 42(10):4802–4820, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muravyev:2001:SCU

[785] Dmitry I. Muravyev. The sufficient conditions under which the S-matrix orthogonal part is real and even. Journal of Mathematical Physics, 42 (10):4821–4826, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Norris:2001:SAO

[786] L. K. Norris. n-symplectic algebra of observables in covariant Lagrangian field theory. Journal of Mathematical Physics, 42(10):4827–4845, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pavon:2001:SMF

[787] Michele Pavon. On the stochastic mechanics of the free relativistic particle. Journal of Mathematical Physics, 42(10):4846–4856, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petz:2001:SQS

[788] Dénes Petz and Milán Mosonyi. Stationary quantum source coding. Journal of Mathematical Physics, 42(10):4857–4864, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ying:2001:SHM

[789] Zu-Jian Ying, You-Quan Li, and Shi-Jian Gu. Symmetries in the Hubbard model with n-fold orbital degeneracy. Journal of Mathematical Physics, 42(10):4865–4872, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Milton:2001:PCC

[790] Graeme W. Milton. Proof of a conjecture on the conductivity of checkerboards. Journal of Mathematical Physics, 42(10):4873–4882, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://scitation.aip.org/content/aip/
journal/jmp/42/10/10.1063/1.1385564.

Kato:2001:GRP

[791] Go Kato and Miki Wadati. Graphical representation of the partition function of a one-dimensional δ-function Bose gas. Journal of Mathematical Physics, 42(10):4883–4893, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2001:LPE

[792] Kai Chen, Bo yu Hou, and Wen li Yang. The Lax pairs for elliptic C_{n} and BC_{n} Ruijsenaars–Schneider models and their spectral curves. Journal of Mathematical Physics, 42(10):4894–4914, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dutour:2001:PDA

[793] Mathieu Dutour. Phase diagram for Abrikosov lattice. Journal of Mathematical Physics, 42(10):4915–4926, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fleury:2001:CMS

[794] M. Fleury. Calogero–Moser systems with periodic and doubly periodic interaction potentials and loop algebras. Journal of Mathematical Physics, 42(10):4927–4937, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2001:RBS

[795] Zixiang Zhou. Relation between the solitons of Yang–Mills–Higgs fields in 2+1 dimensional Minkowski space–time and anti-de Sitter space–time. Journal of Mathematical Physics, 42(10):4938–4946, October 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feroze:2001:CPS

[796] Tooba Feroze, Asghar Qadir, and M. Ziad. The classification of plane symmetric spacetimes by isometries. Journal of Mathematical Physics, 42(10):4947–4955, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:2001:GEC

[797] L. Herrera and N. O. Santos. On the gravitomagnetic effects in cylindrically symmetric space–times. Journal of Mathematical Physics, 42(10): 4956–4962, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2001:RSH

[798] Hongya Liu and Paul S. Wesson. Radiating sources in higher-dimensional gravity. Journal of Mathematical Physics, 42(10):4963–4970, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gasaneo:2001:HHF

[799] G. Gasaneo, F. D. Colavecchia, S. Otranto, and C. R. Garibotti. The hyperbolic–hypergeometric functions. Journal of Mathematical Physics, 42(10):4971–4983, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hernandez-Bermejo:2001:NSJ

[800] Benito Hernández-Bermejo. New solutions of the Jacobi equations for three-dimensional Poisson structures. Journal of Mathematical Physics, 42(10):4984–4996, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2001:RRF

[801] Xufeng Liu and Changpu Sun. Representations of Ures(sl_{2}) via restricted q-Fock spaces. Journal of Mathematical Physics, 42(10):4997–5005, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lyakhovsky:2001:CFS

[802] Vladimir D. Lyakhovsky, Alexander Stolin, and Petr P. Kulish. Chains of Frobenius subalgebras of so(M) and the corresponding twists. Journal of Mathematical Physics, 42(10):5006–5019, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nishino:2001:AAM

[803] Akinori Nishino and Yasushi Komori. An algebraic approach to Macdonald–Koornwinder polynomials: Rodrigues-type formula and inner product identity. Journal of Mathematical Physics, 42(10):5020– 5046, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Athavan:2001:QTCa

[804] N. Athavan, P. O. Fröman, N. Fröman, and M. Lakshmanan. Quantal two-center Coulomb problem treated by means of the phase-integral method. I. General theory. Journal of Mathematical Physics, 42(11): 5051–5076, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Athavan:2001:QTCb

[805] N. Athavan, M. Lakshmanan, and N. Fröman. Quantal two-center Coulomb problem treated by means of the phase-integral method. II. Quantization conditions in the symmetric case expressed in terms of complete elliptic integrals. Numerical illustration. Journal of Mathematical Physics, 42(11):5077–5095, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Athavan:2001:QTCc

[806] N. Athavan, M. Lakshmanan, and N. Fröman. Quantal two-center Coulomb problem treated by means of the phase-integral method. III. Quantization conditions in the general case expressed in terms of complete elliptic integrals. Numerical illustration. Journal of Mathematical Physics, 42(11):5096–5115, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2001:MRA

[807] F. Bagarello. Multi-resolution analysis and fractional quantum Hall effect: An equivalence result. Journal of Mathematical Physics, 42(11): 5116–5129, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boya:2001:BPH

[808] Luis J. Boya, Askold M. Perelomov, and Mariano Santander. Berry phase in homogeneous Kähler manifolds with linear Hamiltonians. Journal of Mathematical Physics, 42(11):5130–5142, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassa:2001:QPS

[809] Antonio Cassa. Quantum physical systems as classical systems. Journal of Mathematical Physics, 42(11):5143–5149, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choe:2001:EMS

[810] Kwangseok Choe. Existence of multivortex solutions in the self-dual Chern–Simons–Higgs theory in a background metric. Journal of Mathematical Physics, 42(11):5150–5162, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cislo:2001:WED

[811] J. Cislo and J. Lopuszański. To what extent do the classical equations of motion determine the quantization scheme? Journal of Mathematical Physics, 42(11):5163–5176, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Florek:2001:TPP

[812] Wojciech Florek. Trions in a periodic potential. Journal of Mathematical Physics, 42(11):5177–5184, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fosco:2001:FSE

[813] C. D. Fosco and N. F. Svaiter. Finite size effects in the anisotropic (λ/4!)(ϕ_{1}^{4} + ϕ_{2}^{4})_{d} model. Journal of Mathematical Physics, 42(11):5185– 5194, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fukui:2001:CIN

[814] Takao Fukui, Sanjeev S. Seahra, and Paul S. Wesson. Cosmological implications of a nonseparable 5D solution of the vacuum Einstein field equations. Journal of Mathematical Physics, 42(11):5195–5201, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ghosh:2001:BTQ

[815] Subir Ghosh. Batalin–Tyutin quantization of the spinning particle model. Journal of Mathematical Physics, 42(11):5202–5211, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gudder:2001:SQM

[816] Stan Gudder and Gabriel Nagy. Sequential quantum measurements. Journal of Mathematical Physics, 42(11):5212–5222, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2001:SEN

[817] B. Guo and G. Yang. Some exact nontrivial global solutions with values in unit sphere for two-dimensional Landau–Lifshitz equations. Journal of Mathematical Physics, 42(11):5223–5227, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2001:EBS

[818] Richard L. Hall, Wolfgang Lucha, and Franz F. Schöberl. Energy bounds for the spinless Salpeter equation. Journal of Mathematical Physics, 42 (11):5228–5237, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langmann:2001:GYM

[819] Edwin Langmann. Generalized Yang–Mills actions from Dirac operator determinants. Journal of Mathematical Physics, 42(11):5238–5256, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pashnev:2001:CES

[820] A. Pashnev and F. Toppan. On the classification of N-extended supersymmetric quantum mechanical systems. Journal of Mathematical Physics, 42(11):5257–5271, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pfeiffer:2001:FDL

[821] Hendryk Pfeiffer. Four-dimensional lattice gauge theory with ribbon categories and the Crane–Yetter state sum. Journal of Mathematical Physics, 42(11):5272–5305, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudolph:2001:NCE

[822] Oliver Rudolph. A new class of entanglement measures. Journal of Mathematical Physics, 42(11):5306–5314, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Russell:2001:SPT

[823] Neil Russell. Superalgebras for the Penning trap. Journal of Mathematical Physics, 42(11):5315–5328, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caseiro:2001:QAA

[824] R. Caseiro, J.-P. Françoise, and R. Sasaki. Quadratic algebra associated with rational Calogero–Moser models. Journal of Mathematical Physics, 42(11):5329–5340, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lafortune:2001:SCD

[825] S. Lafortune, S. Tremblay, and P. Winternitz. Symmetry classification of diatomic molecular chains. Journal of Mathematical Physics, 42(11): 5341–5357, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samoilenko:2001:SPF

[826] Anatoliy M. Samoilenko, Anatoliy K. Prykarpatsky, and Ufuk Taneri. On the structure of Picard–Fuchs type equations for Liouville–Arnold integrable Hamiltonian system on cotangent phase spaces. Journal of Mathematical Physics, 42(11):5358–5370, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chidichimo:2001:AET

[827] Marita C. Chidichimo and Michael D. Thorsley. Asymptotic expansion of _{2}F_{1}(a, b; c; x) in terms of confluent hypergeometric functions. Journal of Mathematical Physics, 42(11):5371–5378, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Craster:2001:CCS

[828] R. V. Craster and Yu. V. Obnosov. Checkerboard composites with separated phases. Journal of Mathematical Physics, 42(11):5379–5388, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Havlicek:2001:RDA

[829] M. Havlíček, A. U. Klimyk, and S. Pos̆ta. Representations of the qdeformed algebra U′(so_{4}). Journal of Mathematical Physics, 42(11):5389– 5416, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lievens:2001:TFD

[830] S. Lievens and J. Van der Jeugt. Transformation formulas for double hypergeometric series related to 9-j coefficients and their basic analogs. Journal of Mathematical Physics, 42(11):5417–5430, November 2001.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parashar:2001:BQD

[831] Deepak Parashar. On the biparametric quantum deformation of GL(2)⊗ GL(1). Journal of Mathematical Physics, 42(11):5431–5443, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Su:2001:WPV

[832] Yucai Su. Weakly primitive vectors of Kac-modules of the Lie superalgebras sl(m/n). Journal of Mathematical Physics, 42(11):5444–5456, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tomoyuki:2001:SSX

[833] Sadakane Tomoyuki. Soliton solutions of XXZ lattice Landau–Lifshitz equation. Journal of Mathematical Physics, 42(11):5457–5471, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zenchuk:2001:FPS

[834] A. I. Zenchuk. Families of particular solutions to multidimensional partial differential equations. Journal of Mathematical Physics, 42(11):5472– 5492, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivashchuk:2001:EHK

[835] V. D. Ivashchuk, S.-W. Kim, and V. N. Melnikov. Erratum: “Hyperbolic Kac–Moody algebra from intersecting p-branes”, [J. Math. Phys. 40, 4072 (1999)]. Journal of Mathematical Physics, 42(11):5493, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [10].

Actor:2001:SQF

[836] A. A. Actor and I. Bender. Scalar quantum field coupled to boundaries and to a background magnetic field. Journal of Mathematical Physics, 42(12):5499–5521, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adams:2001:CLF

[837] David H. Adams. On the continuum limit of fermionic topological charge in lattice gauge theory. Journal of Mathematical Physics, 42(12):5522– 5533, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aspinwall:2001:SNR

[838] Paul S. Aspinwall. Some navigation rules for D-brane monodromy. Journal of Mathematical Physics, 42(12):5534–5552, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bashir:2001:MPS

[839] A. Bashir, R. Delbourgo, and M. L. Roberts. Multidimensional phase space and sunset diagrams. Journal of Mathematical Physics, 42(12): 5553–5564, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dias:2001:GWW

[840] Nuno Costa Dias and João Nuno Prata. Generalized Weyl–Wigner map and Vey quantum mechanics. Journal of Mathematical Physics, 42(12): 5565–5579, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Floratos:2001:SSS

[841] I. Floratos and B. M. A. G. Piette. Spherically symmetric solutions of the sixth order SU(N) Skyrme models. Journal of Mathematical Physics, 42(12):5580–5595, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hainzl:2001:GCS

[842] Christian Hainzl. Gradient corrections for semiclassical theories of atoms in strong magnetic fields. Journal of Mathematical Physics, 42(12):5596– 5625, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hupfer:2001:UBD

[843] Thomas Hupfer, Hajo Leschke, and Simone Warzel. Upper bounds on the density of states of single Landau levels broadened by Gaussian random potentials. Journal of Mathematical Physics, 42(12):5626–5641, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karazija:2001:IRE

[844] R. Karazija and V. Jonauskas. Investigation of the relativistic equivalent Hamiltonian in the LS coupling scheme. Journal of Mathematical Physics, 42(12):5642–5651, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:2001:SER

[845] Avinash Khare and Uday Sukhatme. Some exact results for mid-band and zero band-gap states of associated Lamé potentials. Journal of Mathematical Physics, 42(12):5652–5664, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Osadchy:2001:HBQ

[846] D. Osadchy and J. E. Avron. Hofstadter butterfly as quantum phase diagram. Journal of Mathematical Physics, 42(12):5665–5671, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shibata:2001:CSP

[847] Junya Shibata and Komajiro Niizeki. Coherent state path integral for the Bloch particle. Journal of Mathematical Physics, 42(12):5672–5686, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsutsui:2001:MSS

[848] Izumi Tsutsui, Tamás Fülöp, and Taksu Cheon. Möbius structure of the spectral space of Schrödinger operators with point interaction. Journal of Mathematical Physics, 42(12):5687–5697, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nam:2001:ACM

[849] Hee-Seok Nam. Asymptotics for the condensate multivortex solutions in the self-dual Chern–Simons CP(1) model. Journal of Mathematical Physics, 42(12):5698–5712, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blencowe:2001:PAO

[850] Miles P. Blencowe and Nicholas C. Koshnick. Partition asymptotics from one-dimensional quantum entropy and energy currents. Journal of Mathematical Physics, 42(12):5713–5717, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feinberg:2001:SRT

[851] Joshua Feinberg, R. Scalettar, and A. Zee. “single ring theorem” and the disk-annulus phase transition. Journal of Mathematical Physics, 42(12):

5718–5740, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Campoamor-Stursberg:2001:DSE

[852] O. R. Campoamor-Stursberg, F. G. Gascon, and D. Peralta-Salas. Dynamical systems embedded into Lie algebras. Journal of Mathematical Physics, 42(12):5741–5752, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kapadia:2001:NRS

[853] Devendra A. Kapadia. Nonintegrable reductions of the self-dual Yang– Mills equations in a metric of plane wave type. Journal of Mathematical Physics, 42(12):5753–5761, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tongas:2001:FIN

[854] A. Tongas, D. Tsoubelis, and P. Xenitidis. A family of integrable nonlinear equations of hyperbolic type. Journal of Mathematical Physics, 42(12):5762–5784, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia-Perciante:2001:KTG

[855] Ana Laura García-Perciante, Alfredo Sandoval-Villalbazo, and L. S. García-Colín. Kaluza’s theory in generalized coordinates. Journal of Mathematical Physics, 42(12):5785–5799, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tomaschitz:2001:EPH

[856] Roman Tomaschitz. Electromagnetic propagators in hyperbolic Robertson–Walker cosmologies. Journal of Mathematical Physics, 42 (12):5800–5831, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blumel:2001:CIB

[857] R. Blümel and Yu. Dabaghian. Combinatorial identities for binary necklaces from exact ray-splitting trace formulas. Journal of Mathematical Physics, 42(12):5832–5839, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Devchand:2001:DS

[858] Chandrashekar Devchand and Jean Nuyts. Democratic supersymmetry. Journal of Mathematical Physics, 42(12):5840–5858, December 2001.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Enoch:2001:SSW

[859] S. Enoch, R. C. McPhedran, N. A. Nicorovici, L. C. Botten, and J. N. Nixon. Sums of spherical waves for lattices, layers, and lines. Journal of Mathematical Physics, 42(12):5859–5870, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerin:2001:NGF

[860] Charles-Antoine Guérin. A note on the generalized fractal dimensions of a probability measure. Journal of Mathematical Physics, 42(12):5871– 5875, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenczewski:2001:FRV

[861] Romuald Lenczewski. Filtered random variables, bialgebras, and convolutions. Journal of Mathematical Physics, 42(12):5876–5903, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molnar:2001:OAS

[862] Lajos Molnár. Order-automorphisms of the set of bounded observables. Journal of Mathematical Physics, 42(12):5904–5909, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nagel:2001:EHF

[863] Bengt Nagel. An expansion of the hypergeometric function in Bessel functions. Journal of Mathematical Physics, 42(12):5910–5914, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shlivinski:2001:TDR

[864] Amir Shlivinski, Ehud Heyman, and Anthony J. Devaney. Time domain radiation by scalar sources: Plane wave to multipole transform. Journal of Mathematical Physics, 42(12):5915–5919, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abdallah:2002:COD

[865] N. Ben Abdallah, P. Degond, and I. M. Gamba. Coupling onedimensional time-dependent classical and quantum transport models.

Journal of Mathematical Physics, 43(1):1–24, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bliokh:2002:GAA

[866] K. Yu. Bliokh. Geometric amplitude, adiabatic invariants, quantization, and strong stability of Hamiltonian systems. Journal of Mathematical Physics, 43(1):25–42, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chiang:2002:PDE

[867] Chun-Ming Chiang and Choon-Lin Ho. Planar Dirac electron in Coulomb and magnetic fields: A Bethe ansatz approach. Journal of Mathematical Physics, 43(1):43–51, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Felbacq:2002:NLH

[868] D. Felbacq. Noncommuting limits in homogenization theory of electromagnetic crystals. Journal of Mathematical Physics, 43(1):52–55, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giachetta:2002:CGQ

[869] G. Giachetta, L. Mangiarotti, and G. Sardanashvily. Covariant geometric quantization of nonrelativistic time-dependent mechanics. Journal of Mathematical Physics, 43(1):56–68, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gill:2002:FRQ

[870] Tepper L. Gill and W. W. Zachary. Foundations for relativistic quantum theory. I. Feynman’s operator calculus and the Dyson conjectures. Journal of Mathematical Physics, 43(1):69–93, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2002:SHO

[871] Richard L. Hall, Nasser Saad, and Attila B. von Keviczky. Spiked harmonic oscillators. Journal of Mathematical Physics, 43(1):94–112, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hatzinikitas:2002:NHO

[872] Agapitos Hatzinikitas and Ioannis Smyrnakis. The noncommutative harmonic oscillator in more than one dimension. Journal of Mathematical Physics, 43(1):113–125, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivanov:2002:STR

[873] D. A. Ivanov. The supersymmetric technique for random-matrix ensembles with zero eigenvalues. Journal of Mathematical Physics, 43(1): 126–153, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ling:2002:SSN

[874] Yi Ling. Supersymmetric spin networks. Journal of Mathematical Physics, 43(1):154–169, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2002:FAM

[875] Q. H. Liu, X. Wang, W. H. Qi, L. P. Fu, and B. Hu. The Fejér average and the mean value of a quantity in a quasiclassical wave packet. Journal of Mathematical Physics, 43(1):170–181, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1066].

Martinetti:2002:DKK

[876] Pierre Martinetti and Raimar Wulkenhaar. Discrete Kaluza–Klein from scalar fluctuations in noncommutative geometry. Journal of Mathematical Physics, 43(1):182–204, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:2002:PHVa

[877] Ali Mostafazadeh. Pseudo-Hermiticity versus P T symmetry: The necessary condition for the reality of the spectrum of a non-Hermitian Hamiltonian. Journal of Mathematical Physics, 43(1):205–214, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Popov:2002:ABS

[878] I. Yu. Popov. Asymptotics of bound states and bands for laterally coupled waveguides and layers. Journal of Mathematical Physics, 43(1): 215–234, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakaji:2002:RLG

[879] A. Sakaji, R. S. Hijjawi, N. Shawagfeh, and J. M. Khalifeh. Remarks on the lattice Green’s function: The Glasser case. Journal of Mathematical Physics, 43(1):235–242, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmidt:2002:IIR

[880] Andreas U. Schmidt. Infinite infrared regularization and a state space for the Heisenberg algebra. Journal of Mathematical Physics, 43(1):243–259, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1067].

Teleman:2002:LIL

[881] Kostake Teleman. Lorentz invariant Lagrangians. Journal of Mathematical Physics, 43(1):260–282, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaisman:2002:FQS

[882] Izu Vaisman. Fedosov quantization on symplectic ringed spaces. Journal of Mathematical Physics, 43(1):283–298, January 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Azizi:2002:PTM

[883] Azizollah Azizi. Planar trajectories in a monopole field. Journal of Mathematical Physics, 43(1):299–317, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamath:2002:PRL

[884] S. G. Kamath. A planar Runge–Lenz vector. Journal of Mathematical Physics, 43(1):318–324, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vignolo:2002:IIK

[885] Stefano Vignolo and Danilo Bruno. Iper-ideal kinetic constraints in continuum mechanics. Journal of Mathematical Physics, 43(1):325–343, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frank:2002:GLB

[886] T. D. Frank. On a general link between anomalous diffusion and nonextensivity. Journal of Mathematical Physics, 43(1):344–350, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Requardt:2002:FOS

[887] Manfred Requardt. Fluctuation operators and spontaneous symmetry breaking. Journal of Mathematical Physics, 43(1):351–372, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alba:2002:DBF

[888] David Alba, Luca Lusanna, and Massimo Pauri. Dynamical body frames, orientation-shape variables and canonical spin bases for the nonrelativistic N-body problem. Journal of Mathematical Physics, 43(1):373–402, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avan:2002:SBR

[889] J. Avan and G. Rollet. Structures in BC_{N} Ruijsenaars–Schneider models. Journal of Mathematical Physics, 43(1):403–416, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cardin:2002:GFS

[890] Franco Cardin. The global finite structure of generic envelope loci for Hamilton–Jacobi equations. Journal of Mathematical Physics, 43(1):417– 430, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:2002:HOT

[891] Manuel F. Rañada and Mariano Santander. On harmonic oscillators on the two-dimensional sphere S^{2} and the hyperbolic plane H^{2}. Journal of Mathematical Physics, 43(1):431–451, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shengfan:2002:ASO

[892] Zhou Shengfan. Attractors for second-order lattice dynamical systems with damping. Journal of Mathematical Physics, 43(1):452–465, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2002:DMI

[893] Yufeng Zhang and Hongqing Zhang. A direct method for integrable couplings of TD hierarchy. Journal of Mathematical Physics, 43(1):466– 472, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Accioly:2002:PTL

[894] Antonio Accioly, Abel Azeredo, and Hatsumi Mukai. Propagator, treelevel unitarity and effective nonrelativistic potential for higher-derivative gravity theories in D dimensions. Journal of Mathematical Physics, 43 (1):473–491, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carminati:2002:PACa

[895] J. Carminati, E. Zakhary, and R. G. McLenaghan. On the problem of algebraic completeness for the invariants of the Riemann tensor. II. Journal of Mathematical Physics, 43(1):492–507, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castellani:2002:SNC

[896] O. C. Castellani and H. P. de Oliveira. On solitons with nonminimally coupled scalar fields. Journal of Mathematical Physics, 43(1):508–518, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delgado:2002:SDS

[897] J. Estevez Delgado and T. Zannias. On the structure of degenerate solutions of the Einstein conformally invariant scalar system. Journal of Mathematical Physics, 43(1):519–553, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Etesi:2002:RTN

[898] Gábor Etesi. A rigidity theorem for nonvacuum initial data. Journal of Mathematical Physics, 43(1):554–562, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giannoni:2002:FPG

[899] F. Giannoni, A. Masiello, and P. Piccione. The Fermat principle in general relativity and applications. Journal of Mathematical Physics, 43 (1):563–596, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hosotani:2002:SBE

[900] Yutaka Hosotani. Scaling behavior in the Einstein–Yang–Mills monopoles and dyons. Journal of Mathematical Physics, 43(1):597–603, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Balantekin:2002:CEO

[901] A. B. Balantekin and P. Cassak. Character expansions for the orthogonal and symplectic groups. Journal of Mathematical Physics, 43(1):604–620, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:2002:AGS

[902] Cewen Cao, Xianguo Geng, and Hongye Wang. Algebro-geometric solution of the 2 + 1 dimensional Burgers equation with a discrete variable. Journal of Mathematical Physics, 43(1):621–643, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cartas-Fuentevilla:2002:SSG

[903] R. Cartas-Fuentevilla. On the symplectic structures for geometrical theories. Journal of Mathematical Physics, 43(1):644–650, January 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunajski:2002:HCF

[904] Maciej Dunajski. Hyper-complex four-manifolds from the Tzitzéica equation. Journal of Mathematical Physics, 43(1):651–658, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Edgar:2002:DDT

[905] S. Brian Edgar and A. Höglund. Dimensionally dependent tensor identities by double antisymmetrization. Journal of Mathematical Physics, 43 (1):659–677, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isenberg:2002:SFW

[906] James Isenberg and Steven L. Liebling. Singularity formation in 2 + 1 wave maps. Journal of Mathematical Physics, 43(1):678–683, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kirkpatrick:2002:UCS

[907] K. A. Kirkpatrick. Uniqueness of a convex sum of products of projectors. Journal of Mathematical Physics, 43(1):684–686, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Varsaie:2002:NCR

[908] S. Varsaie. New cohomology ring for supermanifolds. Journal of Mathematical Physics, 43(1):687–691, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassinelli:2002:CLT

[909] Gianni Cassinelli, Ernsesto De Vito, Pekka Lahti, and Juha-Pekka Pellonpää. Covariant localizations in the torus and the phase observables. Journal of Mathematical Physics, 43(2):693–704, February 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castagnino:2002:GTA

[910] Mario A. Castagnino, Jorge Gueron, and Adolfo R. Ordóñez. Global time asymmetry as a consequence of a wave packets theorem. Journal of Mathematical Physics, 43(2):705–713, February 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

ElKinani:2002:GCI

[911] A. H. El Kinani and M. Daoud. Generalized coherent and intelligent states for exact solvable quantum systems. Journal of Mathematical Physics, 43(2):714–733, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2002:SPC

[912] Kei Inoue, Masanori Ohya, and Igor V. Volovich. Semiclassical properties and chaos degree for the quantum Baker’s map. Journal of Mathematical Physics, 43(2):734–755, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karasev:2002:SAQ

[913] M. V. Karasev and T. A. Osborn. Symplectic areas, quantization, and dynamics in electromagnetic fields. Journal of Mathematical Physics, 43 (2):756–788, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khrennikov:2002:EFO

[914] Andrei Khrennikov. Ensemble fluctuations and the origin of quantum probabilistic rule. Journal of Mathematical Physics, 43(2):789–802, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2002:GMG

[915] Kwang Ik Kim and Zuhan Liu. Global minimizer for the Ginzburg– Landau functional of an inhomogeneous superconductor. Journal of Mathematical Physics, 43(2):803–817, February 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kopf:2002:SQS

[916] Tomás̆ Kopf and Mario Paschke. A spectral quadruple for de Sitter space. Journal of Mathematical Physics, 43(2):818–846, February 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lipan:2002:BQE

[917] Ovidiu Lipan and Constantin Rasinariu. Baxter T–Q equation for shape invariant potentials. The finite-gap potentials case. Journal of Mathematical Physics, 43(2):847–865, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Retamal:2002:RDN

[918] Juan C. Retamal and Lautaro Vergara. Regularizing divergences in the von Neumann entropy. Journal of Mathematical Physics, 43(2):866–871, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sako:2002:NCF

[919] Akifumi Sako, Shin-Ichiro Kuroki, and Tomomi Ishikawa. Noncommutative cohomological field theory and GMS soliton. Journal of Mathematical Physics, 43(2):872–896, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:2002:COD

[920] Bei Zeng and Jin-Yan Zeng. Closure of orbits and dynamical symmetry of screened Coulomb potential and isotropic harmonic oscillator. Journal of Mathematical Physics, 43(2):897–903, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meson:2002:KLG

[921] Alejandro M. Mesón and Fernando Vericat. On the Kolmogorov-like generalization of Tsallis entropy, correlation entropies and multifractal analysis. Journal of Mathematical Physics, 43(2):904–917, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palmer:2002:SDA

[922] John Palmer. Short distance asymptotics of Ising correlations. Journal of Mathematical Physics, 43(2):918–953, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ballesteros:2002:CHS

[923] Ángel Ballesteros and Orlando Ragnisco. Classical Hamiltonian systems with sl(2) coalgebra symmetry and their integrable deformations. Journal of Mathematical Physics, 43(2):954–969, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:2002:STD

[924] E. G. Kalnins, J. M. Kress, and P. Winternitz. Superintegrability in a two-dimensional space of nonconstant curvature. Journal of Mathematical Physics, 43(2):970–983, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsuno:2002:DSG

[925] Yoshimasa Matsuno. Dark soliton generation for the intermediate nonlinear Schrödinger equation. Journal of Mathematical Physics, 43(2): 984–1007, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tam:2002:RSD

[926] Hon-Wah Tam, Xing-Biao Hu, and Xian-Min Qian. Remarks on several 2 + 1 dimensional lattices. Journal of Mathematical Physics, 43(2):1008– 1017, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

BezerradeMello:2002:ESG

[927] E. R. Bezerra de Mello. Euclidean scalar Green function in a higher dimensional global monopole space–time. Journal of Mathematical Physics, 43(2):1018–1028, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivanov:2002:ICG

[928] B. V. Ivanov. Integrable cases of gravitating static isothermal fluid spheres. Journal of Mathematical Physics, 43(2):1029–1043, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boiti:2002:IST

[929] M. Boiti, F. Pempinelli, A. K. Pogrebkov, and B. Prinari. Inverse scattering theory of the heat equation for a perturbed one-soliton potential. Journal of Mathematical Physics, 43(2):1044–1062, February 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dubin:2002:AOW

[930] D. A. Dubin and M. A. Hennings. Angle operators on weighted Bergman spaces. Journal of Mathematical Physics, 43(2):1063–1073, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fagnola:2002:SPQ

[931] Franco Fagnola and Rolando Rebolledo. Subharmonic projections for a quantum Markov semigroup. Journal of Mathematical Physics, 43(2): 1074–1082, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Havlicek:2002:AFG

[932] Miloslav Havlíček, Jiří Patera, Edita Pelantová, and Jiří Tolar. Automorphisms of the fine grading of sl(n, C) associated with the generalized Pauli matrices. Journal of Mathematical Physics, 43(2):1083–1094, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leble:2002:RRD

[933] S. B. Leble and A. V. Yurov. Reduction restrictions of Darboux and Laplace transformations for the Goursat equation. Journal of Mathematical Physics, 43(2):1095–1105, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2002:APH

[934] Yezhou Li and Yuzan He. On analytic properties of higher analogs of the second Painlevé equation. Journal of Mathematical Physics, 43(2): 1106–1115, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2002:SIS

[935] Xianwen Zhang. L^{p} spectral independence of the streaming operator. Journal of Mathematical Physics, 43(2):1116–1126, February 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Fa:2002:EEE

[936] Kwok Sau Fa and E. K. Lenzi. Erratum: “Exact equation of state for 2dimensional gravitating system within Tsallis statistical mechanics”, [J. Math. Phys. 42, 1148–1153 (2001)]. Journal of Mathematical Physics, 43(2):1127, February 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [567].

Alhaidari:2002:RMT

[937] A. D. Alhaidari. The relativistic J-matrix theory of scattering: an analytic solution. Journal of Mathematical Physics, 43(3):1129–1135, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Azizi:2002:SOH

[938] Azizollah Azizi. Stable orbits in a Higgs-Monopole field. Journal of Mathematical Physics, 43(3):1136–1151, March 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2002:QML

[939] P. Exner and K. Nĕmcová. Quantum mechanics of layers with a finite number of point perturbations. Journal of Mathematical Physics, 43(3): 1152–1184, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hainzl:2002:SLL

[940] Christian Hainzl. Semiclassics in the lowest Landau band. Journal of Mathematical Physics, 43(3):1185–1210, March 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2002:CSS

[941] Brian C. Hall and Jeffrey J. Mitchell. Coherent states on spheres. Journal of Mathematical Physics, 43(3):1211–1236, March 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1948] and comment [1954].

Hall:2002:RBS

[942] Richard L. Hall, Wolfgang Lucha, and Franz F. Schöberl. Relativistic N-boson systems bound by oscillator pair potentials. Journal of Mathematical Physics, 43(3):1237–1246, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1421].

King:2002:MCN

[943] Christopher King. Maximization of capacity and l_{p} norms for some product channels. Journal of Mathematical Physics, 43(3):1247–1260, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2002:PPN

[944] Zuhan Liu. Pinning phenomena near the lower critical field for superconductor. Journal of Mathematical Physics, 43(3):1261–1272, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nenciu:2002:APT

[945] G. Nenciu. On asymptotic perturbation theory for quantum mechanics: Almost invariant subspaces and gauge invariant magnetic perturbation theory. Journal of Mathematical Physics, 43(3):1273–1298, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pellonpaa:2002:SCP

[946] Juha-Pekka Pellonpää. On the structure of covariant phase observables. Journal of Mathematical Physics, 43(3):1299–1308, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rodriguez:2002:QSE

[947] Miguel A. Rodríguez and Pavel Winternitz. Quantum superintegrability and exact solvability in n dimensions. Journal of Mathematical Physics, 43(3):1309–1322, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sen:2002:FBQ

[948] R. N. Sen and G. L. Sewell. Fiber bundles in quantum physics. Journal of Mathematical Physics, 43(3):1323–1339, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Toprak:2002:LLS

[949] Erdal Toprak and O. Teoman Turgut. Large N limit of SO(N) scalar gauge theory. Journal of Mathematical Physics, 43(3):1340–1352, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vargas:2002:QCA

[950] Jose G. Vargas and Douglas G. Torr. Quantum Clifford algebra from classical differential geometry. Journal of Mathematical Physics, 43(3): 1353–1364, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhalij:2002:SPE

[951] Alexander Zhalij. On separable Pauli equations. Journal of Mathematical Physics, 43(3):1365–1389, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rojas:2002:HTE

[952] Onofre Rojas, S. M. de Souza, and M. T. Thomaz. High temperature expansion for a chain model. Journal of Mathematical Physics, 43(3): 1390–1407, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:2002:BHF

[953] Wen-Xiu Ma. A bi-Hamiltonian formulation for triangular systems by perturbations. Journal of Mathematical Physics, 43(3):1408–1421, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McLenaghan:2002:GIC

[954] Raymond G. McLenaghan, Roman G. Smirnov, and Dennis The. Group invariant classification of separable Hamiltonian systems in the Euclidean plane and the O(4)-symmetric Yang–Mills theories of Yatsun. Journal of Mathematical Physics, 43(3):1422–1440, March 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nutku:2002:MLI

[955] Y. Nutku and M. V. Pavlov. Multi-Lagrangians for integrable systems. Journal of Mathematical Physics, 43(3):1441–1459, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shchesnovich:2002:PTN

[956] V. S. Shchesnovich. Perturbation theory for nearly integrable multicomponent nonlinear PDEs. Journal of Mathematical Physics, 43(3): 1460–1486, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chauvineau:2002:EFR

[957] Bertrand Chauvineau. Euclidean Friedman–Robertson–Walker cosmology in Brans–Dicke-like theories of gravity. Journal of Mathematical Physics, 43(3):1487–1492, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1096].

Heinzle:2002:RDS

[958] J. Mark Heinzle and Roland Steinbauer. Remarks on the distributional Schwarzschild geometry. Journal of Mathematical Physics, 43(3):1493– 1508, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krisch:2002:ATA

[959] J. P. Krisch and E. N. Glass. Adding twist to anisotropic fluids. Journal of Mathematical Physics, 43(3):1509–1518, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nozari:2002:MGM

[960] Kourosh Nozari and Reza Mansouri. The methods of gluing manifolds in general relativity. Journal of Mathematical Physics, 43(3):1519–1535, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pravdova:2002:BRS

[961] A. Pravdová and V. Pravda. Boost–rotation symmetric vacuum space– times with spinning sources. Journal of Mathematical Physics, 43(3): 1536–1546, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alisauskas:2002:MSF

[962] Sigitas Alis̆auskas. The multiple sum formulas for 12-j coefficients of SU(2) and u_{q}(2). Journal of Mathematical Physics, 43(3):1547–1568, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakrabarti:2002:MBE

[963] A. Chakrabarti. Modified braid equations for SO_{q}(3) and noncommutative spaces. Journal of Mathematical Physics, 43(3):1569–1583, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forni:2002:USC

[964] D. M. Forni, M. S. Iriondo, C. N. Kozameh, and M. F. Parisi. Understanding singularities in Cartan’s and null surface formulation geometric structures. Journal of Mathematical Physics, 43(3):1584–1597, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Legault:2002:SSO

[965] Stéphane R. Legault and Thomas B. A. Senior. Solution of a second order difference equation using the bilinear relations of Riemann. Journal of Mathematical Physics, 43(3):1598–1621, March 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Llibre:2002:IAS

[966] Jaume Llibre and Xiang Zhang. Invariant algebraic surfaces of the Lorenz system. Journal of Mathematical Physics, 43(3):1622–1645, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palev:2002:JGQ

[967] T. D. Palev, N. I. Stoilova, and J. Van der Jeugt. Jacobson generators of the quantum superalgebra U_{q}[sl(n + 1|m)] and Fock representations. Journal of Mathematical Physics, 43(3):1646–1663, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiper:2002:ACM

[968] E. V. Tsiper. Analytic Coulomb matrix elements in the lowest Landau level in disk geometry. Journal of Mathematical Physics, 43(3):1664– 1667, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yildiz:2002:BO

[969] A. Yildiz. Braided oscillators. Journal of Mathematical Physics, 43(3): 1668–1674, March 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alba:2002:CMR

[970] David Alba, Luca Lusanna, and Massimo Pauri. Centers of mass and rotational kinematics for the relativistic N-body problem in the restframe instant form. Journal of Mathematical Physics, 43(4):1677–1727, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernido:2002:PIB

[971] Christopher C. Bernido and M. Victoria Carpio-Bernido. Path integrals for boundaries and topological constraints: A white noise functional approach. Journal of Mathematical Physics, 43(4):1728–1736, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freidel:2002:FSP

[972] Laurent Freidel and Kirill Krasnov. The fuzzy sphere ⋆-product and spin networks. Journal of Mathematical Physics, 43(4):1737–1754, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiroshima:2002:OEP

[973] Fumio Hiroshima. Observable effects and parametrized scaling limits of a model in nonrelativistic quantum electrodynamics. Journal of Mathematical Physics, 43(4):1755–1795, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kijowski:2002:GLG

[974] J. Kijowski and G. Rudolph. On the Gauss law and global charge for quantum chromodynamics. Journal of Mathematical Physics, 43(4): 1796–1808, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shvedov:2002:RPT

[975] O. Yu. Shvedov. Renormalization of Poincaré transformations in Hamiltonian semiclassical field theory. Journal of Mathematical Physics, 43 (4):1809–1843, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sigal:2002:SPH

[976] I. M. Sigal, A. Soffer, and L. Zielinski. On the spectral properties of Hamiltonians without conservation of the particle number. Journal of Mathematical Physics, 43(4):1844–1855, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wospakrik:2002:SHM

[977] H. J. Wospakrik and W. J. Zakrzewski. SU(N) skyrmions from harmonic maps S^{2} ⟶ Gr(2, N). Journal of Mathematical Physics, 43(4):1856– 1874, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palmer:2002:PAM

[978] David R. Palmer. A parabolic approximation method with application to global wave propagation. Journal of Mathematical Physics, 43(4):1875– 1905, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Verhoeven:2002:IGH

[979] C. Verhoeven, M. Musette, and R. Conte. Integration of a generalized Hénon–Heiles Hamiltonian. Journal of Mathematical Physics, 43 (4):1906–1915, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Acosta:2002:GSG

[980] José Acosta and Harold Blas. Generalized sine-Gordon/massive Thirring models and soliton/particle correspondences. Journal of Mathematical Physics, 43(4):1916–1937, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2002:LRS

[981] Chunli Chen and Yishen Li. The Lund–Regge surface and its motion’s evolution equation. Journal of Mathematical Physics, 43(4):1938–1955, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2002:CMK

[982] Deng yuan Chen. k-constraint for the modified Kadomtsev–Petviashvili system. Journal of Mathematical Physics, 43(4):1956–1965, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Halburd:2002:SFR

[983] R. G. Halburd. Shear-free relativistic fluids and the absence of movable branch points. Journal of Mathematical Physics, 43(4):1966–1979, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ganguly:2002:ALV

[984] Asish Ganguly. Associated Lamé and various other new classes of elliptic potentials from sl(2, R) and related orthogonal polynomials. Journal of Mathematical Physics, 43(4):1980–1999, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakashima:2002:IMF

[985] Toshiki Nakashima. Irreducible modules of finite dimensional quantum algebras of type A at roots of unity. Journal of Mathematical Physics, 43 (4):2000–2014, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Robinson:2002:RGH

[986] D. C. Robinson. The real geometry of holomorphic four-metrics. Journal of Mathematical Physics, 43(4):2015–2028, April 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2002:QES

[987] R. B. Zhang. Quantum enveloping superalgebras and link invariants. Journal of Mathematical Physics, 43(4):2029–2048, April 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Altafini:2002:CQM

[988] Claudio Altafini. Controllability of quantum mechanical systems by root space decomposition of SU(N). Journal of Mathematical Physics, 43 (5):2051–2062, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

AlvarezM:2002:GCS

[989] Nibaldo Alvarez M. and Véronique Hussin. Generalized coherent and squeezed states based on the h(1) ⊕ su(2) algebra. Journal of Mathematical Physics, 43(5):2063–2096, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barnum:2002:RQD

[990] H. Barnum and E. Knill. Reversing quantum dynamics with near-optimal quantum and classical fidelity. Journal of Mathematical Physics, 43(5): 2097–2106, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boscain:2002:OCL

[991] Ugo Boscain, Grégoire Charlot, Jean-Paul Gauthier, Stéphane Guérin, and Hans-Rudolf Jauslin. Optimal control in laser-induced population transfer for two- and three-level quantum systems. Journal of Mathematical Physics, 43(5):2107–2132, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Demircioglu:2002:TFS

[992] B. Demircioğlu, S. Kuru, M. Önder, and A. Verçin. Two families of superintegrable and isospectral potentials in two dimensions. Journal of Mathematical Physics, 43(5):2133–2150, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:2002:GBC

[993] P. Exner, P. St’ovícek, and P. Vytras. Generalized boundary conditions for the Aharonov–Bohm effect combined with a homogeneous magnetic field. Journal of Mathematical Physics, 43(5):2151–2168, May 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Froman:2002:PAM

[994] Per Olof Fröman and Kilian Larsson. Phase-amplitude method for numerically exact solution of the differential equations of the two-center Coulomb problem. Journal of Mathematical Physics, 43(5):2169–2179, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goldin:2002:GTB

[995] Gerald A. Goldin and Vladimir M. Shtelen. On gauge transformations of Bäcklund type and higher order nonlinear Schrödinger equations. Journal of Mathematical Physics, 43(5):2180–2186, May 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jencova:2002:QIG

[996] Anna Jencová. Quantum information geometry and standard purification. Journal of Mathematical Physics, 43(5):2187–2201, May 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimov:2002:EEE

[997] A. B. Klimov. Exact evolution equations for SU(2) quasidistribution functions. Journal of Mathematical Physics, 43(5):2202–2213, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nonnenmacher:2002:CQM

[998] S. Nonnenmacher and M. R. Zirnbauer. Det–Det correlations for quantum maps: Dual pair and saddle-point analyses. Journal of Mathematical Physics, 43(5):2214–2240, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakaguchi:2002:CSA

[999] Fuminori Sakaguchi and Masahito Hayashi. Coherent states and annihilation–creation operators associated with the irreducible unitary representations of SU(∞, ∞). Journal of Mathematical Physics, 43(5): 2241–2248, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Villegas-Blas:2002:BTC

[1000] Carlos Villegas-Blas. The Bargmann transform and canonical transformations. Journal of Mathematical Physics, 43(5):2249–2283, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagrov:2002:NSR

[1001] V. G. Bagrov, M. C. Baldiotti, D. M. Gitman, and I. V. Shirokov. New solutions of relativistic wave equations in magnetic fields and longitudinal fields. Journal of Mathematical Physics, 43(5):2284–2305, May 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Helein:2002:FDH

[1002] Frédéric Hélein and Joseph Kouneiher. Finite dimensional Hamiltonian formalism for gauge and quantum field theories. Journal of Mathematical Physics, 43(5):2306–2347, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huffel:2002:QYM

[1003] Helmuth Hüffel. Quantizing Yang–Mills theory on a two-point space. Journal of Mathematical Physics, 43(5):2348–2354, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2002:SDC

[1004] Seongtag Kim and Yoonbai Kim. Self-dual Chern–Simons vortices on Riemann surfaces. Journal of Mathematical Physics, 43(5):2355–2362, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oliynyk:2002:LEP

[1005] Todd A. Oliynyk and H. P. Künzle. Local existence proofs for the boundary value problem for static spherically symmetric Einstein–Yang–Mills fields with compact gauge groups. Journal of Mathematical Physics, 43 (5):2363–2393, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strauss:2002:LPS

[1006] Y. Strauss and L. P. Horwitz. Lax–Phillips scattering theory of a relativistic quantum field theoretical Lee–Friedrichs model and Lee–Oehme– Yang–Wu phenomenology. Journal of Mathematical Physics, 43(5):2394– 2422, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wesson:2002:HDD

[1007] Paul S. Wesson. On higher-dimensional dynamics. Journal of Mathematical Physics, 43(5):2423–2438, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Henkel:2002:GPMa

[1008] Oliver Henkel. Global prescribed mean curvature foliations in cosmological space–times. I. Journal of Mathematical Physics, 43(5):2439–2465, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Henkel:2002:GPMb

[1009] Oliver Henkel. Global prescribed mean curvature foliations in cosmological space–times. II. Journal of Mathematical Physics, 43(5):2466–2485, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vacaru:2002:ECB

[1010] Sergiu I. Vacaru and D. Singleton. Ellipsoidal, cylindrical, bipolar and toroidal wormholes in 5D gravity. Journal of Mathematical Physics, 43 (5):2486–2504, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crampin:2002:BQH

[1011] M. Crampin and W. Sarlet. Bi-quasi-Hamiltonian systems. Journal of Mathematical Physics, 43(5):2505–2517, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:2002:MEA

[1012] Jiu Ding and Lawrence R. Mead. Maximum entropy approximation for Lyapunov exponents of chaotic maps. Journal of Mathematical Physics, 43(5):2518–2522, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manas:2002:INI

[1013] Manuel Mañas. From integrable nets to integrable lattices. Journal of Mathematical Physics, 43(5):2523–2546, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nissimov:2002:SSI

[1014] E. Nissimov and S. Pacheva. Symmetries of supersymmetric integrable hierarchies of KP type. Journal of Mathematical Physics, 43(5):2547– 2586, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vakhnenko:2002:SPD

[1015] Oleksiy O. Vakhnenko. Solitons in parametrically driven discrete nonlinear Schrödinger systems with the exploding range of intersite interactions. Journal of Mathematical Physics, 43(5):2587–2605, May 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fel:2002:PPD

[1016] Leonid G. Fel. Piezoelectricity and Piezomagnetism: Duality in twodimensional checkerboards. Journal of Mathematical Physics, 43(5): 2606–2609, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Willis:2002:NAS

[1017] Barton L. Willis. A new approach to static force-free electromagnetic fields. Journal of Mathematical Physics, 43(5):2610–2615, May 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blomker:2002:EDS

[1018] Dirk Blömker, Jinqiao Duan, and Thomas Wanner. Enstrophy dynamics of stochastically forced large-scale geophysical flows. Journal of Mathematical Physics, 43(5):2616–2626, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roshani:2002:ADP

[1019] Farinaz Roshani and Mohammad Khorrami. Annihilation-diffusion processes: An exactly solvable model. Journal of Mathematical Physics, 43 (5):2627–2635, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seaton:2002:DMM

[1020] K. A. Seaton and M. T. Batchelor. The dilute A_{4} model, the E_{7} mass spectrum and the tricritical Ising model. Journal of Mathematical Physics, 43(5):2636–2653, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shiino:2002:NFP

[1021] Masatoshi Shiino. Nonlinear Fokker–Planck equation exhibiting bifurcation phenomena and generalized thermostatistics. Journal of Mathematical Physics, 43(5):2654–2669, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garoni:2002:LFE

[1022] T. M. Garoni and N. E. Frankel. Lévy flights: Exact results and asymptotics beyond all orders. Journal of Mathematical Physics, 43(5):2670– 2689, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gu:2002:PDO

[1023] Qing Gu and Deguang Han. Phases for dyadic orthonormal wavelets. Journal of Mathematical Physics, 43(5):2690–2706, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guhr:2002:RCCa

[1024] Thomas Guhr and Heiner Kohler. Recursive construction for a class of radial functions. I. Ordinary space. Journal of Mathematical Physics, 43 (5):2707–2740, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guhr:2002:RCCb

[1025] Thomas Guhr and Heiner Kohler. Recursive construction for a class of radial functions. II. Superspace. Journal of Mathematical Physics, 43 (5):2741–2769, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kastler:2002:SLT

[1026] Daniel Kastler. A second look at U_{q} ( S l 2) at third root of unity. Journal of Mathematical Physics, 43(5):2770–2790, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazorchuk:2002:QLA

[1027] Volodymyr Mazorchuk. On quasicrystal Lie algebras. Journal of Mathematical Physics, 43(5):2791–2801, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McPhedran:2002:SBS

[1028] R. C. McPhedran and N. A. Nicorovici. Static Bloch sums for the square array. Journal of Mathematical Physics, 43(5):2802–2813, May 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:2002:PHVb

[1029] Ali Mostafazadeh. Pseudo-Hermiticity versus P T -symmetry. II. A complete characterization of non-Hermitian Hamiltonians with a real spectrum. Journal of Mathematical Physics, 43(5):2814–2816, May 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pulmannova:2002:CDE

[1030] Sylvia Pulmannová. Compatibility and decompositions of effects. Journal of Mathematical Physics, 43(5):2817–2830, May 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sosov:2002:CSS

[1031] Yuriy Sosov and Constantine E. Theodosiou. On the complete solution of the Sturm–Liouville problem (d^{2}X/dx^{2})+λ^{2}X = 0 over a closed interval. Journal of Mathematical Physics, 43(5):2831–2843, May 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Albeverio:2002:PSF

[1032] S. Albeverio, G. Guatteri, and S. Mazzucchi. Phase space Feynman path integrals. Journal of Mathematical Physics, 43(6):2847–2857, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buchstaber:2002:MSE

[1033] V. M. Buchstaber, J. C. Eilbeck, V. Z. Enolskii, D. V. Leykin, and M. Salerno. Multidimensional Schrödinger equations with Abelian potentials. Journal of Mathematical Physics, 43(6):2858–2881, June 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Giachetta:2002:GQM

[1034] G. Giachetta, L. Mangiarotti, and G. Sardanashvily. Geometric quantization of mechanical systems with time-dependent parameters. Journal of Mathematical Physics, 43(6):2882–2894, June 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gu:2002:QTB

[1035] Xiao-Yan Gu, Bin Duan, and Zhong-Qi Ma. Quantum three-body system in D dimensions. Journal of Mathematical Physics, 43(6):2895–2906, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwai:2002:RQSa

[1036] Toshihiro Iwai and Toru Hirose. The reduction of a quantum system of three identical particles on a plane. Journal of Mathematical Physics, 43 (6):2907–2926, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwai:2002:RQSb

[1037] Toshihiro Iwai and Toru Hirose. Reduction of quantum systems with symmetry, continuous and discrete. Journal of Mathematical Physics, 43 (6):2927–2947, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klopp:2002:ILT

[1038] Frédéric Klopp. Internal Lifshitz tails for Schrödinger operators with random potentials. Journal of Mathematical Physics, 43(6):2948–2958, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vlasov:2002:NTU

[1039] Alexander Yu. Vlasov. Noncommutative tori and universal sets of nonbinary quantum gates. Journal of Mathematical Physics, 43(6):2959–2964, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anselmi:2002:NIA

[1040] D. Anselmi. A note on the improvement ambiguity of the stress tensor and the critical limits of correlation functions. Journal of Mathematical Physics, 43(6):2965–2977, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cotaescu:2002:HDP

[1041] Ion I. Cotaescu and Mihai Visinescu. Hierarchy of Dirac, Pauli, and Klein–Gordon conserved operators in Taub–NUT background. Journal of Mathematical Physics, 43(6):2978–2987, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konechny:2002:SGL

[1042] Anatoly Konechny and O. Teoman Turgut. Super-Grassmannian and large N limit of quantum field theory with bosons and fermions. Journal of Mathematical Physics, 43(6):2988–3010, June 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendels:2002:SCP

[1043] E. Mendels. Structure conserving parametrization of Feynman diagrams. Journal of Mathematical Physics, 43(6):3011–3035, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Noltingk:2002:CHT

[1044] Duncan Noltingk. Classical history theory of vector fields. Journal of Mathematical Physics, 43(6):3036–3052, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Savvidou:2002:PIC

[1045] Ntina Savvidou. Poincaré invariance for continuous-time histories. Journal of Mathematical Physics, 43(6):3053–3073, June 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Toprak:2002:LLG

[1046] Erdal Toprak and O. Teoman Turgut. Large N limit of SO(N) gauge theory of fermions and Bosons. Journal of Mathematical Physics, 43 (6):3074–3096, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dahia:2002:ESTa

[1047] F. Dahia and C. Romero. The embedding of space–times in five dimensions with nondegenerate Ricci tensor. Journal of Mathematical Physics, 43(6):3097–3106, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaszak:2002:NLI

[1048] Maciej Blaszak and Wen-Xiu Ma. New Liouville integrable noncanonical Hamiltonian systems from the AKNS spectral problem. Journal of Mathematical Physics, 43(6):3107–3123, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foth:2002:BPS

[1049] Philip Foth. Bruhat Poisson structure on CP^{n} and integrable systems. Journal of Mathematical Physics, 43(6):3124–3132, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sahadevan:2002:CMK

[1050] R. Sahadevan and N. Kannagi. The coupled modified Korteweg–de Vries equations: Similarity reduction, Lie–Bäcklund symmetries and integrability. Journal of Mathematical Physics, 43(6):3133–3146, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fatibene:2002:GSM

[1051] L. Fatibene, M. Ferraris, M. Francaviglia, and R. G. McLenaghan. Generalized symmetries in mechanics and field theories. Journal of Mathematical Physics, 43(6):3147–3161, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Margetis:2002:RHE

[1052] Dionisios Margetis. Radiation of horizontal electric dipole on large dielectric sphere. Journal of Mathematical Physics, 43(6):3162–3201, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nunez:2002:PVH

[1053] Manuel Núñez. Plasma velocity in hydromagnetic dynamos. Journal of Mathematical Physics, 43(6):3202–3206, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Santoprete:2002:SPS

[1054] Manuele Santoprete. Symmetric periodic solutions of the anisotropic Manev problem. Journal of Mathematical Physics, 43(6):3207–3219, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Segev:2002:MIA

[1055] Reuven Segev. Metric-independent analysis of the stress-energy tensor. Journal of Mathematical Physics, 43(6):3220–3231, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goudon:2002:DVS

[1056] Thierry Goudon and Antoine Mellet. Discrete version of the SHE asymptotics: multigroup neutron transport equations. Journal of Mathematical Physics, 43(6):3232–3260, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Korepin:2002:ISV

[1057] V. Korepin and P. Zinn-Justin. Inhomogeneous six-vertex model with domain wall boundary conditions and Bethe ansatz. Journal of Mathematical Physics, 43(6):3261–3267, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pyatov:2002:CDA

[1058] P. N. Pyatov and R. Twarock. Construction of diffusion algebras. Journal of Mathematical Physics, 43(6):3268–3279, June 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2002:ADA

[1059] F. Bagarello and C. Trapani. Algebraic dynamics in O^{∗}-algebras: A perturbative approach. Journal of Mathematical Physics, 43(6):3280– 3292, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davies:2002:SHB

[1060] Ian M. Davies, Aubrey Truman, and Huaizhong Zhao. Stochastic heat and Burgers equations and their singularities. I. Geometrical properties. Journal of Mathematical Physics, 43(6):3293–3328, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duplij:2002:ROC

[1061] Steven Duplij and Wladyslaw Marcinek. Regular obstructed categories and topological quantum field theory. Journal of Mathematical Physics, 43(6):3329–3341, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gorin:2002:IMO

[1062] T. Gorin. Integrals of monomials over the orthogonal group. Journal of Mathematical Physics, 43(6):3342–3351, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:2002:WRS

[1063] A. M. Grundland and W. J. Zakrzewski. The Weierstrass representation for surfaces immersed into R^{8} and CP^{2} maps. Journal of Mathematical Physics, 43(6):3352–3362, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moch:2002:NSE

[1064] Sven Moch, Peter Uwer, and Stefan Weinzierl. Nested sums, expansion of transcendental functions, and multiscale multiloop integrals. Journal of Mathematical Physics, 43(6):3363–3386, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pogosyan:2002:SVS

[1065] G. S. Pogosyan and P. Winternitz. Separation of variables and subgroup bases on n-dimensional hyperboloids. Journal of Mathematical Physics, 43(6):3387–3410, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2002:EFA

[1066] Q. H. Liu, X. Wang, W. H. Qi, L. P. Fu, and B. Hu. Erratum: “The Fejér average and the mean value of a quantity in a quasiclassical wave packet”, [J. Math. Phys. 43, 170 (2002)]. Journal of Mathematical Physics, 43(6): 3411, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [875].

Schmidt:2002:EII

[1067] Andreas U. Schmidt. Erratum: “Infinite infrared regularization and a state space for the Heisenberg algebra”, [J. Math. Phys. 43, 243 (2002)]. Journal of Mathematical Physics, 43(6):3412, June 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [880].

Avron:2002:TEC

[1068] J. E. Avron, A. Elgart, G. M. Graf, and L. Sadun. Time-energy coherent states and adiabatic scattering. Journal of Mathematical Physics, 43 (7):3415–3424, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGuise:2002:CSR

[1069] Hubert de Guise and Marco Bertola. Coherent state realizations of su(n + 1) on the n-torus. Journal of Mathematical Physics, 43(7):3425– 3444, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guan:2002:IVO

[1070] X.-W. Guan, A. Foerster, J. Links, H.-Q Zhou, A. Prestes Tonel, and R. H. McKenzie. Integrable variant of the one-dimensional Hubbard model. Journal of Mathematical Physics, 43(7):3445–3457, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mestel:2002:RAC

[1071] B. D. Mestel and A. H. Osbaldestin. Renormalization analysis of correlation properties in a quasiperiodically forced two-level system. Journal of Mathematical Physics, 43(7):3458–3483, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sato:2002:FSQ

[1072] Masatoshi Sato and Toshiaki Tanaka. N -fold supersymmetry in quantum mechanics—analyses of particular models. Journal of Mathematical Physics, 43(7):3484–3510, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spehner:2002:QJD

[1073] D. Spehner and M. Orszag. Quantum jump dynamics in cavity QED. Journal of Mathematical Physics, 43(7):3511–3537, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tempesta:2002:QMR

[1074] P. Tempesta, E. Alfinito, R. A. Leo, and G. Soliani. Quantum models related to fouled Hamiltonians of the harmonic oscillator. Journal of Mathematical Physics, 43(7):3538–3553, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colladay:2002:RSL

[1075] Don Colladay and Patrick McDonald. Redefining spinors in Lorentzviolating quantum electrodynamics. Journal of Mathematical Physics, 43(7):3554–3564, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohrdieck:2002:PSS

[1076] Stephan Mohrdieck. Phase space structure and short distance behavior of local quantum field theories. Journal of Mathematical Physics, 43 (7):3565–3574, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sneddon:2002:CAP

[1077] G. E. Sneddon. Conditions for the alignment of the principal null directions of two Weyl-like tensors. Journal of Mathematical Physics, 43 (7):3575–3585, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chacon:2002:MFB

[1078] R. Chacón and A. Martínez García-Hoz. Modifying fractal basin boundaries by reshaping periodic terms. Journal of Mathematical Physics, 43 (7):3586–3591, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:2002:CSI

[1079] E. G. Kalnins, J. M. Kress, W. Miller, Jr., and G. S. Pogosyan. Complete sets of invariants for dynamical systems that admit a separation of variables. Journal of Mathematical Physics, 43(7):3592–3609, July 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimek:2002:CLC

[1080] M. Klimek. Conservation laws for a class of nonlinear equations with variable coefficients on discrete and noncommutative spaces. Journal of Mathematical Physics, 43(7):3610–3635, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soloviev:2002:BVHa

[1081] Vladimir O. Soloviev. Boundary values as Hamiltonian variables. II. Graded structures. Journal of Mathematical Physics, 43(7):3636–3654, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soloviev:2002:BVHb

[1082] Vladimir O. Soloviev. Boundary values as Hamiltonian variables. III. Ideal fluid with a free surface. Journal of Mathematical Physics, 43(7): 3655–3675, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fel:2002:KTA

[1083] Leonid G. Fel. On Keller theorem for anisotropic media. Journal of Mathematical Physics, 43(7):3676–3677, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haghighipour:2002:RDP

[1084] Nader Haghighipour. Resonance dynamics and partial averaging in a restricted three-body system. Journal of Mathematical Physics, 43(7): 3678–3694, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Batisheva:2002:MPT

[1085] J. G. Batisheva, V. V. Vedenyapin, and S. I. Kuchanov. A mathematical problem of the theory of gelation. Journal of Mathematical Physics, 43 (7):3695–3703, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerra:2002:QRC

[1086] Francesco Guerra and Fabio Lucio Toninelli. Quadratic replica coupling in the Sherrington–Kirkpatrick mean field spin glass model. Journal of Mathematical Physics, 43(7):3704–3716, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:2002:WFL

[1087] Tuncay Aktosun and James H. Rose. Wave focusing on the line. Journal of Mathematical Physics, 43(7):3717–3745, July 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoniou:2002:RSC

[1088] I. Antoniou, S. A. Shkarin, and E. Yarevsky. The resonance spectrum of the cusp map in the space of analytic functions. Journal of Mathematical Physics, 43(7):3746–3758, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adesi:2002:ACH

[1089] Vittorio Barone Adesi and Sergio Zerbini. Analytic continuation of the Hurwitz zeta function with physical application. Journal of Mathematical Physics, 43(7):3759–3765, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grantcharov:2002:RHS

[1090] Gueo Grantcharov, George Papadopoulos, and Yat Sun Poon. Reduction of HKT-structures. Journal of Mathematical Physics, 43(7):3766–3782, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Itoh:2002:CMM

[1091] Mitsuhiro Itoh. Contact metric 5-manifolds, CR twistor spaces and integrability. Journal of Mathematical Physics, 43(7):3783–3797, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1278].

Khare:2002:CII

[1092] Avinash Khare and Uday Sukhatme. Cyclic identities involving Jacobi elliptic functions. Journal of Mathematical Physics, 43(7):3798–3806, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konopelchenko:2002:DSI

[1093] B. Konopelchenko and L. Martinez Alonso. Dispersionless scalar integrable hierarchies, Whitham hierarchy, and the quasiclassical ¯∂-dressing method. Journal of Mathematical Physics, 43(7):3807–3823, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lievens:2002:CCC

[1094] S. Lievens and J. Van der Jeugt. 3nj-coefficients of su(1, 1) as connection coefficients between orthogonal polynomials in n variables. Journal of Mathematical Physics, 43(7):3824–3849, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palev:2002:JGF

[1095] T. D. Palev and J. Van der Jeugt. Jacobson generators, Fock representations and statistics of sl(n + 1). Journal of Mathematical Physics, 43 (7):3850–3873, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chauvineau:2002:EEF

[1096] Bertrand Chauvineau. Erratum: “Euclidean Friedman–Robertson– Walker cosmology in Brans–Dicke-like theories of gravity”, [J. Math. Phys. 43, 1487 (2002)]. Journal of Mathematical Physics, 43(7):3874, July 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [957].

Aschbacher:2002:SBR

[1097] W. H. Aschbacher, J. Fröhlich, G. M. Graf, K. Schnee, and M. Troyer. Symmetry breaking regime in the nonlinear Hartree equation. Journal of Mathematical Physics, 43(8):3879–3891, August 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dittrich:2002:BSS

[1098] J. Dittrich and J. Kríz. Bound states in straight quantum waveguides with combined boundary conditions. Journal of Mathematical Physics, 43(8):3892–3915, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Emary:2002:BTE

[1099] C. Emary and R. F. Bishop. Bogoliubov transformations and exact isolated solutions for simple nonadiabatic Hamiltonians. Journal of Mathematical Physics, 43(8):3916–3926, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feinberg:2002:DOF

[1100] Joshua Feinberg. The Dirac operator in a fermion bag background in 1+1 dimensions and generalized supersymmetric quantum mechanics. Journal of Mathematical Physics, 43(8):3927–3936, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krajewska:2002:SIS

[1101] K. Krajewska and J. Z. Kaminski. Stabilization of impurity states in crossed magnetic and electric fields. Journal of Mathematical Physics, 43(8):3937–3943, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:2002:PHVc

[1102] Ali Mostafazadeh. Pseudo-Hermiticity versus P T -symmetry III: Equivalence of pseudo-Hermiticity and the presence of antilinear symmetries. Journal of Mathematical Physics, 43(8):3944–3951, August 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sogut:2002:ELW

[1103] K. Sogut, A. Havare, and I. Acikgoz. Energy levels and wave functions of vector bosons in a homogeneous magnetic field. Journal of Mathematical Physics, 43(8):3952–3962, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ichinose:2002:PRD

[1104] Takashi Ichinose and Brian Jefferies. The propagator of the radial Dirac equation. Journal of Mathematical Physics, 43(8):3963–3983, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2002:PSS

[1105] Stephen C. Anco and Roh S. Tung. Properties of the symplectic structure of general relativity for spatially bounded space–time regions. Journal of Mathematical Physics, 43(8):3984–4019, August 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1756].

Carminati:2002:PACb

[1106] J. Carminati and E. Zakhary. On the problem of algebraic completeness for the invariants of the Riemann tensor. III. Journal of Mathematical Physics, 43(8):4020–4034, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guha:2002:ANM

[1107] Partha Guha. Applications of Nambu mechanics to systems of hydrodynamical type. Journal of Mathematical Physics, 43(8):4035–4040, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karlovini:2002:UTQ

[1108] Max Karlovini, Giuseppe Pucacco, Kjell Rosquist, and Lars Samuelsson. A unified treatment of quartic invariants at fixed and arbitrary energy. Journal of Mathematical Physics, 43(8):4041–4059, August 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2002:MGM

[1109] Tingting Liu and Mengzhao Qin. Multisymplectic geometry and multisymplectic Preissman scheme for the KP equation. Journal of Mathematical Physics, 43(8):4060–4077, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:2002:DCA

[1110] Sen yue Lou, Chun li Chen, and Xiao yan Tang. (2 + 1)-dimensional (M + N)-component AKNS system: Painlevé integrability, infinitely many symmetries, similarity reductions and exact solutions. Journal of Mathematical Physics, 43(8):4078–4109, August 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amiet:2002:CHO

[1111] Jean-Pierre Amiet and Stefan Weigert. Commensurate harmonic oscillators: Classical symmetries. Journal of Mathematical Physics, 43(8): 4110–4126, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shabanov:2002:LEB

[1112] Sergei V. Shabanov. On a low energy bound in a class of chiral field theories with solitons. Journal of Mathematical Physics, 43(8):4127– 4134, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGier:2002:XSC

[1113] J. de Gier, M. T. Batchelor, B. Nienhuis, and S. Mitra. The XXZ spin chain at Δ = −1/2: Bethe roots, symmetric functions, and determinants. Journal of Mathematical Physics, 43(8):4135–4146, August 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petitjean:2002:CM

[1114] Michel Petitjean. Chiral mixtures. Journal of Mathematical Physics, 43 (8):4147–4157, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baumgartel:2002:TDC

[1115] Hellmut Baumgärtel, Matthias Jurke, and Fernando Lledó. Twisted duality of the CAR-algebra. Journal of Mathematical Physics, 43(8): 4158–4179, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Begin:2002:FBF

[1116] L. Bégin, C. Cummins, L. Lapointe, and P. Mathieu. Fusion bases as facets of polytopes. Journal of Mathematical Physics, 43(8):4180–4201, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:2002:BLT

[1117] Carl M. Bender, Axel Pelster, and Florian Weissbach. Boundary-layer theory, strong-coupling series, and large-order behavior. Journal of Mathematical Physics, 43(8):4202–4220, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leznov:2002:GZS

[1118] A. N. Leznov. The Gelfand–Zeitlin selection rules and representations of quantum A_{n}^{q} algebras. Journal of Mathematical Physics, 43(8):4221– 4233, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruss:2002:CE

[1119] Dagmar Bruß. Characterizing entanglement. Journal of Mathematical Physics, 43(9):4237–4251, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Donald:2002:UTE

[1120] Matthew J. Donald, Michal Horodecki, and Oliver Rudolph. The uniqueness theorem for entanglement measures. Journal of Mathematical Physics, 43(9):4252–4272, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meyer:2002:GEM

[1121] David A. Meyer and Nolan R. Wallach. Global entanglement in multiparticle systems. Journal of Mathematical Physics, 43(9):4273–4278, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schumacher:2002:EPQ

[1122] Benjamin Schumacher and Michael D. Westmoreland. Entanglement and perfect quantum error correction. Journal of Mathematical Physics, 43 (9):4279–4285, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Terhal:2002:EP

[1123] Barbara M. Terhal, Michal Horodecki, Debbie W. Leung, and David P. DiVincenzo. The entanglement of purification. Journal of Mathematical Physics, 43(9):4286–4298, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vollbrecht:2002:CET

[1124] Karl Gerd H. Vollbrecht and Michael M. Wolf. Conditional entropies and their relation to entanglement criteria. Journal of Mathematical Physics, 43(9):4299–4306, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wootters:2002:PTE

[1125] William K. Wootters. Parallel transport in an entangled ring. Journal of Mathematical Physics, 43(9):4307–4325, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holevo:2002:EAC

[1126] A. S. Holevo. On entanglement-assisted classical capacity. Journal of Mathematical Physics, 43(9):4326–4333, September 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shor:2002:ACC

[1127] Peter W. Shor. Additivity of the classical capacity of entanglementbreaking quantum channels. Journal of Mathematical Physics, 43(9): 4334–4340, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Winter:2002:SPQ

[1128] Andreas Winter. Scalable programmable quantum gates and a new aspect of the additivity problem for the classical capacity of quantum channels. Journal of Mathematical Physics, 43(9):4341–4352, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Werner:2002:CAC

[1129] R. F. Werner and A. S. Holevo. Counterexample to an additivity conjecture for output purity of quantum channels. Journal of Mathematical Physics, 43(9):4353–4357, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruskai:2002:IQE

[1130] Mary Beth Ruskai. Inequalities for quantum entropy: A review with conditions for equality. Journal of Mathematical Physics, 43(9):4358– 4375, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1945].

Zalka:2002:QOC

[1131] Christof Zalka and Eleanor Rieffel. Quantum operations that cannot be implemented using a small mixed environment. Journal of Mathematical Physics, 43(9):4376–4381, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamada:2002:LBQ

[1132] Mitsuru Hamada. A lower bound on the quantum capacity of channels with correlated errors. Journal of Mathematical Physics, 43(9):4382– 4390, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsumoto:2002:LBQ

[1133] Ryutaroh Matsumoto and Tomohiko Uyematsu. Lower bound for the quantum capacity of a discrete memoryless quantum channel. Journal of Mathematical Physics, 43(9):4391–4403, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hayden:2002:TQC

[1134] Patrick Hayden, Richard Jozsa, and Andreas Winter. Trading quantum for classical resources in quantum data compression. Journal of Mathematical Physics, 43(9):4404–4444, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harrow:2002:EDA

[1135] Aram W. Harrow, Benjamin Recht, and Isaac L. Chuang. Efficient discrete approximations of quantum gates. Journal of Mathematical Physics, 43(9):4445–4451, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dennis:2002:TQM

[1136] Eric Dennis, Alexei Kitaev, Andrew Landahl, and John Preskill. Topological quantum memory. Journal of Mathematical Physics, 43(9):4452– 4505, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2002:QP

[1137] L.-A. Wu and D. A. Lidar. Qubits as parafermions. Journal of Mathematical Physics, 43(9):4506–4525, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brandt:2002:SCF

[1138] Howard E. Brandt. Secrecy capacity in the four-state protocol of quantum key distribution. Journal of Mathematical Physics, 43(9):4526–4530, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gea-Banacloche:2002:HMQ

[1139] Julio Gea-Banacloche. Hiding messages in quantum data. Journal of Mathematical Physics, 43(9):4531–4536, September 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caves:2002:UQS

[1140] Carlton M. Caves, Christopher A. Fuchs, and Rüdiger Schack. Unknown quantum states: The quantum de Finetti representation. Journal of Mathematical Physics, 43(9):4537–4559, September 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1951].

Mermin:2002:CSA

[1141] N. David Mermin. Compatibility of state assignments. Journal of Mathematical Physics, 43(9):4560–4566, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Astaburuaga:2002:BSP

[1142] M. Angélica Astaburuaga, Pablo Covian, and Claudio Fernández. Behavior of the survival probability in some one-dimensional problems. Journal of Mathematical Physics, 43(10):4571–4581, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:2002:RHS

[1143] Myung-Hoon Chung and Jae-Hoon Kwon. Reduction of the Hilbert space in strongly correlated systems. Journal of Mathematical Physics, 43(10): 4582–4591, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dayi:2002:HEN

[1144] Ömer F. Dayi and Ahmed Jellal. Hall effect in noncommutative coordinates. Journal of Mathematical Physics, 43(10):4592–4601, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1676].

Dias:2002:WFB

[1145] Nuno Costa Dias and João Nuno Prata. Wigner functions with boundaries. Journal of Mathematical Physics, 43(10):4602–4627, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hou:2002:DEQ

[1146] Boyu Hou, Kangjie Shi, Ruihong Yue, and Shaoyou Zhao. The dynamical elliptic quantum Gaudin models and their solutions. Journal of Mathematical Physics, 43(10):4628–4640, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

King:2002:AUQ

[1147] Christopher King. Additivity for unital qubit channels. Journal of Mathematical Physics, 43(10):4641–4653, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kobayashi:2002:MSH

[1148] M. Kobayashi. Method of solving the homogeneous Bloch equation. Journal of Mathematical Physics, 43(10):4654–4667, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Littlejohn:2002:SGD

[1149] Robert G. Littlejohn and Paul Wright. Semiclassical generalization of the Darboux–Christoffel formula. Journal of Mathematical Physics, 43(10): 4668–4680, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parfitt:2002:TDH

[1150] D. G. W. Parfitt and M. E. Portnoi. The two-dimensional hydrogen atom revisited. Journal of Mathematical Physics, 43(10):4681–4691, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanWonderen:2002:UJC

[1151] A. J. van Wonderen and K. Lendi. Unification of the Jaynes–Cummings model and Planck’s radiation law. Journal of Mathematical Physics, 43(10):4692–4720, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bierenbaum:2002:IRF

[1152] Isabella Bierenbaum, Richard Kreckel, and Dirk Kreimer. On the invariance of residues of Feynman graphs. Journal of Mathematical Physics, 43(10):4721–4740, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galiautdinov:2002:CCD

[1153] Andrei A. Galiautdinov and David R. Finkelstein. Chronon corrections to the Dirac equation. Journal of Mathematical Physics, 43(10):4741–4752, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gottschalk:2002:CVT

[1154] Hanno Gottschalk. Complex velocity transformations and the Bisognano–Wichmann theorem for quantum fields acting on Krein spaces. Journal of Mathematical Physics, 43(10):4753–4769, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ilyenko:2002:TRN

[1155] Kostyantin Ilyenko. Twistor representation of null two-surfaces. Journal of Mathematical Physics, 43(10):4770–4789, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Misra:2002:PNC

[1156] Aalok Misra. The peculiarity of a negative coordinate axis in dyonic solutions of noncommutative \ = △ super Yang–Mills. Journal of Mathematical Physics, 43(10):4790–4800, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Noltingk:2002:BRS

[1157] Duncan Noltingk. Becchi–Rouet–Stora–Tyutin quantization of histories electrodynamics. Journal of Mathematical Physics, 43(10):4801–4818, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Siemens:2002:CSS

[1158] Xavier Siemens and Ken D. Olum. Chiral superconducting strings and Nambu–Goto strings in arbitrary dimensions. Journal of Mathematical Physics, 43(10):4819–4827, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Turgut:2002:TDC

[1159] O. Teoman Turgut. On three-dimensional coupled bosons. Journal of Mathematical Physics, 43(10):4828–4838, October 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buchner:2002:PC

[1160] K. Buchner. 1/r-potential without charge. Journal of Mathematical Physics, 43(10):4839–4848, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coley:2002:IMT

[1161] Alan A. Coley and R. J. van den Hoogen. Inhomogeneous M-theory cosmologies. Journal of Mathematical Physics, 43(10):4849–4860, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flores:2002:TMG

[1162] José Luis Flores and Miguel Sánchez. A topological method for geodesic connectedness of space–times: Outer Kerr space–time. Journal of Mathematical Physics, 43(10):4861–4885, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gogberashvili:2002:EBS

[1163] Merab Gogberashvili. Embedding of the brane into six dimensions. Journal of Mathematical Physics, 43(10):4886–4888, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:2002:AGF

[1164] L. Herrera, J. Martin, and J. Ospino. Anisotropic geodesic fluid spheres in general relativity. Journal of Mathematical Physics, 43(10):4889–4897, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kohlprath:2002:STG

[1165] Emmanuel Kohlprath. Schwinger terms in gravitation in two dimensions as a consequence of the gravitational anomaly. Journal of Mathematical Physics, 43(10):4898–4903, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2002:TDS

[1166] Hongya Liu and Qiuyang Zhang. Time-dependent solution for a star immersed in a background radiation. Journal of Mathematical Physics, 43(10):4904–4908, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1542].

Villalba:2002:SVE

[1167] Víctor M. Villalba and Esteban Isasi Catalá. Separation of variables and exact solution of the Klein–Gordon and Dirac equations in an open universe. Journal of Mathematical Physics, 43(10):4909–4920, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dye:2002:BKK

[1168] J. M. Dye and A. Parker. A bidirectional Kaup–Kupershmidt equation and directionally dependent solitons. Journal of Mathematical Physics, 43(10):4921–4949, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Li:2002:NIS

[1169] Yishen Li and Wen-Xiu Ma. A nonconfocal involutive system and constrained flows associated with the MKdV_{−−} equation. Journal of Mathematical Physics, 43(10):4950–4962, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saito:2002:HFA

[1170] Satoru Saito, Akira Shudo, Jun ichi Yamamoto, and Katsuhiko Yoshida. Hamiltonian flow associated with a two-dimensional map. Journal of Mathematical Physics, 43(10):4963–4965, October 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaisman:2002:HSF

[1171] Izu Vaisman. Hamiltonian structures on foliations. Journal of Mathematical Physics, 43(10):4966–4977, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yan:2002:FNI

[1172] Zhenya Yan and Hongqing Zhang. A family of new integrable couplings with two arbitrary functions of TC hierarchy. Journal of Mathematical Physics, 43(10):4978–4986, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2002:EFD

[1173] Xiang Zhang. Exponential factors and Darbouxian first integrals of the Lorenz system. Journal of Mathematical Physics, 43(10):4987–5001, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2002:LIF

[1174] Zixiang Zhou. Liouville integrability of the finite dimensional Hamiltonian systems derived from principal chiral field. Journal of Mathematical Physics, 43(10):5002–5012, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fiorani:2002:GQT

[1175] E. Fiorani, G. Giachetta, and G. Sardanashvily. Geometric quantization of time-dependent completely integrable Hamiltonian systems. Journal of Mathematical Physics, 43(10):5013–5025, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rivera:2002:FSR

[1176] R. Rivera and D. Villarroel. A family of solutions with radiation reaction and retarded interactions for two charges in classical electrodynamics. Journal of Mathematical Physics, 43(10):5026–5044, October 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weston:2002:WFL

[1177] Vaughan H. Weston and B. Lars G. Jonsson. Wave front layer stripping approach to inverse scattering for the wave equation. Journal of Mathematical Physics, 43(10):5045–5059, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kato:2002:DCT

[1178] Go Kato and Miki Wadati. Direct calculation of thermodynamic quantities for the Heisenberg model. Journal of Mathematical Physics, 43(10): 5060–5078, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2002:EFE

[1179] Qiuping A. Wang and Alain Le Méhauté. Extensive form of equilibrium nonextensive statistics. Journal of Mathematical Physics, 43(10):5079– 5089, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garoni:2002:DLF

[1180] T. M. Garoni and N. E. Frankel. d-dimensional Lévy flights: Exact and asymptotic. Journal of Mathematical Physics, 43(10):5090–5107, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Louck:2002:NPA

[1181] James D. Louck. A new parametrization and all integral realizations of the Lorentz group. Journal of Mathematical Physics, 43(10):5108–5134, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prosen:2002:IMS

[1182] T. Prosen, T. H. Seligman, and H. A. Weidenmüller. Integration over matrix spaces with unique invariant measures. Journal of Mathematical Physics, 43(10):5135–5144, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

RauschdeTraubenberg:2002:FDL

[1183] M. Rausch de Traubenberg and M. J. Slupinski. Finite-dimensional Lie algebras of order F. Journal of Mathematical Physics, 43(10):5145–5160, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smolyanov:2002:HFP

[1184] O. G. Smolyanov, A. G. Tokarev, and A. Truman. Hamiltonian Feynman path integrals via the Chernoff formula. Journal of Mathematical Physics, 43(10):5161–5171, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:2002:MEP

[1185] G. F. Torres del Castillo and F. J. Hernández-Moreno. Multipole expansion of a plane wave. Journal of Mathematical Physics, 43(10):5172–5178, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benenti:2002:RCBa

[1186] S. Benenti, C. Chanu, and G. Rastelli. Remarks on the connection between the additive separation of the Hamilton–Jacobi equation and the multiplicative separation of the Schrödinger equation. I. The completeness and Robertson conditions. Journal of Mathematical Physics, 43(11): 5183–5222, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benenti:2002:RCBb

[1187] S. Benenti, C. Chanu, and G. Rastelli. Remarks on the connection between the additive separation of the Hamilton–Jacobi equation and the multiplicative separation of the Schrödinger equation. II. First integrals and symmetry operators. Journal of Mathematical Physics, 43(11):5223– 5253, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brody:2002:ESQ

[1188] Dorje C. Brody and Lane P. Hughston. Efficient simulation of quantum state reduction. Journal of Mathematical Physics, 43(11):5254–5261, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chaturvedi:2002:SCMa

[1189] S. Chaturvedi and N. Mukunda. The Schwinger SU(3) construction. I. Multiplicity problem and relation to induced representations. Journal of Mathematical Physics, 43(11):5262–5277, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chaturvedi:2002:SCMb

[1190] S. Chaturvedi and N. Mukunda. The Schwinger SU(3) construction. II. Relations between Heisenberg–Weyl and SU(3) coherent states. Journal of Mathematical Physics, 43(11):5278–5309, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ganguly:2002:NCQ

[1191] Asish Ganguly. New classes of quasi-solvable potentials, their exactly solvable limit, and related orthogonal polynomials. Journal of Mathematical Physics, 43(11):5310–5324, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaiser:2002:DCD

[1192] Hans-Christoph Kaiser, Hagen Neidhardt, and Joachim Rehberg. Density and current of a dissipative Schrödinger operator. Journal of Mathematical Physics, 43(11):5325–5350, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mathur:2002:SCS

[1193] Manu Mathur and H. S. Mani. SU( N ) coherent states. Journal of Mathematical Physics, 43(11):5351–5364, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melgaard:2002:BSS

[1194] Michael Melgaard. On bound states for systems of weakly coupled Schrödinger equations in one space dimension. Journal of Mathematical Physics, 43(11):5365–5385, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Posilicano:2002:SCF

[1195] Andrea Posilicano and Stefania Ugolini. Scattering into cones and flux across surfaces in quantum mechanics: A pathwise probabilistic approach. Journal of Mathematical Physics, 43(11):5386–5399, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:2002:CST

[1196] D. J. Rowe and J. Repka. Coherent state triplets and their inner products. Journal of Mathematical Physics, 43(11):5400–5438, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Solombrino:2002:WPH

[1197] L. Solombrino. Weak pseudo-Hermiticity and antilinear commutant. Journal of Mathematical Physics, 43(11):5439–5445, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vacchini:2002:QOV

[1198] Bassano Vacchini. Quantum optical versus quantum Brownian motion master equation in terms of covariance and equilibrium properties. Journal of Mathematical Physics, 43(11):5446–5458, November 2002. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yildiz:2002:QGC

[1199] A. Yildiz. Quantum group covariance and the braided structure of deformed oscillators. Journal of Mathematical Physics, 43(11):5459–5469, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chae:2002:GEC

[1200] Dongho Chae and Myeongju Chae. The global existence in the Cauchy problem of the Maxwell–Chern–Simons–Higgs system. Journal of Mathematical Physics, 43(11):5470–5482, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haba:2002:SIE

[1201] Z. Haba. Scale invariant Euclidean field theory in any dimension. Journal of Mathematical Physics, 43(11):5483–5492, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isham:2002:FOH

[1202] C. J. Isham and K. Savvidou. The foliation operator in history quantum field theory. Journal of Mathematical Physics, 43(11):5493–5513, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strohmaier:2002:MAQ

[1203] Alexander Strohmaier, Rainer Verch, and Manfred Wollenberg. Microlocal analysis of quantum fields on curved space–times: Analytic wave front sets and Reeh–Schlieder theorems. Journal of Mathematical Physics, 43 (11):5514–5530, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2002:CHB

[1204] Stephen C. Anco and Roh S. Tung. Covariant Hamiltonian boundary conditions in General Relativity for spatially bounded space–time regions. Journal of Mathematical Physics, 43(11):5531–5566, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1757].

Coley:2002:MDI

[1205] A. A. Coley, G. S. Hall, A. J. Keane, and B. O. J. Tupper. The maximum dimension of the inheriting algebra in perfect fluid space–times. Journal of Mathematical Physics, 43(11):5567–5577, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:2002:WCS

[1206] Simonetta Frittelli and A. O. Petters. Wavefronts, caustic sheets, and caustic surfing in gravitational lensing. Journal of Mathematical Physics, 43(11):5578–5611, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zapata:2002:CSF

[1207] José A. Zapata. Continuum spin foam model for 3d gravity. Journal of Mathematical Physics, 43(11):5612–5623, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bliokh:2002:GBG

[1208] K. Yu. Bliokh. Generalization of Berry’s geometric phase, equivalence of the Hamiltonian nature, quantizability and strong stability of linear oscillatory systems, and conservation of adiabatic invariants. Journal of Mathematical Physics, 43(11):5624–5636, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Komori:2002:ERO

[1209] Yasushi Komori. Elliptic Ruijsenaars operators and functional equations. Journal of Mathematical Physics, 43(11):5637–5653, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarlet:2002:LAS

[1210] W. Sarlet, T. Mestdag, and E. Martínez. Lie algebroid structures on a class of affine bundles. Journal of Mathematical Physics, 43(11):5654– 5674, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Veselov:2002:GCM

[1211] A. P. Veselov. On generalizations of the Calogero–Moser–Sutherland quantum problem and WDVV equations. Journal of Mathematical Physics, 43(11):5675–5682, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Athanasiadis:2002:SRP

[1212] C. Athanasiadis, P. A. Martin, A. Spyropoulos, and I. G. Stratis. Scattering relations for point sources: Acoustic and electromagnetic waves. Journal of Mathematical Physics, 43(11):5683–5697, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fortunato:2002:BIT

[1213] D. Fortunato, L. Orsina, and L. Pisani. Born–Infeld type equations for electrostatic fields. Journal of Mathematical Physics, 43(11):5698–5706, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Laurence:2002:ORV

[1214] Peter Laurence and Edward W. Stredulinsky. Optimal regularity in a variational problem for current sheets in ideal magnetohydrodynamics. Journal of Mathematical Physics, 43(11):5707–5719, November 2002.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhattacharya:2002:ISS

[1215] S. Bhattacharya and J. Blawzdziewicz. Image system for Stokes-flow singularity between two parallel planar walls. Journal of Mathematical Physics, 43(11):5720–5731, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garajeu:2002:BAG

[1216] Daniela Garajeu. Bethe ansatz for the Gaudin model and its relation with Knizhnik–Zamolodchikov equations. Journal of Mathematical Physics, 43(11):5732–5756, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bandelloni:2002:ACD

[1217] G. Bandelloni and S. Lazzarini. W_{∞}-algebras in n complex dimensions and Kodaira–Spencer deformations: A symplectic approach. Journal of Mathematical Physics, 43(11):5757–5772, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biswas:2002:ASI

[1218] Indranil Biswas. Analog of secondary invariants for holomorphic vector bundles. Journal of Mathematical Physics, 43(11):5773–5781, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borstnik:2002:HGS

[1219] N. Mankoc Borstnik and H. B. Nielsen. How to generate spinor representations in any dimension in terms of projection and nilpotent operators. Journal of Mathematical Physics, 43(11):5782–5803, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dahia:2002:ESTb

[1220] F. Dahia and C. Romero. The embedding of the space–time in five dimensions: An extension of the Campbell–Magaard theorem. Journal of Mathematical Physics, 43(11):5804–5814, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeLeo:2002:QEP

[1221] Stefano De Leo, Giuseppe Scolarici, and Luigi Solombrino. Quaternionic eigenvalue problem. Journal of Mathematical Physics, 43(11):5815–5829, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dumitriu:2002:MMB

[1222] Ioana Dumitriu and Alan Edelman. Matrix models for beta ensembles. Journal of Mathematical Physics, 43(11):5830–5847, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:2002:MDG

[1223] Y. Brihaye, B. Hartmann, and D. H. Tchrakian. Monopoles and dyons in SO(3) gauged Skyrme models: Corrigendum. Journal of Mathematical Physics, 43(11):5848–5851, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [694].

Alonso:2002:WFC

[1224] Miguel Angel Alonso, George S. Pogosyan, and Kurt Bernardo Wolf. Wigner functions for curved spaces. I. On hyperboloids. Journal of Mathematical Physics, 43(12):5857–5871, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arias:2002:FPQ

[1225] A. Arias, A. Gheondea, and S. Gudder. Fixed points of quantum operations. Journal of Mathematical Physics, 43(12):5872–5881, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dias:2002:CC

[1226] Nuno Costa Dias. Classicality criteria. Journal of Mathematical Physics, 43(12):5882–5901, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gravel:2002:STO

[1227] Simon Gravel and Pavel Winternitz. Superintegrability with third-order integrals in quantum and classical mechanics. Journal of Mathematical Physics, 43(12):5902–5912, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:2002:CPS

[1228] Richard L. Hall, Wolfgang Lucha, and Franz F. Schöberl. Convexity and potential sums for Salpeter-type Hamiltonians. Journal of Mathematical Physics, 43(12):5913–5925, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanimura:2002:MTG

[1229] Shogo Tanimura. Magnetic translation groups in an n-dimensional torus and their representations. Journal of Mathematical Physics, 43(12):5926– 5948, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2002:TAF

[1230] Tai Tsun Wu and Ming Lun Yu. Theory and application of Fermi pseudopotential in one dimension. Journal of Mathematical Physics, 43(12): 5949–5976, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ying:2002:EBA

[1231] Zu-Jian Ying, You-Quan Li, Shi-Jian Gu, and Bin Chen. Extension of Bethe ansatz to multiple occupancies for one-dimensional SU(4) fermions with δ-function interaction. Journal of Mathematical Physics, 43(12): 5977–5986, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barkallil:2002:RWZ

[1232] A. Barkallil, G. Barnich, and C. Schomblond. Results on the Wess– Zumino consistency condition for arbitrary Lie algebras. Journal of Mathematical Physics, 43(12):5987–6015, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grebeniuk:2002:BBS

[1233] M. A. Grebeniuk, V. D. Ivashchuk, and S.-W. Kim. Black-brane solution for C_{2} algebra. Journal of Mathematical Physics, 43(12):6016–6023, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Polyzou:2002:CPR

[1234] W. N. Polyzou. Cluster properties in relativistic quantum mechanics of N-particle systems. Journal of Mathematical Physics, 43(12):6024–6063, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shabani:2002:ESM

[1235] J. Shabani and A. Vyabandi. Exactly solvable models of relativistic δ-sphere interactions in quantum mechanics. Journal of Mathematical Physics, 43(12):6064–6084, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smyrnakis:2002:BSE

[1236] Ioannis Smyrnakis. Boundary states, extended symmetry algebra, and module structure for certain rational torus models. Journal of Mathematical Physics, 43(12):6085–6095, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bronnikov:2002:STG

[1237] K. A. Bronnikov. Scalar-tensor gravity and conformal continuations. Journal of Mathematical Physics, 43(12):6096–6115, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brunelli:2002:NEN

[1238] J. C. Brunelli and G. A. T. F. da Costa. On the nonlocal equations and nonlocal charges associated with the Harry Dym hierarchy. Journal of Mathematical Physics, 43(12):6116–6128, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosenhaus:2002:ISC

[1239] V. Rosenhaus. Infinite symmetries and conservation laws. Journal of Mathematical Physics, 43(12):6129–6150, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tzirtzilakis:2002:SLS

[1240] E. Tzirtzilakis, V. Marinakis, C. Apokis, and T. Bountis. Soliton-like solutions of higher order wave equations of the Korteweg–de Vries type. Journal of Mathematical Physics, 43(12):6151–6165, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lundmark:2002:DNE

[1241] Hans Lundmark and Stefan Rauch-Wojciechowski. Driven Newton equations and separable time-dependent potentials. Journal of Mathematical Physics, 43(12):6166–6194, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanPutten:2002:UMD

[1242] Maurice H. P. M. van Putten. Uniqueness in MHD in divergence form: Right nullvectors and well-posedness. Journal of Mathematical Physics, 43(12):6195–6208, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Auil:2002:FPD

[1243] F. Auil. Four-particle decay of the Bethe–Salpeter kernel in the hightemperature Ising model. Journal of Mathematical Physics, 43(12):6209– 6223, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerra:2002:CLT

[1244] Francesco Guerra and Fabio Lucio Toninelli. Central limit theorem for fluctuations in the high temperature region of the Sherrington– Kirkpatrick spin glass model. Journal of Mathematical Physics, 43(12): 6224–6237, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnaudon:2002:DRN

[1245] D. Arnaudon, A. Chakrabarti, V. K. Dobrev, and S. G. Mihov. Duality and representations for new exotic bialgebras. Journal of Mathematical Physics, 43(12):6238–6264, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borisov:2002:GCT

[1246] D. Borisov, P. Exner, and R. Gadyl’shin. Geometric coupling thresholds in a two-dimensional strip. Journal of Mathematical Physics, 43(12): 6265–6278, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brink:2002:CSR

[1247] Lars Brink, Abu M. Khan, Pierre Ramond, and Xiaozhen Xiong. Continuous spin representations of the Poincaré and super-Poincaré groups. Journal of Mathematical Physics, 43(12):6279–6295, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dai:2002:AES

[1248] Lianrong Dai, Feng Pan, and J. P. Draayer. Analytical expressions for some multiplicity-free isoscalar factors of S_{f} ⊃ S_{f−1}. Journal of Mathematical Physics, 43(12):6296–6306, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fiore:2002:MRQ

[1249] G. Fiore, M. Maceda, and J. Madore. Metrics on the real quantum plane. Journal of Mathematical Physics, 43(12):6307–6324, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hioe:2002:CNS

[1250] F. T. Hioe. N coupled nonlinear Schrödinger equations: Special set and applications to N = 3. Journal of Mathematical Physics, 43(12): 6325–6338, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1369].

Law:2002:RWD

[1251] Peter R. Law. De Rham–Wu decomposition of holomorphic Riemannian manifolds. Journal of Mathematical Physics, 43(12):6339–6342, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:2002:PHC

[1252] Ali Mostafazadeh. Pseudo-Hermiticity for a class of nondiagonalizable Hamiltonians. Journal of Mathematical Physics, 43(12):6343–6352, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1310].

Patera:2002:EFG

[1253] Jirí Patera, Edita Pelantová, and Milena Svobodová. The eight fine gradings of sl(4, C) and o(6, C). Journal of Mathematical Physics, 43 (12):6353–6378, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garidi:2002:CMV

[1254] T. Garidi, J.-P. Gazeau, and M. V. Takook. Comment on: “ ‘Massive’ vector field in de Sitter space”, [J. Math. Phys. 41, 5920 (2000)]. Journal of Mathematical Physics, 43(12):6379, December 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [345].

Shabani:2002:NSP

[1255] J. Shabani and A. Vyabandi. A note on a series of papers on relativistic δ-sphere interactions in quantum mechanics published by M. N. Hounkonnou and G. Y. H. Avossevou in the Journal of Mathematical Physics. Journal of Mathematical Physics, 43(12):6380–6384, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [16, 103, 104, 496].

Higuchi:2002:EST

[1256] Atsushi Higuchi. Erratum: “Symmetric tensor spherical harmonics on the N-sphere and their application to the de Sitter group SO(N, 1)”, [J. Math. Phys. 28, 1553 (1987)]. Journal of Mathematical Physics, 43 (12):6385, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1].

Bentosela:2003:SWP

[1257] F. Bentosela, Ph. Briet, and L. Pastur. On the spectral and wave propagation properties of the surface Maryland model. Journal of Mathematical Physics, 44(1):1–35, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dehesa:2003:AIE

[1258] J. S. Dehesa, A. Martínez-Finkelshtein, and V. N. Sorokin. Asymptotics of information entropies of some Toda-like potentials. Journal of Mathematical Physics, 44(1):36–47, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garg:2003:SCS

[1259] Anupam Garg, Evgueny Kochetov, Kee-Su Park, and Michael Stone. Spin coherent-state path integrals and the instanton calculus. Journal of Mathematical Physics, 44(1):48–70, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hashimoto:2003:CLT

[1260] Yukihiro Hashimoto, Akihito Hora, and Nobuaki Obata. Central limit theorems for large graphs: Method of quantum decomposition. Journal of Mathematical Physics, 44(1):71–88, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inahama:2003:ESS

[1261] Yuzuru Inahama and Shin ichi Shirai. The essential spectrum of Schrödinger operators with asymptotically constant magnetic fields on the Poincaré upper-half plane. Journal of Mathematical Physics, 44(1): 89–106, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Batista:2003:NGA

[1262] Eliezer Batista and Shahn Majid. Noncommutative geometry of angular momentum space U(SU(∈)). Journal of Mathematical Physics, 44(1): 107–137, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferrari:2003:CSF

[1263] Franco Ferrari. Chern–Simons field theories with non-semisimple gauge group of symmetry. Journal of Mathematical Physics, 44(1):138–145, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grinchuk:2003:PIE

[1264] A. V. Grinchuk and E. A. Ushakov. Path integrals evaluation in twodimensional de Sitter space. Journal of Mathematical Physics, 44(1): 146–155, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marino:2003:TVS

[1265] Marcos Mariño and Ricardo Schiappa. Towards vacuum superstring field theory: The supersliver. Journal of Mathematical Physics, 44(1):156– 187, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pennisi:2003:EFM

[1266] S. Pennisi. An exact fluid model for relativistic electron beams. Journal of Mathematical Physics, 44(1):188–197, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Turgut:2003:TDC

[1267] O. Teoman Turgut. On two dimensional coupled bosons and fermions. Journal of Mathematical Physics, 44(1):198–219, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dassios:2003:MEG

[1268] George Dassios and Fotini Kariotou. Magnetoencephalography in ellipsoidal geometry. Journal of Mathematical Physics, 44(1):220–241, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maccari:2003:INI

[1269] Attilio Maccari. The investigation into new integrable systems of equations in 2 + 1-dimensions. Journal of Mathematical Physics, 44(1):242– 250, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strachan:2003:DMR

[1270] Ian A. B. Strachan. Deformations of the Monge/Riemann hierarchy and approximately integrable systems. Journal of Mathematical Physics, 44 (1):251–262, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bottger:2003:CSH

[1271] Wulf Böttger, Henning Wissowski, and Hans A. Kastrup. Constructing solutions of Hamilton–Jacobi equations for 2D fields with one component by means of Bäcklund transformations. Journal of Mathematical Physics, 44(1):263–301, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnaudon:2003:MPS

[1272] D. Arnaudon, J. Avan, N. Crampé, L. Frappat, and E. Ragoucy. Rmatrix presentation for super-Yangians Y (osp(m|2n)). Journal of Mathematical Physics, 44(1):302–308, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:2003:PBF

[1273] Carl M. Bender, Dorje C. Brody, and Bernhard K. Meister. On powers of Bessel functions. Journal of Mathematical Physics, 44(1):309–314, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeWit:2003:PBW

[1274] David De Wit. A Poincaré–Birkhoff–Witt commutator lemma for U_{q}[gl(m|n)]. Journal of Mathematical Physics, 44(1):315–327, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:2003:CGAa

[1275] A. M. Grundland and W. J. Zakrzewski. On certain geometric aspects of CP^{N} harmonic maps. Journal of Mathematical Physics, 44(1):328–337, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inoue:2003:ELV

[1276] Rei Inoue. The extended Lotka–Volterra lattice and affine Jacobi varieties of spectral curves. Journal of Mathematical Physics, 44(1):338–351, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wurm:2003:FTL

[1277] Alexander Wurm, Nurit Krausz, Cécile DeWitt-Morette, and Marcus Berg. Fourier transforms of Lorentz invariant functions. Journal of Mathematical Physics, 44(1):352–365, January 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Itoh:2003:ECM

[1278] Mitsuhiro Itoh. Erratum: “Contact metric 5-manifolds, CR twistor spaces and integrability,”, [J. Math. Phys. 43, 3783 (2002)]. Journal of Mathematical Physics, 44(1):366–368, January 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1091].

Bruning:2003:SCM

[1279] J. Brüning and V. A. Geyler. Scattering on compact manifolds with infinitely thin horns. Journal of Mathematical Physics, 44(2):371–405, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chadan:2003:BSO

[1280] K. Chadan, N. N. Khuri, A. Martin, and Tai Tsun Wu. Bound states in one and two spatial dimensions. Journal of Mathematical Physics, 44(2): 406–422, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durr:2003:FAS

[1281] D. Dürr and P. Pickl. Flux-across-surfaces theorem for a Dirac particle. Journal of Mathematical Physics, 44(2):423–465, February 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heinonen:2003:CPD

[1282] Teiko Heinonen, Pekka Lahti, and Juha-Pekka Pellonpää. Covariant phase difference observables in quantum mechanics. Journal of Mathematical Physics, 44(2):466–479, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Horowski:2003:SIS

[1283] Maciej Horowski, Anatol Odzijewicz, and Agnieszka Tereszkiewicz. Some integrable systems in nonlinear quantum optics. Journal of Mathematical Physics, 44(2):480–506, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kammerer:2003:WMC

[1284] Clotilde Fermanian Kammerer and Caroline Lasser. Wigner measures and codimension two crossings. Journal of Mathematical Physics, 44(2): 507–527, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leung:2003:CPR

[1285] Debbie W. Leung. Choi’s proof as a recipe for quantum process tomography. Journal of Mathematical Physics, 44(2):528–533, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Havel:2003:RPC

[1286] Timothy F. Havel. Robust procedures for converting among Lindblad, Kraus and matrix representations of quantum dynamical semigroups. Journal of Mathematical Physics, 44(2):534–557, February 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Read:2003:NAB

[1287] N. Read. Non-Abelian braid statistics versus projective permutation statistics. Journal of Mathematical Physics, 44(2):558–563, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Watanabe:2003:CBD

[1288] Teppei Watanabe. A connection between distributivity and locality in compound P-lattices. Journal of Mathematical Physics, 44(2):564–569, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Contino:2003:DRS

[1289] Roberto Contino and Andrea Gambassi. On dimensional regularization of sums. Journal of Mathematical Physics, 44(2):570–587, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DaRold:2003:IFD

[1290] L. Da Rold, C. D. Fosco, and A. López. Interacting fermions and domain wall defects in 2+1 dimensions. Journal of Mathematical Physics, 44(2): 588–598, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Edery:2003:CER

[1291] Ariel Edery. Casimir energy of a relativistic perfect fluid confined to a D-dimensional hypercube. Journal of Mathematical Physics, 44(2): 599–610, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grandou:2003:PMS

[1292] T. Grandou. Proof of a mass singularity free property in high temperature QCD. Journal of Mathematical Physics, 44(2):611–640, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jezierski:2003:GCN

[1293] Jacek Jezierski. Geometry of crossing null shells. Journal of Mathematical Physics, 44(2):641–661, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendy:2003:SSP

[1294] J. E. B. Mendy. Scalar and spin-(1/2) particle creation in gravitational and constant electric field backgrounds. Journal of Mathematical Physics, 44(2):662–675, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kulish:2003:TGM

[1295] P. P. Kulish and N. Manojlovic. Trigonometric osp(1|2) Gaudin model. Journal of Mathematical Physics, 44(2):676–700, February 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qiao:2003:CNE

[1296] Zhijun Qiao, Cewen Cao, and Walter Strampp. Category of nonlinear evolution equations, algebraic structure, and r-matrix. Journal of Mathematical Physics, 44(2):701–722, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bahn:2003:DFS

[1297] Changsoo Bahn, Chul Ki Ko, and Yong Moon Park. Dirichlet forms and symmetric Markovian semigroups on CCR algebras with respect to quasi-free states. Journal of Mathematical Physics, 44(2):723–753, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyallian:2003:FGR

[1298] Carina Boyallian, Victor G. Kac, and Jose I. Liberati. Finite growth representations of infinite Lie conformal algebras. Journal of Mathematical Physics, 44(2):754–770, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Campoamor-Stursberg:2003:ISS

[1299] Rutwig Campoamor-Stursberg. On the invariants of some solvable rigid Lie algebras. Journal of Mathematical Physics, 44(2):771–784, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakrabarti:2003:MBE

[1300] A. Chakrabarti and R. Chakrabarti. Modified braid equations, Baxterizations and noncommutative spaces for the quantum groups GL_{q} ( N ), SO_{q} ( N ), and Sp_{q} ( N ). Journal of Mathematical Physics, 44(2): 785–812, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:2003:CGAb

[1301] A. M. Grundland and W. J. Zakrzewski. On certain geometric aspects of CP^{N} harmonic maps. Journal of Mathematical Physics, 44(2):813–822, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See publisher’s note [1420].

Iorgov:2003:LOH

[1302] N. Z. Iorgov and A. U. Klimyk. A Laplace operator and harmonics on the quantum complex vector space. Journal of Mathematical Physics, 44 (2):823–848, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khachatourian:2003:SRA

[1303] Armik V. M. Khachatourian and Anders O. Wistrom. A sum rule for associated Legendre polynomials with spherical triangles. Journal of Mathematical Physics, 44(2):849–852, February 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1399].

Mazzolo:2003:PDD

[1304] Alain Mazzolo. Probability density distribution of random line segments inside a convex body: Application to random media. Journal of Mathematical Physics, 44(2):853–863, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parashar:2003:IRU

[1305] Preeti Parashar. Induced representations of U_{q}so(5). Journal of Mathematical Physics, 44(2):864–881, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sahadevan:2003:BCB

[1306] R. Sahadevan and S. Khousalya. Belov–Chaltikian and Blaszak– Marciniak lattice equations: Recursion operators and factorization. Journal of Mathematical Physics, 44(2):882–898, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sing:2003:KLP

[1307] Bernd Sing. Kolakoski-(2m, 2n) are limit-periodic model sets. Journal of Mathematical Physics, 44(2):899–912, February 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Topalov:2003:GCI

[1308] Peter Topalov. Geodesic compatibility and integrability of geodesic flows. Journal of Mathematical Physics, 44(2):913–929, February 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wickramasekara:2003:RSR

[1309] S. Wickramasekara and A. Bohm. Representation of semigroups in rigged Hilbert spaces: Subsemigroups of the Weyl–Heisenberg group. Journal of Mathematical Physics, 44(2):930–942, February 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:2003:EPH

[1310] Ali Mostafazadeh. Erratum: “Pseudo-Hermiticity for a class of nondiagonalizable Hamiltonians”, [J. Math. Phys. 43, 6343 (2002)]. Journal of Mathematical Physics, 44(2):943, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1252].

Carvalho:2003:CER

[1311] Túlio O. Carvalho and César R. de Oliveira. Critical energies in random palindrome models. Journal of Mathematical Physics, 44(3):945–961, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cizek:2003:NST

[1312] J. Cízek, J. Zamastil, and L. Skála. New summation technique for rapidly divergent perturbation series. Hydrogen atom in magnetic field. Journal of Mathematical Physics, 44(3):962–968, March 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molnar:2003:PMC

[1313] Lajos Molnár and Werner Timmermann. Preserving the measure of compatibility between quantum states. Journal of Mathematical Physics, 44 (3):969–973, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:2003:PHG

[1314] Ali Mostafazadeh. Pseudo-Hermiticity and generalized P T - and CPT - symmetries. Journal of Mathematical Physics, 44(3):974–989, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shimada:2003:RCS

[1315] Shin ichi Shimada. Resolvent convergence of sphere interactions to point interactions. Journal of Mathematical Physics, 44(3):990–1005, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:2003:GTY

[1316] Stephen C. Anco. Gauge theories of Yang–Mills vector fields coupled to antisymmetric tensor fields. Journal of Mathematical Physics, 44(3): 1006–1043, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brevik:2003:TPP

[1317] Iver Brevik, Andrei A. Bytsenko, and Roger Sollie. Thermodynamic properties of the 2N-piece relativistic string. Journal of Mathematical Physics, 44(3):1044–1055, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cartas-Fuentevilla:2003:AOG

[1318] R. Cartas-Fuentevilla. Adjoint operators, gauge invariant perturbations, and covariant symplectic form for black holes in string theory. Journal of Mathematical Physics, 44(3):1056–1070, March 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marini:2003:SSS

[1319] Antonella Marini and Lorenzo Sadun. Spherically symmetric solutions of a boundary value problem for monopoles. Journal of Mathematical Physics, 44(3):1071–1083, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2003:MCS

[1320] S. Q. Wu and X. Cai. Massive complex scalar field in the Kerr–Sen geometry: Exact solution of wave equation and Hawking radiation. Journal of Mathematical Physics, 44(3):1084–1088, March 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belgiorno:2003:QHT

[1321] F. Belgiorno. Quasi-homogeneous thermodynamics and black holes. Journal of Mathematical Physics, 44(3):1089–1128, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernal:2003:LGN

[1322] Antonio N. Bernal and Miguel Sánchez. Leibnizian, Galilean and Newtonian structures of space–time. Journal of Mathematical Physics, 44(3): 1129–1149, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferapontov:2003:RTH

[1323] E. V. Ferapontov and M. V. Pavlov. Reciprocal transformations of Hamiltonian operators of hydrodynamic type: Nonlocal Hamiltonian formalism for linearly degenerate systems. Journal of Mathematical Physics, 44(3):1150–1172, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seiler:2003:IAP

[1324] Werner M. Seiler. Involution analysis of the partial differential equations characterizing Hamiltonian vector fields. Journal of Mathematical Physics, 44(3):1173–1182, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yurov:2003:DSC

[1325] A. V. Yurov. Discrete symmetry’s chains and links between integrable equations. Journal of Mathematical Physics, 44(3):1183–1201, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Imaikin:2003:SSK

[1326] Valery Imaikin, Alexander Komech, and Peter A. Markowich. Scattering of solitons of the Klein–Gordon equation coupled to a classical particle. Journal of Mathematical Physics, 44(3):1202–1217, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nunez:2003:TED

[1327] Manuel Núñez. Transport of energy in dissipative advection phenomena. Journal of Mathematical Physics, 44(3):1218–1225, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Predescu:2003:PAM

[1328] Cristian Predescu. The partial averaging method. Journal of Mathematical Physics, 44(3):1226–1239, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rubinstein:2003:LGL

[1329] J. Rubinstein and P. Sternberg. Limits of a Ginzburg–Landau model with codimension-one defects. Journal of Mathematical Physics, 44(3): 1240–1251, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Briot:2003:TSY

[1330] C. Briot and E. Ragoucy. Twisted super-Yangians and their representations. Journal of Mathematical Physics, 44(3):1252–1275, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:2003:DYB

[1331] Heng Fan, Boyu Hou, Kangjie Shi, Ruihong Yue, and Shaoyou Zhao. The dynamical Yang–Baxter relation and the minimal representation of the elliptic quantum group. Journal of Mathematical Physics, 44(3): 1276–1296, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fiore:2003:UBT

[1332] Gaetano Fiore, Harold Steinacker, and Julius Wess. Unbraiding the braided tensor product. Journal of Mathematical Physics, 44(3):1297– 1321, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freidel:2003:SNN

[1333] Laurent Freidel and Etera R. Livine. Spin networks for noncompact groups. Journal of Mathematical Physics, 44(3):1322–1356, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:2003:OMF

[1334] Sen-Zhong Huang. An operator method for finding exact solutions to vector Korteweg–de Vries equations. Journal of Mathematical Physics, 44 (3):1357–1388, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:2003:PDA

[1335] Hoil Kim and Chang-Yeong Lee. N-point deformation of algebraic K3 surfaces. Journal of Mathematical Physics, 44(3):1389–1395, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Murata:2003:RSD

[1336] Mikio Murata, Hidetaka Sakai, and Jin Yoneda. Riccati solutions of discrete Painlevé equations with Weyl group symmetry of type E_{8}^{(1)}. Journal of Mathematical Physics, 44(3):1396–1414, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Murugesh:2003:GCL

[1337] S. Murugesh and Radha Balakrishnan. Geometric characterizations of the Lamb equation. Journal of Mathematical Physics, 44(3):1415–1435, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Noakes:2003:NCL

[1338] Lyle Noakes. Null cubics and Lie quadratics. Journal of Mathematical Physics, 44(3):1436–1448, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tarasov:2003:WGI

[1339] V. F. Tarasov. W. Gordon’s integral (1929) and its representations by means of Appell’s functions F_{2}, F_{1}, and F_{3}. Journal of Mathematical Physics, 44(3):1449–1452, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parfitt:2003:ATD

[1340] D. G. W. Parfitt and M. E. Portnoi. Addendum: “The two-dimensional hydrogen atom revisited”, [J. Math. Phys. 43, 4681 (2002)]. Journal of Mathematical Physics, 44(3):1453, March 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abdel-Aty:2003:IEQ

[1341] Mahmoud Abdel-Aty. An investigation of entanglement and quasiprobability distribution in a generalized Jaynes–Cummings model. Journal of Mathematical Physics, 44(4):1457–1471, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alonso:2003:WFC

[1342] Miguel Angel Alonso, George S. Pogosyan, and Kurt Bernardo Wolf. Wigner functions for curved spaces. II. On spheres. Journal of Mathematical Physics, 44(4):1472–1489, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arellano-Balderas:2003:BSP

[1343] Salvador Arellano-Balderas and Valeri V. Kucherenko. On the boson spectrum of the N particle Schrödinger operator with periodic binary potential (dense matter). Journal of Mathematical Physics, 44(4):1490– 1511, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ay:2003:DVD

[1344] Nihat Ay and Wilderich Tuschmann. Duality versus dual flatness in quantum information geometry. Journal of Mathematical Physics, 44 (4):1512–1518, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:2003:MAG

[1345] F. Bagarello. Multiresolution analysis generated by a seed function. Journal of Mathematical Physics, 44(4):1519–1534, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barker:2003:CQS

[1346] Laurence Barker. Continuum quantum systems as limits of discrete quantum systems. IV. Affine canonical transforms. Journal of Mathematical Physics, 44(4):1535–1553, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brau:2003:ULL

[1347] Fabian Brau and Francesco Calogero. Upper and lower limits for the number of S-wave bound states in an attractive potential. Journal of Mathematical Physics, 44(4):1554–1575, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:2003:MAE

[1348] Yves Brihaye and Betti Hartmann. Multiple algebraisations of an elliptic Calogero–Sutherland model. Journal of Mathematical Physics, 44(4): 1576–1583, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gruber:2003:PMS

[1349] Michael J. Gruber. Positive measure spectrum for Schrödinger operators with periodic magnetic fields. Journal of Mathematical Physics, 44(4): 1584–1595, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Olson:2003:WFC

[1350] Jeffrey D. Olson and Jonathan Lee Mace. Wave function confinement via transfer matrix methods. Journal of Mathematical Physics, 44(4):1596– 1624, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samaj:2003:BBQ

[1351] L. Samaj, J. K. Percus, and P. Kalinay. Boundary behavior of quantum Green’s functions. Journal of Mathematical Physics, 44(4):1625–1637, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brzezniak:2003:RCI

[1352] Zdzislaw Brzezniak and Brian Jefferies. Renormalization of Coulomb interactions for the 1D Dirac equation. Journal of Mathematical Physics, 44(4):1638–1659, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jonauskas:2003:GRB

[1353] V. Jonauskas and R. Karazija. General relations between radial integrals in nonrelativistic and relativistic calculation schemes. Journal of Mathematical Physics, 44(4):1660–1665, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ash:2003:MPC

[1354] Avner Ash and Patrick McDonald. Moment problems and the causal set approach to quantum gravity. Journal of Mathematical Physics, 44 (4):1666–1678, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gariel:2003:NAS

[1355] J. Gariel, G. Marcilhacy, and N. O. Santos. A new approach of the stationary axisymmetric vacuum S(A) solutions. Journal of Mathematical Physics, 44(4):1679–1687, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2003:TVH

[1356] Jing-Bo Chen, Han-Ying Guo, and Ke Wu. Total variation in Hamiltonian formalism and symplectic-energy integrators. Journal of Mathematical Physics, 44(4):1688–1702, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalkanli:2003:IKK

[1357] Ayse Karasu Kalkanli, Sergei Yu. Sakovich, and Ísmet Yurdusen. Integrability of Kersten–Krasil’shchik coupled KdV–mKdV equations: singularity analysis and Lax pair. Journal of Mathematical Physics, 44(4): 1703–1708, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Massa:2003:LTA

[1358] Enrico Massa, Enrico Pagani, and Stefano Vignolo. Legendre transformation and analytical mechanics: A geometric approach. Journal of Mathematical Physics, 44(4):1709–1722, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abad:2003:AEQ

[1359] J. Abad and J. Sesma. Asymptotic expansion of the quasiconfluent hypergeometric function. Journal of Mathematical Physics, 44(4):1723–1729, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arik:2003:NPU

[1360] M. Arik and B. T. Kaynak. Noncommutative phase and unitarization of GL_{p,q}(2). Journal of Mathematical Physics, 44(4):1730–1734, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benedetto:2003:BLT

[1361] John J. Benedetto, Wojciech Czaja, and Andrei Ya. Maltsev. The Balian–Low theorem for the symplectic form on R^{2d}. Journal of Mathematical Physics, 44(4):1735–1750, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakravarty:2003:FIG

[1362] S. Chakravarty and R. Halburd. First integrals of a generalized Darboux– Halphen system. Journal of Mathematical Physics, 44(4):1751–1762, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cieslinski:2003:DCE

[1363] Jan L. Cieslinski, Marek Czachor, and Nikolai V. Ustinov. Darboux covariant equations of von Neumann type and their generalizations. Journal of Mathematical Physics, 44(4):1763–1780, April 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakis:2003:DGG

[1364] Aristophanes Dimakis and Folkert Müller-Hoissen. Differential geometry of group lattices. Journal of Mathematical Physics, 44(4):1781–1821, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:2003:CIJ

[1365] Avinash Khare, Arul Lakshminarayan, and Uday Sukhatme. Cyclic identities for Jacobi elliptic and related functions. Journal of Mathematical Physics, 44(4):1822–1841, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Najar:2003:LTR

[1366] Hatem Najar. Lifshitz tails for random acoustic operators. Journal of Mathematical Physics, 44(4):1842–1867, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rasmussen:2003:PHG

[1367] Jørgen Rasmussen. N -point and higher-genus osp(1|2) fusion. Journal of Mathematical Physics, 44(4):1868–1881, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vasilchuk:2003:ADC

[1368] V. Vasilchuk. On the asymptotic distribution of the commutator and anticommutator of random matrices. Journal of Mathematical Physics, 44 (4):1882–1908, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hioe:2003:ECN

[1369] F. T. Hioe. Erratum: “N coupled nonlinear Schrödinger equations: Special set and applications to N = 3”, [J. Math. Phys. 43, 6325 (2002)]. Journal of Mathematical Physics, 44(4):1909, April 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1250].

Aunola:2003:DHO

[1370] M. Aunola. The discretized harmonic oscillator: Mathieu functions and a new class of generalized Hermite polynomials. Journal of Mathematical Physics, 44(5):1913–1936, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barata:2003:PAD

[1371] J. C. A. Barata and D. A. Cortez. Perturbative analysis of dynamical localization. Journal of Mathematical Physics, 44(5):1937–1960, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2003:IBE

[1372] Thomas Chen, Vitali Vougalter, and Semjon A. Vugalter. The increase of binding energy and enhanced binding in nonrelativistic QED. Journal of Mathematical Physics, 44(5):1961–1970, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujiie:2003:BWF

[1373] Setsuro Fujiié and Thierry Ramond. Breit–Wigner formula at barrier tops. Journal of Mathematical Physics, 44(5):1971–1983, May 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giachetta:2003:BHP

[1374] G. Giachetta, L. Mangiarotti, and G. Sardanashvily. Bi-Hamiltonian partially integrable systems. Journal of Mathematical Physics, 44(5): 1984–1997, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heinonen:2003:NPQ

[1375] Teiko Heinonen, Pekka Lahti, Juha-Pekka Pellonpää, Sylvia Pulmannova, and Kari Ylinen. The norm-1-property of a quantum observable. Journal of Mathematical Physics, 44(5):1998–2008, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:2003:RPI

[1376] Yisong Yang. Rigorous proof of isotope effect by Bardeen–Cooper– Schrieffer theory. Journal of Mathematical Physics, 44(5):2009–2025, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caliceti:2003:DBS

[1377] Emanuela Caliceti. Distributional Borel summability for vacuum polarization by an external electric field. Journal of Mathematical Physics, 44 (5):2026–2036, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mund:2003:MLM

[1378] Jens Mund. Modular localization of massive particles with “any” spin in d = 2 + 1. Journal of Mathematical Physics, 44(5):2037–2057, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:2003:TET

[1379] A. G. Smirnov. Towards Euclidean theory of infrared singular quantum fields. Journal of Mathematical Physics, 44(5):2058–2076, May 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Izumiya:2003:SFM

[1380] S. Izumiya and S. Janeczko. A symplectic framework for multiplane gravitational lensing. Journal of Mathematical Physics, 44(5):2077–2093, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Robinson:2003:GFE

[1381] D. C. Robinson. Generalized forms and Einstein’s equations. Journal of Mathematical Physics, 44(5):2094–2110, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marcelli:2003:LPS

[1382] M. Marcelli and M. C. Nucci. Lie point symmetries and first integrals: The Kowalevski top. Journal of Mathematical Physics, 44(5):2111–2132, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dacol:2003:WSW

[1383] Dalcio K. Dacol and Dilip G. Roy. Wave scattering in waveguides. Journal of Mathematical Physics, 44(5):2133–2148, May 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:2003:HOT

[1384] Manuel F. Rañada and Mariano Santander. On harmonic oscillators on the two-dimensional sphere S^{2} and the hyperbolic plane H^{2}. II. Journal of Mathematical Physics, 44(5):2149–2167, May 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yoshida:2003:SBF

[1385] Z. Yoshida, S. Ohsaki, A. Ito, and S. M. Mahajan. Stability of Beltrami flows. Journal of Mathematical Physics, 44(5):2168–2178, May 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenzi:2003:ADF

[1386] E. K. Lenzi, R. S. Mendes, Kwok Sau Fa, L. C. Malacarne, and L. R. da Silva. Anomalous diffusion: Fractional Fokker–Planck equation and its solutions. Journal of Mathematical Physics, 44(5):2179–2185, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bakic:2003:WTC

[1387] Damir Bakic and Boris Guljas. Wigner’s theorem in a class of Hilbert C^{∗}-modules. Journal of Mathematical Physics, 44(5):2186–2191, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyallian:2003:RST

[1388] Carina Boyallian and José I. Liberati. Representations of a symplectic type subalgebra of W_{∞}. Journal of Mathematical Physics, 44(5):2192– 2205, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brouwer:2003:BAM

[1389] R. M. Brouwer. A bicategorical approach to Morita equivalence for von Neumann algebras. Journal of Mathematical Physics, 44(5):2206–2214, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carvalho:2003:NRT

[1390] M. Carvalho and A. L. Oliveira. A new realization for the d = 2 topological algebra. Journal of Mathematical Physics, 44(5):2215–2223, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeLeo:2003:SSQ

[1391] Stefano De Leo and Gisele C. Ducati. Solving simple quaternionic differential equations. Journal of Mathematical Physics, 44(5):2224–2233, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doliwa:2003:GDK

[1392] Adam Doliwa. Geometric discretization of the Koenigs nets. Journal of Mathematical Physics, 44(5):2234–2249, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durand:2003:FOSa

[1393] Loyal Durand. Fractional operators and special functions. I. Bessel functions. Journal of Mathematical Physics, 44(5):2250–2265, May 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Durand:2003:FOSb

[1394] Loyal Durand. Fractional operators and special functions. II. Legendre functions. Journal of Mathematical Physics, 44(5):2266–2292, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geronimo:2003:CIA

[1395] Jeffrey S. Geronimo and Henri Navelet. On certain integrals that appear in conformal field theory. Journal of Mathematical Physics, 44(5):2293– 2319, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khusnutdinov:2003:UAE

[1396] Nail R. Khusnutdinov. On the uniform asymptotic expansion of the Legendre functions. Journal of Mathematical Physics, 44(5):2320–2330, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kobayashi:2003:TSI

[1397] M. Kobayashi. Toward solving the inhomogeneous Bloch equation. Journal of Mathematical Physics, 44(5):2331–2341, May 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leznov:2003:DTM

[1398] A. N. Leznov and R. Torres-Cordoba. Discrete transformation for the matrix three-wave problem in three dimensional space. Journal of Mathematical Physics, 44(5):2342–2352, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khachatourian:2003:ESR

[1399] Armik V. M. Khachatourian and Anders O. Wistrom. Erratum: “A sum rule for associated Legendre polynomials with spherical triangles”, [J. Math. Phys. 44, 849 (2003)]. Journal of Mathematical Physics, 44(5): 2353, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1303].

Romera:2003:EHE

[1400] E. Romera, J. C. Angulo, and J. S. Dehesa. Erratum: “The Hausdorff entropic moment problem”, [J. Math. Phys. 42, 2309 (2001)]. Journal of Mathematical Physics, 44(5):2354, May 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [635].

Altafini:2003:CPF

[1401] Claudio Altafini. Controllability properties for finite dimensional quantum Markovian master equations. Journal of Mathematical Physics, 44 (6):2357–2372, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baro:2003:DST

[1402] M. Baro and H. Neidhardt. Dissipative Schrödinger-type operators as a model for generation and recombination. Journal of Mathematical Physics, 44(6):2373–2401, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benatti:2003:BSE

[1403] Fabio Benatti, H. Narnhofer, and A. Uhlmann. Broken symmetries in the entanglement of formation. Journal of Mathematical Physics, 44 (6):2402–2415, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DelCastillo:2003:LBS

[1404] Pierre Del Castillo. Lower bound for the superheating field in the weak-κ limit: The general case. Journal of Mathematical Physics, 44(6):2416– 2450, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inahama:2003:SPP

[1405] Yuzuru Inahama and Shin ichi Shirai. Spectral properties of Pauli operators on the Poincaré upper-half plane. Journal of Mathematical Physics, 44(6):2451–2462, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karle:2003:GPG

[1406] Regina Karle and Jiannis Pachos. Geometrical phases for the G (4,2) Grassmannian manifold. Journal of Mathematical Physics, 44(6):2463– 2470, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khrennikov:2003:CVQ

[1407] Andrei Khrennikov. Contextual viewpoint to quantum stochastics. Journal of Mathematical Physics, 44(6):2471–2478, June 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mukunda:2003:GCS

[1408] N. Mukunda, Arvind, S. Chaturvedi, and R. Simon. Generalized coherent states and the diagonal representation for operators. Journal of Mathematical Physics, 44(6):2479–2506, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belhaj:2003:NOK

[1409] A. Belhaj, J. J. Manjarín, and P. Resco. On noncommutative orbifolds of K3 surfaces. Journal of Mathematical Physics, 44(6):2507–2520, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salmela:2003:AMS

[1410] Antti Salmela. An algebraic method for solving the SU(3) Gauss law. Journal of Mathematical Physics, 44(6):2521–2533, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chamseddine:2003:IAN

[1411] A. H. Chamseddine. An invariant action for noncommutative gravity in four dimensions. Journal of Mathematical Physics, 44(6):2534–2541, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimitrov:2003:CAE

[1412] Bogdan G. Dimitrov. Cubic algebraic equations in gravity theory, parametrization with the Weierstrass function and nonarithmetic theory of algebraic equations. Journal of Mathematical Physics, 44(6):2542– 2578, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cerkaski:2003:APM

[1413] M. Cerkaski. Application of Poisson maps on coadjoint orbits of Sp(6, R) group to many body dynamics. Journal of Mathematical Physics, 44(6): 2579–2595, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dudnikova:2003:CSE

[1414] T. V. Dudnikova, A. I. Komech, and H. Spohn. On the convergence to statistical equilibrium for harmonic crystals. Journal of Mathematical Physics, 44(6):2596–2620, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golovko:2003:HER

[1415] V. A. Golovko. Hierarchy of equations for reduced density matrices at thermodynamic equilibrium with account taken of the spin of particles. Journal of Mathematical Physics, 44(6):2621–2638, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:2003:SNC

[1416] Xing-Bin Pan. Superconductivity near critical temperature. Journal of Mathematical Physics, 44(6):2639–2678, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Conidis:2003:TOC

[1417] Chris Conidis and Joe Repka. Tensor operators and constructing indecomposable representations of semidirect product groups. Journal of Mathematical Physics, 44(6):2679–2691, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Curtright:2003:EDN

[1418] Thomas Curtright and David Fairlie. Extra dimensions and nonlinear equations. Journal of Mathematical Physics, 44(6):2692–2703, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:2003:WTP

[1419] A. M. Grundland, P. Tempesta, and P. Winternitz. Weak transversality and partially invariant solutions. Journal of Mathematical Physics, 44 (6):2704–2722, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:2003:PNC

[1420] A. M. Grundland and W. J. Zakrzewski. Publisher’s note: “On certain geometric aspects of CP^{N} harmonic maps”, [J. Math. Phys. 44, 813 (2003)]. Journal of Mathematical Physics, 44(6):2723, June 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1301].

Hall:2003:ERB

[1421] Richard L. Hall, Wolfgang Lucha, and Franz F. Schöberl. Erratum: “Relativistic N-boson systems bound by oscillator pair potentials”, [J. Math. Phys. 43, 1237 (2002)]. Journal of Mathematical Physics, 44(6): 2724–2725, June 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [942].

Auberson:2003:BIF

[1422] G. Auberson, G. Mahoux, S. M. Roy, and Virendra Singh. Bell inequalities in four dimensional phase space and the three marginal theorem. Journal of Mathematical Physics, 44(7):2729–2747, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belov:2003:HOT

[1423] S. M. Belov and A. V. Rybkin. Higher order trace formulas of the Buslaev–Faddeev-type for the half-line Schrödinger operator with longrange potentials. Journal of Mathematical Physics, 44(7):2748–2761, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roux:2003:SMS

[1424] Ph. Roux and D. Yafaev. The scattering matrix for the Schrödinger operator with a long-range electromagnetic potential. Journal of Mathematical Physics, 44(7):2762–2786, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruan:2003:GRS

[1425] Dong Ruan and Weicheng Huang. General realization of N = 4 supersymmetric quantum mechanics and its applications. Journal of Mathematical Physics, 44(7):2787–2805, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Toloza:2003:EAE

[1426] J. H. Toloza. Exponentially accurate error estimates of quasiclassical eigenvalues. II. Several dimensions. Journal of Mathematical Physics, 44 (7):2806–2838, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dmitrasinovic:2003:DSB

[1427] V. Dmitrasinovic and Issei Nakamura. Dynamical symmetry breaking and the Nambu–Goldstone theorem in the Gaussian wave functional approximation. Journal of Mathematical Physics, 44(7):2839–2852, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doran:2003:PTS

[1428] Charles F. Doran and Michael Faux. A “periodic table” for supersymmetric M-theory compactifications. Journal of Mathematical Physics, 44 (7):2853–2873, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:2003:GIS

[1429] A. M. Grundland, A. J. Hariton, and V. Hussin. Group-invariant solutions of relativistic and nonrelativistic models in field theory. Journal of Mathematical Physics, 44(7):2874–2890, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pfeiffer:2003:EDT

[1430] Hendryk Pfeiffer. Exact duality transformations for sigma models and gauge theories. Journal of Mathematical Physics, 44(7):2891–2938, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakagawa:2003:ERG

[1431] Hironobu Sakagawa. Entropic repulsion for a Gaussian lattice field with certain finite range interaction. Journal of Mathematical Physics, 44 (7):2939–2951, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silenko:2003:FWT

[1432] Alexander J. Silenko. Foldy–Wouthuysen transformation for relativistic particles in external fields. Journal of Mathematical Physics, 44(7):2952– 2966, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tamura:2003:RCN

[1433] Hideo Tamura. Resolvent convergence in norm for Dirac operator with Aharonov–Bohm field. Journal of Mathematical Physics, 44(7):2967– 2993, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anderson:2003:RLM

[1434] Michael T. Anderson. Regularity for Lorentz metrics under curvature bounds. Journal of Mathematical Physics, 44(7):2994–3012, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dolan:2003:EDS

[1435] P. Dolan and A. Gerber. Exterior differential systems, Janet–Riquier theory and the Riemann–Lanczos problems in two, three, and four dimensions. Journal of Mathematical Physics, 44(7):3013–3034, July 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dolan:2003:WLE

[1436] P. Dolan and A. Gerber. The Weyl–Lanczos equations and the Lanczos wave equation in four dimensions as systems in involution. Journal of Mathematical Physics, 44(7):3035–3045, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krisch:2003:STT

[1437] J. P. Krisch and E. N. Glass. A space–time in toroidal coordinates. Journal of Mathematical Physics, 44(7):3046–3058, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gerber:2003:PMC

[1438] A. Gerber. Prolongation methods and Cartan characters for the three-dimensional Riemann–Lanczos problem. Journal of Mathematical Physics, 44(7):3059–3070, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McLachlan:2003:AEC

[1439] Robert I. McLachlan and Brett Ryland. The algebraic entropy of classical mechanics. Journal of Mathematical Physics, 44(7):3071–3087, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foursov:2003:TCC

[1440] Mikhail V. Foursov. Towards the complete classification of homogeneous two-component integrable equations. Journal of Mathematical Physics, 44(7):3088–3096, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gunther:2003:ISM

[1441] Uwe Günther and Frank Stefani. Isospectrality of spherical MHD dynamo operators: Pseudo-Hermiticity and a no-go theorem. Journal of Mathematical Physics, 44(7):3097–3111, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kobayashi:2003:SSI

[1442] M. Kobayashi. Special solution of the inhomogeneous Bloch equation. Journal of Mathematical Physics, 44(7):3112–3122, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koca:2003:QRR

[1443] Mehmet Koca, Ramazan Koç, and Muataz Al-Barwani. Quaternionic roots of SO(8), SO(9), F_{4} and the related Weyl groups. Journal of Mathematical Physics, 44(7):3123–3140, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vignolo:2003:GDR

[1444] Stefano Vignolo. A geometric description of Routh’s procedure. Addendum: “Legendre transformation and analytical mechanics: A geometric approach”, [J. Math. Phys. 44, 1709 (2003)]. Journal of Mathematical Physics, 44(7):3141–3144, July 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ablowitz:2003:ISR

[1445] M. J. Ablowitz, S. Chakravarty, and R. G. Halburd. Integrable systems and reductions of the self-dual Yang–Mills equations. Journal of Mathematical Physics, 44(8):3147–3173, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grinevich:2003:TPR

[1446] P. G. Grinevich and S. P. Novikov. Topological phenomena in the real periodic sine-Gordon theory. Journal of Mathematical Physics, 44(8): 3174–3184, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

BoutetdeMonvel:2003:GAS

[1447] Anne Boutet de Monvel and Vladimir Kotlyarov. Generation of asymptotic solitons of the nonlinear Schrödinger equation by boundary data. Journal of Mathematical Physics, 44(8):3185–3215, August 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sabatier:2003:LES

[1448] P. C. Sabatier. Lax equations scattering and KdV. Journal of Mathematical Physics, 44(8):3216–3225, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fokas:2003:DSE

[1449] A. S. Fokas. The Davey–Stewartson I equation on the quarter plane with homogeneous Dirichlet boundary conditions. Journal of Mathematical Physics, 44(8):3226–3244, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kac:2003:CKH

[1450] V. G. Kac and J. W. van de Leur. The n-component KP hierarchy and representation theory. Journal of Mathematical Physics, 44(8):3245– 3293, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alonso:2003:ASS

[1451] Luis Martínez Alonso and Manuel Mañas. Additional symmetries and solutions of the dispersionless KP hierarchy. Journal of Mathematical Physics, 44(8):3294–3308, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boiti:2003:ERI

[1452] M. Boiti, F. Pempinelli, A. K. Pogrebkov, and B. Prinari. Extended resolvent and inverse scattering with an application to KPI. Journal of Mathematical Physics, 44(8):3309–3340, August 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rogers:2003:NIR

[1453] C. Rogers and W. K. Schief. Novel integrable reductions in nonlinear continuum mechanics via geometric constraints. Journal of Mathematical Physics, 44(8):3341–3369, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:2003:HMW

[1454] A. M. Grundland and W. J. Zakrzewski. CP^{N−1} harmonic maps and the Weierstrass problem. Journal of Mathematical Physics, 44(8):3370–3382, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dancer:2003:PAR

[1455] Andrew Dancer and McKenzie Y. Wang. Painlevé analysis of the Ricciflat ordinary differential equations associated with Aloff–Wallach spaces and U(1)-bundles over Fano products. Journal of Mathematical Physics, 44(8):3383–3406, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Neugebauer:2003:PRG

[1456] G. Neugebauer and R. Meinel. Progress in relativistic gravitational theory using the inverse scattering method. Journal of Mathematical Physics, 44(8):3407–3429, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunajski:2003:TTH

[1457] Maciej Dunajski and Lionel J. Mason. Twistor theory of hyper-Kähler metrics with hidden symmetries. Journal of Mathematical Physics, 44(8): 3430–3454, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agafonov:2003:HCP

[1458] S. I. Agafonov and A. I. Bobenko. Hexagonal circle patterns with constant intersection angles and discrete Painlevé and Riccati equations.

Journal of Mathematical Physics, 44(8):3455–3469, August 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haskins:2003:GAL

[1459] M. Haskins and J. M. Speight. The geodesic approximation for lump dynamics and coercivity of the Hessian for harmonic maps. Journal of Mathematical Physics, 44(8):3470–3494, August 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baptista:2003:DVN

[1460] J. M. Baptista and N. S. Manton. The dynamics of vortices on S^{2} near the Bradlow limit. Journal of Mathematical Physics, 44(8):3495–3508, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tong:2003:MSN

[1461] David Tong. The moduli space of noncommutative vortices. Journal of Mathematical Physics, 44(8):3509–3516, August 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Murray:2003:NMM

[1462] Michael K. Murray and Michael A. Singer. A note on monopole moduli spaces. Journal of Mathematical Physics, 44(8):3517–3531, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Battye:2003:PSF

[1463] Richard A. Battye, Gary W. Gibbons, Paulina Rychenkova, and Paul M. Sutcliffe. Polyhedral scattering of fundamental monopoles. Journal of Mathematical Physics, 44(8):3532–3542, August 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Battye:2003:IS

[1464] Richard A. Battye, Conor J. Houghton, and Paul M. Sutcliffe. Icosahedral Skyrmions. Journal of Mathematical Physics, 44(8):3543–3554, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ward:2003:TQS

[1465] R. S. Ward. Topological Q-solitons. Journal of Mathematical Physics, 44 (8):3555–3561, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioannidou:2003:QBP

[1466] T. A. Ioannidou and N. D. Vlachos. On the Q-ball profile function. Journal of Mathematical Physics, 44(8):3562–3568, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hikami:2003:TQI

[1467] Kazuhiro Hikami and Miki Wadati. Topics in quantum integrable systems. Journal of Mathematical Physics, 44(8):3569–3594, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mezincescu:2003:FMS

[1468] G. Andrei Mezincescu and Luca Mezincescu. Factorization method and the supersymmetric monopole harmonics. Journal of Mathematical Physics, 44(8):3595–3606, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goldhaber:2003:FCD

[1469] Alfred Scharff Goldhaber. Fractional charge definitions and conditions. Journal of Mathematical Physics, 44(8):3607–3618, August 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klinkhamer:2003:SSF

[1470] Frans R. Klinkhamer and Christian Rupp. Sphalerons, spectral flow, and anomalies. Journal of Mathematical Physics, 44(8):3619–3639, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davies:2003:MAA

[1471] N. Michael Davies, Timothy J. Hollowood, and Valentin V. Khoze. Monopoles, affine algebras and the gluino condensate. Journal of Mathematical Physics, 44(8):3640–3656, August 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baik:2003:PRC

[1472] Jinho Baik, Percy Deift, and Eugene Strahov. Products and ratios of characteristic polynomials of random Hermitian matrices. Journal of Mathematical Physics, 44(8):3657–3670, August 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andai:2003:MRM

[1473] Attila Andai. Monotone Riemannian metrics on density matrices with non-monotone scalar curvature. Journal of Mathematical Physics, 44(9): 3675–3688, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brizhik:2003:SLE

[1474] L. S. Brizhik, A. A. Eremko, B. M. A. G. Piette, and W. J. Zakrzewski. Spontaneous localization of electrons in two-dimensional lattices within the adiabatic approximation. Journal of Mathematical Physics, 44(9): 3689–3697, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsko:2003:IHH

[1475] Andrei G. Bytsko. On integrable Hamiltonians for higher spin XXZ chain. Journal of Mathematical Physics, 44(9):3698–3717, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chruscinski:2003:QMD

[1476] Dariusz Chruściński. Quantum mechanics of damped systems. Journal of Mathematical Physics, 44(9):3718–3733, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferrari:2003:EES

[1477] Christian Ferrari and Nicolas Macris. Extended edge states in finite Hall systems. Journal of Mathematical Physics, 44(9):3734–3751, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gibilisco:2003:WYI

[1478] Paolo Gibilisco and Tommaso Isola. Wigner–Yanase information on quantum state space: The geometric approach. Journal of Mathematical Physics, 44(9):3752–3762, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gu:2003:QFB

[1479] Xiao-Yan Gu, Zhong-Qi Ma, and Jian-Qiang Sun. Quantum four-body system in D dimensions. Journal of Mathematical Physics, 44(9):3763– 3774, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okubo:2003:GST

[1480] Susumu Okubo. Global symmetries of time-dependent Schrödinger equations. Journal of Mathematical Physics, 44(9):3775–3799, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakrabarti:2003:NTM

[1481] A. Chakrabarti. Nonlinear transforms of momenta and Planck scale limit. Journal of Mathematical Physics, 44(9):3800–3808, September 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coquereaux:2003:DQS

[1482] R. Coquereaux and G. Schieber. Determination of quantum symmetries for higher ADE systems from the modular T matrix. Journal of Mathematical Physics, 44(9):3809–3837, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garidi:2003:MSF

[1483] T. Garidi, J-P. Gazeau, and M. V. Takook. “massive” spin-2 field in de Sitter space. Journal of Mathematical Physics, 44(9):3838–3862, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lazzarini:2003:IEG

[1484] Serge Lazzarini and José M. Gracia-Bondía. Improved Epstein–Glaser renormalization. II. Lorentz invariant framework. Journal of Mathematical Physics, 44(9):3863–3875, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miskovic:2003:DSI

[1485] Olivera Miskovic and Jorge Zanelli. Dynamical structure of irregular constrained systems. Journal of Mathematical Physics, 44(9):3876–3887, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calabrese:2003:DIP

[1486] Gioel Calabrese and Olivier Sarbach. Detecting ill posed boundary conditions in general relativity. Journal of Mathematical Physics, 44(9): 3888–3899, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miritzis:2003:DSA

[1487] John Miritzis. Dynamical system approach to FRW models in higherorder gravity theories. Journal of Mathematical Physics, 44(9):3900– 3910, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Georgieva:2003:GVP

[1488] Bogdana Georgieva, Ronald Guenther, and Theodore Bodurov. Generalized variational principle of Herglotz for several independent variables. First Noether-type theorem. Journal of Mathematical Physics, 44 (9):3911–3927, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:2003:TCG

[1489] Jingsong He, Yishen Li, and Yi Cheng. Two choices of the gauge transformation for the AKNS hierarchy through the constrained KP hierarchy. Journal of Mathematical Physics, 44(9):3928–3960, September 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sahadevan:2003:THF

[1490] R. Sahadevan and S. Khousalya. Tri-Hamiltonian formulation for certain integrable lattice equations. Journal of Mathematical Physics, 44(9): 3961–3978, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takasaki:2003:SCH

[1491] Kanehisa Takasaki. Spectral curve and Hamiltonian structure of isomonodromic SU(2) Calogero–Gaudin system. Journal of Mathematical Physics, 44(9):3979–3999, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tang:2003:EMV

[1492] Xiao-Yan Tang and Sen-Yue Lou. Extended multilinear variable separation approach and multivalued localized excitations for some (2 + 1)- dimensional integrable systems. Journal of Mathematical Physics, 44(9): 4000–4025, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Afanasiev:2003:VCT

[1493] G. N. Afanasiev, V. G. Kartavenko, and Yu. P. Stepanovsky. Vavilov– Cherenkov and transition radiations on the dielectric and metallic spheres. Journal of Mathematical Physics, 44(9):4026–4056, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Corngold:2003:CTC

[1494] Noel Corngold. Classical transport of charged particles in a magnetic field. Journal of Mathematical Physics, 44(9):4057–4077, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jimenez-Lara:2003:TBP

[1495] Lidia Jiménez-Lara and Eduardo Piña. The three-body problem with an inverse square law potential. Journal of Mathematical Physics, 44 (9):4078–4089, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leach:2003:EBC

[1496] P. G. L. Leach, K. Andriopoulos, and M. C. Nucci. The Ermanno– Bernoulli constants and representations of the complete symmetry group of the Kepler problem. Journal of Mathematical Physics, 44(9):4090– 4106, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nucci:2003:LIS

[1497] M. C. Nucci. Lorenz integrable system moves à la Poinsot. Journal of Mathematical Physics, 44(9):4107–4118, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Emamizadeh:2003:ABC

[1498] B. Emamizadeh and M. H. Mehrabi. Asymptotic behavior of the centroids of a family of vortices. Journal of Mathematical Physics, 44(9): 4119–4133, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pavlov:2003:IHC

[1499] Maxim V. Pavlov. Integrable hydrodynamic chains. Journal of Mathematical Physics, 44(9):4134–4156, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:2003:RMA

[1500] P. J. Forrester, N. E. Frankel, and T. M. Garoni. Random matrix averages and the impenetrable Bose gas in Dirichlet and Neumann boundary conditions. Journal of Mathematical Physics, 44(9):4157–4175, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2003:EPF

[1501] Da-Quan Jiang, Min Qian, and Fu-Xi Zhang. Entropy production fluctuations of finite Markov chains. Journal of Mathematical Physics, 44 (9):4176–4188, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kato:2003:DCF

[1502] Go Kato and Miki Wadati. A direct calculation of the free energy from the Bethe ansatz equation for the Heisenberg model. Journal of Mathematical Physics, 44(9):4189–4200, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chillingworth:2003:SPF

[1503] David Chillingworth and Martin Golubitsky. Symmetry and pattern formation for a planar layer of nematic liquid crystal. Journal of Mathematical Physics, 44(9):4201–4219, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakis:2003:RGB

[1504] Aristophanes Dimakis and Folkert Müller-Hoissen. Riemannian geometry of bicovariant group lattices. Journal of Mathematical Physics, 44(9): 4220–4259, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fashandi:2003:CWT

[1505] M. Fashandi, R. A. Kamyabi Gol, A. Niknam, and M. A. Pourabdollah. Continuous wavelet transform on a special homogeneous space. Journal of Mathematical Physics, 44(9):4260–4266, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guhr:2003:AGT

[1506] Thomas Guhr and Heiner Kohler. Angular Gelfand–Tzetlin coordinates for the supergroup UOSp(k_{1}/2k_{2}). Journal of Mathematical Physics, 44 (9):4267–4286, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hofmann:2003:CSP

[1507] Gerald Hofmann. On the construction of self-polar and self-polar Hilbertian norms on Minkowski space of dimension 2. Journal of Mathematical Physics, 44(9):4287–4307, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Malmendier:2003:EEE

[1508] Andreas Malmendier. The eigenvalue equation on the Eguchi–Hanson space. Journal of Mathematical Physics, 44(9):4308–4343, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shklyarov:2003:BTQ

[1509] Dmitry Shklyarov and Genkai Zhang. Berezin transform on the quantum unit ball. Journal of Mathematical Physics, 44(9):4344–4373, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:2003:ACC

[1510] G. F. Torres del Castillo and L. F. Gómez-Ceballos. Algebraic classification of the curvature of three-dimensional manifolds with indefinite metric. Journal of Mathematical Physics, 44(9):4374–4380, September 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borg:2003:PAS

[1511] J. L. Borg and J. V. Pulé. Pauli approximations to the self-adjoint extensions of the Aharonov–Bohm Hamiltonian. Journal of Mathematical Physics, 44(10):4385–4410, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwai:2003:SRM

[1512] Toshihiro Iwai and Hidetaka Yamaoka. Stratified reduction of manybody kinetic energy operators. Journal of Mathematical Physics, 44(10): 4411–4435, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Richard:2003:DDC

[1513] Christoph Richard. Dense Dirac combs in Euclidean space with pure point diffraction. Journal of Mathematical Physics, 44(10):4436–4449, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scolarici:2003:PHN

[1514] G. Scolarici and L. Solombrino. On the pseudo-Hermitian nondiagonalizable Hamiltonians. Journal of Mathematical Physics, 44(10):4450–4459, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silbergleit:2003:PFS

[1515] Alexander Silbergleit, Ilya Mandel, and Ilya Nemenman. Potential and field singularity at a surface point charge. Journal of Mathematical Physics, 44(10):4460–4466, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dong:2003:GTA

[1516] Shi-Hai Dong, Guo-Hua Sun, and Dusan Popov. Group theory approach to the Dirac equation with a Coulomb plus scalar potential in D + 1 dimensions. Journal of Mathematical Physics, 44(10):4467–4479, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fewster:2003:QWE

[1517] Christopher J. Fewster and Michael J. Pfenning. A quantum weak energy inequality for spin-one fields in curved space–time. Journal of Mathematical Physics, 44(10):4480–4513, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ochiai:2003:DDP

[1518] Tomoshiro Ochiai. A direct derivation of polynomial invariants from perturbative Chern–Simons gauge theory. Journal of Mathematical Physics, 44(10):4514–4526, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Popov:2003:SWM

[1519] Alexander D. Popov, Armen G. Sergeev, and Martin Wolf. Seiberg– Witten monopole equations on noncommutative R^{4}. Journal of Mathematical Physics, 44(10):4527–4554, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sparber:2003:SAM

[1520] C. Sparber and P. Markowich. Semiclassical asymptotics for the Maxwell–Dirac system. Journal of Mathematical Physics, 44(10):4555– 4572, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1944].

Geng:2003:AGC

[1521] Xianguo Geng, H. H. Dai, and Cewen Cao. Algebro–geometric constructions of the discrete Ablowitz–Ladik flows and applications. Journal of Mathematical Physics, 44(10):4573–4588, October 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kundu:2003:USG

[1522] Anjan Kundu. Unifying scheme for generating discrete integrable systems including inhomogeneous and hybrid models. Journal of Mathematical Physics, 44(10):4589–4603, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shchesnovich:2003:GSM

[1523] Valery S. Shchesnovich and Jianke Yang. General soliton matrices in the Riemann–Hilbert problem for integrable nonlinear equations. Journal of Mathematical Physics, 44(10):4604–4639, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:2003:IDZ

[1524] Derchyi Wu. Isomonodromy deformations for the ZS-AKNS system with quadratic spectral variables. Journal of Mathematical Physics, 44(10): 4640–4651, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bustamante:2003:MLH

[1525] Miguel D. Bustamante and Sergio A. Hojman. Multi-Lagrangians, hereditary operators and Lax pairs for the Korteweg–de Vries positive and negative hierarchies. Journal of Mathematical Physics, 44(10):4652–4671, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurses:2003:LWP

[1526] Metin Gürses and Özgür Sarioğlu. Liénard–Wiechert potentials in even dimensions. Journal of Mathematical Physics, 44(10):4672–4680, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jiang:2003:GKF

[1527] Da-Quan Jiang and Fu-Xi Zhang. The Green–Kubo formula and power spectrum of reversible Markov processes. Journal of Mathematical Physics, 44(10):4681–4689, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:2003:SMB

[1528] Huan-Qiang Zhou, Jon Links, Mark D. Gould, and Ross H. McKenzie. Solvable models of Bose–Einstein condensates: A new algebraic Bethe ansatz scheme. Journal of Mathematical Physics, 44(10):4690–4701, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Almeida:2003:VCI

[1529] R. Almeida, A. Kumpera, and J. L. Rubin. On the variational cohomology of g-invariant foliations. Journal of Mathematical Physics, 44(10): 4702–4712, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Balachandran:2003:DOC

[1530] A. P. Balachandran, Giorgio Immirzi, Joohan Lee, and Peter Presnajder. Dirac operators on coset spaces. Journal of Mathematical Physics, 44(10):4713–4735, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blohmann:2003:CRQ

[1531] Christian Blohmann. Covariant realization of quantum spaces as star products by Drinfeld twists. Journal of Mathematical Physics, 44(10): 4736–4755, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brunelli:2003:SEH

[1532] J. C. Brunelli, Ashok Das, and Ziemowit Popowicz. Supersymmetric extensions of the Harry Dym hierarchy. Journal of Mathematical Physics, 44(10):4756–4767, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassinelli:2003:POV

[1533] G. Cassinelli, E. De Vito, and A. Toigo. Positive operator valued measures covariant with respect to an irreducible representation. Journal of Mathematical Physics, 44(10):4768–4775, October 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dalton:2003:RHE

[1534] J. Dalton and B. D. Mestel. Renormalization for the Harper equation for quadratic irrationals. Journal of Mathematical Physics, 44(10):4776– 4783, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

ElBaz:2003:DCG

[1535] M. El Baz, A. El Hassouni, Y. Hassouni, and E. H. Zakkari. n = 3 differential calculus and gauge theory on a reduced quantum plane. Journal of Mathematical Physics, 44(10):4784–4794, October 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leitner:2003:IKS

[1536] Felipe Leitner. Imaginary Killing spinors in Lorentzian geometry. Journal of Mathematical Physics, 44(10):4795–4806, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lippert:2003:MMJ

[1537] Ross A. Lippert. A matrix model for the β-Jacobi ensemble. Journal of Mathematical Physics, 44(10):4807–4816, October 2003. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borstnik:2003:HGF

[1538] N. Mankoc Borstnik and H. B. Nielsen. How to generate families of spinors. Journal of Mathematical Physics, 44(10):4817–4827, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okumura:2003:IAS

[1539] Shoji Okumura. The indefinite anti-self-dual metrics and the Painlevé equations. Journal of Mathematical Physics, 44(10):4828–4838, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramos:2003:DWS

[1540] M. P. M. Ramos, E. G. L. R. Vaz, and J. Carot. Double warped space–times. Journal of Mathematical Physics, 44(10):4839–4865, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borchers:2003:FLT

[1541] H. J. Borchers and Ch. Borchers. The first law of thermodynamics and Dyson–Lieb’s perpetuum mobile. Journal of Mathematical Physics, 44 (10):4866–4867, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Govinder:2003:CTD

[1542] K. S. Govinder and M. Govender. Comment on “Time-dependent solution for a star immersed in a background radiation”, [H. Liu and Q. Zhang, J. Math. Phys. 43, 4904 (2002)]. Journal of Mathematical Physics, 44(10):4868–4870, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1166].

Aktosun:2003:RPR

[1543] Tuncay Aktosun and Vassilis G. Papanicolaou. Recovery of a potential from the ratio of reflection and transmission coefficients. Journal of Mathematical Physics, 44(11):4875–4883, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassam-Chenai:2003:HAI

[1544] Patrick Cassam-Chenaï and Frédéric Patras. The Hopf algebra of identical, fermionic particle systems — fundamental concepts and properties. Journal of Mathematical Physics, 44(11):4884–4906, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cresson:2003:SCS

[1545] Jacky Cresson. Scale calculus and the Schrödinger equation. Journal of Mathematical Physics, 44(11):4907–4938, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DellAntonio:2003:D

[1546] Gianfausto Dell’Antonio. On decoherence. Journal of Mathematical Physics, 44(11):4939–4956, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nunez:2003:SRE

[1547] J. Fernández Núñez, W. García Fuertes, and A. M. Perelomov. Some results on the eigenfunctions of the quantum trigonometric Calogero– Sutherland model related to the Lie algebra D_{4}. Journal of Mathematical Physics, 44(11):4957–4974, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klopp:2003:NAL

[1548] Frédéric Klopp and Shu Nakamura. A note on Anderson localization for the random hopping model. Journal of Mathematical Physics, 44 (11):4975–4980, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mardoyan:2003:GMK

[1549] Levon Mardoyan. The generalized MIC-Kepler system. Journal of Mathematical Physics, 44(11):4981–4987, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakano:2003:RMF

[1550] Fumihiko Nakano and Yuji Nomura. Random magnetic fields on line graphs. Journal of Mathematical Physics, 44(11):4988–5002, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Raginsky:2003:RND

[1551] Maxim Raginsky. Radon–Nikodym derivatives of quantum operations. Journal of Mathematical Physics, 44(11):5003–5020, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saad:2003:PEC

[1552] Nasser Saad, Richard L. Hall, and Attila B. von Keviczky. Perturbation expansions for a class of singular potentials. Journal of Mathematical Physics, 44(11):5021–5041, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakamoto:2003:EFA

[1553] Makoto Sakamoto and Shogo Tanimura. An extension of Fourier analysis for the n-torus in the magnetic field and its application to spectral analysis of the magnetic Laplacian. Journal of Mathematical Physics, 44 (11):5042–5069, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thirulogasanthar:2003:CVC

[1554] K. Thirulogasanthar and S. Twareque Ali. A class of vector coherent states defined over matrix domains. Journal of Mathematical Physics, 44(11):5070–5083, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:2003:PSM

[1555] S. Zhang and A. Vourdas. Phase space methods for particles on a circle. Journal of Mathematical Physics, 44(11):5084–5094, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loffelholz:2003:MST

[1556] J. Löffelholz, G. Morchio, and F. Strocchi. Mathematical structure of the temporal gauge in quantum electrodynamics. Journal of Mathematical Physics, 44(11):5095–5107, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anderson:2003:EST

[1557] E. Anderson, F. Dahia, J. E. Lidsey, and C. Romero. Embeddings in space–times sourced by scalar fields. Journal of Mathematical Physics, 44(11):5108–5119, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Balakin:2003:PPP

[1558] Alexander B. Balakin, Veronika R. Kurbanova, and Winfried Zimdahl. Parametric phenomena of the particle dynamics in a periodic gravitational wave field. Journal of Mathematical Physics, 44(11):5120–5140, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sharif:2003:SEM

[1559] M. Sharif. Symmetries of the energy–momentum tensor of spherically symmetric Lorentzian manifolds. Journal of Mathematical Physics, 44 (11):5141–5158, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:2003:FST

[1560] F. Calogero, J.-P. Françoise, and A. Guillot. A further solvable threebody problem in the plane. Journal of Mathematical Physics, 44(11): 5159–5165, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aldaya:2003:GPR

[1561] V. Aldaya, J. L. Jaramillo, and J. Guerrero. Gauge principle revisited: Towards a unification of space–time and internal gauge interactions. Journal of Mathematical Physics, 44(11):5166–5184, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:2003:ESW

[1562] Xiaofeng Liu and Houyu Jia. Existence of suitable weak solutions of complex Ginzburg–Landau equations and properties of the set of singular points. Journal of Mathematical Physics, 44(11):5185–5193, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anteneodo:2003:MNM

[1563] Celia Anteneodo and Constantino Tsallis. Multiplicative noise: A mechanism leading to nonextensive statistical mechanics. Journal of Mathematical Physics, 44(11):5194–5203, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:2003:GPB

[1564] M. Chen. On the generalized problem of the Boltzmann equation and the moment method in kinetic theory of gases. Journal of Mathematical Physics, 44(11):5204–5211, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Field:2003:SDS

[1565] Timothy R. Field and Robert J. A. Tough. Stochastic dynamics of the scattering amplitude generating K-distributed noise. Journal of Mathematical Physics, 44(11):5212–5223, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tutschka:2003:ISA

[1566] Christian Tutschka. Interacting squares in arbitrary external field. Journal of Mathematical Physics, 44(11):5224–5242, November 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adam:2003:RSE

[1567] C. Adam and J. Sánchez-Guillén. Ring-shaped exact Hopf solitons. Journal of Mathematical Physics, 44(11):5243–5249, November 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:2003:WHA

[1568] N. Aizawa and P. S. Isaac. Weak Hopf algebras corresponding to U_{q} [ sl_{n} ]. Journal of Mathematical Physics, 44(11):5250–5267, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alexandrov:2003:GFT

[1569] A. Alexandrov. Givental formula in terms of Virasoro operators. Journal of Mathematical Physics, 44(11):5268–5278, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Botero:2003:GPM

[1570] Alonso Botero. Geometric phase and modulus relations for probability amplitudes as functions on complex parameter spaces. Journal of Mathematical Physics, 44(11):5279–5295, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alcaras:2003:PIB

[1571] J. A. Castilho Alcarás, J. Tambergs, T. Krasta, J. Ruza, and O. Katkevicius. Plethysms and interacting boson models. Journal of Mathematical Physics, 44(11):5296–5319, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakrabarti:2003:CFD

[1572] A. Chakrabarti. Canonical factorization and diagonalization of Baxterized braid matrices: Explicit constructions and applications. Journal of Mathematical Physics, 44(11):5320–5349, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clarkson:2003:FPE

[1573] Peter A. Clarkson. The fourth Painlevé equation and associated special polynomials. Journal of Mathematical Physics, 44(11):5350–5374, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Edgar:2003:ECR

[1574] S. Brian Edgar. On effective constraints for the Riemann–Lanczos system of equations. Journal of Mathematical Physics, 44(11):5375–5385, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarizadeh:2003:HRC

[1575] M. A. Jafarizadeh and S. Behnia. Hierarchy of random chaotic maps with an invariant measure. Journal of Mathematical Physics, 44(11): 5386–5400, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:2003:TFR

[1576] Min Ho Lee. Tau functions and residues. Journal of Mathematical Physics, 44(11):5401–5409, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sadun:2003:TSI

[1577] Lorenzo Sadun. Tiling spaces are inverse limits. Journal of Mathematical Physics, 44(11):5410–5414, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scharnhorst:2003:GIE

[1578] K. Scharnhorst. A Grassmann integral equation. Journal of Mathematical Physics, 44(11):5415–5449, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yamane:2003:CEM

[1579] Hiroyuki Yamane. A central extension of U_{q}sl(2|2)^{(1)} and R-matrices with a new parameter. Journal of Mathematical Physics, 44(11):5450– 5455, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aldaya:2003:ECC

[1580] V. Aldaya and J. Guerrero. Erratum: “Canonical coherent states for the relativistic harmonic oscillator”, [J. Math. Phys. 36, 3191 (1995)]. Journal of Mathematical Physics, 44(11):5456, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [4].

Beneduci:2003:SRU

[1581] Roberto Beneduci and Giuseppe Nisticò. Sharp reconstruction of unsharp quantum observables. Journal of Mathematical Physics, 44(12):5461– 5473, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weigert:2003:LTC

[1582] Stefan Weigert and Paul Busch. Lüders theorem for coherent-state POVMs. Journal of Mathematical Physics, 44(12):5474–5486, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ahmedov:2003:CFB

[1583] H. Ahmedov and I. H. Duru. Casimir force between surfaces close to each other. Journal of Mathematical Physics, 44(12):5487–5503, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ayan:2003:NLM

[1584] Ali Ayan and O. Teoman Turgut. On the nonrelativistic Lee model. Journal of Mathematical Physics, 44(12):5504–5516, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hannabuss:2003:BCF

[1585] K. C. Hannabuss and M. Semplice. Boundary conformal fields and Tomita–Takesaki theory. Journal of Mathematical Physics, 44(12):5517– 5529, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leandre:2003:SWZ

[1586] Rémi Léandre. Stochastic Wess–Zumino–Novikov–Witten model on the torus. Journal of Mathematical Physics, 44(12):5530–5568, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morchio:2003:CDE

[1587] G. Morchio and F. Strocchi. Charge density and electric charge in quantum electrodynamics. Journal of Mathematical Physics, 44(12):5569– 5587, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Requardt:2003:GRG

[1588] Manfred Requardt. A geometric renormalization group in discrete quantum space–time. Journal of Mathematical Physics, 44(12):5588–5615, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chae:2003:ESG

[1589] Dongho Chae and Kwangseok Choe. Existence of the self-graviting Chern–Simons vortices. Journal of Mathematical Physics, 44(12):5616– 5636, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:2003:GSE

[1590] A. Das, A. DeBenedictis, and N. Tariq. General solutions of Einstein’s spherically symmetric gravitational equations with junction conditions. Journal of Mathematical Physics, 44(12):5637–5655, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:2003:DSD

[1591] Ya-Jun Gao. Double structures and double symmetries for the general symplectic gravity models. Journal of Mathematical Physics, 44(12): 5656–5663, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seahra:2003:UEF

[1592] Sanjeev S. Seahra and Paul S. Wesson. Universes encircling fivedimensional black holes. Journal of Mathematical Physics, 44(12):5664– 5680, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tian:2003:SVN

[1593] Guihua Tian and Zhao Zheng. The second variation of a null geodesic. Journal of Mathematical Physics, 44(12):5681–5687, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ay:2003:HEO

[1594] Ahmet Ay, Metin Gürses, and Kostyantyn Zheltukhin. Hamiltonian equations in R^{3}. Journal of Mathematical Physics, 44(12):5688–5705, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Caudrelier:2003:LPS

[1595] V. Caudrelier and E. Ragoucy. Lax pair and super-Yangian symmetry of the nonlinear super-Schrödinger equation. Journal of Mathematical Physics, 44(12):5706–5732, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choudhury:2003:ICT

[1596] S. Roy Choudhury. Integrability characteristics of two-dimensional generalizations of NLS type equations. Journal of Mathematical Physics, 44 (12):5733–5750, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grunberg:2003:CGM

[1597] Daniel B. Grünberg. N = 4 characters in Gepner models, orbits and elliptic genera. Journal of Mathematical Physics, 44(12):5751–5792, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2003:NLA

[1598] Fukui Guo and Yufeng Zhang. A new loop algebra and a corresponding integrable hierarchy, as well as its integrable coupling. Journal of Mathematical Physics, 44(12):5793–5803, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jia:2003:KEG

[1599] Qiuli Jia, Shengfan Zhou, and Fuqi Yin. Kolmogorov entropy of global attractor for dissipative lattice dynamical systems. Journal of Mathematical Physics, 44(12):5804–5810, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:2003:SSD

[1600] E. G. Kalnins, J. M. Kress, W. Miller, Jr., and P. Winternitz. Superintegrable systems in Darboux spaces. Journal of Mathematical Physics, 44(12):5811–5848, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamvissis:2003:SNS

[1601] Spyridon Kamvissis. Semiclassical nonlinear Schrödinger on the half line. Journal of Mathematical Physics, 44(12):5849–5868, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:2003:VSL

[1602] S. Y. Lou, C. Rogers, and W. K. Schief. Virasoro structure and localized excitations of the LKR system. Journal of Mathematical Physics, 44(12): 5869–5887, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boulahoual:2003:NGF

[1603] A. Boulahoual and M. B. Sedra. Noncommutative geometry framework and the Feynman’s proof of Maxwell equations. Journal of Mathematical Physics, 44(12):5888–5901, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deotto:2003:HSSa

[1604] E. Deotto, E. Gozzi, and D. Mauro. Hilbert space structure in classical mechanics. I. Journal of Mathematical Physics, 44(12):5902–5936, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deotto:2003:HSSb

[1605] E. Deotto, E. Gozzi, and D. Mauro. Hilbert space structure in classical mechanics. II. Journal of Mathematical Physics, 44(12):5937–5957, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Efroimsky:2003:GSB

[1606] Michael Efroimsky and Peter Goldreich. Gauge symmetry of the Nbody problem in the Hamilton–Jacobi approach. Journal of Mathematical Physics, 44(12):5958–5977, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:2003:VDM

[1607] Han-Ying Guo and Ke Wu. On variations in discrete mechanics and field theory. Journal of Mathematical Physics, 44(12):5978–6004, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fannes:2003:EGS

[1608] M. Fannes, B. Haegeman, and M. Mosonyi. Entropy growth of shiftinvariant states on a quantum spin chain. Journal of Mathematical Physics, 44(12):6005–6019, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marucho:2003:GBT

[1609] Marcelo Marucho and Gustavo A. Carri. On the generalized Borel transform and its application to the statistical mechanics of macromolecules. Journal of Mathematical Physics, 44(12):6020–6031, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tonel:2003:IGS

[1610] Arlei Prestes Tonel, Angela Foerster, Katrina Hibberd, and Jon Links. Integrable generalized spin ladder models based on the SU(1|3) and SU(3|1) algebras. Journal of Mathematical Physics, 44(12):6032–6039, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zuyev:2003:EPT

[1611] S. Zuyev and J. Quintanilla. Estimation of percolation thresholds via percolation in inhomogeneous media. Journal of Mathematical Physics, 44(12):6040–6046, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alekseevsky:2003:YMC

[1612] Dmitri V. Alekseevsky, Vicente Cortés, and Chandrashekar Devchand. Yang–Mills connections over manifolds with Grassmann structure. Journal of Mathematical Physics, 44(12):6047–6076, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersson:2003:FBM

[1613] E. Andersson. On the M-function and Borg–Marchenko theorems for vector-valued Sturm–Liouville equations. Journal of Mathematical Physics, 44(12):6077–6100, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anzaldo-Meneses:2003:GDS

[1614] A. Anzaldo-Meneses and F. Monroy-Pérez. Goursat distribution and sub-Riemannian structures. Journal of Mathematical Physics, 44(12): 6101–6111, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aubert:2003:IIU

[1615] S. Aubert and C. S. Lam. Invariant integration over the unitary group. Journal of Mathematical Physics, 44(12):6112–6131, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chae:2003:GES

[1616] Dongho Chae and Hyungjin Huh. Global existence for small initial data in the Born–Infeld equations. Journal of Mathematical Physics, 44(12): 6132–6139, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Edgar:2003:ONR

[1617] S. Brian Edgar and Ola Wingbrant. Old and new results for superenergy tensors from dimensionally dependent tensor identities. Journal of Mathematical Physics, 44(12):6140–6159, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kudryashov:2003:APE

[1618] Nikolai A. Kudryashov. Amalgamations of the Painlevé equations. Journal of Mathematical Physics, 44(12):6160–6178, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lievens:2003:MOP

[1619] S. Lievens and J. Van der Jeugt. Multivariable orthogonal polynomials associated with tensor products of the oscillator algebra ⌊(∞). Journal of Mathematical Physics, 44(12):6179–6194, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mazzolo:2003:PCL

[1620] Alain Mazzolo, Benoît Roesslinger, and Wilfried Gille. Properties of chord length distributions of nonconvex bodies. Journal of Mathematical Physics, 44(12):6195–6208, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Otte:2003:DIO

[1621] Peter Otte. Deficiency indices of operator polynomials in creation and annihilation operators. Journal of Mathematical Physics, 44(12):6209– 6222, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Torre:2003:HRW

[1622] C. G. Torre. The helically-reduced wave equation as a symmetric-positive system. Journal of Mathematical Physics, 44(12):6223–6232, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1950].

Almeida:2004:SSV

[1623] J. Almeida and J. Teixeira. The space of state vectors: A hyperfinite approach. Journal of Mathematical Physics, 45(1):1–20, January 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/p1_s1.

Baro:2004:DSP

[1624] M. Baro, H.-Chr. Kaiser, H. Neidhardt, and J. Rehberg. Dissipative Schrödinger–Poisson systems. Journal of Mathematical Physics, 45(1): 21–43, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i1/p21_s1.

Federbush:2004:PEQ

[1625] Paul Federbush. A polymer expansion for the quantum Heisenberg ferromagnet wave function. Journal of Mathematical Physics, 45(1):44– 50, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i1/p44_s1.

Geyler:2004:POA

[1626] V. A. Geyler and P. St’ovícek. On the Pauli operator for the Aharonov– Bohm effect with two solenoids. Journal of Mathematical Physics, 45(1):

51–75, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i1/p51_s1.

Giachetta:2004:NHO

[1627] G. Giachetta, L. Mangiarotti, and G. Sardanashvily. Nonadiabatic holonomy operators in classical and quantum completely integrable systems. Journal of Mathematical Physics, 45(1):76–86, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/p76_s1.

Hartmann:2004:WCS

[1628] Lorenz Hartmann and John R. Klauder. Weak coherent state path integrals. Journal of Mathematical Physics, 45(1):87–99, January 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/p87_s1.

Mitra:2004:EAF

[1629] Saibal Mitra. Expansions about free-fermion models. Journal of Mathematical Physics, 45(1):100–106, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i1/p100_s1.

Moretti:2004:QSS

[1630] P. Moretti, M. Lantieri, and L. Cianchi. Quantum study of the spin inversion. Journal of Mathematical Physics, 45(1):107–113, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i1/p107_s1.

Mukunda:2004:WDQ

[1631] N. Mukunda, Arvind, S. Chaturvedi, and R. Simon. Wigner distributions and quantum mechanics on Lie groups: The case of the regular representation. Journal of Mathematical Physics, 45(1):114–148, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i1/p114_s1.

Albeverio:2004:LAC

[1632] Sergio Albeverio, Maria Simonetta Bernabei, and Xian Yin Zhou. Lattice approximations and continuum limits of ϕ_{2}^{4}-quantum fields. Journal of Mathematical Physics, 45(1):149–178, January 2004. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i1/p149_s1.

Caicedo:2004:ELS

[1633] M. I. Caicedo and N. F. Svaiter. Effective Lagrangians for scalar fields and finite size effects in field theory. Journal of Mathematical Physics, 45(1):179–196, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i1/p179_s1.

Didenko:2004:FFD

[1634] V. E. Didenko and M. A. Vasiliev. Free field dynamics in the generalized AdS (super)space. Journal of Mathematical Physics, 45(1):197– 215, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i1/p197_s1.

Herrera-Aguilar:2004:STE

[1635] Alfredo Herrera-Aguilar and Oleg V. Kechkin. String theory extensions of Einstein–Maxwell fields: The stationary case. Journal of Mathematical Physics, 45(1):216–229, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i1/p216_s1.

Moretti:2004:HSR

[1636] Valter Moretti and Nicola Pinamonti. Holography and SL(2, R) symmetry in 2D Rindler space–time. Journal of Mathematical Physics, 45(1):230–256, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i1/p230_s1.

Moretti:2004:VAC

[1637] Valter Moretti and Nicola Pinamonti. Virasoro algebra with central charge c = 1 on the horizon of a two-dimensional-Rindler space–time. Journal of Mathematical Physics, 45(1):257–284, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/p257_s1.

Nieto:2004:SCB

[1638] J. A. Nieto. Searching for a connection between matroid theory and string theory. Journal of Mathematical Physics, 45(1):285–301, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i1/p285_s1.

Paturyan:2004:MAS

[1639] Vanush Paturyan and D. H. Tchrakian. Monopole-antimonopole solutions of the Skyrmed SU(2) Yang–Mills–Higgs model. Journal of Mathematical Physics, 45(1):302–309, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i1/p302_s1.

Dolan:2004:WLR

[1640] P. Dolan and A. Gerber. The Weyl–Lanczos relations and the fourdimensional Lanczos tensor wave equation and some symmetry generators. Journal of Mathematical Physics, 45(1):310–326, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/
p310_s1.

vandenBrink:2004:WAR

[1641] Alec Maassen van den Brink. WKB analysis of the Regge–Wheeler equation down in the frequency plane. Journal of Mathematical Physics, 45(1):327–335, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i1/p327_s1.

Dorodnitsyn:2004:CSL

[1642] Vladimir Dorodnitsyn, Roman Kozlov, and Pavel Winternitz. Continuous symmetries of Lagrangians and exact solutions of discrete equations. Journal of Mathematical Physics, 45(1):336–359, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/p336_s1.

Echeverria-Enriquez:2004:LHU

[1643] Arturo Echeverría-Enríquez, Carlos López, Jesús Marín-Solano, Miguel C. Muñoz-Lecanda, and Narciso Román-Roy. Lagrangian–Hamiltonian unified formalism for field theory. Journal of Mathematical Physics, 45(1):360–380, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i1/p360_s1.

Bogoyavlenskij:2004:UEM

[1644] Oleg I. Bogoyavlenskij. Unsteady equipartition MHD solutions. Journal of Mathematical Physics, 45(1):381–390, January 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i1/p381_s1.

Caspers:2004:CBS

[1645] W. J. Caspers, A. Wal, M. Labuz, M. Kuźma, and T. Lulek. Continuity of Bethe solutions with respect to chain length N and winding numbers λ_{l}. Journal of Mathematical Physics, 45(1):391–399, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/p391_s1.

Pechen:2004:QSE

[1646] A. N. Pechen. Quantum stochastic equation for a test particle interacting with a dilute Bose gas. Journal of Mathematical Physics, 45(1):400– 417, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i1/p400_s1.

Cassinelli:2004:POV

[1647] G. Cassinelli, E. De Vito, and A. Toigo. Positive operator valued measures covariant with respect to an Abelian group. Journal of Mathematical Physics, 45(1):418–433, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i1/p418_s1.

Crampe:2004:HSY

[1648] N. Crampé. Hopf structure of the Yangian Y (sl_{n}) in the Drinfel’d realization. Journal of Mathematical Physics, 45(1):434–447, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i1/p434_s1.

Ji:2004:LSA

[1649] Xiaoda Ji, Chunli Chen, Jin E. Zhang, and Yishen Li. Lie symmetry analysis and some new exact solutions of the Wu–Zhang equation. Journal of Mathematical Physics, 45(1):448–460, January 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i1/p448_s1.

Kim:2004:TFN

[1650] Hoil Kim and Chang-Yeong Lee. Theta functions on noncommutative T^{4}. Journal of Mathematical Physics, 45(1):461–474, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/p461_s1.

Kobayashi:2004:ISS

[1651] M. Kobayashi. Import of a set of solutions of the homogeneous Bloch equation. Journal of Mathematical Physics, 45(1):475–485, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/
p475_s1.

Kobayashi:2004:TGS

[1652] M. Kobayashi. Towards a general solution of the inhomogeneous Bloch equation. Journal of Mathematical Physics, 45(1):486–494, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/
p486_s1.

Nagel:2004:CEG

[1653] Bengt Nagel. Confluence expansions of the generalized hypergeometric function. Journal of Mathematical Physics, 45(1):495–508, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i1/p495_s1.

Su:2004:LDC

[1654] Yucai Su. Low dimensional cohomology of general conformal algebras gc_{N}. Journal of Mathematical Physics, 45(1):509–524, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i1/
p509_s1.

Su:2004:CDS

[1655] Yucai Su, Kaiming Zhao, and Linsheng Zhu. Classification of derivationsimple color algebras related to locally finite derivations. Journal of Mathematical Physics, 45(1):525–536, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i1/p525_s1.

vanSuijlekom:2004:NLC

[1656] W. D. van Suijlekom. The noncommutative Lorentzian cylinder as an isospectral deformation. Journal of Mathematical Physics, 45(1):537– 556, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i1/p537_s1.

Hou:2004:ERS

[1657] B. Y. Hou, R. Sasaki, and W.-L. Yang. Eigenvalues of Ruijsenaars– Schneider models associated with A_{n−1} root system in Bethe ansatz formalism. Journal of Mathematical Physics, 45(2):559–575, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i2/p559_s1.

Lin:2004:EPT

[1658] Wen-Wei Lin and Shih-Feng Shieh. Eigenvalue problems of a twodimensional Schrödinger operator with nonparabolic effective mass. Journal of Mathematical Physics, 45(2):576–584, February 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i2/p576_
s1.

Swanson:2004:TEN

[1659] Mark S. Swanson. Transition elements for a non-Hermitian quadratic Hamiltonian. Journal of Mathematical Physics, 45(2):585–601, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i2/p585_s1.

Cartas-Fuentevilla:2004:ECF

[1660] R. Cartas-Fuentevilla. The Euler characteristic and the first Chern number in the covariant phase space formulation of string theory. Journal of Mathematical Physics, 45(2):602–608, February 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i2/p602_s1.

Ko:2004:CFQ

[1661] Chul Ki Ko and Yong Moon Park. Construction of a family of quantum Ornstein–Uhlenbeck semigroups. Journal of Mathematical Physics, 45(2):609–627, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i2/p609_s1.

Neto:2004:EMM

[1662] A. Francisco Neto, P. A. Faria da Veiga, and M. O’Carroll. Existence of mesons and mass splitting in strong coupling lattice quantum chromodynamics. Journal of Mathematical Physics, 45(2):628–641, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i2/p628_s1.

Choi:2004:WPA

[1663] Jaedong Choi and Soon-Tae Hong. Warped product approach to universe with nonsmooth scale factor. Journal of Mathematical Physics, 45(2):642–651, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i2/p642_s1.

Ferrando:2004:CTS

[1664] Joan Josep Ferrando and Juan Antonio Sáez. On the classification of type D space–times. Journal of Mathematical Physics, 45(2):652–667, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i2/p652_s1.

Noundjeu:2004:EID

[1665] P. Noundjeu, N. Noutchegueme, and A. D. Rendall. Existence of initial data satisfying the constraints for the spherically symmetric Einstein– Vlasov–Maxwell system. Journal of Mathematical Physics, 45(2):668– 676, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i2/p668_s1.

Pareja:2004:RSD

[1666] M. J. Pareja. Relativistic stars in differential rotation: bounds on the dragging rate and on the rotational energy. Journal of Mathematical Physics, 45(2):677–695, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i2/p677_s1.

Brezhnev:2004:ESG

[1667] Yurii V. Brezhnev. Elliptic solitons and Gröbner bases. Journal of Mathematical Physics, 45(2):696–712, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i2/p696_s1.

Turhan:2004:CNC

[1668] Refik Turhan. A class of nonautonomous coupled KdV systems. Journal of Mathematical Physics, 45(2):713–724, February 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i2/p713_s1.

Garcia-Godinez:2004:GHJa

[1669] Patricia García-Godínez, Ezra Ted Newman, and Gilberto Silva-Ortigoza. 2-geometries and the Hamilton–Jacobi equation. Journal of Mathematical Physics, 45(2):725–735, February 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i2/p725_s1.

Vlad:2004:ABC

[1670] Marcel O. Vlad, Manuel G. Velarde, and John Ross. Analogies between colored Lévy noise and random channel approach to disordered kinetics. Journal of Mathematical Physics, 45(2):736–760, February 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i2/p736_
s1.

Becker:2004:IPW

[1671] Peter A. Becker. On the integration of products of Whittaker functions with respect to the second index. Journal of Mathematical Physics, 45(2):761–773, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i2/p761_s1.

Carron:2004:TNQ

[1672] G. Carron, P. Exner, and D. Krejcirík. Topologically nontrivial quantum layers. Journal of Mathematical Physics, 45(2):774–784, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i2/p774_s1.

Maillard:2004:SEO

[1673] Jean-Marie Maillard. Star exponentials for any ordering of the elements of the inhomogeneous symplectic Lie algebra. Journal of Mathematical Physics, 45(2):785–794, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i2/p785_s1.

Matsuno:2004:SNA

[1674] Yoshimasa Matsuno. A system of nonlinear algebraic equations connected with the multisoliton solution of the Benjamin–Ono equation. Journal of Mathematical Physics, 45(2):795–802, February 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i2/p795_
s1.

Panayotounakos:2004:REA

[1675] Dimitrios E. Panayotounakos and Nikolaos V. Sotiropoulos. On the reduction and the existence of approximate analytic solutions of some basic nonlinear ODEs in mathematical physics and nonlinear mechanics. Journal of Mathematical Physics, 45(2):803–826, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i2/p803_s1.

Dayi:2004:EHE

[1676] Ömer F. Dayi and Ahmed Jellal. Erratum: “Hall effect in noncommutative coordinates”, [J. Math. Phys. 43, 4592 (2002)]. Journal of Mathematical Physics, 45(2):827, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i2/p827_s1. See [1144].

Christandl:2004:SEA

[1677] Matthias Christandl and Andreas Winter. “squashed entanglement”: An additive entanglement measure. Journal of Mathematical Physics, 45(3): 829–840, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i3/p829_s1.

Chruscinski:2004:QMD

[1678] Dariusz Chruściński. Quantum mechanics of damped systems. II. Damping and parabolic potential barrier. Journal of Mathematical Physics, 45(3):841–854, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i3/p841_s1.

Dai:2004:GAS

[1679] Dao-Qing Dai, Bin Han, and Rong-Qing Jia. Galerkin analysis for Schrödinger equation by wavelets. Journal of Mathematical Physics, 45(3):855–869, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i3/p855_s1.

Datta:2004:RMA

[1680] Nilanjana Datta and Hervé Kunz. A random matrix approach to the crossover of energy-level statistics from Wigner to Poisson. Journal of Mathematical Physics, 45(3):870–886, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i3/p870_s1.

Dias:2004:TDT

[1681] Nuno Costa Dias and João Nuno Prata. Time dependent transformations in deformation quantization. Journal of Mathematical Physics, 45(3): 887–901, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i3/p887_s1.

Khrennikov:2004:CAQ

[1682] Andrei Khrennikov. Contextual approach to quantum mechanics and the theory of the fundamental prespace. Journal of Mathematical Physics, 45(3):902–921, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i3/p902_s1.

Martin:2004:BST

[1683] André Martin and Tai Tsun Wu. Bound states in two spatial dimensions in the noncentral case. Journal of Mathematical Physics, 45(3):922– 931, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i3/p922_s1.

Mostafazadeh:2004:POP

[1684] Ali Mostafazadeh. Pseudounitary operators and pseudounitary quantum dynamics. Journal of Mathematical Physics, 45(3):932–946, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i3/p932_s1.

Pengpan:2004:ERE

[1685] Teparksorn Pengpan. Equal rank embedding and its related construction to superconformal field theories. Journal of Mathematical Physics, 45(3): 947–958, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i3/p947_s1.

Svetlichny:2004:LSM

[1686] George Svetlichny. On linearity of separating multiparticle differential Schrödinger operators for identical particles. Journal of Mathematical Physics, 45(3):959–964, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i3/p959_s1.

Ahmedov:2004:CEI

[1687] H. Ahmedov and I. H. Duru. Casimir energy inside a cavity with triangular cross section. Journal of Mathematical Physics, 45(3):965–977, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i3/p965_s1.

Deng:2004:SSN

[1688] Hui Deng, Bo-Yu Hou, Kang-Jie Shi, Zhan-Ying Yang, and Rui-Hong Yue. Soliton solutions on noncommutative orbifold T^{2}/Z_{4}. Journal of Mathematical Physics, 45(3):978–995, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i3/p978_s1.

Melas:2004:BGG

[1689] Evangelos Melas. The BMS group and generalized gravitational instantons. Journal of Mathematical Physics, 45(3):996–1002, March 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i3/p996_
s1.

Gravel:2004:HSC

[1690] Simon Gravel. Hamiltonians separable in Cartesian coordinates and third-order integrals of motion. Journal of Mathematical Physics, 45 (3):1003–1019, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i3/p1003_s1.

Lou:2004:EAO

[1691] S. Y. Lou and Xiao-Yan Tang. Equations of arbitrary order invariant under the Kadomtsev–Petviashvili symmetry group. Journal of Mathematical Physics, 45(3):1020–1030, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i3/p1020_s1.

Saito:2004:HNF

[1692] Satoru Saito, Noriko Saitoh, and Katsuhiko Yoshida. Hyperelliptic Nambu flow associated with integrable maps. Journal of Mathematical Physics, 45(3):1031–1041, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i3/p1031_s1.

deLeon:2004:GIN

[1693] M. de León, D. Martín de Diego, and A. Santamaría-Merino. Geometric integrators and nonholonomic mechanics. Journal of Mathematical Physics, 45(3):1042–1064, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i3/p1042_s1.

Maciejewski:2004:NSP

[1694] Andrzej J. Maciejewski and Maria Przybylska. Nonintegrability of the Suslov problem. Journal of Mathematical Physics, 45(3):1065– 1078, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i3/p1065_s1.

McLenaghan:2004:ECT

[1695] Raymond G. McLenaghan, Roman G. Smirnov, and Dennis The. An extension of the classical theory of algebraic invariants to pseudo-Riemannian geometry and Hamiltonian mechanics. Journal of Mathematical Physics, 45(3):1079–1120, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i3/p1079_s1.

Ganikhodjaev:2004:PMC

[1696] Nasir Ganikhodjaev. The Potts model on Z^{d} with countable set of spin values. Journal of Mathematical Physics, 45(3):1121–1127, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i3/p1121_s1.

Maltsev:2004:QFT

[1697] Andrei Ya. Maltsev. Quasiperiodic functions theory and the superlattice potentials for a two-dimensional electron gas. Journal of Mathematical Physics, 45(3):1128–1149, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i3/p1128_s1.

Arik:2004:DTP

[1698] M. Arik and A. Peker-Dobie. Decomposition of tensor product representations of the unitary matrix quantum group SU_{q}(2). Journal of Mathematical Physics, 45(3):1150–1155, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i3/p1150_s1.

deMello:2004:IUR

[1699] R. O. de Mello and V. O. Rivelles. The irreducible unitary representations of the extended Poincaré group in (1 + 1) dimensions. Journal of Mathematical Physics, 45(3):1156–1167, March 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i3/p1156_s1.

Elizalde:2004:MAZ

[1700] E. Elizalde and M. Tierz. Multiplicative anomaly and zeta factorization. Journal of Mathematical Physics, 45(3):1168–1179, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i3/p1168_s1.

Gerry:2004:CGP

[1701] Christopher C. Gerry. On the Clebsch–Gordan problem for SU(1, 1): Coupling nonstandard representations. Journal of Mathematical Physics, 45(3):1180–1190, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i3/p1180_s1.

Lafortune:2004:SCA

[1702] S. Lafortune and A. Goriely. Singularity confinement and algebraic integrability. Journal of Mathematical Physics, 45(3):1191–1208, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i3/p1191_s1.

Weniger:2004:MPN

[1703] Ernst Joachim Weniger. Mathematical properties of a new Levin-type sequence transformation introduced by Cízek, Zamastil, and Skála. I. Algebraic theory. Journal of Mathematical Physics, 45(3):1209–1246, March 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i3/p1209_s1.

Baran:2004:SPR

[1704] S. A. Baran and G. A. Kerimov. On solvable potentials related to SO(2, 2). II. Natanzon potentials. Journal of Mathematical Physics, 45(4):1249–1266, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1249_s1.

Bruning:2004:SPS

[1705] J. Brüning, V. Geyler, and I. Lobanov. Spectral properties of a shortrange impurity in a quantum dot. Journal of Mathematical Physics, 45(4):1267–1290, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1267_s1.

Durr:2004:DTP

[1706] Detlef Dürr, Rodolfo Figari, and Alessandro Teta. Decoherence in a two-particle model. Journal of Mathematical Physics, 45(4):1291–1309, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i4/p1291_s1.

Giraud:2004:CSM

[1707] B. G. Giraud and A. Weiguny. Continuum singularities of a mean-field theory of collisions. Journal of Mathematical Physics, 45(4):1310–1331, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i4/p1310_s1.

Harrison:2004:QIG

[1708] J. M. Harrison and J. M. Robbins. Quantum indistinguishability from general representations of SU(2n). Journal of Mathematical Physics, 45(4):1332–1358, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1332_s1.

Linton:2004:BSCa

[1709] C. M. Linton and K. Ratcliffe. Bound states in coupled guides. I. Two dimensions. Journal of Mathematical Physics, 45(4):1359–1379, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i4/p1359_s1.

Linton:2004:BSCb

[1710] C. M. Linton and K. Ratcliffe. Bound states in coupled guides. II. Three dimensions. Journal of Mathematical Physics, 45(4):1380–1393, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i4/p1380_s1.

Mantoiu:2004:MWC

[1711] Marius Mantoiu and Radu Purice. The magnetic Weyl calculus. Journal of Mathematical Physics, 45(4):1394–1417, April 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i4/p1394_s1.

Rybkin:2004:ACN

[1712] Alexei Rybkin. On the absolutely continuous and negative discrete spectra of Schrödinger operators on the line with locally integrable globally square summable potentials. Journal of Mathematical Physics, 45(4):1418–1425, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1418_s1.

Saenz:2004:QMS

[1713] A. W. Sáenz. Quantum-mechanical scattering in exterior domains with impenetrable periodic boundaries and short-range potentials. Journal of Mathematical Physics, 45(4):1426–1446, April 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i4/p1426_s1. See er-

ratum [1949].

Franco:2004:SFT

[1714] Daniel H. T. Franco and Caio M. M. Polito. Supersymmetric fieldtheoretic models on a supermanifold. Journal of Mathematical Physics, 45(4):1447–1473, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1447_s1.

Terekidi:2004:VDR

[1715] Andrei G. Terekidi and Jurij W. Darewych. Variational derivation of relativistic fermion-antifermion wave equations in QED. Journal of Mathematical Physics, 45(4):1474–1496, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i4/p1474_s1.

Haesen:2004:IES

[1716] Stefan Haesen and Leopold Verstraelen. Ideally embedded space–times. Journal of Mathematical Physics, 45(4):1497–1510, April 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i4/p1497_

s1. See erratum [1955].

Park:2004:HSV

[1717] Eun Kyung Park and Pyung Seong Kwon. Homotopy structure of 5 d vacua. Journal of Mathematical Physics, 45(4):1511–1517, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i4/
p1511_s1.

Sharif:2004:CSP

[1718] M. Sharif. Classification of static plane symmetric space–times according to their matter collineations. Journal of Mathematical Physics, 45(4):1518–1531, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1518_s1.

Sharif:2004:SEM

[1719] M. Sharif. Symmetries of the energy-momentum tensor of cylindrically symmetric static space–times. Journal of Mathematical Physics, 45(4):1532–1560, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1532_s1.

Guha:2004:GFS

[1720] Partha Guha. Geodesic flow on (super-) Bott–Virasoro group and Harry Dym family. Journal of Mathematical Physics, 45(4):1561–1570, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i4/p1561_s1.

Golovko:2004:SFL

[1721] V. A. Golovko. Superfluidity of a Fermi liquid from the viewpoint of a hierarchy of equations for reduced density matrices. Journal of Mathematical Physics, 45(4):1571–1592, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i4/p1571_s1.

Kutoviy:2004:EGS

[1722] O. V. Kutoviy and A. L. Rebenko. Existence of Gibbs state for continuous gas with many-body interaction. Journal of Mathematical Physics, 45(4):1593–1605, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1593_s1.

Vandevenne:2004:ESB

[1723] Lieselot Vandevenne, André Verbeure, and Valentin A. Zagrebnov. Equilibrium states for the Bose gas. Journal of Mathematical Physics, 45(4):1606–1622, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i4/p1606_s1.

Aizawa:2004:NGS

[1724] N. Aizawa and R. Chakrabarti. Noncommutative geometry of super-Jordanian OSp_{h}(2/1) covariant quantum space. Journal of Mathematical Physics, 45(4):1623–1638, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i4/p1623_s1.

Guo:2004:EUG

[1725] Boling Guo and Yongqian Han. Existence and uniqueness of global solution of the Hasegawa–Mima equation. Journal of Mathematical Physics, 45(4):1639–1647, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i4/p1639_s1.

Khorunzhy:2004:EDL

[1726] O. Khorunzhy, M. Shcherbina, and V. Vengerovsky. Eigenvalue distribution of large weighted random graphs. Journal of Mathematical Physics, 45(4):1648–1672, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i4/p1648_s1.

Lim:2004:DAS

[1727] A. E. K. Lim and J. Carminati. The determination of all syzygies for the dependent polynomial invariants of the Riemann tensor. I. Pure Ricci and pure Weyl invariants. Journal of Mathematical Physics, 45(4):1673– 1698, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i4/p1673_s1.

Sacks:2004:IPC

[1728] Paul Sacks. An inverse problem in coupled mode theory. Journal of Mathematical Physics, 45(4):1699–1710, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i4/p1699_s1.

Stoyanov:2004:CTD

[1729] B. J. Stoyanov. Comment on “Two-dimensional, highly directive currents on large circular loops”, [J. Math. Phys. 41, 6130 (2000)]. Journal of Mathematical Physics, 45(4):1711–1712, April 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i4/p1711_s1. See [359].

Andrea:2004:SES

[1730] S. Andrea, A. Restuccia, and A. Sotomayor. Supersymmetric exact sequence, heat kernel and super Korteweg–de Vries hierarchy. Journal of Mathematical Physics, 45(5):1715–1729, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i5/p1715_s1.

Munteanu:2004:GFC

[1731] Florian Munteanu, Angel M. Rey, and Modesto Salgado. The Günther’s formalism in classical field theory: momentum map and reduction. Journal of Mathematical Physics, 45(5):1730–1751, May 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i5/p1730_s1.

Wu:2004:TNT

[1732] Wei Wu and Ling-An Wu. Two new types of quantum states generated via higher powers of Bogoliubov’s transformation. Journal of Mathematical Physics, 45(5):1752–1761, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i5/p1752_s1.

Ma:2004:ISB

[1733] Tian Ma and Shouhong Wang. Interior structural bifurcation and separation of 2D incompressible flows. Journal of Mathematical Physics, 45(5): 1762–1776, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1762_s1.

Kastler:2004:ESG

[1734] Daniel Kastler. On the exterior structure of graphs. Journal of Mathematical Physics, 45(5):1777–1786, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i5/p1777_s1.

Jencova:2004:GDD

[1735] Anna Jencová. Geodesic distances on density matrices. Journal of Mathematical Physics, 45(5):1787–1794, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i5/p1787_s1.

Watanabe:2004:LOS

[1736] Teppei Watanabe. Locality and orthomodular structure of compound systems. Journal of Mathematical Physics, 45(5):1795–1803, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i5/
p1795_s1.

Dadzie:2004:ASK

[1737] S. Kokou Dadzie and J. Gilbert Méolans. Anisotropic scattering kernel: Generalized and modified Maxwell boundary conditions. Journal of Mathematical Physics, 45(5):1804–1819, May 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i5/p1804_s1.

Bayramoglu:2004:DSN

[1738] Mehmet Bayramoğlu, Fatih Tasçi, and Djafar Zeynalov. On the discrete spectrum of non-self-adjoint Schrödinger differential equation with an operator coefficient. Journal of Mathematical Physics, 45(5):1820– 1825, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1820_s1.

Kondratiev:2004:DDC

[1739] Yuri G. Kondratiev, Alexei L. Rebenko, and Michael Röckner. On diffusion dynamics for continuous systems with singular superstable interaction. Journal of Mathematical Physics, 45(5):1826–1848, May 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i5/p1826_
s1.

Sanchez:2004:GGC

[1740] Alberto Sánchez, Alfredo Macías, and Hernando Quevedo. Generating Gowdy cosmological models. Journal of Mathematical Physics, 45(5): 1849–1858, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1849_s1.

Pletyukhov:2004:EPC

[1741] Mikhail Pletyukhov. On the extra phase correction to the semiclassical spin coherent-state propagator. Journal of Mathematical Physics, 45(5): 1859–1872, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1859_s1.

Gavrilov:2004:GFD

[1742] S. P. Gavrilov, D. M. Gitman, and A. A. Smirnov. Green functions of the Dirac equation with magnetic-solenoid field. Journal of Mathematical Physics, 45(5):1873–1886, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i5/p1873_s1.

Moses:2004:RIR

[1743] Harry E. Moses. The role of the irreducible representations of the Poincaré group in solving Maxwell’s equations. Journal of Mathematical Physics, 45(5):1887–1918, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i5/p1887_s1.

Louko:2004:GAO

[1744] Jorma Louko and Alberto Molgado. Group averaging in the (p, q) oscillator representation of SL(2, R). Journal of Mathematical Physics, 45(5): 1919–1943, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1919_s1.

Soloviev:2004:GFC

[1745] M. A. Soloviev. On the generalized function calculus for infrared and ultraviolet singular quantum fields. Journal of Mathematical Physics, 45(5):1944–1958, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i5/p1944_s1.

Berube:2004:ISQ

[1746] Josée Bérubé and Pavel Winternitz. Integrable and superintegrable quantum systems in a magnetic field. Journal of Mathematical Physics, 45(5): 1959–1973, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1959_s1.

Yamano:2004:SCM

[1747] Takuya Yamano. A statistical complexity measure with nonextensive entropy and quasi-multiplicativity. Journal of Mathematical Physics, 45(5): 1974–1987, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1974_s1.

Nash:2004:ABF

[1748] Patrick L. Nash and Rafael López-Mobilia. Associated Bessel functions and the discrete approximation of the free-particle time evolution operator in cylindrical coordinates. Journal of Mathematical Physics, 45(5): 1988–1993, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p1988_s1.

Pantoja:2004:CSD

[1749] N. R. Pantoja, H. Rago, and R. O. Rodríguez. Curvature singularity of the distributional Bañados, Teitelboim, and Zanelli black hole geometry. Journal of Mathematical Physics, 45(5):1994–2002, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i5/p1994_s1.

Forrester:2004:AGF

[1750] P. J. Forrester and N. E. Frankel. Applications and generalizations of Fisher–Hartwig asymptotics. Journal of Mathematical Physics, 45(5): 2003–2028, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i5/p2003_s1.

Pinaud:2004:AAS

[1751] Olivier Pinaud. Adiabatic approximation of the Schrödinger–Poisson system with a partial confinement: The stationary case. Journal of Mathematical Physics, 45(5):2029–2050, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i5/p2029_s1.

Guerrero:2004:GCR

[1752] J. Guerrero, J. L. Jaramillo, and V. Aldaya. Group-cohomology refinement to classify G-symplectic manifolds. Journal of Mathematical Physics, 45(5):2051–2072, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i5/p2051_s1.

Grillo:2004:IMT

[1753] S. Grillo and H. Montani. Integrable mixing of A_{n−1} type vertex models. Journal of Mathematical Physics, 45(5):2073–2089, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i5/p2073_s1.

Hudson:2004:DPI

[1754] R. L. Hudson and S. Pulmannová. Double product integrals and Enriquez quantization of Lie bialgebras I: The quasitriangular identities. Journal of Mathematical Physics, 45(5):2090–2105, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v45/i5/p2090_s1. See comment

[1755].

Enriquez:2004:CDP

[1755] B. Enriquez. Comment on “Double product integrals and Enriquez quantization of Lie bialgebras I: The quasitriangular identities”, Hudson and Pulmannová, [J. Math. Phys. 45, 2090 (2004)]. Journal of Mathematical Physics, 45(5):2106–2107, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i5/p2106_s1. See [1754].

Anco:2004:EPS

[1756] Stephen C. Anco and Roh S. Tung. Erratum: “Properties of the symplectic structure of general relativity for spatially bounded space–time regions”, [J. Math. Phys. 43, 3984 (2002)]. Journal of Mathematical Physics, 45(5):2108, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i5/p2108_s1. See [1105].

Anco:2004:ECH

[1757] Stephen C. Anco and Roh S. Tung. Erratum: “Covariant Hamiltonian boundary conditions in General Relativity for spatially bounded space– time regions”, [J. Math. Phys. 43, 5531 (2002)]. Journal of Mathematical Physics, 45(5):2109, May 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i5/p2109_s1. See [1204].

Fernandez-Jambrina:2004:NSF

[1758] L. Fernández-Jambrina and L. M. González-Romero. Nonsingular G_{2} stiff fluid cosmologies. Journal of Mathematical Physics, 45(6):2113– 2123, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i6/p2113_s1.

Paul:2004:ARS

[1759] Samir K. Paul and Siddhartha Sen. The appearance of the resolved singular hypersurface in the classical phase space of the Lie group SU(n). Journal of Mathematical Physics, 45(6):2124–2133, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2124_s1.

Yarotsky:2004:PGS

[1760] D. A. Yarotsky. Perturbations of ground states in weakly interacting quantum spin systems. Journal of Mathematical Physics, 45(6):2134– 2152, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i6/p2134_s1.

Burch:2004:HE

[1761] Aidan Burch. Histories electromagnetism. Journal of Mathematical Physics, 45(6):2153–2170, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i6/p2153_s1.

Renes:2004:SIC

[1762] Joseph M. Renes, Robin Blume-Kohout, A. J. Scott, and Carlton M. Caves. Symmetric informationally complete quantum measurements. Journal of Mathematical Physics, 45(6):2171–2180, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2171_s1.

Schurmann:2004:TWS

[1763] H. W. Schürmann and V. S. Serov. Traveling wave solutions of a generalized modified Kadomtsev–Petviashvili equation. Journal of Mathematical Physics, 45(6):2181–2187, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i6/p2181_s1.

Patera:2004:FGM

[1764] Jirí Patera, Edita Pelantová, and Milena Svobodová. Fine gradings of o(4, C). Journal of Mathematical Physics, 45(6):2188–2198, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/
p2188_s1.

Bruning:2004:RQF

[1765] E. Brüning and S. Nagamachi. Relativistic quantum field theory with a fundamental length. Journal of Mathematical Physics, 45(6):2199– 2231, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i6/p2199_s1.

Melamed:2004:PSG

[1766] T. Melamed. Phase-space Green’s functions for modeling time-harmonic scattering from smooth inhomogeneous objects. Journal of Mathematical Physics, 45(6):2232–2247, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i6/p2232_s1.

Roberts:2004:QLE

[1767] Mark D. Roberts. Quantizing the line element field. Journal of Mathematical Physics, 45(6):2248–2257, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i6/p2248_s1.

Alhaidari:2004:RDM

[1768] A. D. Alhaidari. The relativistic Dirac–Morse Green’s function. Journal of Mathematical Physics, 45(6):2258–2265, June 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2258_s1.

Calogero:2004:TISa

[1769] Francesco Calogero. A technique to identify solvable dynamical systems, and a solvable generalization of the goldfish many-body problem. Journal of Mathematical Physics, 45(6):2266–2279, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2266_s1.

Gungor:2004:SCK

[1770] F. Güngör, V. I. Lahno, and R. Z. Zhdanov. Symmetry classification of KdV-type nonlinear evolution equations. Journal of Mathematical Physics, 45(6):2280–2313, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i6/p2280_s1.

ElBaz:2004:DCC

[1771] M. El Baz, A. El Hassouni, Y. Hassouni, and E. H. Zakkari. d^{3} = 0, d^{2} = 0 differential calculi on certain noncommutative (super) spaces. Journal of Mathematical Physics, 45(6):2314–2322, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2314_s1.

Bena:2004:SLS

[1772] Ioana Bena, Avinash Khare, and Avadh Saxena. Soliton lattice and single soliton solutions of the associated Lamé and Lamé potentials. Journal of Mathematical Physics, 45(6):2323–2337, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i6/p2323_s1.

Sakovich:2004:BLT

[1773] S. Yu. Sakovich. On bosonic limits of two recent supersymmetric extensions of the Harry Dym hierarchy. Journal of Mathematical Physics, 45(6):2338–2342, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i6/p2338_s1.

Haesen:2004:CPS

[1774] Stefan Haesen and Leopold Verstraelen. Classification of the pseudosymmetric space–times. Journal of Mathematical Physics, 45(6):2343–2346, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i6/p2343_s1.

Ori:2004:CSF

[1775] Amos Ori and Eran Rosenthal. Calculation of the self force using the extended-object approach. Journal of Mathematical Physics, 45(6):2347– 2364, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i6/p2347_s1.

Ferapontov:2004:HRM

[1776] E. V. Ferapontov and K. R. Khusnutdinova. Hydrodynamic reductions of multidimensional dispersionless PDEs: The test for integrability. Journal of Mathematical Physics, 45(6):2365–2377, June 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2365_s1.

Freidel:2004:CFT

[1777] Laurent Freidel and Kirill Krasnov. 2D conformal field theories and holography. Journal of Mathematical Physics, 45(6):2378–2404, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i6/p2378_s1.

Coll:2004:LBS

[1778] Bartolomé Coll and Joan Josep Ferrando. On the Leibniz bracket, the Schouten bracket and the Laplacian. Journal of Mathematical Physics, 45(6):2405–2410, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i6/p2405_s1.

Howard:2004:SSO

[1779] I. A. Howard, A. Minguzzi, N. H. March, and M. P. Tosi. Slater sum for the one-dimensional sech^{2} potential in relation to the kinetic energy density. Journal of Mathematical Physics, 45(6):2411–2419, June 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2411_
s1.

Aksenova:2004:CCF

[1780] E. V. Aksenova, V. P. Romanov, and A. Yu. Valkov. Calculation of correlation function of the director fluctuations in cholesteric liquid crystals by WKB method. Journal of Mathematical Physics, 45(6):2420– 2446, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i6/p2420_s1.

Bullock:2004:CDQ

[1781] Stephen S. Bullock and Gavin K. Brennen. Canonical decompositions of n-qubit quantum computations and concurrence. Journal of Mathematical Physics, 45(6):2447–2467, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i6/p2447_s1.

Atoyan:2004:PCF

[1782] A. Atoyan and J. Patera. Properties of continuous Fourier extension of the discrete cosine transform and its multidimensional generalization. Journal of Mathematical Physics, 45(6):2468–2491, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2468_s1.

Nagi:2004:SPF

[1783] Jasbir Nagi. Superconformal primary fields on a graded Riemann sphere. Journal of Mathematical Physics, 45(6):2492–2514, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i6/p2492_s1.

Surya:2004:CST

[1784] Sumati Surya. Cyclic statistics in three dimensions. Journal of Mathematical Physics, 45(6):2515–2525, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i6/p2515_s1.

Carmeli:2004:PMO

[1785] Claudio Carmeli, Teiko Heinonen, and Alessandro Toigo. Position and momentum observables on R and on R^{3}. Journal of Mathematical Physics, 45(6):2526–2539, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i6/p2526_s1.

Garcia-Godinez:2004:GHJb

[1786] Patricia García-Godínez, Ezra Ted Newman, and Gilberto Silva-Ortigoza. 3-geometries and the Hamilton–Jacobi equation. Journal of Mathematical Physics, 45(7):2543–2559, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i7/p2543_s1.

McPhedran:2004:DAP

[1787] R. C. McPhedran, G. H. Smith, N. A. Nicorovici, and L. C. Botten. Distributive and analytic properties of lattice sums. Journal of Mathematical Physics, 45(7):2560–2578, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i7/p2560_s1.

Wreszinski:2004:EBO

[1788] Walter F. Wreszinski, Luiz A. Manzoni, and Oscar Bolina. Existence of the Bogoliubov S(g) operator for the (: ϕ^{4}: )_{2} quantum field theory.

Journal of Mathematical Physics, 45(7):2579–2593, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i7/p2579_s1.

Maerten:2004:KID

[1789] Daniel Maerten. Killing initial data revisited. Journal of Mathematical Physics, 45(7):2594–2599, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i7/p2594_s1.

Gregoratti:2004:QEC

[1790] M. Gregoratti and R. F. Werner. On quantum error-correction by classical feedback in discrete time. Journal of Mathematical Physics, 45(7): 2600–2612, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2600_s1.

Isozaki:2004:ISP

[1791] Hiroshi Isozaki, Hideo Nakazawa, and Gunther Uhlmann. Inverse scattering problem in nuclear physics — optical model. Journal of Mathematical Physics, 45(7):2613–2632, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i7/p2613_s1.

Brunelli:2004:IHD

[1792] J. C. Brunelli and Ashok Das. On an integrable hierarchy derived from the isentropic gas dynamics. Journal of Mathematical Physics, 45(7): 2633–2645, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2633_s1.

Brunelli:2004:DHD

[1793] J. C. Brunelli, Ashok Das, and Ziemowit Popowicz. Deformed Harry Dym and Hunter–Zheng equations. Journal of Mathematical Physics, 45(7):2646–2655, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i7/p2646_s1.

Weinzierl:2004:EAH

[1794] Stefan Weinzierl. Expansion around half-integer values, binomial sums, and inverse binomial sums. Journal of Mathematical Physics, 45(7):2656– 2673, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2656_s1.

Zamastil:2004:UAE

[1795] J. Zamastil, J. Cízek, M. Kalhous, L. Skála, and M. Simánek. The use of so(2, 1) algebra for the evaluation of atomic integrals: The study of two-electron atoms. Journal of Mathematical Physics, 45(7):2674– 2693, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2674_s1. See erratum [1947].

Thirulogasanthar:2004:TSC

[1796] K. Thirulogasanthar, Nasser Saad, and Attila B. von Keviczky. Temporally stable coherent states for a free magnetic Schrödinger operator. Journal of Mathematical Physics, 45(7):2694–2717, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i7/p2694_s1.

Yan:2004:LCH

[1797] Xiaodong Yan. On limits to convective heat transport at infinite Prandtl number with or without rotation. Journal of Mathematical Physics, 45(7):2718–2743, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i7/p2718_s1.

Debbasch:2004:DPC

[1798] F. Debbasch. A diffusion process in curved space–time. Journal of Mathematical Physics, 45(7):2744–2760, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i7/p2744_s1.

Rowe:2004:SHB

[1799] D. J. Rowe, P. S. Turner, and J. Repka. Spherical harmonics and basic coupling coefficients for the group SO(5) in an SO(3) basis. Journal of Mathematical Physics, 45(7):2761–2784, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i7/p2761_s1.

Cendra:2004:GCP

[1800] Hernán Cendra, Alberto Ibort, Manuel de León, and David Martín de Diego. A generalization of Chetaev’s principle for a class of higher order nonholonomic constraints. Journal of Mathematical Physics, 45(7):

2785–2801, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2785_s1.

Fairbairn:2004:SHS

[1801] Winston Fairbairn and Carlo Rovelli. Separable Hilbert space in loop quantum gravity. Journal of Mathematical Physics, 45(7):2802–2814, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i7/p2802_s1.

Almog:2004:LSS

[1802] Y. Almog. The loss of stability of surface superconductivity. Journal of Mathematical Physics, 45(7):2815–2832, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i7/p2815_s1.

Boumenir:2004:ECT

[1803] A. Boumenir and Vu Kim Tuan. Existence and construction of the transmutation operator. Journal of Mathematical Physics, 45(7):2833–2843, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i7/p2833_s1.

Salmela:2004:CPE

[1804] Antti Salmela. Covariant Poisson equation with compact Lie algebras. Journal of Mathematical Physics, 45(7):2844–2863, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i7/p2844_s1.

Guha:2004:AHB

[1805] Partha Guha. AKS hierarchy and bi-Hamiltonian geometry of Gelfand– Zakharevich type. Journal of Mathematical Physics, 45(7):2864–2884, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i7/p2864_s1.

Binz:2004:CUW

[1806] Ernst Binz, Reinhard Honegger, and Alfred Rieckers. Construction and uniqueness of the C^{∗}–Weyl algebra over a general pre-symplectic space. Journal of Mathematical Physics, 45(7):2885–2907, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i7/p2885_s1.

Rodrigues:2004:ADH

[1807] Waldyr A. Rodrigues. Algebraic and Dirac–Hestenes spinors and spinor fields. Journal of Mathematical Physics, 45(7):2908–2944, July 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i7/p2908_
s1.

Mosna:2004:BAD

[1808] Ricardo A. Mosna and Waldyr A. Rodrigues. The bundles of algebraic and Dirac–Hestenes spinor fields. Journal of Mathematical Physics, 45(7):2945–2966, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i7/p2945_s1.

Lu:2004:BCD

[1809] Lu Lu, Charles R. Doering, and Friedrich H. Busse. Bounds on convection driven by internal heating. Journal of Mathematical Physics, 45(7):2967– 2986, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2967_s1.

Chadan:2004:ESC

[1810] K. Chadan and R. Kobayashi. Erratum: “A sufficient condition for the existence of bound states in a potential”, [J. Math. Phys. 38, 4900 (1997)]. Journal of Mathematical Physics, 45(7):2987, July 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i7/p2987_

s1. See [5].

Dmitrasinovic:2004:ETQ

[1811] V. Dmitrasinović. Erratum: “Three-quark exchange operators, crossing matrices and Fierz transformations in SU(2) and SU(3)”, [J. Math. Phys. 42, 991 (2001)]. Journal of Mathematical Physics, 45(7):2988– 2989, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2988_s1. See [556].

Zeitsch:2004:SOR

[1812] Peter J. Zeitsch. Symmetry operators for Riemann’s method. Journal of Mathematical Physics, 45(8):2993–3000, August 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i8/p2993_s1.

Zeitsch:2004:RFG

[1813] Peter J. Zeitsch. Riemann functions and the group E(1, 1). Journal of Mathematical Physics, 45(8):3001–3018, August 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3001_s1.

Aubert:2004:IGT

[1814] S. Aubert and C. S. Lam. Invariant and group theoretical integrations over the U(n) group. Journal of Mathematical Physics, 45(8):3019– 3039, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i8/p3019_s1.

Lesage:2004:STG

[1815] Frédéric Lesage and Jørgen Rasmussen. SLE-type growth processes and the Yang–Lee singularity. Journal of Mathematical Physics, 45(8):3040– 3048, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i8/p3040_s1.

Popovych:2004:GCD

[1816] Roman O. Popovych, Nataliya M. Ivanova, and Homayoon Eshraghi. Group classification of (1 + 1)-dimensional Schrödinger equations with potentials and power nonlinearities. Journal of Mathematical Physics, 45(8):3049–3057, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i8/p3049_s1.

Katori:2004:SMV

[1817] Makoto Katori and Hideki Tanemura. Symmetry of matrix-valued stochastic processes and noncolliding diffusion particle systems. Journal of Mathematical Physics, 45(8):3058–3085, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3058_s1.

Hall:2004:RBS

[1818] Richard L. Hall, Wolfgang Lucha, and Franz F. Schöberl. Relativistic N-boson systems bound by pair potentials V (r_{ij}) = g(r_{ij}^{2}). Journal of Mathematical Physics, 45(8):3086–3094, August 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3086_s1.

Alvarez:2004:LCD

[1819] Gabriel Álvarez. Langer–Cherry derivation of the multi-instanton expansion for the symmetric double well. Journal of Mathematical Physics, 45(8):3095–3108, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i8/p3095_s1.

Ullrich:2004:RQF

[1820] Peter Ullrich. On the restriction of quantum fields to a lightlike surface. Journal of Mathematical Physics, 45(8):3109–3145, August 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3109_
s1.

Kojima:2004:DRE

[1821] Takeo Kojima and Hitoshi Konno. The Drinfeld realization of the elliptic quantum group B_{q,λ}(A_{2}^{(2)}). Journal of Mathematical Physics, 45(8):3146–3179, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i8/p3146_s1.

Galapon:2004:STA

[1822] Eric A. Galapon. Shouldn’t there be an antithesis to quantization? Journal of Mathematical Physics, 45(8):3180–3215, August 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3180_

s1. See erratum [1952].

Toft:2004:MUE

[1823] Joachim Toft. Minimization under entropy conditions, with applications in lower bound problems. Journal of Mathematical Physics, 45(8):3216– 3227, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i8/p3216_s1.

Colladay:2004:YMI

[1824] Don Colladay and Patrick McDonald. Yang–Mills instantons with Lorentz violation. Journal of Mathematical Physics, 45(8):3228–3238, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i8/p3228_s1.

Grundland:2004:PIS

[1825] A. M. Grundland and A. J. Hariton. Partially invariant solutions of models obtained from the Nambu–Goto action. Journal of Mathematical Physics, 45(8):3239–3265, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v45/i8/p3239_s1.

Li:2004:QDC

[1826] Fang Li and Yao-Zhong Zhang. Quantum doubles from a class of noncocommutative weak Hopf algebras. Journal of Mathematical Physics, 45(8):3266–3281, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i8/p3266_s1.

Bonetto:2004:ASM

[1827] F. Bonetto, P. Falco, and A. Giuliani. Analyticity of the SRB measure of a lattice of coupled Anosov diffeomorphisms of the torus. Journal of Mathematical Physics, 45(8):3282–3309, August 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3282_s1.

Gruber:2004:RFQ

[1828] Christian Gruber and Valentin A. Zagrebnov. Random-field quantum spherical ferroelectric model. Journal of Mathematical Physics, 45(8):3310–3321, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i8/p3310_s1.

Rao:2004:PCM

[1829] S. Eswara Rao. Partial classification of modules for Lie-algebra of diffeomorphisms of d-dimensional torus. Journal of Mathematical Physics, 45(8):3322–3333, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i8/p3322_s1.

Jiang:2004:AEE

[1830] Tongsong Jiang. An algorithm for eigenvalues and eigenvectors of quaternion matrices in quaternionic quantum mechanics. Journal of Mathematical Physics, 45(8):3334–3338, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i8/p3334_s1.

Naber:2004:TFS

[1831] Mark Naber. Time fractional Schrödinger equation. Journal of Mathematical Physics, 45(8):3339–3352, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i8/p3339_s1.

Ciccariello:2004:LFT

[1832] Salvino Ciccariello. The Lerch function and the thermodynamical functions of the ideal quantum gases. Journal of Mathematical Physics, 45(8):3353–3361, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i8/p3353_s1.

Gabardo:2004:BLP

[1833] Jean-Pierre Gabardo and Deguang Han. Balian–Low phenomenon for subspace Gabor frames. Journal of Mathematical Physics, 45(8):3362– 3378, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i8/p3362_s1.

Pareja:2004:BDR

[1834] M. J. Pareja. Bounds on the dragging rate of slowly and differentially rotating relativistic stars. Journal of Mathematical Physics, 45(8):3379– 3398, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i8/p3379_s1.

Lehnert:2004:DTW

[1835] Ralf Lehnert. Dirac theory within the Standard-Model Extension. Journal of Mathematical Physics, 45(8):3399–3412, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3399_s1.

Brodlie:2004:NCT

[1836] Alastair Brodlie. Nonlinear canonical transformations in classical and quantum mechanics. Journal of Mathematical Physics, 45(8):3413– 3431, August 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i8/p3413_s1.

Daoud:2004:ARB

[1837] M. Daoud. Analytic representations based on su(3) coherent states and Robertson intelligent states. Journal of Mathematical Physics, 45 (9):3435–3443, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i9/p3435_s1.

Lenzi:2004:SFN

[1838] E. K. Lenzi, R. S. Mendes, Kwok Sau Fa, L. R. da Silva, and L. S. Lucena. Solutions for a fractional nonlinear diffusion equation: Spatial time dependent diffusion coefficient and external forces. Journal of Mathematical Physics, 45(9):3444–3452, September 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i9/p3444_s1.

Bruneau:2004:HEA

[1839] Vincent Bruneau and Georgi D. Raikov. High energy asymptotics of the magnetic spectral shift function. Journal of Mathematical Physics, 45 (9):3453–3461, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i9/p3453_s1.

Li:2004:GPA

[1840] Hongyan Li, Shengfan Zhou, and Fuqi Yin. Global periodic attractor for strongly damped wave equations with time-periodic driving force. Journal of Mathematical Physics, 45(9):3462–3467, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i9/p3462_s1.

Boillat:2004:EMW

[1841] Guy Boillat and Tommaso Ruggeri. Energy momentum, wave velocities and characteristic shocks in Euler’s variational equations with application to the Born–Infeld theory. Journal of Mathematical Physics, 45(9):3468– 3478, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i9/p3468_s1.

Ma:2004:IBV

[1842] Shixiang Ma and Shijin Ding. On the initial boundary value problem for a shallow water equation. Journal of Mathematical Physics, 45(9):3479– 3497, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i9/p3479_s1.

G:2004:AAF

[1843] J. Fernando Barbero G., Guillermo A. Mena Marugán, and Eduardo J. S. Villaseñor. Asymptotic analysis of field commutators for Einstein–Rosen gravitational waves. Journal of Mathematical Physics, 45(9):3498–3532, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i9/p3498_s1.

Amaya-Tapia:2004:IRO

[1844] A. Amaya-Tapia, G. Gasaneo, S. Ovchinnikov, J. H. Macek, and S. Y. Larsen. Integral representation of one-dimensional three particle scattering for δ function interactions. Journal of Mathematical Physics, 45 (9):3533–3545, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i9/p3533_s1.

Flores:2004:GSR

[1845] J. L. Flores, Y. Parra, and U. Percoco. On the general structure of Ricci collineations for type B warped space–times. Journal of Mathematical Physics, 45(9):3546–3557, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i9/p3546_s1.

Morosi:2004:TT

[1846] Carlo Morosi and Livio Pizzocchero. On a theorem by Treves. Journal of Mathematical Physics, 45(9):3558–3564, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i9/p3558_s1.

Pule:2004:CPB

[1847] Joseph V. Pulé and Valentin A. Zagrebnov. The canonical perfect Bose gas in Casimir boxes. Journal of Mathematical Physics, 45(9):3565–3583, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i9/p3565_s1.

Stremler:2004:EPM

[1848] Mark A. Stremler. Evaluation of phase-modulated lattice sums. Journal of Mathematical Physics, 45(9):3584–3589, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i9/p3584_s1.

Leach:2004:RCM

[1849] P. G. L. Leach and M. C. Nucci. Reduction of the classical MICZ–Kepler problem to a two-dimensional linear isotropic harmonic oscillator. Journal of Mathematical Physics, 45(9):3590–3604, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i9/p3590_s1.

Jorgensen:2004:EPD

[1850] Palle E. T. Jorgensen. The existence problem for dynamics of dissipative systems in quantum probability. Journal of Mathematical Physics, 45 (9):3605–3619, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i9/p3605_s1.

DAriano:2004:ECQ

[1851] Giacomo Mauro D’Ariano. Extremal covariant quantum operations and positive operator valued measures. Journal of Mathematical Physics, 45 (9):3620–3635, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i9/p3620_s1.

Guhr:2004:DSH

[1852] Thomas Guhr and Heiner Kohler. Derivation of the supersymmetric Harish-Chandra integral for UOSp(k_{1}/2k_{2}). Journal of Mathematical Physics, 45(9):3636–3644, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i9/p3636_s1.

Ganikhodjaev:2004:IMT

[1853] N. N. Ganikhodjaev, C. H. Pah, and M. R. B. Wahiddin. An Ising model with three competing interactions on a Cayley tree. Journal of Mathematical Physics, 45(9):3645–3658, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i9/p3645_s1.

Turko:2004:FSU

[1854] Ludwik Turko. Finite size universe or perfect squash problem. Journal of Mathematical Physics, 45(9):3659–3675, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i9/p3659_s1.

Nenciu:2004:SLM

[1855] Gheorghe Nenciu and Vania Sordoni. Semiclassical limit for multistate Klein–Gordon systems: almost invariant subspaces, and scattering theory. Journal of Mathematical Physics, 45(9):3676–3696, September 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i9/
p3676_s1.

Majid:2004:IQG

[1856] Shahn Majid. Introduction: Quantum groups and noncommutative geometry. Journal of Mathematical Physics, 45(10):3701–3702, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i10/p3701_s1.

Bonneau:2004:QGD

[1857] Philippe Bonneau, Murray Gerstenhaber, Anthony Giaquinto, and Daniel Sternheimer. Quantum groups and deformation quantization: Explicit approaches and implicit aspects. Journal of Mathematical Physics, 45(10):3703–3741, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i10/p3703_s1.

Soibelman:2004:MSN

[1858] Yan Soibelman. Mirror symmetry and noncommutative geometry of A_{∞}- categories. Journal of Mathematical Physics, 45(10):3742–3757, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i10/p3742_s1.

Ebrahimi-Fard:2004:IRL

[1859] Kurusch Ebrahimi-Fard, Li Guo, and Dirk Kreimer. Integrable renormalization I: The ladder case. Journal of Mathematical Physics, 45(10):3758– 3769, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i10/p3758_s1.

Goldin:2004:FSN

[1860] Gerald A. Goldin and Shahn Majid. On the Fock space for nonrelativistic anyon fields and braided tensor products. Journal of Mathematical Physics, 45(10):3770–3787, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i10/p3770_s1.

Dobrev:2004:CIE

[1861] V. K. Dobrev and S. T. Petrov. q-conformal invariant equations and q-plane wave solutions. Journal of Mathematical Physics, 45(10):3788– 3799, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i10/p3788_s1.

Bershtein:2004:DPS

[1862] Olga Bershtein. Degenerate principal series of quantum Harish–Chandra modules. Journal of Mathematical Physics, 45(10):3800–3827, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i10/p3800_s1.

Gateva-Ivanova:2004:CAS

[1863] Tatiana Gateva-Ivanova. A combinatorial approach to the set-theoretic solutions of the Yang–Baxter equation. Journal of Mathematical Physics, 45(10):3828–3858, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i10/p3828_s1.

Joyce:2004:RLA

[1864] William P. Joyce. Recoupling Lie algebra and universal ω-algebra. Journal of Mathematical Physics, 45(10):3859–3877, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i10/p3859_s1.

Dixon:2004:DAF

[1865] Geoffrey Dixon. Division algebras: Family replication. Journal of Mathematical Physics, 45(10):3878–3882, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i10/p3878_s1.

Akrami:2004:BCC

[1866] S. E. Akrami and S. Majid. Braided cyclic cocycles and nonassociative geometry. Journal of Mathematical Physics, 45(10):3883–3911, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i10/p3883_s1.

Panaite:2004:QHA

[1867] Florin Panaite and Freddy Van Oystaeyen. Quasi-Hopf algebras and representations of octonions and other quasialgebras. Journal of Mathematical Physics, 45(10):3912–3929, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i10/p3912_s1.

Street:2004:FMP

[1868] Ross Street. Frobenius monads and pseudomonoids. Journal of Mathematical Physics, 45(10):3930–3948, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i10/p3930_s1.

Girelli:2004:HGT

[1869] Florian Girelli and Hendryk Pfeiffer. Higher gauge theory-differential versus integral formulation. Journal of Mathematical Physics, 45(10):3949– 3971, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i10/p3949_s1.

Bressler:2004:HMS

[1870] Paul Bressler and Yan Soibelman. Homological mirror symmetry, deformation quantization and noncommutative geometry. Journal of Mathematical Physics, 45(10):3972–3982, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i10/p3972_s1.

Maldonado-Mercado:2004:DQS

[1871] Cesar Maldonado-Mercado. Deformation quantization in singular spaces. Journal of Mathematical Physics, 45(10):3983–3994, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i10/
p3983_s1.

Androulidakis:2004:CEP

[1872] Iakovos Androulidakis. Classification of extensions of principal bundles and transitive Lie groupoids with prescribed kernel and cokernel. Journal of Mathematical Physics, 45(10):3995–4012, October 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i10/p3995_s1.

Adam:2004:HMS

[1873] C. Adam. Hopf maps as static solutions of the complex eikonal equation. Journal of Mathematical Physics, 45(11):4017–4024, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i11/p4017_s1.

Khamraev:2004:AMC

[1874] Murod Khamraev and Farruh Mukhamedov. On P -adic λ-model on the Cayley tree. Journal of Mathematical Physics, 45(11):4025–4034, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i11/p4025_s1.

Rudolph:2004:EQS

[1875] Oliver Rudolph. On extremal quantum states of composite systems with fixed marginals. Journal of Mathematical Physics, 45(11):4035–4041, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i11/p4035_s1.

Calogero:2004:NLR

[1876] Simone Calogero. The Newtonian limit of the relativistic Boltzmann equation. Journal of Mathematical Physics, 45(11):4042–4052, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i11/p4042_s1.

Blennow:2004:ESS

[1877] Mattias Blennow and Tommy Ohlsson. Exact series solution to the two flavor neutrino oscillation problem in matter. Journal of Mathematical Physics, 45(11):4053–4063, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i11/p4053_s1.

Desheng:2004:ABS

[1878] Yang Desheng. The asymptotic behavior of the stochastic Ginzburg– Landau equation with multiplicative noise. Journal of Mathematical Physics, 45(11):4064–4076, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i11/p4064_s1.

Levi:2004:UCD

[1879] Decio Levi, Piergiulio Tempesta, and Pavel Winternitz. Umbral calculus, difference equations and the discrete Schrödinger equation. Journal of Mathematical Physics, 45(11):4077–4105, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/p4077_s1.

Ren:2004:SSR

[1880] Xiaofeng Ren and Juncheng Wei. Stability of spot and ring solutions of the diblock copolymer equation. Journal of Mathematical Physics, 45 (11):4106–4133, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4106_s1.

Caponio:2004:AST

[1881] Erasmo Caponio and Antonio Masiello. The Avez–Seifert theorem for the relativistic Lorentz force equation. Journal of Mathematical Physics, 45(11):4134–4140, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4134_s1.

Smirnov:2004:CJI

[1882] Roman G. Smirnov and Jin Yue. Covariants, joint invariants and the problem of equivalence in the invariant theory of Killing tensors defined in pseudo-Riemannian spaces of constant curvature. Journal of Mathematical Physics, 45(11):4141–4163, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i11/p4141_s1.

Richard:2004:PDO

[1883] Serge Richard and Rafael Tiedra de Aldecoa. On perturbations of Dirac operators with variable magnetic field of constant direction. Journal of Mathematical Physics, 45(11):4164–4173, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/p4164_s1.

Catto:2004:EBR

[1884] Isabelle Catto, Pavel Exner, and Christian Hainzl. Enhanced binding revisited for a spinless particle in nonrelativistic QED. Journal of Mathematical Physics, 45(11):4174–4185, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i11/p4174_s1.

Gallo:2004:TDR

[1885] Emanuel Gallo. Two-dimensional Riemannian and Lorentzian geometries from second-order ODE’s. Journal of Mathematical Physics, 45 (11):4186–4190, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4186_s1.

Qadir:2004:CMC

[1886] Asghar Qadir and K. Saifullah. Comments on matter collineations of plane symmetric, cylindrically symmetric, and spherically symmetric space–times. Journal of Mathematical Physics, 45(11):4191–4192, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i11/p4191_s1.

Sharif:2004:ASE

[1887] M. Sharif. Addendum: Symmetries of the energy-momentum tensor. Journal of Mathematical Physics, 45(11):4193–4195, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/
p4193_s1.

Arias:2004:ASQ

[1888] Alvaro Arias and Stan Gudder. Almost sharp quantum effects. Journal of Mathematical Physics, 45(11):4196–4206, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/p4196_s1.

Arik:2004:IQG

[1889] M. Arik and A. Baykal. Inhomogeneous quantum groups for particle algebras. Journal of Mathematical Physics, 45(11):4207–4217, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i11/p4207_s1.

Jiang:2004:AQL

[1890] Tongsong Jiang. An algorithm for quaternionic linear equations in quaternionic quantum theory. Journal of Mathematical Physics, 45 (11):4218–4222, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4218_s1.

Gravanis:2004:IHD

[1891] Elias Gravanis and Steven Willison. Intersecting hypersurfaces in dimensionally continued topological density gravitation. Journal of Mathematical Physics, 45(11):4223–4238, November 2004. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i11/p4223_s1.

Beckermann:2004:AUB

[1892] B. Beckermann, A. Martínez-Finkelshtein, E. A. Rakhmanov, and F. Wielonsky. Asymptotic upper bounds for the entropy of orthogonal polynomials in the Szegő class. Journal of Mathematical Physics, 45 (11):4239–4254, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4239_s1.

Chabanov:2004:RCS

[1893] V. M. Chabanov. Recovering the M-channel Sturm–Liouville operator from M + 1 spectra. Journal of Mathematical Physics, 45(11):4255– 4260, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i11/p4255_s1.

Rothstein:2004:SWP

[1894] Ivan Rothstein. Semiclassical wave-packet scattering in one and two dimensions. Journal of Mathematical Physics, 45(11):4261–4270, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i11/p4261_s1.

Abdel-Aty:2004:ASE

[1895] Mahmoud Abdel-Aty and Abdel-Shafy F. Obada. Analytic solution for entangled two-qubit in a cavity field. Journal of Mathematical Physics, 45(11):4271–4281, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4271_s1.

Corona-Corona:2004:WJS

[1896] Gulmaro Corona-Corona. A Wronskian of Jost solutions. Journal of Mathematical Physics, 45(11):4282–4287, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/p4282_s1.

Sankovich:2004:BCS

[1897] D. P. Sankovich. On the Bose condensation in some model of a nonideal Bose gas. Journal of Mathematical Physics, 45(11):4288–4300, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i11/p4288_s1.

Yang:2004:SDR

[1898] W.-L. Yang and R. Sasaki. Solution of the dual reflection equation for A_{n−1}^{(1)} solid-on-solid model. Journal of Mathematical Physics, 45 (11):4301–4309, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4301_s1.

Karatsuba:2004:ASO

[1899] Ekatherina A. Karatsuba. Approximation of sums of oscillating summands in certain physical problems. Journal of Mathematical Physics, 45(11):4310–4321, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i11/p4310_s1.

Schlingemann:2004:ESC

[1900] Dirk-M. Schlingemann. Error syndrome calculation for graph codes on a one-way quantum computer: Towards a quantum memory. Journal of Mathematical Physics, 45(11):4322–4333, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/p4322_s1.

Noakes:2004:NNL

[1901] Lyle Noakes. Non-null Lie quadratics in E^{3}. Journal of Mathematical Physics, 45(11):4334–4351, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i11/p4334_s1.

Klimcik:2004:D

[1902] Ctirad Klimcík. q-deformation of z → (αz+β)/(γz+δ). Journal of Mathematical Physics, 45(11):4352–4359, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i11/p4352_s1.

Yu:2004:GCS

[1903] Hongjun Yu. Global classical solution of the Vlasov–Maxwell–Landau system near Maxwellians. Journal of Mathematical Physics, 45(11):4360– 4376, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i11/p4360_s1.

Jauslin:2004:UQL

[1904] H. R. Jauslin, O. Sapin, S. Guérin, and W. F. Wreszinski. Upper quantum Lyapunov exponent and parametric oscillators. Journal of Mathematical Physics, 45(11):4377–4385, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i11/p4377_s1.

Jordan:2004:FSN

[1905] Andrew N. Jordan, Eugene V. Sukhorukov, and Sebastian Pilgram. Fluctuation statistics in networks: A stochastic path integral approach. Journal of Mathematical Physics, 45(11):4386–4417, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/
p4386_s1.

Znojil:2004:FSS

[1906] Miloslav Znojil. Fragile P T -symmetry in a solvable model. Journal of Mathematical Physics, 45(11):4418–4430, November 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i11/p4418_s1.

Chiribella:2004:ECP

[1907] Giulio Chiribella and Giacomo Mauro D’Ariano. Extremal covariant positive operator valued measures. Journal of Mathematical Physics, 45 (12):4435–4447, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i12/p4435_s1.

Vignolo:2004:NGL

[1908] Stefano Vignolo and Roberto Cianci. A new geometrical look at gravity coupled with Yang–Mills fields. Journal of Mathematical Physics, 45 (12):4448–4463, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i12/p4448_s1.

deBuyl:2004:HKM

[1909] Sophie de Buyl and Christiane Schomblond. Hyperbolic Kac Moody algebras and Einstein billiards. Journal of Mathematical Physics, 45 (12):4464–4492, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i12/p4464_s1.

Borg:2004:LTR

[1910] J. L. Borg and J. V. Pulé. Lifshits tails for random smooth magnetic vortices. Journal of Mathematical Physics, 45(12):4493–4505, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4493_s1.

Maccari:2004:CSR

[1911] Attilio Maccari. Coherent solutions for relativistic vectorial fields. Journal of Mathematical Physics, 45(12):4506–4514, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/p4506_s1.

Napsuciale:2004:ASP

[1912] Mauro Napsuciale and Mariana Kirchbach. Avoiding superluminal propagation of higher spin waves via projectors onto W^{2} invariant subspaces. Journal of Mathematical Physics, 45(12):4515–4523, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/
p4515_s1.

Svaiter:2004:FSE

[1913] N. F. Svaiter. Finite size effects in thermal field theory. Journal of Mathematical Physics, 45(12):4524–4538, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i12/p4524_s1.

Song:2004:CG

[1914] Jun S. Song and Yun S. Song. On a conjecture of Givental. Journal of Mathematical Physics, 45(12):4539–4550, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/p4539_s1.

Giblin:2004:STS

[1915] John T. Giblin and Andrew D. Hwang. Space–time slices and surfaces of revolution. Journal of Mathematical Physics, 45(12):4551–4559, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4551_s1.

Falomir:2004:RSF

[1916] H. Falomir, M. A. Muschietti, and P. A. G. Pisani. On the resolvent and spectral functions of a second order differential operator with a regular singularity. Journal of Mathematical Physics, 45(12):4560–4577, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4560_s1.

Skrypnyk:2004:DLA

[1917] T. Skrypnyk. Deformations of loop algebras and integrable systems: hierarchies of integrable equations. Journal of Mathematical Physics, 45 (12):4578–4595, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i12/p4578_s1.

Majid:2004:MQR

[1918] S. Majid and E. Raineri. Moduli of quantum Riemannian geometries on ≤ 4 points. Journal of Mathematical Physics, 45(12):4596–4627, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4596_s1.

Yoshimura:2004:NNN

[1919] Kazuyuki Yoshimura and Ken Umeno. Nonintegrability of nonhomogeneous nonlinear lattices. Journal of Mathematical Physics, 45(12):4628– 4639, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i12/p4628_s1.

Bahns:2004:ICN

[1920] Dorothea Bahns. The invariant charges of the Nambu–Goto string and canonical quantization. Journal of Mathematical Physics, 45(12):4640– 4660, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i12/p4640_s1.

Calogero:2004:TISb

[1921] Francesco Calogero. A technique to identify solvable dynamical systems, and another solvable extension of the goldfish many-body problem. Journal of Mathematical Physics, 45(12):4661–4678, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/p4661_s1.

Girelli:2004:ABD

[1922] Florian Girelli, Thomas Krajewski, and Pierre Martinetti. An algebraic Birkhoff decomposition for the continuous renormalization group. Journal of Mathematical Physics, 45(12):4679–4697, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/p4679_s1.

Cardoso:2004:AQF

[1923] Vitor Cardoso, José Natário, and Ricardo Schiappa. Asymptotic quasinormal frequencies for black holes in nonasymptotically flat space–times. Journal of Mathematical Physics, 45(12):4698–4713, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/
p4698_s1.

Niestegge:2004:CSR

[1924] Gerd Niestegge. Composite systems and the role of the complex numbers in quantum mechanics. Journal of Mathematical Physics, 45(12):4714– 4725, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i12/p4714_s1. See erratum [1953].

Dehnen:2004:SSS

[1925] H. Dehnen and V. D. Ivashchuk. On spherically symmetric solutions with horizon in model with multicomponent anisotropic fluid. Journal of Mathematical Physics, 45(12):4726–4736, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/p4726_s1.

Morosi:2004:TTA

[1926] Carlo Morosi and Livio Pizzocchero. On the Treves theorem for the Ablowitz–Kaup–Newell–Segur equation. Journal of Mathematical Physics, 45(12):4737–4753, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i12/p4737_s1.

Grandou:2004:AIH

[1927] T. Grandou. Angular intricacies in hot gauge field theories. Journal of Mathematical Physics, 45(12):4754–4763, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/p4754_s1.

Sultana:2004:CKH

[1928] J. Sultana and C. C. Dyer. Conformal Killing horizons. Journal of Mathematical Physics, 45(12):4764–4776, December 2004. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i12/p4764_s1.

Zhang:2004:PEN

[1929] Keke Zhang, Xinhao Liao, and Paul Earnshaw. The Poincaré equation: A new polynomial and its unusual properties. Journal of Mathematical Physics, 45(12):4777–4790, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i12/p4777_s1.

Raab:2004:ANQ

[1930] A. Raab. An approach to nonstandard quantum mechanics. Journal of Mathematical Physics, 45(12):4791–4809, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v45/i12/p4791_s1.

Buchholz:2004:SQS

[1931] Detlev Buchholz and Stephen J. Summers. Stable quantum systems in anti-de Sitter space: Causality, independence, and spectral properties. Journal of Mathematical Physics, 45(12):4810–4831, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4810_s1.

Auberson:2004:MDD

[1932] G. Auberson, G. Mahoux, S. M. Roy, and Virendra Singh. Marginal distributions in (2N)-dimensional phase space and the quantum (N + 1) marginal theorem. Journal of Mathematical Physics, 45(12):4832– 4854, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i12/p4832_s1.

Briand:2004:MST

[1933] Emmanuel Briand, Jean-Gabriel Luque, Jean-Yves Thibon, and Frank Verstraete. The moduli space of three-qutrit states. Journal of Mathematical Physics, 45(12):4855–4867, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i12/p4855_s1.

Furuichi:2004:FPT

[1934] S. Furuichi, K. Yanagi, and K. Kuriyama. Fundamental properties of Tsallis relative entropy. Journal of Mathematical Physics, 45(12):4868– 4877, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i12/p4868_s1.

Kim:2004:CBG

[1935] Jeong-Ah Kim and Dong-Uy Shin. Crystal bases and generalized Lascoux–Leclerc–Thibon (LLT) algorithm for the quantum affine algebra U_{q}(C_{n}(1)). Journal of Mathematical Physics, 45(12):4878–4895, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4878_s1.

Tanimoto:2004:HAL

[1936] Masayuki Tanimoto. Harmonic analysis of linear fields on the nilgeometric cosmological model. Journal of Mathematical Physics, 45(12):4896– 4919, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/

jmapaq/v45/i12/p4896_s1. See publisher’s note [1946].

Halvorson:2004:RPA

[1937] Hans Halvorson. Remote preparation of arbitrary ensembles and quantum bit commitment. Journal of Mathematical Physics, 45(12):4920– 4931, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i12/p4920_s1.

Contopoulos:2004:CCT

[1938] G. Contopoulos and M. Harsoula. Chaos in the case of two fixed black holes. Journal of Mathematical Physics, 45(12):4932–4956, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4932_s1.

Esposti:2004:FFU

[1939] Mirko Degli Esposti and Andreas Knauf. On the form factor for the unitary group. Journal of Mathematical Physics, 45(12):4957–4979, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v45/i12/p4957_s1.

Ritter:2004:DNT

[1940] William Gordon Ritter. Description of noncommutative theories and matrix models by Wightman functions. Journal of Mathematical Physics, 45(12):4980–5002, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i12/p4980_s1.

Iochum:2004:CIA

[1941] Bruno Iochum, Thomas Schücker, and Christoph Stephan. On a classification of irreducible almost commutative geometries. Journal of Mathematical Physics, 45(12):5003–5041, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i12/p5003_s1.

Mestel:2004:GMR

[1942] B. D. Mestel and A. H. Osbaldestin. Golden mean renormalization for a generalized Harper equation: The Ketoja–Satija orchid. Journal of Mathematical Physics, 45(12):5042–5075, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i12/p5042_s1.

Giacobbe:2004:MQR

[1943] A. Giacobbe, R. H. Cushman, D. A. Sadovskií, and B. I. Zhilinskií. Monodromy of the quantum 1: 1: 2 resonant swing spring. Journal of Mathematical Physics, 45(12):5076–5100, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i12/p5076_s1.

Sparber:2004:ESA

[1944] C. Sparber and P. Markowich. Erratum: “Semiclassical asymptotics for the Maxwell–Dirac system”, [J. Math. Phys. 44, 4555 (2003)]. Journal of Mathematical Physics, 45(12):5101, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v45/i12/p5101_s1. See [1520].

Ruskai:2005:EIQ

[1945] Mary Beth Ruskai. Erratum: “Inequalities for quantum entropy: A review with conditions for equality”, [J. Math. Phys. 43, 4358 (2002)]. Journal of Mathematical Physics, 46(1):019901, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/

p019901_s1. See [1130].

Tanimoto:2005:PNH

[1946] Masayuki Tanimoto. Publisher’s note: “Harmonic analysis of linear fields on the nilgeometric cosmological model”, [J. Math. Phys. 45, 4896 (2004)]. Journal of Mathematical Physics, 46(1):019902, January 2005.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/

p019902_s1. See [1936].

Zamastil:2005:EUA

[1947] J. Zamastil, J. Cízek, M. Kalhous, L. Skála, and M. Simánek. Erratum: “The use of so(2, 1) algebra for the evaluation of atomic integrals: The study of two-electron atoms”, [J. Math. Phys. 45, 2674 (2004)]. Journal of Mathematical Physics, 46(3):039901, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p039901_s1. See [1795].

Hall:2005:ECS

[1948] Brian C. Hall and Jeffrey J. Mitchell. Erratum: “Coherent states on spheres”, [J. Math. Phys. 43, 1211 (2002)]. Journal of Mathematical Physics, 46(5):059901, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p059901_s1. See [941, 1954].

Saenz:2005:EQM

[1949] A. W. Sáenz. Erratum: “Quantum-mechanical scattering in exterior domains with impenetrable boundaries and short-range potentials”, [J. Math. Phys. 45, 1426 (2004)]. Journal of Mathematical Physics, 46(6): 069901, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p069901_s1. See [1713].

Torre:2007:EHR

[1950] C. G. Torre. Erratum: “The helically-reduced wave equation as a symmetric-positive system”, [J. Math. Phys. 44, 6223 (2003)]. Journal of Mathematical Physics, 48(2):029901, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i2/p029901_s1. See [1622].

Caves:2008:EUQ

[1951] Carlton M. Caves, Christopher A. Fuchs, and Rüdiger Schack. Erratum: “Unknown quantum states: The quantum de Finetti representation”, [J. Math. Phys. 44, 4537 (2002)]. Journal of Mathematical Physics, 49(1): 019902, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p019902_s1. See [1140].

Galapon:2008:EST

[1952] Eric. A. Galapon. Erratum: “Shouldn’t there be an antithesis to quantization?”, [J. Math. Phys. 45, 3180 (2004)]. Journal of Mathematical Physics, 49(9):099902, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i9/p099902_s1. See [1822].

Niestegge:2010:ECS

[1953] Gerd Niestegge. Erratum: “Composite systems and the role of the complex numbers in quantum mechanics”, [J. Math. Phys. 45, 4714 (2004)]. Journal of Mathematical Physics, 51(2):029901, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/

p029901_s1. See [1924].

Rabeie:2011:CCS

[1954] A. Rabeie and Z. Jalilian. Comment on “Coherent states on spheres”, [J. Math. Phys. 43, 1211 (2002)]. Journal of Mathematical Physics, 52(8): 084101, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p084101_s1. See [941, 1948].

Haesen:2014:EIE

[1955] Stefan Haesen and Leopold Verstraelen. Erratum: “Ideally embedded space–times” [J. Math. Phys. 45, 1497 (2004)]. Journal of Mathematical Physics, 55(4):049902, April 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1716].

Taillebois:2020:CQT

[1956] E. R. F. Taillebois and A. T. Avelar. Comment on “Quantum time and spatial localization: An analysis of the Hegerfeldt paradox” [j. math. phys. 41, 6093 (2000)]. Journal of Mathematical Physics, 61(10):104101, October 2020. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [357].